A Bibliography of Publications in IEEE Computer, 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 December 2021 Version 1.92

Title word cross-reference

+ [455]. $1 [2363]. 2 [1831]. 3 [5, 100, 229, 352, 409, 820, 1129, 1157, 1214, 1299, 1407, 1487, 1831, 2183, 2361, 2417]. $50 [238]. × [1750].

1 [1133, 1562]. 10th [1028]. 12 [786, 930, 2320]. 1968 [2736]. 1992 [2151, 2286, 2316]. 1993 [2664]. 19th [341]. 19th-Century [341]. 1B [1972].

2 [1590]. 2.0 [55, 121, 196]. 2.5D [2267]. 2010 [88, 199]. 2011 [317]. 2012 [461, 648]. 2013 [1135]. 2014 [1023, 1061, 1731]. 2015 [1237, 1292, 1302, 1338, 1340, 1358, 1377, 1391, 1413, 1427, 1527, 1604, 2136]. 2016 [1613, 1627, 1629, 1715, 1761, 1912, 1942, 2274, 2275, 2389]. 2017 [1993, 2191]. 2018 [2471, 2480, 2587, 2695]. 2019 [2591, 2606, 2617, 2619, 2656, 2665, 2681, 2710, 2712, 2747, 2858, 2860, 2864, 2922, 3079]. 2020 [3112]. 2022 [1470]. 21st [261, 341, 465, 569, 812, 958, 1375, 1694, 1940, 1957, 2476]. 21st-Century [341]. 25 [2433]. 29th [1362].

3.0 [15]. 30th [1638]. 31 [1633]. 31st [2022]. 39th [1569].

4.0 [2249]. 4G [268].

5 [60]. 5.0 [2653, 3067]. 50 [2716]. 50th [2153]. 5G [2935].

64-Bit [1070]. 6500 [2232].

754 [3108].

802 [1223]. 8086 [2231].

99 [2788].

A-Changin [1576]. above [1298]. Abroad [757]. Absence [1730]. Abstractions [2129]. Abundance [1753]. Abundant [1750]. Abundant-Data [1750]. Abuse [2572]. Abuse-Forming [2572]. Academia [241]. Academic [261, 465, 665]. Academy [1405]. Accelerate [2592]. Accelerated [2625]. Accelerating [2872]. Acceptance [1872]. Access [108, 153, 593, 637, 1139, 1159, 1264, 1344, 1448, 1455, 1486, 1513, 1550, 1593, 2134, 2262, 2329, 2336]. Accessibility [310, 858, 2939]. Accessible [869]. Accessing [269]. Accidents [1511]. Accomplishments [2003]. Accountability [76, 893, 2648]. Accounting [282]. Accredited [2815]. Accurate [681, 2674]. Achieve [783, 2025, 2057, 2315, 2397]. Achieving [241, 980]. ACM [373, 1846, 2150, 2185, 2771]. ACM/IEEE [2771]. Across [72, 395, 2299, 2519, 2560, 2840]. Act [694, 3110]. Action [192, 417, 615, 2334, 2436]. Actionable [2842]. Actions [3067]. Activating [1647]. Active [1126, 3061]. Activities [301, 1960]. Activity [185, 684, 763, 964, 1691, 2158, 2596, 2998]. Activity-Driven [185]. Activity-Enriched [684]. Actually [97]. Ad [1177]. Ada [2377]. Adapting [90, 456, 1279, 1401, 1753]. Adaptive [680, 1447]. Adding [108, 2500]. Address [141, 375, 643, 919, 1253, 1630, 2016, 2366, 2678, 2980]. Addressable [2537]. Addresses [1070]. Addressing [35, 629, 784, 930]. Adopter [271]. Adopters [270]. Adoption [387, 962, 1042, 2438]. Adopts [2447]. Advanced [1160]. Advances [1437, 2261, 2892, 2956]. Advancing [155, 194, 718]. Advantage [1366]. Adventures [1275]. Adversarial [3005]. Advertisement [971, 996, 1025, 1028, 1034, 1053, 1055, 1080, 1081, 1083, 1108, 1110, 1124, 1139, 1140, 1146, 1149, 1169, 1171, 1173, 1182, 1186, 1194, 1228, 1245, 1256, 1288, 1291, 1292, 1295, 1296, 1302, 1306, 1320, 1321, 1323– 1325, 1327, 1328, 1331, 1338, 1340, 1342, 1344, 1348, 1354, 1358, 1362, 1366, 1367, 1370, 1374, 1377, 1379, 1383, 1467, 1495, 1498, 1509, 1525, 1531, 1545, 1560, 1569, 1572, 1577, 1629, 1633, 1635, 1636, 1642, 1648, 1688, 1690, 1693, 1698, 1704, 1715, 1720, 1724, 1735, 1737, 1739, 1752]. Advertisement [1440, 1455, 1456, 1471, 1486–1488, 1503, 1513– 1515, 1527, 1546, 1547, 1558, 1567, 1573, 1574, 1580, 1582, 1602, 1603, 1608, 1610, 1613, 1617, 1621, 1627, 1638, 1645, 1664, 1666, 1674, 1675, 1679, 1683, 1700, 1702, 1707, 1722, 1741, 1742, 1767–1769, 1774, 1798, 1799, 1813, 1817, 1825, 1827, 1838, 1852, 1855, 1856, 1865, 1881, 1883– 1885, 1890, 1906, 1912, 1917, 1930, 1933, 1935, 1936, 1941, 1942, 1948, 1961, 1963, 1967, 1970, 1980, 2024, 2025, 2045, 2055, 2087, 2110, 2112, 2113]. Advertisement [1264, 1267, 1271, 1404, 1407, 1800, 1828, 1857, 1904, 1913, 1919, 1951, 1981, 1982, 2053, 2145, 2150, 2154, 2164, 2167, 2169, 2172– 2174, 2192, 2214]. Advertising [568, 2763]. Advice [3000]. Advise [3000]. Aerial [100]. Affect [1020]. Affective [2260, 2612]. Affordability [551]. Affordable [1641]. Africa [1261]. Afrofuturism [1877]. after [1009]. Aftershock [2725]. Again [751, 2188, 3042]. Against [436, 506, 527, 538, 1315, 2267]. Agarwal [1680]. Age [76, 1033, 1386, 1943, 2575, 2642, 2696, 2737, 2759, 3098]. Agency [535]. Agenda [990, 1927, 2383]. Agent [2236]. Agent-Based [2236]. Agents [2668]. Aggregate [1649]. Agile [111, 1899]. Agility [91, 1313]. Agitation [3063]. Ago [4, 21, 42, 77, 96, 115, 131, 152, 171, 189, 212, 228, 247, 267, 287, 307, 325, 363, 385, 408, 428, 452, 539, 561, 583, 606, 649, 672, 695, 716, 739, 758, 779, 809, 833, 857, 878, 901, 926, 949, 1037, 1091, 1117, 1148, 1179, 1207, 1235, 1262, 1300, 1333, 1372, 1408, 1435, 1461, 1493, 1521, 1552, 1611, 1646, 1681, 1712, 1746, 1777, 1806, 1833, 1889, 1921, 1953, 1986, 2029, 2061, 2091, 2119, 2202, 2229, 2257, 2349, 2378, 2406, 2434, 2464, 2492, 2516, 2538, 2561, 2613, 2641, 3018, 3053]. Ago [59, 475, 496, 519, 628, 975, 1005, 1062, 1583, 1863, 2716, 2736, 2753, 2779, 2808, 2838, 2865, 2896, 2928, 2958, 2990, 3085]. Ahead [906, 2155, 2419]. Ahmed [1725]. AI [2102, 2416, 2694, 2698, 2762, 3005, 3037, 3067]. AI-Enabled [3005]. Air [1562, 1590, 2718]. Aisle [2299]. Akron [1240]. Akyildiz [296]. Alan [581]. Alexa [2392]. Algebra [1615]. Algorithm [632, 1642, 1643]. Algorithmic [437]. Algorithms [525, 1612, 1616, 2037, 2787]. Alice [1581]. Aligning [1337]. Alleviating [2638]. Allocation [889]. Allowed [386]. Also [2746]. Alt [2247]. Alt-News [2247]. Alterations [938]. Altered [486]. Alternative [291, 637]. Alternatives [391]. Altitude [2323]. Alzheimer [302]. Am [688, 817]. Ambient [219, 2295]. Ambient-Powered [2295]. Amdahl [530, 1011, 1589, 1808, 2289]. Amendment [2015]. America [462, 694]. American [1789]. among [2210]. Amplify [2423]. Analog [425]. Analyses [1403]. Analysis [69, 159, 421, 528, 720, 842, 844, 884, 910, 1134, 1184, 1518, 1837, 2264]. Analytic [636]. Analytica [2601, 2672]. Analytics [457, 663, 673, 740, 810, 822, 846, 882–887, 916, 1267, 1281, 1320, 1615, 1653, 1779, 1786, 1870, 1871, 1982, 1989, 2297, 2414, 2595, 2758]. Analyze [2097, 2235]. Analyzing [46, 817, 1017, 1163]. Anant [1680]. Ancient [69]. and16 [42]. Andrew [1003, 1205]. Android [1185, 1215, 1926, 1966, 2570]. Anil [1609]. Animal [1213, 3026]. Animation [1387]. Annals [2791, 2825, 2826, 2847, 2868, 2913, 2932, 2961, 3045]. Annealing [2904]. Annealing-Based [2904]. Anniversary [440, 2153]. Annotation [1087]. Announces [106, 731]. Annual [1569, 2650]. Anomalies [2387]. Anonymous [1839]. Answer [620]. Answering [2159]. Antarctic [1275]. Anthropomorphic [2693]. Antifragility [2785]. Antivaccination [2443]. Antivirus [277, 1721]. Ants [575]. Anxiety [1845]. Any [387, 446, 2242]. Anything [446, 1192]. Anytime [822]. Anywhere [610, 822]. AOSD [24]. Apache [670]. API [902]. Apoptotic [235]. App [381, 403, 1104, 1563, 2414, 2523]. Appearance [359]. Applicants [773]. Application [315, 364, 386, 506, 939, 954, 1046, 1068, 1069, 1183, 1381, 1651, 2301, 2356, 3066]. Application-Aware [3066]. Application-Oriented [1651]. Application-Tailored [954]. Applications [47, 121, 213, 219, 271, 310, 328, 431, 460, 543, 545, 554, 590, 635, 659, 727, 766, 1160, 1204, 1276, 1530, 1564, 1569, 1612, 1721, 1805, 1897, 2068, 2211, 2563, 2789, 2792, 2849, 2866, 2909, 2971, 3022]. Applied [2704, 2875]. Applying [345, 764, 900, 987, 1526, 1814, 3095]. Apprehend [320]. Approach [6, 117, 316, 395, 540, 747, 786, 818, 954, 976, 1076, 1965, 2002, 2274, 2436, 2447, 2718, 2757, 2818, 2842, 2848, 2901]. Approaches [302, 455, 810, 2100, 2133, 2815]. Appropriated [110]. Approximate [2293]. Apps [405, 2570]. AppStore [1187]. Aragorn [2470]. Archaeological [351, 354]. Archaeology [349, 353]. Arching [290]. Architect [199, 1899]. Architecting [275, 1754, 2540]. Architectural [184, 1246, 1620, 2100]. Architecturally [47]. Architecture [83, 123, 256, 316, 616, 724, 849, 1011, 1381, 1654, 1678, 2074, 2236, 2244, 2721, 2845, 3012, 3068]. Architecture-Based [316]. Architectures [635, 1612, 1618, 2353, 2625, 2661]. Archiving [79]. Arduino [1035]. Area [814, 837]. Arithmetic [2031]. Arm [288, 1557].

Armed [641]. Arms [133]. Army [2479]. Arpanet [3021]. Array [1273]. Arrhythmia [3064]. Art [604, 1234, 1710, 1811, 2121, 3093]. Articles [2755, 2782, 2784, 2786, 2806, 2807, 2821, 2822, 2839, 2851, 2853, 2870, 2897, 2903, 2907, 2936, 2969, 2971, 2973, 2975, 2997, 3025, 3041, 3072, 3084]. Artifacts [2519]. Artificial [2276, 2692, 2696, 2974, 3031, 3064, 3065, 3095]. Arts [931]. Ascends [1135]. Asia [586]. Ask [560]. Asking [1532]. Aspect [24]. Aspect-Oriented [24]. Aspects [206, 322, 1620]. Assess [1786]. Assessing [2238, 3069]. Assessment [1479, 2047, 2239, 2240]. Assessments [1284, 2616]. Assets [1107]. Assimilation [1714]. Assistant [2181, 2544]. Assistants [448, 2474]. Assisted [354, 589, 1157, 2760]. Assistive [1125, 2541]. Association [2972]. Assurance [155, 157, 338, 3006]. Assuring [3006]. Astonishing [2404]. Astroinformatics [1272]. Astronomy [1268, 1276]. Astrophysics [1719]. Asynchronous [955]. At-Risk [1019]. Athletes [2543]. Atomic [727]. Atomic-Level [727]. Atrial [3062]. Attack [453, 641, 1497, 2721]. Attack-Resilient [2721]. Attacks [527, 1217, 1432, 1841, 1850, 2387, 2566]. Attention [805, 1587, 1686, 2209]. Attention-Aware [1686]. Attentive [1790]. Attenuated [1901]. Attitudes [2616]. Attracting [1314]. Attribute [1448, 2647]. Attribute-Based [1448, 2647]. Attributes [108]. Attribution [2217]. Audience [1941, 2542, 2565]. Audio [1160]. Audiovisual [589]. Auditing [55]. Auditor [55]. Augment [2183]. Augmented [609–614, 1045, 2080]. Augmenting [2937]. August [2664]. Authenticated [1162]. Authenticating [1624]. Authentication [25, 708, 1532, 1728, 2597]. Autism [594]. Autistic [2612]. Auto [47]. Auto-Pipe [47]. Automated [448, 991, 1121, 1619, 2703].

Automatic [486, 619, 1272, 1382].

Automating [231]. Automobile [588]. Automotive [315]. Autonomic [9, 223]. Autonomous [2103, 2190, 2276, 2353, 2466, 2469]. Autonomy [914, 3006]. Autos [357]. Autoscaling [1663]. Availability [500, 1555, 2163, 2303]. Avalanche [353, 3033]. Average [1159]. Averting [298]. Avizienis [621]. Avoid [166, 605, 2573]. Avoided [2419]. Avoiding [1317]. Award [32, 51, 68, 88, 106, 199, 257, 296, 317, 336, 399, 419, 438, 439, 441, 596, 621, 1237, 1263, 1274, 1613, 1650, 1690, 1700, 1881, 1912, 1917, 1942, 1967, 1993, 1998, 2150, 2185, 2191, 2220, 2343, 2363, 2587, 2591, 2593, 2606, 2617, 2619, 2634, 2645, 2656, 2665, 2681, 2683, 2695, 2697, 2710, 2712, 2768, 2771, 2797, 2993]. Awards [278, 513, 731, 1956, 1970, 1984, 2032, 2137, 2167, 2192, 2208, 2214, 2363, 2567, 2599, 2608, 2632, 2663, 2691, 2976, 3089]. Aware [178, 1075, 1119, 1414, 1564, 1584, 1585, 1620, 1663, 1686, 1710, 1848, 1965, 1978, 2267, 2329, 2647, 3066]. Awareness [313, 565, 728, 814, 847, 863, 1586, 1587]. Away [2237].

B. [2768, 2797]. Babbage [2606, 2617, 2681, 2710]. Back [450, 458, 1061, 2130, 2669]. Backed [2322]. Backup [2330]. Bad [246, 2767]. Badly [1692]. Baker [756]. Balanced [2101]. Balancing [1122]. Ballooning [2323]. Band [2728]. Bands [1557]. Banking [2163, 2382]. Banzi [1035]. Bar [829, 1771]. Barcodes [230]. Barns [20]. Barrett [399]. Barriers [784]. Based [11, 62, 108, 127, 144, 178, 185, 235, 316, 357, 591, 593, 611, 615, 636, 679, 784, 906, 988, 1046, 1059, 1125, 1220, 1448, 1656, 1665, 1755, 1828, 1857, 1965, 2159, 2184, 2209, 2236, 2319, 2387, 2409, 2527, 2564, 2597, 2598, 2647, 2668, 2722, 2742, 2877, 2904, 2905, 3064, 3067]. Bases [2136]. Basic [2843]. Batteries [823].

Battery [50, 960, 1216]. Battle [288, 615, 2132]. Battlefield [2719]. Be [240, 404, 634, 1683, 1767, 1798, 1809, 2419, 2427, 2455, 2483, 2510, 2553, 2582, 2682, 2711]. Beat [2776]. Beauty [769, 1128]. Becoming [2875]. Bed [2603]. Been [209, 405]. Beeping [1557]. before [829]. Begin [2543]. Beginning [1012]. Behavior [359, 1095]. Behavioral [2385]. Behaviors [2544]. behind [2504]. Behlendorf [670]. Being [1122, 2094, 2337, 2653]. Bell [854]. Benefit [2746]. Benefits [1587, 2393, 2629]. Bertrand [88, 647]. Bespoke [915]. Best [557, 671, 1331, 1370, 1559, 1974, 2370, 2395, 2428, 2456, 2485, 2512, 2578, 2614, 2700]. Better [75, 1769, 2024, 2086, 2113, 2126, 2143, 2169, 2173, 2186, 2187, 2194, 2224, 2253, 2269, 2309, 2314, 2348, 2725, 3067]. Between [464, 785, 1216, 1243, 1907, 2008, 2324]. Beyond [260, 352, 542, 543, 929, 982, 1021, 1281, 1523, 1704, 1713, 1749, 1778, 1789, 2078, 2297, 2320, 2522, 2626, 2935]. BI [464]. Bias [160]. Biclusters [916]. Big [663, 673, 696, 740, 822, 846, 860, 862–865, 927, 1044, 1131, 1178, 1267, 1307, 1315, 1320, 1324, 1363, 1369, 1396, 1411, 1427, 1463, 1465, 1466, 1469, 1595, 1714, 1814, 1856, 1866, 1869, 1870, 1884, 1885, 1935, 1981, 2041, 2297, 2452, 2489, 2571, 2699, 2702, 2737, 2764, 2780, 2812, 2841, 2884, 2900, 2940, 2999, 3036, 3056, 3098, 3111]. BigData [1358, 1377]. BigDataService [1340]. Binarized [2873]. Binary [910]. Binds [3099]. Bioacoustics [2184]. Bioacoustics-Based [2184]. Biobotic [1896]. Biochips [1923, 1996]. Bioinformatics [590, 738]. Biologically [2974]. Biology [522]. Biometric [25, 29, 708, 938, 1281, 1609]. Biometrically [799]. Biometrics [404, 847, 1346]. Biopharmaceutical [2388]. Biosignals [1101]. Birman [68]. BIS [3042]. Bit [1070]. Bitcoin [2384, 2522]. Bits [1849]. Bixplorer [916]. Black [1839]. Black-Box [1839]. Blackout [615]. Blame [574, 2298].

Bletchley [581]. Blindfolded [1015]. Bloat [402]. Blockchain [2379, 2380, 2382, 2522, 2698, 2729, 2744, 2877, 2996, 3083, 3103]. Blockchain-Based [2877]. Blockchain-Enabled [3083]. BlockPy [2262]. Blocks [854]. Blood [1014]. Blood-Flow [1014]. Blue [199]. Blueprint [2934]. Board [2773, 2833, 2889, 2916, 2949, 3010, 3048, 3074]. Boards [2799, 2953]. Bob [831]. Bodies [1270, 1829]. Body [185, 1128, 2184, 2408]. Boeing [2814]. Bogus [1873]. Bombings [991]. Bond [654]. Bonding [1260]. Book [1003]. Books [3108]. Bookshelf [31, 74, 87, 124, 142, 180, 201, 220, 237, 279, 398]. Boolean [1874]. Boosting [1309, 2072]. Boot [1215, 1966]. Borderline [1506]. Borenstein [1147]. Boston [991]. Bot [1929]. Botnets [818, 843, 2189, 2333]. Bots [531, 3002]. Bottleneck [634]. Boundaries [2040, 2102]. Bowel [2598]. Box [339, 701, 888, 1839]. Boxes [2902]. BPM [464]. BPO [459]. Bracelets [2246]. Braille [1776, 2105]. Brain [221, 251, 543, 1686, 2072, 2073, 2186, 2848]. Brain-Computer [543, 1686]. Brain-Inspired [2072]. Brain-Network [2073]. Brains [425]. Branch [2566]. Brand [2549]. Brand-Equity [2549]. Breaking [574]. Breakthrough [879]. Breathing [2597]. Breathing-Based [2597]. Brexit [2495]. Brian [670]. Bribes [757]. Bridge [464, 881]. Bridges [1932]. Bridging [234, 1243, 2324, 3097]. Brief [694, 715, 2574]. Briefs [327, 347, 366, 388, 410, 478, 498, 521, 541, 563, 585, 608, 630, 651, 674, 697, 717, 741, 760, 781, 811, 835, 859, 880, 903, 928, 951, 977, 1007, 1038, 1063, 1092, 1118, 1151, 1180, 1209, 1236, 1266, 1301, 1334, 1373]. Bring [1188, 1327, 1328, 1366, 1367]. Bringing [409, 740, 1261, 2393, 2618]. Brings [497]. Broadening [787]. Broader [1941, 2542, 2565]. Brokers [1094, 1932]. Bronze [2650]. Browser [636]. Browsing [314, 2588]. Bruce [1804, 1862, 2524]. Buffer [436]. Buffer-Overflow [436]. Buffers [1217]. Bug [688]. Build [592, 1727]. Building [231, 253, 272, 333, 509, 571, 675, 729, 777, 807, 826, 854, 1035, 1115, 1132, 1350, 1484, 1519, 1784, 1862, 1876, 2187, 2646]. Bundeswehr [2053]. Bus [362]. Business [67, 116, 528, 627, 671, 1016, 1324, 1363, 1396, 1418, 1588, 1845, 3003]. Business-Critical [1588]. Businesses [959, 2005]. Busted [1019]. Buzz [665]. BYOD [1434].

C [685, 924]. C/C [685]. Cables [341]. Cache [954]. Calendar [14, 33, 52, 162, 200, 239, 337, 376, 400, 442, 463, 485, 514, 534, 556, 578, 600, 622, 644, 666, 690, 711, 732, 774, 804, 828, 851, 874, 896, 920, 943, 966, 994, 1026, 1051, 1078, 1106, 1137, 1168, 1198, 1225, 1254, 1287, 1318, 1359, 1390, 1422, 1450, 1482, 1508, 1541, 1566, 1598, 1631, 1669, 1699, 1733, 1762, 1794, 1822, 1851, 1879, 1911, 1945, 1976, 2018, 2050, 2081, 2108, 2139, 2168, 2219, 2248, 2278, 2305, 2340]. Call [33, 200, 239, 376, 400, 442, 463, 485, 514, 556, 578, 600, 622, 644, 666, 690, 711, 732, 774, 804, 828, 851, 874, 896, 920, 943, 966, 994, 1026, 1051, 1078, 1106, 1137, 1168, 1198, 1225, 1237, 1254, 1256, 1274, 1278, 1287, 1295, 1318, 1359, 1390, 1422, 1425, 1450, 1453, 1474, 1482, 1498, 1508, 1525, 1527, 1541, 1558, 1566, 1577, 1598, 1604, 1631, 1635, 1636, 1650, 1661, 1669, 1672, 1688, 1690, 1698, 1699, 1733, 1735, 1762, 1794, 1822, 1851, 1879, 1911, 1945, 1976, 2018, 2050, 2081, 2108]. Call [14, 52, 162, 337, 534, 1825, 1838, 1865, 1912, 1936, 1942, 1967, 1970, 1984, 1998, 2032, 2110, 2124, 2139, 2168, 2191, 2192, 2208, 2219, 2220, 2248, 2278, 2305, 2326, 2340, 2343, 2491, 2599, 2632, 2663, 2691, 2755, 2782, 2784, 2786, 2806, 2807, 2821, 2822, 2839, 2851, 2853, 2870, 2897, 2903, 2907, 2936, 2946, 2969, 2971, 2973, 2975, 2976, 2984, 2995, 2997, 3025, 3028, 3034, 3041, 3072, 3084]. Calls [1836]. Cambridge [2601, 2672]. Camera [44, 1152–1158, 1472, 1501, 1585]. Cameras [913, 1155, 2878, 3024]. Can [66, 202, 240, 404, 412, 446, 575, 584, 634, 771, 783, 1016, 1020, 1134, 1555, 1727, 2126, 2169, 2206, 2260, 2273, 2392, 2419, 2505, 2651, 2878, 3027, 3033]. Cancer [120, 525]. Candidate [1868]. Candidates [141, 375, 643, 919, 1253, 1630, 2014, 2016, 2365, 2366, 2677, 2678, 2979, 2980]. Can’t [592]. CAP [499–504]. Capabilities [1307, 2790]. Capacity [1465, 2699]. Capes [1350]. Capital [39, 605, 1242]. Capstone [1786]. Captcha [255]. Capturing [99, 101, 684]. Car [1166, 2149]. Carbon [7, 231]. Carbon-Constrained [7]. Cardiac [2098]. Care [865]. Career [1289, 1319, 1360, 1392, 1423, 1451, 1483, 1510, 1529, 1542, 1568, 1599, 1602, 1608, 1632, 1670, 1675, 1701, 1722, 1736, 1764, 1768, 1795, 1799, 1823, 1827, 1855, 1883, 1890, 1914, 1933, 1946, 1963, 1977, 2013, 2025, 2051, 2055, 2057, 2082, 2085, 2087, 2112, 2114, 2142, 2145, 2172, 2175, 2193, 2196, 2223, 2252, 2308, 2315, 2331, 2371, 2397, 2399, 2881]. Careers [2, 1004, 2322]. Caregiver [3063]. Carrying [1403]. Cars [2190, 2465, 2740]. CASAS [888]. Case [17, 193, 592, 701, 934, 991, 1351, 1782, 1905, 2100, 2616, 2620, 2721, 2944, 3062]. Cases [1492, 1520]. Cash [650]. Catalyst [843]. Catastrophes [1317]. Categorizing [105, 1814]. Category [636]. Category-Based [636]. Causal [1754]. Causality [421]. Cause [1232]. CCEM [1135]. CDC [2232]. CE2016 [1667]. Celebration [1446, 1577, 1635, 1672]. Cell [16, 1968]. Cellular [2412]. Censoring [56]. Center [902, 1683, 1767, 1798, 1809, 2427, 2455, 2483, 2510, 2553, 2582, 2682, 2711, 3044, 3104]. Centered [121, 764, 1385, 1964]. Centers [125]. Centimeter [2158]. Centimeter-Scale [2158]. Central [2650]. Centric [233, 737, 824, 2237, 2742]. Century [261, 341, 465, 569, 812, 958, 1375, 1694, 1940, 1957, 2476]. Cerf [715]. Certain [446]. Certainty [258]. Certificates [453].

Certification [2388, 2576, 2854]. CES [996, 1292, 1302, 1338]. Chain [392, 1990, 1994, 1996, 2527]. Chaining [1339, 2101, 2943]. Challenge [79, 223, 254, 469, 597, 629, 887, 927, 937, 938, 1929, 2274, 2319, 3095]. Challenge-Based [2319]. Challenges [89, 330, 355, 403, 434, 477, 706, 721, 722, 766, 793, 838, 860, 905, 930, 1014, 1071, 1384, 1524, 1591, 1623, 1713, 1780, 1811, 2036, 2211, 2242, 2318, 2388, 2467, 2562, 2598, 2602, 2701, 2740, 2758, 2968]. Challenging [648]. Chance [2776]. Change [57, 90, 297, 331, 407, 908, 1170, 1349, 1400, 1401, 1747]. Changed [802]. Changers [825, 1389]. Changes [172, 794, 2206, 2330, 2460]. Changin [1576]. Changing [837, 1048, 1529, 2019, 2244]. Channel [1260, 1550]. Channels [1579]. Chaos [58, 143]. Chapters [293]. Characteristics [338]. Characterization [1968]. Characterizing [820]. Charles [2606, 2617, 2681, 2710]. Charting [431]. Chasing [2106]. Chasm [290]. Checking [1403, 1999]. Chemistry [2266]. Chemistry-Inspired [2266]. Chen [278]. Chicken [362]. Children [2612]. China [627, 705, 1311, 2481, 2742]. Chinese [936]. Chip [955, 1756, 1994, 2231]. Chipmakers [288]. Chips [6, 23, 249, 1342, 1739, 2259, 2361]. Chirping [1557]. Chocolate [2362]. Choice [849, 1116, 2385]. Choose [2087, 2114, 2145, 2175, 2196]. Choreographed [1065]. Choreographing [2128]. Chuck [438]. Chutzpah [1446]. Cinefile [636]. Circuit [2267, 2318]. Circuitry [251]. Circuits [175]. Cirrus [64]. Cities [329, 330, 333, 1310, 1414, 2647, 2737, 2742]. Citizens [2618]. City [332, 460, 1311, 2741, 2743]. Civil [1017, 2761]. Class [1087]. Classical [2746]. Classifier [1691]. Classifying [176]. Classroom [614, 929, 932]. Clearance [2419]. Cleve [439]. Client [907]. Clients [894]. Climate [57, 935, 1349]. Clock [1725]. Close [523]. Closer [387]. Closing [1239]. Cloud [64, 66, 70, 177, 270–272, 274– 276, 303, 346, 371, 372, 393, 416, 454, 480, 497, 503, 570, 616, 702, 724, 727, 748, 797, 894, 939, 962, 988, 990, 998, 1013, 1046, 1053, 1080, 1081, 1108, 1135, 1140, 1145, 1157, 1196, 1204, 1227, 1246, 1298, 1312, 1321, 1335, 1384, 1415, 1442, 1444, 1473, 1518, 1536, 1564, 1661, 1663, 1698, 1726, 1807, 1808, 1810, 1811, 1816, 1844, 2042, 2047, 2107, 2160, 2236, 2243, 2258, 2303, 2362, 2424, 2449, 2469, 2475, 2491, 2500, 2595, 2962, 3004, 3026, 3064, 3083, 3104]. Cloud-Assisted [1157]. Cloud-Based [988, 1046, 3064]. Cloud-Computing [346]. Cloud-Integrated [2236]. Cloud-Scale [275]. Cloudebug [1222]. Clouds [65, 834, 846, 1298, 1351, 1696]. Clustering [2387]. Clusters [433, 2122]. Coaches [2543]. Coaching [2539, 2540, 2666]. Coda [2046]. Code [511, 910, 1217, 1553, 1818, 2566, 2783]. Code.org [917]. Code4GoodPBC [2324]. Codes [63, 2673, 2823]. Codesign [431–434, 960, 1618]. Coding [1160, 2263]. Cognition [2937]. Cognitive [964, 1550, 1684, 1842, 1869, 2541, 2869, 3094]. Coherence [954]. Coherent [952, 2498]. Cold [2350]. Cold-War [2350]. Collaboration [155, 768, 777, 1846, 2447, 2743, 2765, 2769]. Collaborative [392, 435, 523, 1087, 1093, 1095– 1097, 2125, 2263, 2381]. Collection [120, 340, 868, 1126]. Collections [103]. Collective [143, 863, 1087, 1114, 1586, 1778, 2387]. College [399, 683, 789, 1018, 1969, 2629, 3106]. Column [1537]. Combat [1212]. Combat-Related [1212]. Combating [818, 1696]. Combination [114]. Combinatorial [1067, 1502, 2076, 2157]. Combines [117]. Combining [1684].

Comes [780]. Comfort [875]. Comfortable [23]. Coming [149, 229, 1943]. Command [2479]. Commerce [213]. Commercial [277]. Commercialization [1238, 1240]. Committee [179]. Committees [2090]. Common [491, 598, 944]. Communication [979, 2184, 2931, 2934]. Communications [429, 569, 655, 1834, 2527]. Communities [335, 2623]. Community [2, 103, 348, 565, 1241, 1969]. Community-Driven [1241]. Comodeling [295]. Companies [97]. Company [510, 948]. Comparative [2521]. Comparing [2813]. Comparison [2239]. Compass [2285]. Compatible [440]. Competition [2394, 2444, 2502, 2650]. Competitions [2322]. Competitive [1366]. Compiler [160]. Complex [524, 813, 1190]. Complexity [640, 2133, 2540, 2783]. Compliance [30, 2970]. Complicated [667]. Component [981, 2129]. Components [1992, 3105]. Comprehensive [86, 1239]. Compressive [1156]. COMPSAC [1525, 1629, 1715, 2064, 2471, 2860]. Computation [66, 234, 253, 350, 414, 1120, 1505, 2034, 2036]. Computational [122, 216, 217, 349, 1403, 1719, 2133]. Computationally [522]. Compute [618]. COMPUTER [140, 374, 461, 642, 689, 918, 971, 985, 1011, 1022, 1023, 1025, 1034, 1055, 1194, 1252, 1274, 1331, 1357, 1370, 1391, 1480, 1490, 1503, 1567, 1582, 1610, 1628, 1650, 1652, 1675, 1683, 1702, 1752, 1761, 1767, 1798, 1809, 1817, 1883, 1890, 1891, 1900, 1912, 1917, 1930, 1933, 1942, 1956, 1961, 1963, 1967, 1985, 1987, 1998, 2007, 2012, 2013, 2020, 2028, 2055, 2060, 2067, 2085, 2087, 2090, 2112, 2114, 2117, 2121, 2142, 2145, 2175, 2191, 2196, 2199, 2220, 2273, 2343, 2351, 2358, 2364, 2391, 2711, 2994]. Computer [138, 342, 373, 465, 1387, 1835, 1898, 2172, 2193, 2223, 2244, 2252, 2308, 2352, 2371, 2399, 2421, 2427, 2437, 2455, 2468, 2480, 2483, 2497, 2510, 2530, 2532, 2536, 2553, 2554, 2559, 2582, 2583, 2593, 2604, 2628, 2634, 2635, 2640, 2645, 2657, 2658, 2675, 2676, 2682, 2683, 2697, 2705, 2707, 2730, 2738, 2747, 2756, 2772, 2773, 2789, 2798, 2799, 2832, 2833, 2849, 2859, 2887, 2888, 2909, 2923, 2933, 2947, 2949, 2950, 2971, 2977, 2981, 2992, 3001, 3009, 3012, 3022, 3047, 3068, 3073, 3080, 3089, 3112, 3113]. Computer [1, 12, 62, 94, 100, 141, 145, 221, 235, 238, 283, 350, 354, 375, 424, 466, 533, 543, 555, 577, 589, 598, 604, 615, 643, 656, 691, 710, 730, 786, 803, 827, 850, 873, 895, 919, 930, 934, 942, 965, 993, 1024, 1050, 1077, 1105, 1125, 1136, 1167, 1197, 1224, 1253, 1286, 1347, 1405, 1481, 1530, 1630, 1667, 1686, 1693, 1801, 1846, 1860, 1881, 1937, 1955, 2016, 2031, 2065, 2109, 2138, 2149, 2294, 2316, 2318, 2327, 2353, 2356, 2366, 2404, 2503, 2589, 2622, 2630, 2661, 2686, 2689, 2724, 2751, 2761, 3079, 3090]. Computer [2, 13, 51, 106, 113, 129, 150, 167, 187, 207, 209, 224, 245, 264, 284, 336, 361, 369, 383, 406, 426, 449, 471, 473, 492, 513, 516, 536, 558, 580, 602, 624, 646, 668, 692, 713, 734, 754, 776, 806, 830, 853, 876, 898, 922, 945, 972, 1001, 1031, 1058, 1086, 1113, 1144, 1176, 1203, 1231, 1259, 1297, 1329, 1368, 1401, 1402, 1431, 1458, 1517, 1549, 1578, 1606, 1640, 1641, 1677, 1709, 1744, 1747, 1775, 1803, 2148, 2179, 2286, 2577, 2678, 2736, 2830, 2831, 2871, 2919, 2980]. Computer [305, 323, 1832, 1859, 1888, 1922, 1954, 2030, 2062, 2092, 2117, 2120, 2151–2153, 2180, 2201, 2256, 2284, 2313, 2375, 2403, 2463, 2558, 2662, 2664, 2752, 2778, 2805, 2836, 2863, 2866, 2894, 2927, 2957, 2976, 2984, 2988, 3015, 3034, 3081]. Computer-Assisted [354, 589]. Computer-Based [235, 615]. Computerization [2941]. Computers [28, 206, 244, 1459, 1496, 1555, 1569, 1903, 1939, 2008, 2462, 2477, 2573, 2746, 2776, 2901, 3000, 3076]. Computing [9, 10, 37, 54, 64, 89, 109, 128, 186, 232, 263, 272, 282, 343, 351, 431, 455, 586, 588, 597, 616, 631, 652, 655, 682, 684, 709, 783, 785, 788, 789, 798, 823, 845, 929, 932, 933, 1008, 1041, 1076, 1080, 1119, 1270, 1273, 1275, 1276, 1298, 1314, 1321, 1384, 1415, 1570, 1627, 1662, 1682, 1698, 1748, 1750, 1751, 1778, 1807, 1808, 2004, 2010, 2042, 2066, 2107, 2133, 2155, 2160, 2244, 2293, 2300, 2350, 2390, 2412, 2414, 2422, 2453, 2518, 2567, 2569, 2608, 2671, 2746, 2869, 2898, 2905, 2951, 2959, 2965, 2982, 2991, 3019, 3026, 3087]. Computing [36, 66, 75, 135, 161, 177, 179, 218, 235, 242, 252, 270, 276, 288, 290, 297, 311, 346, 371, 372, 379, 416, 479, 508, 517, 587, 607, 664, 696, 748, 768, 782, 784, 791, 792, 797, 844, 854, 899, 924, 931, 962, 998, 1053, 1089, 1128, 1220, 1268, 1335, 1477, 1536, 1553, 1587, 1594, 1614, 1620, 1661, 1718, 1749, 1842, 1846, 1870, 1893, 1894, 1902, 2033, 2035, 2038, 2039, 2041, 2072, 2156, 2178, 2258, 2260, 2303, 2354, 2357, 2410, 2424, 2439, 2484, 2491, 2590, 2596, 2602, 2612, 2658, 2904, 2931, 2933, 2937, 2962, 3016, 3040, 3083]. Computing [2511, 2766, 2786, 2791, 2816, 2817, 2821, 2822, 2825, 2826, 2844, 2847, 2853, 2868, 2870, 2879, 2893, 2903, 2913, 2930, 2932, 2961, 2975, 2997, 3028, 3039, 3045, 3070, 3082, 3088]. Concept [107]. Concepts [184, 1356, 1388]. Concerns [141, 375, 643, 919, 1253, 1630, 2016, 2366, 2678, 2910, 2980]. Concurrent [1159]. Condition [1732]. Conduct [63]. Conference [852, 1227, 1569, 1872]. Conferences [1146, 1852, 1891, 2216]. Confidence [1780]. Confident [170]. Confidentiality [393]. Configurability [2328]. Configuration [1926]. Configurations [680, 1663]. Conflict [2210]. Conflicts [122]. Congress [1028, 1358, 1377, 1427, 2858, 2864, 2922]. Connect [136, 1756, 2310, 2345, 2429, 2551, 2580]. Connected [762, 1149, 1221, 1471, 2096, 2386, 2442, 2627, 2687]. Connecting [933, 2341]. Connection [533, 710, 827, 850, 873, 895, 942, 965, 993, 1024, 1050, 1077, 1105, 1136, 1167, 1197, 1224, 1286, 1357, 1421, 1449, 1507, 1565, 1597, 1668, 1697, 1734, 1763, 1793, 1821, 1910, 1975, 2049, 2166, 2279, 2304, 2339, 2367, 2451, 2506, 2529, 2550, 2579, 2631, 2654, 2679, 2708, 2745, 2748, 2885, 3008, 3071]. Connections [2306, 2920, 2945]. Connectivity [2623, 2792]. Consequences [864]. Considerations [861, 2162]. Considered [1872]. Consistency [500, 502, 2519]. Consistent [504]. Console [262]. Consoles [772]. Consortium [2381, 2658, 2933]. Constitution [2015]. Constrained [7, 2597]. Constraint [680]. Constructing [2718]. Consumer [1155, 1689, 3037]. Consumers [959]. Consumption [357, 820, 1163, 1216]. Consumption-Based [357]. Contact [1164]. Conte [1023]. Content [215, 737, 1841, 2128, 2454, 2537]. Content-Centric [737]. Contents [970, 1000, 1030, 1057, 1085, 1112, 1143, 1175, 1202, 1230, 1258, 1398, 1429, 2027, 2059, 2089, 2116, 2147, 2177, 2198, 2226, 2254, 2282, 2311, 2346, 2373, 2401, 2430, 2458, 2486, 2513, 2533, 2555, 2584, 2609, 2636, 2659, 2684, 2713, 2733, 2749, 2774, 2800, 2802, 2834, 2861, 2890, 2925, 2954, 2985, 3013, 3051, 3077]. Contests [3029]. Context [540, 763, 805, 1075, 1120, 1624, 2292]. Context-Aware [1075]. Continually [2330]. Continuous [1002, 1728]. Contract [2722]. Contract-Based [2722]. Contracts [49, 1090, 2383]. Contrasting [2813]. Contribute [2273]. Contribution [2337]. Control [108, 153, 593, 958, 984, 1341, 1448, 1689, 2070, 2134, 2329, 2564, 2721]. Controlled [1908]. Controlling [356, 466, 601]. Controls [866, 894, 974, 1848]. Controversies [2300]. Controversy [1972]. Converged [629]. Conversational [2668]. Conversations [899, 1089, 1947]. Conversion [2295]. Converts [133]. Cookies [941]. COOL [1342, 1739]. Cooler [23]. Coordinating [3002]. Coordination [105]. Copernican [940]. Coproduction [1960]. Core [37, 458, 1679, 1800, 1904, 2131]. Corp. [1379, 1666]. Corpora [507].

Corporate [862, 1036, 1241, 2972]. Corporation [1913]. Corrections [35]. Corrupted [1388]. Cortically [2038]. Cosmetic [1164]. Cost [862, 1018, 1444, 2323, 3026]. counterfeiters [1992]. Countermeasures [2301]. Coupled [2038]. Couples [2210]. Course [2477]. Coursera [626]. Courses [455, 2121]. Court [1206]. Cover [968, 997, 1054, 1082, 1109, 1141, 1172, 1200, 1394, 1426, 1454, 1485, 1512, 1544, 1571, 1601, 1637, 1673, 1706, 1740, 1766, 1797, 1826, 1854, 1882, 1918, 1949, 1979, 2021, 2054, 2084, 2111, 2141, 2171, 2222, 2250, 2281, 2307, 2342, 2369, 2398, 2426, 2482, 2509, 2531, 2552, 2581, 2605, 2633, 2655, 2680, 2709, 2731, 2732, 2770, 2796, 2828, 2829, 2857, 2886, 2921, 2948, 2952, 2983, 3011, 3049, 3075]. Cows [1561]. CPaaS.io [2743]. CPS [2721]. CPUs [45]. Crashes [2419]. Crawling [1161]. Crawls [2620]. Cray [199, 438, 2137, 2167, 2214, 2567, 2608]. Crazy [1657]. Create [2625]. Creating [472, 686, 933, 1233, 1277, 1339, 1444, 2965]. Creation [1347]. Creativity [37]. Credentials [2647]. Creep [1626]. Crime [149, 574]. Criminals [320]. Crisis [35, 298, 552, 1788, 1895]. Critical [81, 83, 85, 814, 958, 1588, 3005]. Cross [527, 659, 1745, 1841, 1864, 2040]. Cross-Domain [1864]. Cross-Platform [659]. Cross-Project [1745]. Cross-Referencing [2040]. Cross-Site [527, 1841]. Crossbars [253]. Crowd [137, 2760]. Crowd-Assisted [2760]. Crowd-Powered [137]. Crowdfunding [801]. Crowdsourcing [328, 447, 752, 1284, 1897, 1932, 1955, 2544, 2740, 2742]. Crowdsourcing-Based [2742]. Crunch [673, 2941]. Cryptocurrencies [2727]. Cryptocurrency [2382]. Cryptographic [1862]. Cryptographically [1505]. Cryptography [2044]. Crystals [253]. CS [182, 293, 419, 484, 731, 773, 1421, 1449, 1470, 1507, 1565, 1597, 1668, 1697, 1734, 1763, 1793, 1821, 1910, 1975, 2049, 2150, 2164, 2166, 2185, 2228, 2273, 2279, 2304, 2339, 2367, 2451, 2506, 2529, 2550, 2568, 2579, 2587, 2606, 2617, 2631, 2654, 2679, 2681, 2695, 2708, 2710, 2715, 2735, 2748, 2771, 2885, 2920, 2945, 3008, 3057, 3071]. CS2013 [1481]. CSDA [336]. CSDP [336]. CSI [2587, 2695]. CSI-IEEE [2587, 2695]. Cuddly [1958]. Cues [1556]. Cultivating [293]. Cultural [352, 382, 457]. Culturally [686]. Culture [744, 2170]. Cup [2788]. Curious [170]. Curiouser [1581]. Current [1816, 2467, 2478, 2760, 2898]. Curricula [1481, 1667, 2814]. Curricular [2811, 2814]. Curriculum [932, 1731, 2162]. Curtail [562]. Customer [709]. Cutting [72, 745, 836]. Cutting-Edge [836]. Cyber [7, 743, 1071, 1389, 1446, 1494, 1499, 1928, 2217, 2234–2237, 2461, 2476, 2548, 2703, 2717, 2722, 2726, 2843, 3101]. Cyber-infrastructure [7]. Cyber-Physical [743, 1071, 2234– 2237, 2461, 2476, 2548, 2703, 2717, 2722, 2726]. Cyberattack [641]. Cyberattacks [2720]. Cyberbullying [401, 2852]. Cybercrime [817]. Cyberentity [816]. Cyberethics [2704]. Cyberfrontier [1925]. Cyberhuman [1538]. Cyberinfrastructure [192, 435, 1990]. Cyberinfrastructures [1243]. Cyberphysical [397, 767, 2943]. Cyberpreparedness [73]. Cybersecuring [2005]. Cybersecurity [38, 959, 1048, 1173, 1245, 1247, 1277, 1412, 1494, 1496, 1500, 1515, 1591, 1645, 1760, 1878, 1909, 1919, 1931, 1951, 2023, 2230, 2472, 2477, 2526, 2562, 2811, 2813–2815, 2818–2820, 3095]. Cyberspace [812, 1317, 1622, 2104, 3033, 3097]. Cyberthreats [2603]. Cybertrust [2642, 2727, 2744]. Cyborgs [531]. Cycle [955, 1419, 1994]. Cycles [97, 1002, 1796]. Cycling [1622]. Czar [1600].

D [5, 100, 229, 352, 409, 820, 1129, 1157, 1214, 1274, 1299, 1407, 1487, 1831, 2183, 2361, 2417, 2593, 2634, 2645, 2683, 2697]. D-Printed [2183]. Dada [1315]. Daily [2183]. Damages [1711]. Dangerous [2335]. Daphne [626]. Dark [1244, 2245, 2335]. Darkly [892]. DARPA [1929]. Darwin [1104]. Data [120, 125, 148, 217, 233, 322, 340, 341, 353, 380, 421, 480, 503, 518, 572, 632, 634, 663, 673, 700, 701, 797, 822, 824, 842, 844, 846, 860–863, 865, 868, 871, 883, 927, 1013, 1015, 1044, 1131, 1178, 1204, 1267, 1307, 1320, 1349, 1369, 1384, 1411, 1427, 1442, 1463, 1464, 1466, 1468, 1469, 1614, 1714, 1745, 1750, 1856, 1866, 1869, 1884, 1885, 1935, 1962, 1981, 2040, 2073, 2074, 2099, 2122, 2207, 2233, 2241, 2262, 2297, 2330, 2405, 2501, 2520, 2595, 2669, 2672, 2699, 2737, 2739, 2757, 2788, 2845, 3026, 3037, 3044, 3098, 3104]. Data [864, 935, 990, 1465, 1595, 1814, 1864, 1870, 2041, 2452, 2489, 2571, 2702, 2764, 2780, 2812, 2841, 2884, 2900, 2940, 2999, 3036, 3056, 3111]. Data-Capacity [2699]. Data-Centric [233, 824]. Data-Driven [861]. Data-Intensive [148, 572, 1204, 1384, 1614, 1962]. Data-Parallel [634]. Data-Science [2262]. Database [502, 1233, 1656, 1868]. Databases [22, 599, 1464]. Datacenter [618, 657, 2297]. Datacenters [724, 1353, 1588, 1812, 2475]. DataPort [3032]. Datasets [1308]. David [1219]. Dawn [15]. Day [349, 1006, 2075]. Days [494]. DBMS [937]. DDoS [2333, 2644]. Deadline [1633]. Dealers [1166]. Dealing [353, 2540]. Death [235, 312]. Debates [340]. Debt [2269]. Debug [723]. Debugging [704]. Decade [2318]. December [2736]. Deception [1315, 1499]. Deceptions [2504]. Deciding [1726]. Decision [2098, 2669, 2848]. Decline [248]. Decode [350]. Deduction [2943]. Deep [2335, 2420, 2590, 2592, 2594– 2596, 2840, 2873, 3033, 3062]. Default [235].

Defeating [795, 1992]. Defects [1559]. Defending [436, 506, 527, 538, 1315]. Defense [143, 1499, 2761]. Defensive [1047, 1315, 1346]. Deficit [1239]. Defined [879, 1335, 1343, 2101, 2499, 2520]. Defining [2122]. Definition [2354]. Degradation [2318]. Degree [1731]. Degrees [2629, 2815]. Déjà [751, 3042]. Delegate [2014, 2365, 2677, 2979]. Delegate/Director [2014, 2365, 2677, 2979]. Delivering [303, 459]. Delivery [843, 1177, 1497, 3096]. Delusion [2006]. Demand [616, 871]. Dementia [2541, 3063]. Democracy [304, 1310, 2495]. Demystified [1305]. Demystify [1840]. Demystifying [1938]. Denial [1432, 1499, 1850]. Denial-of-Service [1432, 1850]. Density [234, 2318, 2560]. Department [925]. Departments [2273]. Dependability [2912]. Dependency [633]. Deployment [1196, 1343, 1651, 2804]. Deployments [613]. Depression [1955]. Descriptions [83]. Deserve [1720]. Design [194, 219, 355, 380, 502, 547, 618, 778, 1006, 1232, 1385, 1530, 1651, 1836, 1957, 1999, 2157, 2231, 2265, 2267, 2276, 2325, 2352, 2540, 2649]. Design-for-Trust [355]. Designed [309]. Designing [3, 125, 494, 635, 915, 1564, 1686, 1776, 1965]. Designs [1475]. Desktop [617]. Destined [1791]. Detect [1217, 2097, 2210]. Detecting [813, 1997, 2566, 2638, 2776, 2964]. Detection [144, 355, 486, 658, 719, 815, 1059, 1684, 1685, 2000, 2387, 2646, 2788, 2892]. Detects [418]. Determining [1232, 2241]. Develop [44, 1962]. Developer [318, 1529]. Developers [1316, 1488, 2963]. Developing [177, 223, 381, 550, 2466, 2729]. Development [24, 67, 97, 98, 291, 403, 405, 422, 481, 659, 939, 986, 1042, 1313, 2269, 2438, 2518, 2741, 2972]. Device [960, 1188, 1780, 2207, 2336, 2441]. Devices [132, 133, 252, 254, 497, 762, 915, 1728, 1966, 2125, 2420, 2597, 2627, 3066].

Diabetic [2206]. Diabolical [1634]. Diagnosis [1955]. Diagonal [912]. Dictate [1760]. Dictator [1971]. dielets [1992]. Dies [70]. Differential [1184]. Difficult [468]. Digital [57, 79, 93, 166, 215, 339, 425, 453, 470, 473, 564, 718, 726, 736, 744, 817, 1019, 1039, 1163, 1248, 1641, 1785, 1844, 1849, 1906, 1923, 1985, 1996, 2105, 2106, 2221, 2246, 2350, 2524, 2562, 2623, 2723, 2726, 2881, 3061, 3097]. Digital-Microfluidic [1923]. Digitization [420]. Dilemma [1971]. Dimension [660, 775, 2500]. Dimensions [1810]. Direction [950]. Directions [1816, 2760]. Director [2014, 2365, 2677, 2979]. Directories [953]. Directoryless [1654]. Disability [1555]. Disaster [767, 1120, 1894, 1897, 1905]. Disaster-Prepared [767]. Discipline [126, 130, 465, 1703, 2504]. Disciplines [72]. Disconnect [866, 1634]. Discover [1331, 1370]. Discovering [460, 507, 537]. Discovery [37, 421]. Discrete [11]. Discrimination [1555]. Discriminatively [1059]. Discussion [1528]. Disease [302, 1134, 1439]. Diseases [524]. Disinformatics [2504, 2728]. Disney [611]. Disorder [1212, 1506]. Display [371, 418, 568, 569, 2105]. Displays [229, 508, 565–567, 837, 1475, 1554, 1776]. Disputes [63]. Disruption [792]. Disruptive [1570, 2698, 3006]. Dissecting [243]. Distance [2263]. Distinct [719]. Distinguished [1237, 1263, 1613, 1993, 2591, 2619, 2656, 2665, 2712]. Distribute [661]. Distributed [138, 502, 1310, 1362, 1550, 1586, 1638, 1850, 1896, 1898, 2022, 2035, 2127, 2241, 2381, 2795]. Distributed-Ledger [2381]. Diverse [47, 788]. Diversity [782, 783, 791, 1878, 2218]. Divide [1243]. Division [2014, 2365, 2677, 2979]. DLT [2729]. DLT/Blockchain [2729]. Dmail [1162]. Do [202, 358, 560, 1016, 1459, 1600, 2479, 2620, 2881, 2963]. Does [989, 2186, 2843]. Dog [2998]. Doing [627, 2295]. Domain [1864, 2330, 2385]. Domestic [459, 2601]. Don’t [2623]. Doors [1937]. Doorway [1165]. Doug [1115]. Douglas [2077]. Down [662]. Downtime [2785]. Dozens [336]. Dragging [2048]. Drastic [90]. Drawing [354]. Dream [823, 1789]. Dreams [2207, 2382]. Dress [1553]. Drive [2149, 2577]. Driven [185, 369, 522, 861, 1241, 2666, 2974]. Driver [111, 525]. Driverless [2740]. Drivers [1337]. Driving [688, 2353, 2465, 2467, 2469, 2703]. Drone [3096]. Dual [2964]. Dual-View [2964]. Dumb [324]. Dumbing [662]. Dynamic [675–679, 813, 953, 1759, 2987, 3066]. Dynamics [460, 1014, 1717, 1719]. Dystopia [1476, 2601]. Dystory [1725].

E-Assistant [2544]. E-Coaching [2539, 2540]. E-Commerce [213]. E-Government [1303–1305, 1784]. E-Health [1305]. E-Learning [2121]. E-mails [186]. E-Participation [1414]. e-Science [92]. E-Textiles [931]. Ear [444]. Early [261, 1433, 1782, 3003]. EARtH [2071]. Easy [1876]. Easy-to-Use [1876]. Eating [2544]. Ebb [507]. Eben [947]. Eckert [621, 2150, 2185, 2771]. Eco [313]. Eco-Feedback [313]. Ecology [1695]. Economic [17]. Economics [2385]. Economies [177]. Economy [273, 1072, 1238, 1242, 1753]. Ecosystem [1494, 2670]. Ecosystems [1187, 1695, 2475]. ECSEE [1724]. Edge [172, 745, 836, 1643, 1902, 1916, 1950, 2156, 2243, 2410, 2412–2414, 2475, 2595, 2937, 2951, 2962, 2965, 2982, 3104]. Edge-to-Cloud [2595]. Editor [40, 208, 265, 285, 499, 522, 1008, 1064, 1093, 1152, 1238, 1375, 1990]. Editors [81, 860, 904, 1119, 1437, 1553, 1612, 1647, 2066, 2123, 2717, 2737]. Educating [765, 788, 2503, 2819]. Education [54, 157, 222, 278, 301, 342, 381, 399, 511, 786, 929–931, 934, 1076, 1247, 1594, 1710, 1846, 1970, 1984, 2032, 2060, 2192, 2208, 2263, 2265, 2266, 2422, 2476, 2526, 2587, 2599, 2632, 2663, 2691, 2695, 2818, 2820]. Educational [198, 2813, 3040]. Educator [870]. edX [1680]. EEG [619, 1684]. Eerie [2450]. Effect [160, 1543]. Effective [1776]. Effectiveness [277, 2472, 3062]. Effects [377, 474, 1716, 3104]. Efficiency [629, 1046, 1220, 1353, 1594, 2068, 2072, 2207]. Efficient [125, 240, 1157, 1643, 1750, 1812, 2066, 2125, 2390]. Effort [422, 1249]. EIC [773, 2754]. EICs [689]. Eiffel [647]. Eight [2268]. Elasticity [1473]. Elbridge [2106]. Elect [141, 375, 461, 643, 730, 919, 1023, 1253, 1391, 1630, 1761, 2016, 2138, 2366, 2480, 2678, 2747, 2980, 3112]. Election [140, 374, 642, 918, 1252, 1628, 2012, 2275, 2364, 2389, 2676, 2725, 2977, 2992, 3001]. Electric [301]. Electrical [341, 2409]. Electronic [181, 251, 574, 698, 699, 701, 1045, 1310, 1992, 2744, 3069]. Electronics [2295]. Elements [767]. Elephant [1504]. Elephant-Sized [1504]. Eliciting [2470]. Eliminating [2040]. Elizabeth [1233]. Elsewhere [2228, 2715, 2735, 3057]. Elusive [1477]. Email [1162]. Embedded [82, 90, 676, 986, 1530, 1788, 1992, 2336, 2590]. Embodied [963, 2036]. Embracing [736, 2460]. Emergence [126, 348, 1041, 2156]. Emergency [309, 1893–1895]. Emerging [29, 1135, 1160, 2493, 3016]. eMMA [2667]. Emotion [2541]. Emotional [2696]. Emotions [2776]. Empirical [30, 277, 868, 936, 1123]. Empower [1367, 2334]. Empowering [119, 871]. Empowerment [3063]. Enable [1719]. Enabled [3005, 3083]. Enables [1837]. Enabling [682, 824, 858, 1042, 1410, 1501, 1655, 1968, 2420, 2742]. Encounters [523]. Encrypted [1836]. End [443, 1163, 1273, 2563]. End-to-End [443, 1163, 1273, 2563]. Ending [175, 1012].

Endless [1705, 1948]. Endoscopy [2598]. Endoscopy-Image-Based [2598]. Endowing [2518]. Enemy [2448]. Energy [18, 61, 66, 89, 125, 214, 301, 313, 396, 443, 530, 654, 657, 750, 798, 1046, 1157, 1163, 1216, 1220, 1353, 1750, 1812, 2066, 2068, 2071, 2125, 2354, 2560, 3066]. Energy-Efficient [125, 1157, 1750, 1812, 2066, 2125]. Energy-Harvesting [654]. Energy-Proportional [89, 2354]. Energy-Saving [214]. Enforced [2063]. Enforcement [663]. Enforcing [511]. Engagement [2128]. Engaging [2528, 2972]. Engine [11, 137, 1687]. Engineer [2318]. Engineering [62, 145, 255, 259, 268, 292, 301, 316, 411, 413– 417, 647, 679, 769, 770, 914, 984, 1043, 1071, 1130, 1213, 1227, 1243, 1274, 1620, 1625, 1667, 1731, 1745, 1779, 2003, 2102, 2136, 2299, 2317, 2444, 2517, 2519, 2966]. Engineers [1004, 2320]. Engines [2328]. English [57]. Enhance [328]. Enhanced [1070, 1641, 2240]. Enhancing [483, 990, 1312, 1518]. Enlightens [2630]. Enough [274, 367, 1974]. Enriched [684]. Ensemble [1714]. Ensuring [1438, 1771]. Enterprise [236, 570, 824, 1094, 1812, 2212]. Enterprise-Wide [824]. Entertain [2376]. Entities [2003]. Entity [671]. Entrepreneur [336, 1881, 1917]. Entrepreneurial [293, 294]. Entropy [1073, 2044]. Entry [1596]. Environment [240, 372, 477, 727, 824, 1361, 1723, 2262]. Environments [457, 767, 768, 3035]. Envisioning [2334, 2448]. Envy [205]. Epidemics [867]. Epsilon [1700, 2363]. Equal [310, 829]. Equifax [2473]. Equity [2103, 2549]. Era [1460, 1717, 1808, 2103, 2247, 2289, 2646, 2881]. Errant [944, 1170, 1511]. Erratum [2762, 2855, 3107]. Error [956, 2475]. Error-Resilient [2475]. Escalating [1018]. ESCO [1309]. Essentials [156]. Estate [2508]. Estimation [422, 1249, 2902].

Ethernet [831, 1009, 3044]. Ethical [775, 2276]. Ethics [511, 745, 1758, 2823]. EU [797, 959, 2743]. Eucalyptus [303]. Europe [24, 1304, 1309]. Evading [1850]. Evaluate [236, 657]. Evaluating [1810]. Evaluation [160, 887, 1123, 2521]. Evasion [843]. Event [144, 725, 2174, 2758]. Events [1465, 2097]. Ever [1974, 2615]. Everyone [420, 683, 858, 3106]. Everything [273, 405, 2287]. Evidence [796, 3067]. Evidence-Based [3067]. Evolution [84, 148, 595, 1102, 1313, 1336, 2107, 2277, 2384, 2643]. Evolutionary [414]. Evolve [85]. Evolving [81–83, 637, 1260, 1988, 2706, 2824]. Exact [1643]. Exactly [1395]. Exaggerated [312]. Example [2477, 3065]. Examples [1276, 1382, 1962]. Exascale [234, 431, 432, 434, 490, 696, 1713, 1988, 2070]. Excellence [1700, 2363, 2567, 2608, 2751, 3079]. Exchange [3026]. Exciting [869]. Execution [1046, 1654]. Executive [1506]. Exhausted [856]. existing [2765]. Expand [1555]. Expanding [1187, 1728, 2692]. Expansion [137]. Expect [1395]. Expected [1395]. Experience [292, 300, 339, 937, 1099, 1284, 1842, 2502]. Experienced [344]. Experiences [332, 1221]. Experiential [744]. Experimental [522]. Experiments [164, 566, 2724]. Expert [3000]. Experts [1083, 1124, 1658, 1679, 1800, 1904, 2131]. Explainable [2694, 2762]. Explained [2914]. Exploiting [2592]. Exploration [415, 883, 1607]. Explore [251, 1704]. Exploring [1932, 2292, 2898]. Exposure [236, 1212]. Express [3099]. Expressing [359]. Expression [48, 1586]. Expressions [1382, 2612]. Expressive [1584, 1585, 1588, 1620]. Extending [1808, 2243]. Extends [1523]. Extensible [1065]. Extensions [1147]. External [1361]. Externalities [1072]. Extract [1466].

Extract-Transform-Load [1466]. Extreme [2413]. Eye [2181, 2876]. Eyeglasses [2183].

F2654hD4 [2046]. Fabric [44, 275].

Fabrication [726]. Face

[28, 79, 280, 297, 837, 913, 991, 2242].

Facebook [664, 2011, 2045, 2672]. Facial [663, 2612]. Facilitate [1536]. Facing [793]. Fact [749]. Factor [1020]. Factorizing [725]. Factors [73, 962, 2096]. Faculty [2273]. Fagin [599]. Failed [591]. Failures [1190, 2449]. Fake [2247, 2389]. Fallout [1581]. False [1971]. Familiar [505, 2867]. Families [676]. Family [1251]. FAQs [1901]. Far [2237, 2996]. Far-Away [2237]. Farewell [1562, 1590]. Farming [3026]. Fashion [3093]. Fast [48, 1966]. Faster [249, 520]. Fatigue [1103]. Fault [591, 1184, 2002]. Faults [2651]. Favorite [2554, 2583, 2635, 2657]. FBI [592]. FCC [2574]. Fear [1353, 2612]. Feast [1255]. Feature [982]. Features [1330]. Federal [1500, 2467]. Federated [724, 1145]. Federation [483, 1536, 2107]. Feed [435]. Feedback [313, 334]. Fellows [1801]. Female [789]. Fernbach [439, 2137, 2167, 2214, 2567, 2608]. Fi [190, 430, 520, 2158]. Fibrillation [3062]. Fiction [749, 848, 1418, 1533, 1742, 1789, 1943, 2009]. Field [2200, 2929]. Fielding [1551]. Fifteen [2140]. Fiftieth [440]. Fight [132]. Fighting [78]. File [719, 2330]. Filtering [56, 2646]. Final [393]. Finally [5, 326, 1937]. Finance [216, 217, 2379, 2382]. Financial [11, 488]. Find [289, 605]. Finding [151, 2093, 2598]. Findings [1787]. Fine [2790]. Fine-Grained [2790]. Fingerprint [27]. Fingerprints [486]. Fingertips [467, 1098]. Finish [2321]. Firefox [1189]. First [350, 419, 613, 856, 1012, 1089, 1985, 2179, 2213, 2540]. FISMA [1285]. Fit [2024, 2025, 2057, 2086, 2087, 2113, 2114, 2143, 2145, 2173, 2175, 2194, 2196, 2224, 2253, 2309, 2314, 2315, 2397]. Fitness [2206]. Fitting [157]. Five [161, 800]. Fixing [1033]. Flagship [473, 2136]. Flash [905]. Flat [163]. Flaws [454, 2093, 2507]. Flesh [137]. Flexible [1475, 2454]. Floating [252, 1924]. Floating-Gate [252]. Floriani [2747]. Florida [2650]. Flow [507, 1014, 1997, 2235, 2267]. Flying [309]. Focus [1169, 1467, 1509, 1531, 1560, 1648, 1737, 1774, 1813, 1980, 2154, 2204]. Foe [3005]. Fog [1536, 1587, 2010, 2104, 2178, 2390, 3026]. Fog-Computing [2390]. Fog-to-Cloud [3026]. Folksonomy [321]. Follow [3110]. Following [792, 1440]. Fong [1233]. Food [2527, 3103]. Footprinting [329]. Force [641]. Forces [38]. Forecasting [308]. Forefront [834]. Foreign [821]. Forensic [320, 721]. Forensics [718, 720, 1844]. Foreseeable [2380]. Forget [2213]. Form [309, 2242]. Formal [8, 219, 259, 1909, 2001, 2718]. Formality [1313]. Format [1641]. Formation [321]. Forming [2572]. Forms [179]. Forty [831]. Forward [26, 210, 450, 1061, 1067, 1325, 1364, 1397, 1415, 1428, 1528, 1602, 1608, 1667, 1675, 1722, 1768, 1799, 1827, 1855, 1883, 1890, 1933, 1963, 2013, 2055, 2085, 2112, 2142, 2172, 2193, 2223, 2252, 2308, 2331, 2371, 2399, 2648]. Forward-Looking [1667]. Foster [317, 1332, 1557, 2325]. Fostering [782]. Foundation [145, 670, 756, 1779]. Foundations [17, 193, 2811]. Four [38, 1787, 2100, 2363, 2815]. Fourth [574, 660]. Frame [2876]. Framework [768, 986, 1065, 1310, 1466, 1814, 2237, 2266, 2387, 2467]. Frameworks [1810, 2813]. Franca [2368]. Fraud [488, 2763]. Free [535, 1557, 1969, 2723, 2914]. FreeBSD [2521]. Frequency [2720, 2987]. Fresh [2527]. Friend [3005]. Friends [489]. Front [902, 968, 997, 1054, 1082, 1109, 1141, 1172, 1200, 1394, 1426, 1454, 1485, 1512, 1544, 1571, 1601, 1637, 1673, 1706, 1740, 1766, 1797, 1826, 1854, 1882, 1918, 1949, 1979, 2021, 2054, 2084, 2111, 2141, 2171, 2222, 2250, 2281, 2307, 2342, 2369, 2398, 2426, 2482, 2509, 2531, 2552, 2581, 2605, 2633, 2655, 2680, 2709, 2731, 2732, 2770, 2796, 2828, 2829, 2857, 2886, 2921, 2948, 2952, 2983, 3011, 3049, 3075]. Frontier [393, 2692, 2706]. Frontiers [1187, 2066, 2238, 2357, 2496]. FTC [1166]. Fuel [1110, 1228, 1354, 1374]. Fueling [1242]. Fujii [1391]. Fulfilling [9]. Full [930, 2044]. Fun [793]. Function [343, 1339, 2101]. Functional [2073]. Fundamentals [655]. Fusion [702, 1864]. Future [3, 54, 90, 128, 144, 186, 225, 230, 254, 281, 311, 329, 405, 472, 479, 491, 520, 570, 572, 743, 770, 893, 909, 955, 1039, 1040, 1043, 1045, 1133, 1195, 1218, 1247, 1248, 1322, 1380, 1642, 1727, 1760, 1791, 1816, 2034, 2080, 2095, 2118, 2187, 2230, 2231, 2271, 2330, 2334, 2405, 2416, 2450, 2474, 2524, 2525, 2528, 2760, 2765, 2769, 2819, 2869, 2934]. Futures [1819]. Futurology [3004].

Gadget [2723]. Gadget-Free [2723]. Gaining [867]. Galaxy [1269, 1916, 1950]. Gallup [1417]. Gamble [1420]. Gambling [1818]. Game [232, 510, 825, 858, 1019, 1074, 1251, 1388, 1389, 2298, 2564]. Game-Model-Based [2564]. Gamers [1316]. Games [686, 820, 1132]. Gang [401]. Gap [234, 464, 2324, 2699]. Gaps [1562, 1590]. Gara [199]. Gate [252, 1755, 1997]. Gate-Level [1997]. Gateway [159]. Gateways [1345]. Gaudiot [1761]. Gauging [1285]. Gen [2503]. Gender [782, 784, 786, 791]. Gender-Based [784]. Gene [199]. General [2845]. Generalized [796]. Generalizing [530, 2336]. Generation [16, 365, 473, 588, 635, 765, 785, 818, 904, 1411, 1444, 1596, 2033, 2320, 2666, 2881, 2935]. Generative [3005]. Generator [334, 1756]. Genesis [2728]. Genetics [1213]. Genome [525, 1044, 1443]. Genomes [526].

Genomics [590]. Geocomputational [1723]. Geography [1452]. Geologic [308]. Gerrymander [2106]. Gestural [545, 950]. Gesture [185]. Gesture-Based [185]. Get [757, 1495, 1545, 1572, 1720]. Getting [559, 912, 3038]. Ghost [2106]. Giant [1325, 1364, 1397, 1428]. Giants [704, 2338]. Gift [2554, 2583, 2635, 2657]. Gigabit [190]. GitHub [1543]. Give [427, 1351, 2554, 2583, 2635, 2657]. Give-and-Take [427]. Given [147]. Glasses [229]. Glimpse [2416]. Glitters [1138, 2472]. Global [64, 490, 819, 1013, 1880, 2170, 2306, 2394, 2425]. Globalizing [1191]. Globally [661, 1162]. Globe [2588]. Globus [1332]. Gluttony [93]. Gnome [359]. Go [70, 313, 683, 759, 927, 1853, 2568, 2960, 3106]. Go/No [2960]. Goal [1477]. goals [2025, 2057, 2315, 2397]. Goes [810, 1013]. Gold [1138, 2472]. Gonzo [1758]. Good [95, 944, 1537, 1974, 2362, 2767, 3020]. Google [101, 294]. Gordon [854]. Got [2507]. Governance [862, 3038]. Government [197, 1303–1305, 1308, 1784, 2048, 2447]. Governments [2737]. GPGPU [634]. GPS [430, 771]. GPU [633]. GPUs [759, 1719]. Grade [1441]. Graduate [174]. Graduates [182]. Graham [441]. Grained [2790]. Grand [1713, 2504, 3095]. Grandest [2602]. Graph [528, 654, 885, 1464, 1615, 1618]. Graphical [1959]. Graphics [1387, 2789, 2849, 2866, 2909, 2971, 3022]. Grassroots [2443]. Gravitational [1719]. Gravity [1270, 3104]. Great [800, 1417, 2652, 2840]. Greatly [312]. Green [53, 165, 205, 402, 573, 652, 655, 657, 798, 893, 1419]. Greene [336]. GreenLight [7]. Greenpower [2321]. Grid [123, 139, 317, 553, 1723, 2624]. Grids [324, 2240, 2478, 2720]. Grier [461]. Ground [559]. Groundwork [329]. Group [1019, 1480]. Growing [112, 499]. Grows [1369]. Growth [2404]. Guest [81, 499, 522, 860, 904, 1008, 1064, 1093, 1119, 1152, 1238, 1255, 1375, 1437, 1553, 1612, 1647, 1990, 2066, 2123, 2717, 2737]. Guidance [2814, 2842]. Guided [1349]. Guidelines [422, 547, 1667, 1731]. Guido [1433, 1460].

H [1972, 2071]. H-1B [1972]. H-EARtH [2071]. Habit [407]. Habitat [100]. Habits [2170]. Hack [800, 1446]. Hackers [132, 872, 2644]. Hacking [2121, 2389, 2478, 2549]. Hafner [877]. Half [1226]. Half-Told [1226]. Hamilton [169]. Hand [466, 1146, 1852, 2165, 2216]. Handcuffs [2246]. Handheld [610]. Handle [700]. Handling [2328]. Hands [133, 344]. happen [1511]. Happened [153, 2001, 2270]. Haptic [2285]. Harbor [1928]. Hard [2444, 3110]. Hardin [1060]. Hardness [2268]. Hardware [147, 176, 295, 355, 432, 954, 960, 1997, 2063, 2267, 2292, 2296, 2330, 2384, 2661, 2892]. Hardware-Enforced [2063]. Hardware/Software [960]. Harlan [1274, 2593, 2634, 2645, 2683, 2697]. Harmful [1872]. Harmonization [2939]. Harmonizing [338]. Harnessing [1101]. Harvesting [654]. Harvey [789]. Hashes [719]. Hashtag [944, 1170, 1511]. Having [793]. Hazards [1072, 2618, 3101]. HCI [826]. Head [1554]. Head-Up [1554]. Headers [1339]. Health [121, 181, 196, 698– 700, 742, 1305, 1347, 1438, 1927, 2094– 2098, 2260, 2539, 2647, 2668, 2701, 2842, 3061]. Health-Active [3061]. Healthcare [118, 119, 701–703, 742, 841, 987, 1153, 1441, 1442, 1468, 2099, 2302, 2332, 2666, 2744]. Healthier [3059]. Healthy [3061]. Heart [209]. Heartbleed [1250]. Heights [1679, 1800, 1904, 2131]. Hell [166]. Help [85, 380, 1019, 2075, 2206, 2543]. Helping [2010, 2322]. Here [260, 753, 1039, 1324, 1363, 1396, 1642]. Heritage [352]. Hescott [278]. Heterogeneity [2292]. Heterogeneous [316, 372, 1415, 2071]. HHS [699]. Hidden [746, 1738, 2776]. Hiding [1416, 3027]. Hierarchical [2721, 2722, 2804]. High [218, 342, 424, 682, 844, 1223, 1270, 1480, 1718, 1723, 1780, 2233, 2303, 2323, 2444, 2567, 2569, 2608, 2658, 2933]. High-Altitude [2323]. High-Availability [2303]. High-Confidence [1780]. High-level [2233]. High-Performance [218, 682, 844, 1270, 1480, 1718, 2569]. High-Throughput [1723]. Higher [157, 222, 983]. Higher-Order [983]. Highlights [1, 94, 113, 129, 150, 167, 187, 207, 224, 245, 264, 284, 361, 383, 406, 426, 449, 471, 492, 516, 536, 558, 580, 602, 624, 646, 668, 692, 713, 734, 754, 776, 806, 830, 853, 876, 898, 922, 945, 972, 1001, 1031, 1058, 1086, 1113, 1144, 1176, 1203, 1231, 1259, 1297, 1329, 1368, 1402, 1431, 1458, 1490, 1517, 1549, 1578, 1606, 1640, 1677, 1709, 1744, 1775, 1803, 1832, 1859, 1888, 1922, 1954, 1987, 2030, 2062, 2092, 2120, 2152, 2180, 2201, 2256, 2313, 2375, 2403]. Highlights [305, 323, 1135, 2284, 2432, 2463, 2490, 2515, 2535, 2558, 2586, 2611, 2639, 2662, 2686, 2752, 2778, 2805, 2836, 2863, 2894, 2927, 2957, 2988, 3015, 3081]. Highly [780]. Hiring [1729]. Hiscox [1887]. Historical [867]. Histories [684]. History [368, 517, 612, 715, 924, 1791, 2791, 2825, 2826, 2847, 2868, 2913, 2932, 2961, 3045, 3108]. Hit [3079]. Hoc [1177]. Hold [134, 462]. Holders [336]. Holding [2781]. Holds [1040]. Holes [1932]. Holistic [618, 656, 956, 2327]. Hollar [924]. Home [888, 1345, 1908, 1959, 2393]. Homes [2097]. Honest [427]. Honors [161, 484, 577, 599]. Horizontal [508]. Horrocks [973]. Hospital [1134]. Hot [757]. House [1028, 1034, 1053, 1055, 1080, 1081, 1083, 1108, 1110, 1124, 1139, 1140, 1146, 1149, 1169, 1171, 1173, 1182, 1186, 1194, 1228, 1245, 1256, 1288, 1295, 1320, 1321, 1331, 1340, 1342, 1344, 1358, 1362, 1370, 1374, 1377, 1383, 1440, 1455, 1467, 1471, 1487, 1503, 1509, 1514, 1515, 1525, 1527, 1531, 1546, 1547, 1558, 1560, 1567, 1573, 1574, 1577, 1582, 1608, 1613, 1617, 1621, 1627, 1629, 1633, 1635, 1636, 1638, 1683, 1688, 1690, 1693, 1698, 1702, 1704, 1707, 1715, 1720, 1722, 1735, 1737, 1739, 1752, 1767, 1768]. House [1292, 1495, 1545, 1572, 1679, 1700, 1774, 1798– 1800, 1813, 1817, 1825, 1827, 1828, 1838, 1852, 1855–1857, 1865, 1881, 1883–1885, 1890, 1904, 1906, 1912, 1917, 1919, 1930, 1933, 1935, 1936, 1941, 1942, 1951, 1961, 1963, 1967, 1970, 1980– 1982, 2045, 2087, 2110, 2112, 2113, 2145, 2150, 2154, 2164, 2167, 2169, 2172– 2174, 2182, 2192, 2214]. Houweling [1115]. However [1641]. HP [203]. HPC [1726, 2068]. HTML [60]. Huawei [3042]. Human [127, 137, 243, 444, 743, 763, 764, 1120, 1385, 1732, 1842, 1964, 2038, 2158, 2184, 2385, 2404, 2418, 2596, 2653, 2694, 2742, 2762, 2876]. Human-Body-Mediated [2184]. Human-Centered [1385, 1964]. Human-Centric [2742]. Human-in-the-Loop [743]. Human-Machine [2038]. Human-Understandable [2694, 2762]. Humanitarian [2422]. Humans [28, 1772, 2008, 2776, 3067]. Hurdle [402]. Hybrid [834, 1196, 1751, 2595, 2740, 2901]. Hygiene [2843]. Hyperbole [1446].

I/O [909]. I/Ocloud [2500]. IaaS [724]. Ian [973, 1332]. Ibis [138]. IBM [146, 249, 367–370, 1792]. IC [1999]. IC2E [1227]. ICT [174]. Idea [246]. Ideas [507]. Identical [358, 799]. Identification [29, 2745]. Identified [454]. Identifies [430]. Identify [2209, 2543]. Identifying [176, 215, 525, 1094, 1624]. Identity [243, 275, 300, 358, 488, 2473]. IEC [2136]. IEEE [140, 141, 374, 375, 419, 642, 643, 689, 918, 919, 971, 985, 998, 1022, 1025, 1028, 1034, 1053, 1055, 1080, 1139, 1182, 1194, 1227, 1252, 1253, 1264, 1274, 1291, 1296, 1306, 1321, 1323, 1331, 1340, 1344, 1358, 1362, 1370, 1377, 1413, 1427, 1455, 1470, 1486, 1503, 1513, 1514, 1567, 1582, 1610, 1628, 1630, 1638, 1650, 1652, 1661, 1664, 1675, 1683, 1702, 1767, 1798, 1801, 1809, 1817, 1891, 1900, 1912, 1942, 1956, 2022, 2121, 2150, 2185, 2587, 2606, 2617, 2658, 2681, 2695, 2710, 2747, 3079, 3112]. IEEE [1883, 1890, 1906, 1933, 1963, 1967, 1985, 1995, 1998, 2007, 2012–2016, 2028, 2055, 2060, 2067, 2085, 2087, 2090, 2112, 2114, 2117, 2142, 2145, 2172, 2175, 2191, 2193, 2196, 2199, 2220, 2223, 2252, 2308, 2343, 2351, 2358, 2364– 2366, 2371, 2391, 2399, 2421, 2427, 2437, 2439, 2452, 2453, 2455, 2468, 2483, 2497, 2510, 2532, 2536, 2553, 2554, 2559, 2582, 2583, 2593, 2604, 2622, 2628, 2634, 2640, 2645, 2675–2678, 2682, 2683, 2689, 2697, 2702, 2705, 2711, 2730, 2738]. IEEE [293, 373, 1135, 1160, 1283, 1792, 1846, 1924, 2136, 2424, 2484, 2489, 2491, 2511, 2571, 2635, 2657, 2671, 2751, 2756, 2771– 2773, 2789, 2791, 2798, 2799, 2825, 2826, 2832, 2833, 2847, 2849, 2858, 2859, 2864, 2868, 2887– 2889, 2907, 2909, 2913, 2916, 2919, 2922, 2923, 2932, 2933, 2947, 2949– 2951, 2953, 2961, 2971, 2972, 2976, 2977, 2979– 2982, 2992, 3001, 3007, 3009, 3010, 3012, 3022, 3025, 3032, 3042, 3043, 3045, 3047, 3048, 3050, 3068, 3073, 3074, 3080, 3082, 3089, 3090, 3108, 3113]. IEEE [1076, 2755, 2764, 2766, 2777, 2780, 2786, 2793, 2794, 2812, 2816, 2817, 2821, 2822, 2837, 2839, 2841, 2844, 2853, 2856, 2870, 2879, 2882– 2884, 2893, 2895, 2897, 2900, 2903, 2915, 2918, 2930, 2936, 2938, 2940, 2942, 2967, 2969, 2973, 2975, 2978, 2989, 2995, 2997, 2999, 3017, 3028, 3030, 3036, 3039, 3041, 3054, 3056, 3058, 3060, 3070, 3072, 3076, 3084, 3086, 3088, 3091, 3109, 3111, 3114]. IEEE-CS [293, 2606, 2617, 2681, 2710]. IEEE/IBM [1792]. IEEE’s [1985]. If [446, 560, 1395]. Ignorance [211]. II [495]. I’m [2109]. Image [576, 2598]. Imagery [1134]. Images [1783, 1968]. Imagination [1110, 1228, 1354, 1374, 1765]. Imagineering [1192]. Imaging [1129, 2073, 2745]. Imagining [479, 1350, 2618]. Immersive [292, 1211, 1214]. Immortal [1849]. Immune [531]. Impact [499, 705, 953, 957, 959, 1008, 1205, 1285, 1401, 1941, 2377, 2380, 2542, 2565, 2845, 2998, 3021]. Impacting [962, 2272]. Impaired [1125, 2181]. Imperative [2417]. Implanted [132]. Implementing [714, 1500, 1618]. Implications [391, 1010, 1591]. Importance [1073, 2327, 2385, 2876]. Important [1492, 1520, 3037]. Impossible [512, 2161]. Impromptu [1895]. Improve [120, 154, 1046, 1684, 1894, 2068, 2264, 2292, 2302]. Improved [2842]. Improvement [1002, 1897]. Improving [280, 786, 841, 1964, 2560]. Improvisation [1915]. In-Memory [673, 1464, 1614]. Inc. [1440, 2011, 2017, 2045]. Included [823]. Inclusion [2724]. Inconsistencies [1232]. Increase [310]. Increasing [551, 789, 1579]. Incremental [1343]. Incubator [1241]. Independent [510, 1691]. Index [1417]. Indie [2390]. Indigenous [705]. Indirect [3035]. Individual [2099, 2209]. Indoor [2674, 2934]. Induction [2943]. Industrial [191, 1671, 2726]. Industrial-Plant [191]. Industrie [2249]. Industry [241, 1385, 1729, 2322, 2380, 2447, 2577, 2630, 2917, 2935, 2972]. Industry-Academia [241]. Industry-Backed [2322]. Inevitable [225, 2699]. Inference [2944]. Inferences [2292]. Inferring [1556, 3035]. InfiniBand [433]. Influences [1096]. Influential [813]. Informatics [121, 1044, 1047, 1442, 1732, 1894]. Information [30, 112, 126, 273, 505, 598, 703, 979, 985, 987, 1022, 1073, 1093, 1095, 1098, 1194, 1416, 1468, 1593, 1625, 1656, 1900, 1907, 1997, 2007, 2235, 2237, 2239, 2351, 2385, 2391, 2493, 2496, 2532, 2559, 2640, 2730, 2738, 2798, 2832, 2859, 2888, 2923, 2947, 2981, 3027, 3047, 3073]. Information-Centric [2237]. Information-Flow [1997, 2235]. Informed [1944, 2673]. Infrastructure [7, 195, 481, 629, 703, 814, 1876, 1898, 2332, 2390, 2527, 2624]. Infrastructures [724, 883, 1965, 2500]. Infringement [538, 1711]. Ingenuity [1504]. Inhabit [1819]. Initiative [2273, 2746, 3023, 3055]. Initiatives [1591, 2569, 2739, 3003]. Injection [795, 1217]. Inner [359, 792]. Innovating [2083]. Innovation [2, 37, 111, 203, 292, 294, 330, 665, 705, 801, 861, 1238, 1239, 1242, 1445, 2265, 2296, 2324, 2381]. Innovations [2850, 3006]. Input [133, 185]. Inria [2350]. Insanity [1729]. Insecure [1835, 1836, 2809]. Insider [961]. Insight [2652]. Insights [867, 1533, 1847, 1909]. Inspired [2037, 2072, 2266]. Inspiring [2320]. Instagram [2616]. Instant [1139, 1264, 1344, 1455, 1486, 1513]. Instruction [1133, 1195, 1959, 2462]. Instructional [1937]. Instrumentation [2336]. Intangible [579]. Integrate [931]. Integrated [175, 295, 333, 2236, 2267, 3103]. Integrating [73, 701, 1155, 1466, 2263]. Integration [204, 657, 983, 986, 1442, 2438, 2698]. Integrity [990, 1356, 2725]. Intel [2231, 2545]. IntellBatt [50]. Intellectual [476, 495, 738, 1116]. Intelligence [34, 116, 219, 348, 884, 1114, 1754, 1955, 2037, 2276, 2692, 2696, 2974, 3031, 3064, 3065, 3095]. Intelligent [378, 532, 1439, 2461, 2474, 2479, 2793, 2882, 2915, 3017, 3058, 3091]. Intelligently [771]. Intensive [148, 572, 676, 728, 865, 1204, 1384, 1614, 1962]. Intent [418]. Intentions [12]. Intents [2964]. Interact [3094]. Interacting [822]. Interaction [110, 185, 524, 542, 545, 546, 840, 911, 1067, 1098, 2038, 2123, 2129, 2404, 2408]. Interactions [544, 963]. Interactive [331, 379, 508, 549, 564, 565, 567, 841, 915, 2233, 2409, 2758]. Interconnected [2440]. Interconnection [234, 1579]. Interconnects [953]. Intercontinental [139]. Interdependent [2241]. Interdevice [2128]. Interdisciplinary [455, 2265, 2818].

Interface

[49, 1875, 1957, 2157, 2551, 2580, 2772]. Interfaces [543, 584, 610, 950, 1021, 1042, 1686, 1790, 1958, 1962, 2407]. International [1227, 1292, 1302, 1338, 1362, 1569, 1638, 2022, 2444, 2502, 2569, 2650]. Internationalists [515]. Internet [172, 269, 314, 394, 453, 571, 587, 615, 761, 762, 765, 766, 816, 877, 1033, 1089, 1123, 1147, 1192, 1261, 1410, 1647, 1649, 1820, 1876, 1905, 1938, 2010, 2074, 2129, 2130, 2132, 2162, 2189, 2266, 2287, 2291, 2293, 2294, 2332, 2390, 2425, 2563, 2594, 2719, 2726, 2846, 2910, 2962, 2968, 3003, 3082, 3083, 3103]. Internet-of-Things [765]. Internetware [587]. Interoperability [1303, 1304, 1307–1309, 2438, 2706]. Interoperable [1414, 2729]. Intersect [1193]. Interventions [2668]. Interview [158, 283]. Intimidation [1020]. Intrinsic [2540]. Introducing [1502]. Introduction [81, 499, 522, 860, 904, 1008, 1064, 1093, 1119, 1152, 1238, 1375, 1437, 1553, 1612, 1647, 1990, 2066, 2123, 2717, 2737]. Introductory [2262]. Intrusion [2646, 2788]. Intrusions [78]. Intuit [1440, 2017]. Invent [2126, 2169]. Invented [405]. Inventing [693]. Inventory [2523]. Invents [694, 832]. Investigating [1095]. Investigation [30, 1839]. Investigations [817]. Investigators [721]. Investing [211]. Invisibility [71]. Invisible [547, 550]. iOS [1416]. IoT [764, 1624, 1651, 1653, 2002, 2187, 2298, 2333, 2360, 2420, 2436, 2500, 2527, 2575, 2576, 2597, 2627, 2642–2644, 2646– 2648, 2674, 2745, 2792, 2965, 3044]. IoT-Based [2527, 2647]. IP [1061, 1070, 1206, 1299, 1406, 1492, 1520, 2267]. iPlant [435]. IPv4 [1070]. IPv6 [387, 843]. IRI [1413]. IRI-2015 [1413]. Iris [26, 1164]. Iron [696]. Irregular [635, 1612, 1614, 1616]. Irregularities [2359]. ISA [446]. Isn’t [664]. ISO [2136]. ISO/IEC/IEEE [2136]. Issue [893, 1688, 1735, 2110, 2124, 2984, 3034]. Issues [768, 1406, 1644, 1835, 2179, 2501, 2760]. ITiP [2064]. ITO [459]. Itself [3033]. ITSy [979].

Jacobson [559, 737]. Jail [1019]. Jain [1609]. January [2151]. Japan [1048, 2739, 2743]. Java [487, 685]. JavaScript [494, 537]. Jimmy [802, 992]. Job [1169, 1309, 1467, 1509, 1531, 1560, 1648, 1737, 1774, 1813, 1980, 2154, 2204, 2370, 2395, 2428, 2456, 2485, 2512, 2578, 2614, 2700, 2799, 2949, 2953, 3048]. Jobs [474, 1417, 1726, 2773, 2833, 2889, 2916, 3010, 3074]. John [924, 1519]. Joint [2587, 2695, 2934]. Jones [2077]. Joseph [1060]. Joules [402]. Journalism [845]. Journey [209, 2118]. Joy [2612]. JQuery [1519]. Judson [1387]. Juliet [685]. July [2316]. June [2286]. Junk [1873]. Just [664, 1052].

K-12 [786]. Kanai [68, 317]. Karlsson [32]. Kasahara [2138]. Katie [877]. KDD [2788]. Keep [132, 1293, 2505]. Keeping [2425]. Keeps [23, 540, 3067]. Ken [68, 2137, 2167, 2214, 2567, 2608]. Kennedy [441, 731, 2137, 2167, 2214, 2567, 2608]. Key [513, 962, 2046, 2096]. Keyboard [542, 544]. Keyboards [382]. Khan [1405]. Kickstarting [2296]. Kill [446]. Killer [2414, 2523]. Kilometre [1273]. Kind [523]. Kinder [913]. Kinect [300, 339]. Kinecting [963]. Kinetic [2701]. Kinetic-Powered [2701]. KISS [987]. Kleinrock [899, 1089]. Know [182, 345, 364, 386, 900, 2824]. Know-How [2824]. Know-What [2824]. KnowLang [1447]. Knowledge [34, 137, 378, 1094, 1447, 1829, 2130, 2344, 2372, 2400, 2411, 2554, 2583, 2635, 2657, 3000]. Known [350]. Koller [626]. Kuck [336].

Labs [203]. LADEE [1539]. LAN [1223]. LAN/MAN [1223]. Land [2644]. Landmarks [2408]. Landscape [270, 677, 1048, 1868]. Landscapes [351]. Language [494, 589, 647, 821, 2093, 2233, 2359, 2377, 2624]. Languages [184, 607, 1191]. LANs [1260]. Large [457, 837, 885, 1502, 1615, 1616, 1751, 1968, 2790, 3064]. Large-Scale [457, 885, 1615, 1616, 1751, 2790, 3064]. Larry [807]. Latency [1465]. Later [567, 704, 2259, 2269, 2433]. Latest [2473]. Launch [2887]. Laundering [1248]. Law [163, 530, 663, 1009, 1011, 1012, 1555, 1589, 1749, 1808, 2188, 2289, 2361, 2648, 2787, 3110]. Laws [1008, 1116]. Lawyer [560]. Layer [980]. Lead [607]. Leadership [2, 1190, 1652, 1671, 2199, 2358]. Leading [1679, 1800, 1904, 2131]. Leap [1325, 1364, 1397, 1428]. Learn [575, 821, 1546, 1573, 1604, 1674, 1707, 1741]. Learned [1275, 1618, 3029]. Learning [62, 302, 424, 557, 589, 764, 933, 1043, 1132, 1491, 1665, 1710, 1786, 1892, 1939, 2080, 2098, 2121, 2261, 2264, 2319, 2325, 2420, 2454, 2502, 2525, 2590, 2592, 2594–2596, 2598, 2621, 2649, 2693, 2741, 2760, 2809, 2873, 2875, 2964, 2966, 2987, 3006, 3033, 3062, 3095, 3098]. Lease [65]. Lecture [1087]. Ledger [2381, 2795]. Legacies [1849]. Legacy [1127, 2048]. Legal [63, 73, 925, 1004, 1644]. Legislating [1692]. Legos [2187]. Leisure [384]. Len [899, 1089]. Lens [2569]. Lenses [1164]. Lerdorf [693]. Lesions [2598]. Less [978, 1495, 1545, 1572]. Lessons [164, 299, 413, 1385, 1618, 3029]. Let [401, 1973, 2376]. Lets [2543]. Letters [40, 168, 208, 265, 285, 493, 603, 625, 669, 755, 855, 923, 946, 1088, 1150, 1181, 1208, 1265, 2557, 2622, 2689, 2830, 2831, 2871, 2919, 3012, 3068]. Level [101, 232, 727, 911, 1356, 1820, 1997, 2233, 2968]. Levels [1704]. Leveraging [120, 1120, 1419, 2873]. Liberated [634]. Libertarian [2382]. Libraries [79]. Library [1906, 1985, 2568]. License [1619, 1783, 2970]. Licenses [2914]. Lie [771, 1226]. Lies [906]. Life [97, 331, 446, 960, 1216, 1327, 1328, 1352, 1366, 1367, 1419, 1994, 2183, 2190, 2205, 2842]. Life-Cycle [1994]. Lifeloggers [2211].

Lifelong [2325]. Lifestyle [2206]. Lifestyles [3059]. Light [249, 289, 2245]. Lights [269, 429]. Lightweight [1899, 2413]. Like [1470, 1641]. Limited [796]. Limits [134, 819, 2000]. Line [259, 354, 677]. Linear [334, 1615]. Lines [675, 678, 679]. Lingua [2368]. Link [528, 1589, 3099]. Linked [1307]. Linkedin [1379, 1666]. Linking [67, 1303, 2814]. List [1718, 2714]. Lists [165]. Lithography [6]. Litter [2572]. Little [1504, 2295]. Live [22, 2263, 2809]. Lives [1417, 2260]. Living [2065, 2178, 3061]. Load [1466]. Loan [897]. Local [1013, 2170]. Localization [591, 681, 1013, 1831]. Locally [661]. Location [430, 810, 1440, 1710, 2674]. Location-Aware [1710]. Locations [459]. Log [2998]. Logic [548, 1874]. Logical [2944]. Logistics [3103]. Loihi [2545]. London [612]. Long [209, 566, 1839, 1898, 2135]. Long-Term [566, 1839]. Look [1470, 1528, 2130, 2935]. Lookahead [1217]. Looking [210, 311, 450, 1061, 1667, 2370, 2395, 2428, 2456, 2485, 2512, 2578, 2614, 2700]. Loop [743, 1772, 3067]. Losing [2543]. Loss [1730]. Lot [2295]. Lots [1134]. Love [712, 1275]. Low [1444, 1872, 2293, 2323, 2419, 2537, 2669, 3026]. Low-Cost [1444, 2323, 3026]. Low-Power [2537]. Lowly [902]. Loyal [209]. Lung [482]. Lying [1322].

M [2666]. M-Coaching [2666]. Mac [2521]. Machine [302, 1120, 1654, 1934, 2038, 2098, 2396, 2525, 2649, 2693, 2760, 2875, 2966, 3006, 3095, 3098]. Machinery [425]. Machines [1939, 2077]. Made [44, 1378]. Magazine [1080, 2117, 2782, 2874, 2880, 2882, 2883, 2895, 2897, 2909, 2915, 2918, 2924, 2938, 2951, 2967, 2969, 2976, 2989, 3030, 3041, 3054, 3058, 3060, 3072, 3082, 3084, 3086, 3091, 3109]. Magazines [1, 94, 113, 129, 150, 167, 187, 207, 224, 245, 264, 284, 361, 383, 406, 426, 449, 471, 492, 516, 536, 558, 580, 602, 624, 646, 668, 692, 713, 734, 754, 776, 806, 830, 853, 876, 898, 922, 945, 972, 1001, 1031, 1058, 1086, 1113, 1144, 1176, 1203, 1231, 1259, 1297, 1329, 1368, 1402, 1431, 1458, 1490, 1517, 1549, 1578, 1606, 1640, 1677, 1709, 1744, 1775, 1803, 1832, 1859, 1888, 1922, 1954, 1987, 2030, 2062, 2092, 2120, 2152, 2180, 2201, 2256, 2313, 2375, 2403]. Magazines [305, 323, 1039, 2284, 2432, 2463, 2490, 2515, 2535, 2558, 2586, 2611, 2639, 2662, 2686, 2752, 2778, 2805, 2836, 2863, 2894, 2927, 2957, 2988, 3015, 3081]. Magic [2246]. Mail [1147]. mails [186]. Maintaining [571, 2470, 2519]. Major [1650, 1912, 1942, 1967, 1998, 2191, 2220, 2343, 2976]. Make [213, 249, 367, 2206]. Maker [2546]. Makes [1537]. Makimoto [1010]. Making [550, 869, 1130, 1375, 1376, 1380, 1660, 2848, 3067, 3094]. Malice [2601]. Malicious [390]. Malware [658, 843, 1126, 1184, 1222, 2000]. MAN [1223]. Manage [680, 2667]. Management [84, 215, 217, 231, 333, 392, 418, 443, 503, 592, 746, 747, 849, 883, 982, 1353, 1439, 1588, 1868, 1894, 1897, 1934, 1994, 2071, 2295, 2405, 2668, 2722, 2783, 3066]. Managing [118, 395, 702, 1204, 1277, 1468, 1480, 1811, 1849, 2238, 2269, 2449, 2508]. Manhattan [992]. Manhunt [819]. Manipulation [1069, 1534]. Manning [529]. Many [1468]. Manycore [2649]. Mapping [1895, 2813, 2905]. Maps [447, 1831]. Marathon [991]. Marginalized [2623]. Market [288, 369, 423, 445, 1223, 1337, 1534]. Marketing [268, 2299]. Markets [306, 324, 1135, 3101]. Markup [2620]. Mashup [1021]. Mashups [665]. Mass [1315]. Masses [116, 480, 740]. Massimo [1035]. Massively [458, 631, 632, 635]. Masthead [969, 999, 1029, 1056, 1084, 1111, 1142, 1174, 1201, 1229, 1257, 1294, 1326, 1365, 1399, 1430, 1457, 1489, 1516, 1548, 1575, 1605, 1639, 1676, 1708, 1743, 1770, 1802, 1830, 1858, 1886, 1920, 1952, 1983, 2026, 2058, 2088, 2115, 2146, 2176, 2197, 2227, 2255, 2283, 2312, 2347, 2374, 2402, 2431, 2459, 2487, 2514, 2534, 2556, 2585, 2610, 2637, 2660, 2685, 2734, 2750, 2775, 2801, 2803, 2835, 2862, 2891, 2926, 2955, 2986, 3014, 3052, 3078]. Matching [27, 48, 320, 576, 1015, 1309, 1618]. Material [79, 2745]. Mathematical [372]. Mathematics [963]. Matter [529, 1355]. Matters [783, 1171, 1186, 1271]. Maturity [91, 256, 1307]. Mauchly [621, 2150, 2185, 2771]. Maximize [2128]. McDowell [296]. Me [1351]. Meanings [2396]. Means [700, 2052]. Measurement [67, 1102]. Measurements [1643, 1840]. Measures [2441]. Measuring [573, 1216, 1417]. Mechanisms [2720]. Media [178, 197, 744, 842, 891, 1017, 1036, 1094, 1848, 1907, 2128, 2495, 2572, 2668, 3027]. Media-based [2668]. Mediaprinting [215]. Mediated [193, 194, 198, 2184]. Medical [132, 543, 752, 938, 1129, 1441, 1780, 2706]. Medical-Grade [1441]. Medication [747, 2667]. Medicine [1437, 1443]. Medium [569, 1898]. Medium-Scale [1898]. Meet [174, 341, 749, 2260]. Meeting [2917]. Meets [411, 414, 416, 665, 709, 1341, 1626, 1759, 2496, 3083]. Megawatts [490]. Meltdown [2573]. Member [1906, 2751, 3079]. Members [13, 484, 1771, 2747]. Membership [971, 1025, 1055, 1171, 1186, 1271, 1675, 1883, 1890, 1933, 1963, 2013, 2024, 2028, 2055, 2085, 2086, 2112, 2113, 2142, 2143, 2172, 2173, 2193, 2194, 2223, 2224, 2252, 2253, 2308, 2309, 2314, 2344, 2371, 2372, 2399, 2400, 2411, 2554, 2583, 2635, 2657]. Memoriam [169]. Memories [2790]. Memory [251, 673, 904–908, 952, 956, 1040, 1159, 1279, 1464, 1614, 1654, 1753, 1934, 2203, 2356, 2498, 2537, 2560, 3099]. Memristive [251]. Memristor [1755]. Memristor-Based [1755]. Mental [130, 1122, 2209]. Mentoring [785]. Mercator [1033]. Merwin [1237, 1263, 1613, 1621, 1652, 1993, 2199, 2358, 2591, 2619, 2656, 2665, 2712, 2993]. meSchup [2440]. Message [2754]. Metaheuristics [2037]. Metamorphic [1931]. Metaphor [617, 1098]. Metascheduling [123]. Metcalfe [831, 1009]. Method [8, 2209, 2239]. Methodological [2359]. Methodology [2240]. Methods [259, 1909, 1964, 2001, 2076, 2235]. Metra [2480]. Metrics [357, 2270]. Meyer [88, 647]. Microelectronics [250]. Microelements [2962]. Microexpressions [2776]. Microfluidic [1923, 1996]. Micropower [2295]. Microprocessor [134, 2231]. Microprocessors [233]. microscopic [1992]. Microscopy [1968]. Middle [227]. Might [596]. Migration [227, 1246, 1654]. Milky [1269]. Millimeter [2934]. Millimeter-Wave [2934]. Million [2872]. Milliseconds [1412]. Mills [88, 1274, 2593, 2634, 2645, 2683, 2697]. Milojicic [730]. MIME [1147]. Mind [964, 1908, 2423]. Mind-Controlled [1908]. Mindset [1804]. Miners [341]. Minimizing [2449]. Mining [55, 181, 380, 464, 684, 687, 867, 1161, 1814]. MINIX [1205]. Minus [1742]. Mirai [2333]. Mirror [691]. Misconceptions [598, 2268]. Misconduct [261]. Misinformation [1386]. Missing [417, 1097, 2570]. Mission [1539, 1626]. Missions [195]. Mist [1587]. Mitchell [756]. Mitigating [1841]. Mitigation [2000]. Mix [2348]. Mobile [16, 66, 104, 213, 226, 290, 311, 326, 328, 371, 403, 445, 497, 554, 650, 658, 742, 759, 772, 820, 825, 976, 1046, 1068, 1081, 1108, 1140, 1183, 1189, 1439, 1441, 1556, 1563, 1721, 1728, 1805, 1927, 2041, 2260, 2441, 2590, 2596, 2666, 2669]. Mobile-Driven [2666]. Mobilization [819]. Model [121, 295, 633, 952, 980, 1232, 1273, 1403, 1559, 1965, 2209, 2381, 2385, 2564]. Model-Based [1965, 2209]. Modeling [86, 122, 236, 372, 652, 654, 655, 1191, 1716, 2068, 2130]. Models [291, 352, 981, 982, 1059, 1240, 1468]. Modern [502, 870, 1460, 2031, 2075]. Modern-Day [2075]. Modernization [1623]. Modularity [402]. Molecular [1717, 2036]. Moler [439]. Moment [940]. Monash [517]. Monetization [808]. Money [109, 1248]. Monitor [1934]. Monitoring [1217, 1837, 2070, 2098, 2125, 2544, 3063]. Months [1412]. MOOC [1824]. MOOCs [961, 2454]. Mood [423]. Moore [32, 163, 1012, 1749, 2188, 2361, 3110]. Moral [1072, 3101]. Morphing [315]. Mosaic [807, 1060]. Motivating [313]. Motivation [2541, 2543]. Move [698, 834, 1602, 1608, 1675, 1722, 1726, 1768, 1799, 1827, 1850, 1855, 1883, 1890, 1933, 1963, 2013, 2055, 2085, 2112, 2142, 2172, 2193, 2223, 2252, 2308, 2331, 2371, 2399]. Movement [2443]. Movements [466, 2876]. Moves [696]. Moving [156, 911, 950, 1067]. Mozilla [756]. MPI [1988]. Mr. [1127]. Much [1532]. Mudd [789]. Mug [320]. Multi [705, 1679, 1800, 1904, 2131, 2453]. Multi-Core [1679, 1800, 1904, 2131]. Multi-national [705]. Multi-Scale [2453]. MultiAmdahl [889]. Multicloud [768]. Multicore [45, 723, 952, 953, 2039, 2071, 2289, 2355]. Multicores [956]. Multidevice [2123]. Multidimensional [2757]. Multidisciplinary [2819]. Multihop [955]. Multilayered [2236]. Multilingual [1593]. Multimedia [135, 215, 1099, 1161, 2794, 2856, 2883, 2918, 2938, 2967, 2989, 3054, 3086]. Multimodal [2210]. Multiple [395]. Multipurpose [567, 1147]. Multisurface [546]. Multitenancy [2413]. Multithreaded [45, 631, 635, 1618]. Multitier [988, 1518]. Multitouching [660]. Multivariate [1607]. München [2053]. Muscle [2409]. Museum [517, 554, 612, 2065]. Museums [467]. Music [455]. Must [1747]. My [359, 792, 1280, 2149, 2164, 2998]. myCS [2182, 2251, 2290, 2457, 2494, 2607]. Myth [548].

N3XT [1750]. Name [95, 2385]. Named [317, 336, 689, 2501]. Names [51, 513]. NAND [905]. Nano [766]. Nanoelectronics [250]. Nanogrids [661]. Nanoscale [254, 1754]. Nanostores [233]. Nanowire [253]. NASA [890]. Nathaniel [1147]. Nation [703]. National [192, 705, 1190, 2446, 2746, 3023]. Natural [612, 1624, 2093]. Nature [1076, 2037]. Nature-Inspired [2037]. Navigating [270, 1868]. Navigation [2285]. Navigational [2842]. NCSA [1060]. Near [229, 1662]. Near-Real-Time [1662]. Need [85, 560, 1600, 1861]. Needs [174, 1122, 2917, 2991]. Negative [1072, 3102]. Neglect [412]. Neonatal [865]. Nerd [792]. Net [1843, 2415, 2574]. Nets [238]. Network [153, 159, 172, 309, 365, 720, 936, 1339, 1341, 1345, 1579, 1607, 1678, 1756, 1839, 1896, 2073, 2101, 2243, 2413, 2564, 2788, 2968, 3024]. Network-Access-Control [153]. Network-Level [2968]. Network-on-Chip [1756]. Network-Usage [172]. Networked [1131, 1907, 2327]. Networking [34, 117, 665, 737, 866, 879, 1335, 2911, 2950, 3007, 3021, 3043, 3050, 3114]. Networks [78, 118, 136, 236, 289, 390, 524, 568, 569, 653, 654, 701, 813, 843, 955, 1003, 1152–1158, 1177, 1312, 1343, 1352, 1381, 1472, 1497, 1501, 1550, 1585, 1755, 1897, 2101, 2127, 2259, 2412, 2501, 2545, 2572, 2597, 2621, 2804, 2873, 3005]. Neumann [704]. Neural [1678, 2545, 2597, 2873]. Neuroimaging [302]. Neuromemristive [2072]. Neurons [2872]. Neuroscience [2569]. Neurotechnology [2186]. Neutering [1843]. Neutrality [1843, 2415, 2574]. Neutrino [1275]. Never [2213]. News [11, 327, 347, 366, 388, 410, 478, 498, 521, 541, 563, 585, 608, 630, 651, 674, 697, 717, 741, 760, 781, 811, 835, 859, 880, 903, 928, 950, 951, 977, 1007, 1038, 1063, 1092, 1118, 1151, 1180, 1209, 1236, 1266, 1301, 1334, 1373, 1409, 1436, 1462, 2247, 2389]. Newspapers [867]. Newton [2652]. Next [16, 18, 80, 232, 298, 365, 473, 511, 588, 635, 761, 785, 818, 839, 904, 911, 917, 937, 938, 1324, 1363, 1389, 1396, 1522, 1596, 1704, 1820, 1867, 1875, 2033, 2318, 2320, 2503, 2666, 2881, 2935]. Next-Gen [2503]. Next-Generation [16, 473, 588, 635, 818, 904, 1596, 2033]. Nexus [2549]. NFMI [2792]. Nigeria [2341]. Nightmare [1443]. Nii [1261]. Niklaus [604]. NIST [703, 2136, 2910]. No [2626, 2912, 2960]. NoCs [2987]. Node [2804]. Node-Deployment [2804]. Noirware [1478]. Nomadic [1870]. Nominations [1237, 1274, 1650, 1690, 1912, 1942, 1967, 1970, 1984, 1998, 2032, 2137, 2191, 2192, 2208, 2214, 2220, 2343, 2567, 2599, 2608, 2632, 2663, 2691, 2976]. Nominees [1838, 1936, 1970, 1984, 2032, 2192, 2208, 2326, 2599, 2632, 2663, 2691]. Non [900, 2560]. Non-US [900]. Non-Volatile [2560]. Nonfunctional [1540]. Nonhumans [2218]. Nonlinear [222]. Nonverbal [1121]. Nonvolatile [907, 3099]. Normal [748]. North [462]. NoSQL [22, 1867]. Notation [537]. Novel [252, 544, 2804]. NP [2268]. NP-Hardness [2268]. NSFNet [559, 1115]. Nuclear [1782]. NUMA [1616, 1934]. Numbers [206, 462]. Numerical [1714]. Nutritionist [2544].

O [909]. Obese [2206]. Obesity [322]. Obfuscated [1184]. Object [420, 537, 1059, 1892]. Objective [2616]. Objects [1624]. Observations [3029]. Obstacles [1155]. Ocloud [2500]. October [1633]. Odd [750, 798]. Oddities [750, 798]. Off [91, 509, 559, 2725]. Off-the-Shelf [509]. Offers [98]. Office [1293, 3044]. Offloading [66]. Oh [2389]. Ohmic [1755]. Oil [144]. Olympic [597]. OMG [1434]. Omitted [160]. Omitted-Variable [160]. On-Body [185]. On-Chip [955]. On-Demand [616, 871]. On-Skin [2407, 2408]. On-Time [1177]. Onboard [23]. One [274, 319, 912, 1325, 1364, 1397, 1428, 1468, 1492, 2069, 2344, 2372, 2400, 2411, 3108]. Online [12, 137, 154, 164, 226, 389, 390, 483, 489, 620, 936, 1065, 1133, 1195, 1222, 1534, 1840, 2443, 2618, 2670, 2763]. Only [502, 591]. Onshore [403]. Opaque [3102]. Open [17, 64, 203, 356, 413, 547, 569, 582, 618, 637, 1021, 1303, 1307, 1308, 1337, 1341, 1472, 1529, 1559, 2140, 2160, 2262, 2296, 2422, 2447, 2729, 2739, 2743, 2850, 2914, 2963, 2970, 3002, 3038, 3105]. OpenFlow [365, 1336]. Opening [1937]. openings [1440]. OpenSpace [2588]. Operating [613, 824, 1781, 2499]. Operation [653]. Operational [1867]. Operationalized [2627]. Operations [191]. Opportunistic [10, 763]. Opportunities [10, 445, 860, 905, 939, 1125, 1155, 1289, 1319, 1360, 1384, 1392, 1423, 1451, 1483, 1510, 1524, 1542, 1568, 1599, 1632, 1670, 1701, 1736, 1764, 1795, 1823, 1914, 1946, 1977, 2051, 2082, 2242, 2300, 2467, 2625, 2701, 2758, 2968]. Opportunity [310, 1381, 2644]. Optimal [889, 1550]. Optimist [490]. Optimization [632, 1687, 2037, 2157, 2901, 2906]. Optimizations [160]. Optimize [960]. Optimized [549]. Optimizing [7, 1615, 1934, 2337]. Option [1641]. Options [1246, 2024, 2086, 2113, 2143, 2173, 2194, 2224, 2253, 2309, 2314]. Optique [1469]. Orchestrating [1196]. Orchestration [1810]. Order [474, 983, 1554]. Organization [1502]. Organizational [1249, 1307]. Organizations [783, 2620]. Organize [1027]. Oriented [24, 256, 678, 1651, 2127]. Origin [391]. Origins [877, 1008]. OSINT [1219]. Osmosis [2962]. Osmotic [2962]. Other [2333]. Otherwise [1837]. OuijAmazon [617]. Oulu [332]. Our [472, 489, 921, 1199, 1716, 1771, 1849, 2118, 2562, 2570, 2881]. Ourselves [1027]. Out-of-Box [339]. Outbreak [1134]. Outcome [120, 1760]. Outcomes [706]. Outlook [1389, 1772, 1781, 2493, 2754]. Output [466]. Outsourcing [459]. Outstanding [555]. Overcoming [254, 504]. Overflow [436]. Overhead [269]. Overload [112]. Overview [694, 1196, 2003, 2574]. OWL [973]. Own [1188].

PaaS [939]. Packets [715, 899, 1089]. Pain [2669]. Pains [112]. Palm [1146, 1852, 2216]. Panopticon [1343]. Paper [1098, 2724]. Paperless [1420]. Papers [1256, 1278, 1295, 1425, 1453, 1474, 1498, 1525, 1636, 1661, 1698, 1865, 1872, 2491, 2946, 2995, 3028]. Paradigm [575, 587, 595, 840, 1103, 1188, 1779, 2038, 2073]. Paradigms [1842, 2033]. ParaHeap [1269]. ParaHeap-k [1269]. Parallel [46, 458, 548, 633, 634, 1362, 1638, 2022]. Parallelism [45]. Parallelization [1272]. Park [581]. Parking [1134]. Parks [611]. Part [476, 495, 502, 750, 772, 798, 825, 967, 1059, 1133, 1195, 1492, 1520, 1562, 1590]. Part-Based [1059]. Participation [193, 194, 196, 198, 787, 789, 1414, 1577, 1635, 1672, 1825]. Parties [395]. Party [1757, 3020]. Passionate [1378]. Passive [1126, 1840]. Password [468]. Passwords [547, 1016]. Past [281, 349, 770, 3090]. Patching [3102]. Patent [345, 364, 386, 538, 648, 808, 1711]. Patents [778, 900]. Path [26, 431, 785, 2035, 2237, 2840]. Pathologies [2547]. Pathways [525, 2820]. Patient [121, 701, 1122]. Patient-Centered [121]. Patients [119]. Patt [419]. Pattern [1015]. Patterns [1244]. Paying [757, 2209]. Payment [213]. Payments [650]. PC [367, 368, 370]. PCs [214, 772]. Peace [1199]. Pearl [1928]. Peculiar [1847]. Peer [248]. Peer-to-Peer [248]. Pending [364]. Penetration [1066, 2575]. People [1472]. Perceived [2998]. Perception [260]. Perceptual [1842]. PerCom [1627]. Performance [218, 232, 434, 456, 682, 820, 844, 1145, 1270, 1312, 1480, 1534, 1718, 1871, 2068, 2069, 2567, 2569, 2608, 2658, 2933]. Performance-Power [2068]. Peril [242]. Perils [594, 1926, 2441]. Periodicals [773]. Pernicious [2600]. Persistent [908, 2203]. Person [619]. Personal [368, 369, 417, 448, 700, 791, 1437, 1438, 2595]. Personality [1459]. Personalized [1871]. Personalizing [2183]. Personas [817]. Perspective [123, 222, 255, 276, 318, 417, 510, 699, 703, 815, 1336, 1355, 2499, 2703, 2719, 2788]. Perspectives [501, 862, 2875]. Persuaders [1738]. Persuasive [2853]. Pervasive [120, 1627, 1790, 1837, 1982, 1989, 2292, 2786, 2821, 2822, 2870, 2903, 2975, 2997, 3028]. Pessimist [490]. Petaflops [1717]. PhD [462, 1535]. Phenomenon [1589]. Philanthropy [2324]. Phishing [2626]. Phone [16, 771, 1836, 2441]. Phones [772, 1185]. Photo [103]. PHP [693]. Phylogenetic [1403]. Physical [743, 1071, 1526, 1820, 2234–2237, 2461, 2476, 2548, 2703, 2717, 2722, 2726, 3097]. Physiological [1682, 1685]. Pi [947, 1700, 2363]. Piazza [777]. Picking [1554]. PII [1166]. Pilots [1311]. Pin [46]. Pioneer [317, 336]. PIPA [615]. Pipe [47]. Pipeline [729, 784, 1535]. Piracy [175, 1696, 2267]. Pirated [2827]. Pitfalls [2301]. PKI [104]. Placement [1812]. Plain [352]. Plan [967]. Plane [1339]. Planes [1341]. Planning [225, 1465]. Plant [191, 1782]. Plants [2388]. Plate [1619]. Plates [1783]. Platform [272, 333, 659, 1536, 1994, 2071, 2440, 2469, 2962]. Platforms [271, 290, 315, 1189, 2743]. Play [733, 772, 825, 858, 983, 1557]. Playdates [638]. PLiM [2294]. Plotting [2477]. Plug [983]. Plug-and-Play [983]. Point [1290, 1924, 2626]. Policies [30, 395, 1016, 2470, 2939]. Policy [73, 391, 705, 861, 866, 1591, 1841, 2134, 2467, 2574, 2739, 3067]. Policymakers [2618].

Politeness [360]. Political [615, 2165]. Politics [1785]. Polling [2275]. Pollyanna [2006]. Pong [1694]. Pooja [777]. Population [1044]. Porous [932]. Portable [3066]. Portfolio [687]. Portfolios [1307]. positions [1440]. PossessedHand [466]. Possibilities [272, 1705, 1948, 2618]. Possible [512]. Post [1717, 1824, 2247, 2823]. Post-MOOC [1824]. Post-Petaflops [1717]. Post-Truth [2823]. Post-Truths [2247]. Postmodern [2728]. Posts [390, 3027]. Posttraumatic [1212]. Postures [706]. Potential [382, 2010, 2036, 2044, 2869]. Power [418, 434, 530, 562, 661, 820, 849, 978, 995, 1101, 1102, 1145, 1353, 1659, 1782, 2068– 2070, 2293, 2318, 2355, 2478, 2537, 2720]. Power-Performance [1145]. Powerball [3046]. Powered [137, 931, 1041, 2295, 2701]. Powerful [6]. Powering [273, 823]. Practical [164, 1847, 2845, 2904, 3096]. Practice [24, 54, 351, 412, 1079, 1193, 2673, 2742, 2814]. Practices [826, 932, 1816]. Practitioners [412]. Pragmatic [2783]. Pre [399]. Pre-College [399]. Precise [482]. Predict [1559]. Predictable [2419]. Predicting [1871, 2207, 2385]. Prediction [11, 178, 1714, 1782, 2275, 2566]. Predictions [3092, 3093, 3100]. Predictive [1982, 1989]. Predictor [423]. Prepare [424, 2310, 2322, 2345, 2429, 2881]. Prepared [767]. preparing [2765]. Present [281, 349, 770, 1045]. Present-Day [349]. Presentations [1527, 1558, 1604, 1607]. Preservation [1814]. Preserving [2688]. President [141, 375, 461, 643, 730, 919, 1023, 1253, 1391, 1630, 1761, 2016, 2138, 2366, 2480, 2678, 2747, 2980, 3112]. President-Elect [141, 375, 461, 643, 730, 919, 1023, 1253, 1391, 1630, 1761, 2016, 2138, 2366, 2480, 2678, 2747, 2980, 3112]. Presidential [484, 2275, 2389]. Presidents [373]. Press [1331, 1370]. Prevent [2651]. Preventing [144, 1250].

Prevention [3063]. Price [11, 1758].

Primary [2720]. Primer [1047]. Priming [784]. Principles [764, 2854]. Print [226, 1641]. Printed [2183]. Printing [1407, 1487]. Prior [1234]. Priorities [192, 198]. Prioritize [1895]. PRISM [892]. Privacy [340, 389, 393, 395, 518, 708, 797, 866, 990, 1016, 1047, 1178, 1182, 1355, 1414, 1438, 1443, 1514, 1664, 1730, 1814, 1816, 1834, 1848, 1927, 1995, 2063, 2099, 2103, 2130, 2616, 2618, 2647, 2672, 2688, 2727, 2777, 2810, 2837, 2846, 2848, 2874, 2880, 2895, 2924, 2942, 2978, 3030, 3060, 3098, 3109]. Privacy-Aware [1414, 2647]. Privacy-Knowledge [2130]. Privacy-Preservation [1814]. Privacy-Preserving [2688]. Private [155, 303, 529, 2595]. Prize [238]. Proactive [1347, 2518]. Probabilistic [69]. Probably [560]. Problem [326, 346, 723, 1013, 1504, 2217, 2356, 2478, 2905]. Problems [505, 562, 1463, 2901]. Process [55, 464, 729, 983]. Processes [338, 549]. Processing [178, 1362, 1638, 2022, 2099, 2385]. Processor [562]. Processors [458, 1501, 2039]. Produce [2186]. Producer [3037]. Product [259, 338, 675, 677–679, 980, 1444, 1534]. Production [2052]. Products [380, 1771]. Profession [19, 128, 186, 470, 1880, 2212]. Professional [174, 794, 932, 1020, 2784, 2806, 2807, 2814, 2851, 2946, 2995]. Professionals [161, 263, 555, 785, 2907, 2973, 3025]. Profiles [2560]. Program [872, 1337, 2972]. Programmable [1222]. Programmed [235, 360]. Programmer [39]. Programmers [1964, 2262]. Programming [127, 184, 595, 607, 634, 647, 680, 869, 1021, 1074, 1251, 1388, 1473, 1649, 1962, 1964, 2127, 2356, 2359, 2377, 2435, 2436, 2440, 2545, 2625, 2650]. Programming-Language [2359]. Programs [13, 46, 146, 174, 704, 1731, 2906]. Progress [562, 714, 1311, 1718, 2441, 2739]. Project [7, 133, 249, 618, 1332, 1665, 1745, 1786, 2213, 2600]. Project-Based [1665]. Projection [611, 1033]. Projection-Based [611]. Projects [356, 1378, 2963, 3002]. Promise [22, 98, 242, 469, 594, 661, 1281, 1902]. Promises [6, 154, 1463]. Promote [2963]. Promoting [1304]. Propagation [796, 1386]. Property [476, 495, 738, 1116]. Proportional [89, 2354]. Proportionality [396]. Proportions [488]. Proposals [1688, 1735, 2110, 2124, 2984, 3034]. Proposed [2015]. Proposes [249]. Prospects [1623]. Prostate [120]. Protecting [738, 1185, 1497, 1716, 1816, 1925, 2099]. Protection [480, 814, 958, 2562, 2672, 2790, 2845]. Protective [2441]. Protein [524]. Protocol [553]. Protocols [1836]. Prototyping [762, 988, 1418, 1533, 2009]. Provenance [3103]. Provides [61]. Providing [894, 937]. Provisioning [372, 871, 2669]. Proxemic [1244]. Proxying [18]. Psst [2505]. Psychology [1620]. Public [155, 566–568, 612, 861, 913, 1154, 1239, 1696, 2620, 2939]. Public-Private [155]. Publication [473, 2136]. Publications [689, 1020, 1045, 1139, 1264, 1344, 1455, 1486, 1513, 1860]. Publish [2620]. Publishing [736, 1039]. Pufferies [2600]. Puncturing [2600]. Pupillometry [1684]. Purdue [1417, 2394]. Purposeful [790]. Pursuing [794]. Push [172]. Pushes [819]. Pushing [500]. Put [1903]. Putting [913, 1044]. Puzzle [62]. Puzzle-Based [62]. Python [1272, 1433, 1460].

Q&A [620]. QoS [178]. QoS-Aware [178]. Quagmire [1757, 2576]. Quake [308]. Qualifications [2472]. Quality [338, 937, 1223, 1284, 1441, 1479, 1656, 1771, 2205, 2842]. Quality-Based [1656]. Quality-of-Experience [1284]. Quality-of-Life [2205]. Quantification [1589]. Quantified [964]. Quantifying [957, 1308]. Quantum [1717, 1765, 2034, 2035, 2525, 2671, 2746, 2898, 2899, 2901, 2902, 2904–2906, 3023, 3046]. Quaynor [1261]. Queries [1656, 2078]. Query [1096, 2964]. Question [2159]. Question-Answering [2159]. Questioning [71]. Questions [1472]. Quick [2277]. Quickly [673].

R&D [156, 705, 1500]. Race [490]. Races [520]. Racing [2321]. Rack [2790]. Rack-Scale [2790]. Radical [2249]. Radio [1550]. Radiotherapy [482]. Radon [1783]. Raising [20, 605, 1771]. Ramakrishna [2768, 2797]. Range [930, 2792]. Ransomware [2448]. Rapid [988, 1042]. Rasmus [693]. Raspberry [947]. Rates [1872]. Rating [1534]. Rau [419, 731, 2768, 2797]. Read [1020]. Reader [186, 1116]. Reading [1045]. Ready [5, 650, 902]. Real [43, 138, 158, 609, 886, 1170, 1441, 1662, 2039, 2414, 2508, 2602]. Real-Time [43, 886, 1441, 2039, 2414]. Real-World [138]. Realism [2899]. Realistic [2902]. Realities [2382]. Reality [548, 609–614, 1211–1213, 2080]. Realize [2010]. Realizing [766]. Really [8, 43, 91, 145, 489, 2520, 2902]. Rearrangement [1614]. Reasoning [378, 981, 1655]. Rebooting [1748, 1939, 2155, 2244, 2573, 2746]. Receives [88, 199, 336, 441]. Recipients [51]. Reciprocal [1960]. Recognition [26, 28, 280, 444, 619, 763, 964, 991, 1164, 1281, 1290, 1609, 1619, 1720, 2242, 2596]. Recognize [1783]. Recognizes [555]. Recognizing [2567, 2608]. Recommendation [1114]. Recommended [780]. Recommender [2543, 2956]. Reconfigurable [955, 1158]. Reconstructing [102]. Reconstruction [103, 1892]. Record [1438]. Records [181, 698–702, 2744]. Recovering [872].

Recovery [1643]. Recruiting [109]. Recurrent [2597]. Recycle [909]. Redefine [2188, 2330]. Redefining [2357, 2757]. Redesigning [526]. Reduce [909]. Reduced [2462]. Reducing [18, 2651]. Reduction [2905]. Reduction-Based [2905]. Redundancy [2159]. Redundancy-Based [2159]. Refactor [909]. Refactoring [2328]. Referencing [2040]. Reflections [566, 791]. Reform [648]. Reforming [174]. Reformulation [1096]. Refreshable [1776]. Register [334]. Regular [48, 1382]. Regulation [2787, 2845]. Regulations [797]. Reimagining [840]. Reinforcement [2987]. Reinventing [1032]. Related [1212]. Relation [84]. Relational [1656]. Relations [709]. Relationship [241, 1907, 2008]. Relevance [1032]. Relevant [465, 686, 1130, 1223]. Reliability [254, 434, 989, 1559, 1589, 1782, 2069, 2318, 2355]. Reloaded [2415]. Remote [371, 459, 1835]. Renaissance [1868]. Rendering [1330]. Rendezvous [328]. Renren [936]. Repeat [1791]. Repeated [955]. Replication [504, 1465]. Report [1311, 1470, 2747]. Reports [312]. Represent [3000, 3027]. Representation [378, 1447, 1491, 3062]. Reproducibility [3055]. Repurposers [623]. Repurposing [1653]. Required [1837]. Requirements [86, 295, 468, 914, 1540, 1837, 2093]. Requires [976]. Rescue [1896]. Research [594, 742, 752, 836, 838, 868, 934, 979, 1094, 1130, 1243, 1811, 1898, 1927, 2303, 2337, 2359, 2383, 2404, 2417, 2724, 2788, 2994]. Researchers [132]. Reservoir [274]. Residual [1559]. Resig [1519]. Resilience [1277, 2719, 2722]. Resiliency [1312, 2717]. Resilient [956, 2475, 2721, 2965]. Resistance [906]. Resistance-Based [906]. Resistant [2443]. Resolution [633]. Resolve [63]. Resource [178, 282, 889, 1564, 1810, 2597, 2877, 2902]. Resource-Aware [1564].

Resource-Constrained [2597]. Resources [1663, 2087, 2114, 2145, 2175, 2196, 2266]. Response [1120, 1893, 1895, 1905, 2720]. Responsible [2477]. Responsive [2074]. REST [1551]. Restandardize [60]. Restructuring [909]. Results [3003, 3102]. Resurrecting [2232]. Retail [297]. Retaining [1314]. Retention [34]. Rethinking [8, 233, 340, 432, 1120, 1496, 1540, 2008, 2795]. Retrieval [576]. Retrospect [593]. Retrospective [367, 2289]. Returning [792]. Reuse [159, 2566]. Reveal [2876]. Revealing [2443]. Reveals [2612]. Reverse [1213]. Reversible [2300]. Review [714]. Revisited [1100, 1623]. Revisiting [2377, 2404, 2498, 3106]. Revitalization [1371]. revitalizing [2765]. Reviving [349]. Revolution [761, 2546, 2548]. Revolutionary [1443, 2754]. Revolutionizing [742]. RF [1041]. RF-Powered [1041]. RFIDiocy [751]. RFIDs [2745]. Rich [1121, 1122]. Richard [1237, 1263, 1613, 1621, 1652, 1993, 2199, 2358, 2591, 2619, 2656, 2665, 2712, 2993]. Right [367, 1660, 2025, 2057, 2315, 2397]. Rights [215]. Rigid [1475]. Rigor [665]. Ring [334]. Rise [93, 2443, 2461, 2522, 2661]. Rising [1049]. Risk [84, 236, 862, 1019, 1277, 1828, 1857, 2047, 2239–2241, 2564]. Risk-Assessment [2240]. Risk-Based [1828, 1857]. Risks [746, 1036, 2238]. RIT [770]. Road [2155, 2444, 2466]. Roadmap [86]. Roadmapping [2630]. Robert [1219]. Robot [149, 802, 1006, 1280, 1592, 2127]. Robotic [745]. Robotics [316, 1759, 3094]. Robots [309, 638, 1350, 1940, 1959, 2135, 3094]. RobotX [2274]. Robustness [255]. Rock [1081, 1108, 1140, 1173, 1245, 1267, 1320, 1404, 1407, 1456, 1487, 1515, 1547, 1574, 1603, 1617, 1645, 1828, 1856, 1857, 1884, 1885, 1919, 1935, 1951, 1981, 1982, 1989, 2023, 2174]. Role [108, 282, 593, 1094, 1239, 1969, 2244, 2412, 2495, 2671, 2788]. Role-Based [108, 593]. Rome [102]. Room [546]. Rooms [613]. Rooted [1185]. RootGuard [1185]. Rosebush [1387]. Rosenblum [336]. Rossum [1433, 1460]. Roulette [2473]. Round [2473]. Roundtable [1528]. Roy [1551]. Rs [397]. Rubble [1896]. Rule [1999, 2069]. Ruling [2574]. Runs [591]. Runtime [49, 469]. Runtimes [290, 1616]. Rural [459]. Russian [2280, 2389].

S [1205]. S7 [1916, 1950]. Sabotage [188]. Safe [132, 173, 1293, 1876]. Safely [700]. Safety [1944, 2103, 2673, 2912, 3005]. Safety-Critical [3005]. Sale [856]. Same [391]. Same-Origin [391]. Sampling [586, 815, 2646]. Samsung [1324, 1325, 1916]. SanDisk [1913]. Sankar [777]. Satellite [1134]. Satisfaction [2488, 2751, 3079]. Save [66, 2260, 3033]. Saving [214, 319]. Say [1395]. SCADA [721]. Scalable [846, 1721, 2035, 2498, 2595]. Scale [275, 457, 656, 702, 885, 1464, 1615, 1616, 1751, 1898, 2158, 2453, 2454, 2790, 3064]. Scaling [134, 632, 1272, 1616, 1748, 2355, 2356, 2987]. Scan [815]. Scanning [230]. SCDOR [800]. Scenario [178]. Scenario-Based [178]. Scene [103]. Schema.org [2620]. Schneier [1804, 1862, 2524]. Scholarship [1076, 1621, 1652, 2199, 2358]. School [424, 2444]. Schools [342]. Science [62, 92, 145, 202, 283, 342, 358, 384, 424, 465, 572, 604, 718, 749, 786, 848, 930, 934, 1347, 1349, 1405, 1418, 1481, 1533, 1742, 1789, 1873, 1898, 1937, 1943, 2009, 2207, 2262, 2273, 2445, 2496, 2615, 2994, 3065, 3093]. Sciences [2350, 2493]. Scientific [145, 159, 193, 840, 1045, 1243, 1713, 1947]. Scientist [2109, 2875]. Scientists [44, 577, 2320, 2503]. Score [2996]. Scott [169]. Scratch [1074]. Screen [229, 409]. Screening [3064]. Screens [331]. Script [69]. Scripting [527, 1841]. SDN [1336]. SE [1731]. Seamless [1710]. Search [15, 43, 107, 137, 154, 445, 575, 1015, 1096, 1097, 1161, 1169, 1467, 1509, 1531, 1560, 1634, 1648, 1687, 1737, 1774, 1813, 1896, 1960, 2004, 2078, 2154, 2204]. Search-Engine [1687]. Searches [1835]. Second [474, 1335, 1411]. Second-Generation [1411]. Second-Order [474]. Secondary [2820]. Secret [848, 2505]. Secretariat [2004]. Secretocracy [1815]. Secrets [1393]. Sector [719]. Secure [97, 173, 392, 605, 871, 1015, 1494, 1497, 1505, 1532, 1876, 1991, 1994, 2332, 2362, 2470, 2627, 2718]. Securely [1660]. Securing [394, 477, 553, 812, 1188, 1805, 2723]. Security [12, 30, 104, 158, 255, 326, 389, 391, 453, 454, 505, 509, 518, 540, 598, 616, 708, 722, 727, 768, 800, 816, 976, 989, 1073, 1103, 1182, 1183, 1189, 1190, 1285, 1313, 1355, 1389, 1514, 1538, 1664, 1787, 1804, 1828, 1841, 1857, 1925– 1927, 1944, 1974, 1990, 1995, 1996, 1999, 2001, 2047, 2076, 2100, 2136, 2189, 2235, 2238, 2239, 2267, 2272, 2277, 2293, 2301, 2501, 2523, 2562– 2564, 2643, 2653, 2661, 2673, 2761, 2777, 2810, 2846, 2874, 2880, 2895, 2912, 2924, 2942– 2944, 2966, 2968, 2978, 3030, 3060, 3098, 3109]. Security-Aware [2267]. Security-Informed [1944, 2673]. See [1268, 2623, 3024]. Seeding [2820]. Seeing [2043]. Seek [3000]. Seeking [882, 1093, 1095]. Seen [735]. Segmentation [2944]. Seitz [438]. Select [3105]. Selecting [1559]. Selection [528, 1114]. Self [680, 964, 1158, 1211, 1270, 1439, 1447, 1584– 1588, 1620, 1663, 1965, 2465, 2467, 2508]. Self-Adaptive [680, 1447]. Self-Aware [1584, 1585, 1620, 1663, 1965]. Self-Awareness [1586, 1587]. Self-Driving [2465, 2467]. Self-Expression [1586]. Self-Expressive [1584, 1585, 1588, 1620]. Self-Gravity [1270]. Self-Management [1439]. Self-Managing [2508]. Self-Reconfigurable [1158]. Semantic [15, 179, 1309, 1466, 1655, 1842, 2436, 2438, 2620]. Semiconductor [2296, 2336]. Sense [2466, 3094]. Sensemaking [1095]. Sensing [185, 597, 1121, 1122, 1244, 2158, 2527, 2759, 2934, 2935, 3035]. Sensor [653, 654, 1691, 1896, 2804]. Sensor-Tolerant [1691]. Sensors [1685, 3035]. Sentient [314, 1943, 2271]. Sentiment [2264]. September [648, 2690]. Sequences [725]. Sequencing [525, 1443]. Serendipitous [2649]. Serious [326, 454]. Serve [13]. Server [319, 396, 1279, 2475]. Servers [125]. Service [49, 154, 256, 273, 333, 483, 678, 679, 682, 709, 1114, 1162, 1237, 1263, 1339, 1377, 1432, 1441, 1595, 1613, 1850, 1993, 2044, 2099, 2101, 2470, 2591, 2619, 2624, 2656, 2665, 2712]. Service-Based [679]. Service-Oriented [256, 678]. Services [127, 136, 204, 270, 416, 459, 894, 988, 1028, 1065, 1066, 1163, 1358, 1427, 1518, 1625, 1771, 1784, 1960, 2647, 2666, 2723, 2858, 2864, 2922]. Serving [790]. Sesame [547]. Set [2462]. Sets [2788]. Seven [166]. Seymour [2137, 2167, 2214, 2567, 2608]. Shadows [2048]. Shannon [2496]. Shape [1842]. Shaping [38, 351, 2360]. Share [601, 2554, 2583, 2635, 2657]. Shared [956, 1221, 1654, 2498]. Shared-Memory [956, 1654]. Sharing [440, 2241, 2520]. Shelf [509]. Shift [334, 1103]. Shifting [3104]. Shin [796]. Shine [1973]. Shining [2245]. Shoe [2902]. Shopbots [80]. Shopping [2181, 2450]. Shorelines [1716]. Short [897, 2792]. Short-Range [2792]. Short-Term [897]. Shortening [1215]. Shots [320]. Should [182, 683, 3106]. Shoulders [704, 2338]. Show [712]. Showcase [1305, 1792]. Shows [1769]. Shull [3112]. Shutdown [1966]. Sidney [2137, 2167, 2214, 2567, 2608]. Siege [1047]. Signage [564]. Signals [2158, 2935]. Significance [1220, 1234]. Significance-Based [1220]. Silence [2450]. Silent [1689]. Silent-Speech [1689]. Silicon [370]. Simple [1589]. Simplicity [111, 978–980, 2944]. Simplifying [2436]. Simulating [139]. Simulation [236, 548, 652, 655, 657, 1719]. Simulator [656, 2327]. SimWare [656]. Single [955, 2297]. Single-Cycle [955]. Site [527, 528, 1841]. Sites [351, 620]. Situ [2873]. Situation [1779, 1978]. Situational [814, 847]. Six [1389, 1533, 2272]. Sized [1504]. Sketches [320]. Skewed [1783]. Skills [1121]. Skin [1875, 2407, 2408]. Slam [2325]. Sleep [2207]. Sleeping [214]. Sleight [2165]. Slow [289]. Small [409, 1270, 2003, 2005, 2357, 2598, 2901, 2902]. Small-Bowel [2598]. Smarr [807]. Smart [119, 324, 329, 332, 553, 652, 747, 767, 888, 955, 1152, 1154, 1155, 1157, 1158, 1187, 1303, 1310, 1311, 1414, 1501, 1960, 2094– 2097, 2125, 2178, 2183, 2237, 2240, 2329, 2332, 2383, 2440, 2647, 2737, 2741–2743, 2878, 3104]. Smart-City [2743]. Smarter [50, 330, 1317, 1663]. Smartphone [659, 722, 1187, 1216, 1501, 1532]. Smartphones [315, 509]. Smartwatches [2246]. Snowden [1127]. Snowflakes [253]. SoC [960]. Social [34, 117, 118, 135, 136, 192–194, 196–198, 206, 236, 242, 348, 390, 510, 664, 665, 709, 792, 819, 842, 866, 891, 936, 1017, 1036, 1043, 1044, 1094, 1099, 1121, 1161, 1352, 1662, 1848, 1897, 1907, 1940, 2041, 2190, 2495, 2572, 2621, 2668, 2759, 2911, 2950, 3007, 3027, 3043, 3050, 3114]. Social-Networks [136]. Social-Participation [193]. Socially [1848, 2477]. Society [1, 94, 113, 129, 140, 141, 150, 167, 187, 207, 224, 238, 245, 264, 284, 361, 374, 375, 383, 406, 426, 449, 461, 471, 492, 516, 533, 536, 555, 558, 573, 577, 580, 599, 602, 624, 642, 643, 646, 668, 689, 692, 710, 713, 730, 734, 754, 776, 806, 830, 853, 876, 1023, 1047, 1274, 1357, 1391, 1480, 1490, 1652, 1761, 1801, 1912, 1942, 1987, 2138, 2191, 2199, 2337, 2358, 2380, 2432, 2480, 2593, 2634, 2645, 2658, 2683, 2697, 2711, 2747, 2887, 2933, 2949, 3089, 3112]. Society [827, 850, 873, 895, 898, 918, 919, 922, 942, 945, 965, 971, 972, 985, 993, 1001, 1022, 1024, 1025, 1031, 1034, 1050, 1055, 1058, 1077, 1086, 1105, 1113, 1136, 1144, 1167, 1176, 1194, 1197, 1203, 1224, 1231, 1252, 1253, 1259, 1286, 1297, 1329, 1368, 1402, 1431, 1458, 1503, 1517, 1549, 1567, 1578, 1582, 1606, 1610, 1628, 1630, 1640, 1650, 1675, 1677, 1683, 1702, 1709, 1744, 1767, 1775, 1798, 1803, 1809, 1817, 1832, 1859, 1860, 1888, 1891, 1900, 1922, 1954, 2030, 2062, 2092, 2120, 2121]. Society [1883, 1890, 1933, 1956, 1963, 1967, 1985, 1998, 2007, 2012, 2013, 2016, 2028, 2055, 2060, 2067, 2085, 2087, 2090, 2112, 2114, 2117, 2142, 2145, 2152, 2172, 2175, 2180, 2193, 2196, 2201, 2220, 2223, 2252, 2256, 2308, 2313, 2343, 2351, 2364, 2366, 2371, 2375, 2391, 2399, 2403, 2421, 2427, 2437, 2455, 2463, 2468, 2483, 2490, 2497, 2510, 2515, 2532, 2535, 2536, 2553, 2554, 2558, 2559, 2582, 2583, 2586, 2604, 2611, 2622, 2628, 2635, 2639, 2640, 2662, 2675, 2676, 2678, 2682, 2686, 2689, 2705, 2730, 2738]. Society [2, 13, 51, 106, 161, 179, 209, 278, 305, 323, 373, 513, 1747, 1846, 2148, 2284, 2577, 2653, 2657, 2751, 2752, 2756, 2772, 2773, 2778, 2798, 2799, 2805, 2830–2833, 2836, 2859, 2863, 2871, 2888, 2894, 2919, 2923, 2927, 2947, 2950, 2957, 2977, 2980, 2981, 2988, 2992, 3001, 3009, 3015, 3047, 3067, 3073, 3079–3081, 3090, 3113]. SocietyConnection [803]. Sociotechnical [195, 639]. Socks [1557]. Socrates [3031]. SoDA [2638]. Soft [404, 504, 956]. Soft-Error-Resilient [956]. Soft-State [504]. Software [24, 35, 67, 83, 90, 127, 144, 145, 155, 157, 158, 259, 277, 295, 298, 316, 369, 378, 402, 405, 411, 413–415, 417, 422, 432, 481, 552, 582, 587, 640, 647, 670, 675–679, 728, 769, 770, 826, 838, 879, 908, 948, 960, 1043, 1064, 1083, 1124, 1130, 1243, 1285, 1313, 1316, 1335, 1343, 1529, 1540, 1559, 1569, 1600, 1623, 1658, 1695, 1696, 1729, 1731, 1745, 1757, 1779, 1780, 1782, 1787, 2003, 2096, 2101, 2129, 2165, 2270, 2272, 2317, 2417, 2422, 2499, 2517, 2523, 2718, 2782, 2785, 2827, 2837, 2845, 2914, 3002, 3029, 3072]. Software [416, 957, 1274, 1788, 2755, 2839, 2897, 2936, 2969, 3041, 3084]. Software-Based [127, 144]. Software-Defined [879, 1335, 1343, 2101, 2499]. Software-Driven [369]. Software-Intensive [676, 728]. Software-Testing [3029]. Software/Hardware [295]. Softwarization [2332]. Solar [1270]. Solicited [2137, 2214, 2567, 2608]. Solid [1040]. Solid-State [1040]. SolidThinking [2056]. Solution [1468, 1994, 2740]. Solutions [371, 766, 1331, 1370, 2096, 2160, 2325]. Solve [723]. Solved [107]. Solves [1504]. Solving [893, 2901]. Some [206, 289, 488]. Someday [41]. Sony [1446]. Sophistry [2165]. SOSE [1296]. Sought [773]. Source [17, 356, 413, 582, 1337, 1341, 1529, 1559, 1818, 2140, 2160, 2296, 2422, 2447, 2850, 2914, 2963, 2970, 3002, 3038, 3105]. Sourcing [2454]. SP7 [714]. Space [331, 415, 612, 1381, 2040, 3097]. Spaces [397, 1154, 2329]. Spam [390]. Sparse [1615]. Special [335, 1688, 1735, 2110, 2124, 2984, 3034]. Specifying [1656]. Spectre [2573]. Spectrum [105]. Speech [1042, 1689, 2125]. Speech-User [1042]. Speed [289, 1861]. Spending [1488]. Spiking [1678, 2545]. Spills [144]. Spirit [114]. Spoofing [1164]. SportPhishing [662]. Sports [460, 597, 681, 3065]. Spotlighting [1445]. Spreaders [813]. Spreadsheet [2040]. SQL [795, 1233, 1656]. Square [1273]. SSWAP [1655]. Stack [909, 2560]. Staff [2714]. Stage [331]. Stages [800]. Stakeholders [2673]. Stand [2338]. Standard [1924, 3108]. Standardization [346, 1335, 1791, 2388, 2917]. Standardizing [973, 1129, 2070, 2276]. Standards [310, 1160, 1223, 1233, 1277, 1283, 1336, 1337, 1419, 1480, 1539, 1625, 1690, 1773, 1847, 1905, 2003, 2212, 2438, 2526, 2528, 2577, 2630, 2729, 2765, 2854, 2939, 2972]. Standing [1008]. Star [2174]. Stars [1081, 1108, 1140, 1173, 1245, 1267, 1320, 1404, 1407, 1456, 1487, 1515, 1547, 1574, 1603, 1617, 1645, 1828, 1856, 1857, 1884, 1885, 1919, 1935, 1951, 1981, 1982, 1989, 2023]. Start [292, 948, 1406]. Start-up [292]. Start-ups [1406]. Started [3038]. Starting [671]. Starts [2352]. Startup [1404, 1456]. State [271, 504, 1020, 1040, 1393, 1424, 1811, 2335, 2898, 3021]. States [2209]. Static [2328]. Statistics [2688]. Status [1556]. Stay [1019, 1149, 1471, 2386, 2442, 2687]. STC [1527, 1558, 1604]. Steady [462]. STEAM [931]. STEAM-Powered [931]. Steampunk [1592]. Steele [1219]. Steering [356]. Steganography [98]. Stem [931, 1100, 1657, 1845, 1968, 2321]. Step [18, 343, 511, 839, 872, 902, 2540]. Steps [166]. Stereoscopic [377]. Steve [474, 1592]. Sticky [395, 1282]. Stimulation [2409]. Stochastic [2906]. Stochasticity [2873]. Stock [11, 423, 2618]. Stop [1275, 1853]. Storage [65, 907, 909, 927, 1869, 2699]. Store [381]. Stories [586]. Storm [1716]. Story [370, 502, 789, 1659]. Storytelling [839]. Strain [2443]. Strangeness [488]. Stratagainst [707]. Strategic [34]. Strategies [500, 959, 1679, 1748, 1800, 1904, 2131, 2360]. Strategy [67, 109, 1500, 2804]. Stratfor [707]. Stray [2935]. Stream [217]. Streaming [47, 497, 1157]. Streamlining [2724]. Streams [886]. Street [101, 218]. Stress [1212, 1685, 2638]. Structural [1932]. Structure [671]. Structures [3035]. Student [293, 1445, 1652, 1700, 1792, 1871, 2199, 2358, 2363, 2454, 2554, 2583, 2635, 2657]. Student-Sourcing [2454]. Students [788, 930, 1193, 2322]. Studies [2100, 2620]. Study [137, 277, 302, 936, 991, 1782, 1905, 2073, 2616, 2721, 2944]. Studying [1269]. Stuxnet [299]. Style [1551]. Subjective [2616]. Subscribe [2810, 2849, 2866, 2874, 2895, 2978, 2982, 3030, 3060, 3109]. Subscription [2777, 2789, 2942]. Substance [1444]. Subway [2481]. Success [586, 2096, 2444, 2996]. Successes [403]. Successful [2274]. Suite [685]. Summer [1280]. Summit [1083, 1124]. Sun [1973]. Supercomputer [2232]. Supercomputers [16]. Supercomputing [1713]. Superconducting [1751]. Superorganisms [639]. Supervised [2987]. Supply [392, 1990, 1994, 1996, 2527]. Supply-Chain [392, 1990, 1994, 1996]. Support [55, 884, 1347, 1653, 1710, 2098, 2666, 2669]. Supporting [565, 2673]. Supremacy [3046]. Supreme [1206]. Surface [1128]. Surfaces [110]. Surgery [745]. Surprises [358]. Surveillance [532, 1020, 1072, 1153, 1154, 1834, 1835, 2441]. Survey [1732]. Survivable [845]. Survive [1928]. Susan [441]. Suspects [991]. Sustain [2206]. Sustainability [2488]. Sustainable [380, 706, 986, 1494, 2439, 2484, 2511, 2766, 2804, 2816, 2817, 2844, 2879, 2893, 2930, 3039, 3070, 3088]. Sustaining [1401, 2361]. Swarm [2037, 2127]. Swarm-Oriented [2127]. Sweet [1908, 2207]. Symantec [1288, 1348, 1580]. Symbiosis [2649]. Symposium [1362, 1638, 2022, 2064]. Synapses [251]. Syndicate [2877]. Synergistic [2038]. Synergy [241]. Synthesis [1382]. Synthetic [3037]. System [148, 430, 440, 502, 531, 592, 676, 708, 723, 777, 958, 1270, 1283, 1438, 1619, 1782, 1898, 1932, 1934, 2069, 2236, 2264, 2518, 2543, 2560, 3063, 3064]. Systemic [786]. Systems [47, 81– 83, 85, 156, 194, 204, 217, 223, 233, 235, 254, 301, 316, 378, 431, 432, 434, 454, 631, 632, 640, 652, 680, 721, 728, 743, 764, 908, 911, 986, 988, 1071, 1075, 1125, 1279, 1411, 1447, 1468, 1584, 1586, 1616, 1620, 1686, 1751, 1781, 1862, 1867, 1868, 1965, 2003, 2048, 2072, 2159, 2234, 2235, 2237, 2241, 2269, 2276, 2330, 2409, 2453, 2461, 2476, 2499, 2540, 2548, 2598, 2648, 2649, 2717, 2721, 2722, 2726, 2785, 2793, 2869, 2882, 2915, 2934, 2943, 2956, 3005, 3017, 3058, 3069, 3091].

T [1551]. Table [970, 1000, 1030, 1057, 1085, 1107, 1112, 1143, 1175, 1202, 1230, 1258, 1398, 1429, 2027, 2059, 2089, 2116, 2147, 2177, 2198, 2226, 2254, 2282, 2311, 2346, 2373, 2401, 2430, 2458, 2486, 2513, 2533, 2555, 2584, 2609, 2636, 2659, 2684, 2713, 2733, 2749, 2774, 2800, 2802, 2834, 2861, 2890, 2925, 2954, 2985, 3013, 3051, 3077]. Tabletops [508]. Tackles [2394]. Tactics [1047, 1315]. Tactile [1330, 2105]. Tagging [2418]. Tailored [954]. Tailoring [1530]. Take [427, 1679, 1800, 1904, 2131, 2277, 2568]. Takes [645, 992, 2650]. Taking [116, 786, 984, 1820, 2618, 2996]. Tale [165]. Tales [24]. Talking [1218]. Taming [1465, 1625, 2627]. Tanenbaum [1003, 1205]. Tangible [584, 911]. Tangled [2389]. Target [719, 2745, 3079]. Targeted [1837]. Targeting [2751]. Targets [191]. Task [633]. Task-Parallel [633]. Tattoo [576]. Tax [1874]. TCHPC [2658]. Teach [1074, 1388]. Teachers [424]. Teaches [3031]. Teaching [579, 604, 626, 1132, 1251, 2042, 2264, 2337]. Team [2394, 2444]. Teams [1786, 2819]. Tech [1969, 2272, 2370, 2395, 2428, 2456, 2485, 2512, 2578, 2614, 2700]. TechIgnite [2144, 2174, 2195, 2225]. Technical [179, 335, 1835, 2090, 2269, 2658, 2751, 2933, 3079]. Technique [61, 1966]. Techniques [632, 720, 1814, 1893, 2987]. Technological [761, 1437, 1915]. Technologies [29, 120, 262, 270, 377, 906, 1279, 1526, 2169, 2205, 2261, 2380, 2423, 2541, 3026]. Technology [35, 105, 119, 126, 153, 193, 194, 197, 198, 280, 291, 300, 304, 351, 470, 550, 584, 594, 780, 801, 845, 950, 979, 987, 1021, 1121, 1125, 1128, 1238, 1337, 1367, 1439, 1644, 1692, 1710, 1748, 1772, 1796, 2126, 2174, 2210, 2242, 2249, 2368, 2379, 2382, 2446, 2630, 2795, 3092, 3093]. Technology-Mediated [193, 194, 198]. tekMountain [1241]. Teleimmersion [1153]. Telematics [2767]. Telemedia [1907]. Telepresence [1214]. Tell [1134]. Ten [345, 364, 386, 900]. Tenor [921]. Term [566, 897, 1839, 1898]. Terms [773]. Territory [451]. Test [685]. Testbed [64, 1222, 1651]. Testing [483, 703, 1064–1068, 1317, 1337, 1502, 1721, 1931, 2076, 2570, 2575, 2627, 3029]. Text [507, 886, 1596]. Textiles [931]. Thanks [2135]. Theater [800]. Theatrical [456]. Theft [488, 2473]. Their [22, 244, 289, 330, 959, 1094, 2540, 2869]. Them [2069, 3099]. Theme [611]. Theorem [499, 501, 1873]. Theoreticians [412]. Theory [193, 412, 899, 1079, 1193, 1349]. Theory-Guided [1349]. Therac [2433]. Therac-25 [2433]. Therapy [1212]. There [36, 91, 367, 620, 2912]. Thespians [574]. Thin [458]. Thin-Core [458]. Thing [1324, 1363, 1396, 2161]. Things [314, 345, 364, 386, 394, 761, 762, 765, 766, 816, 900, 1410, 1524, 1526, 1647, 1649, 1660, 1876, 1905, 1938, 2010, 2074, 2129, 2130, 2132, 2162, 2189, 2291, 2293, 2294, 2332, 2390, 2440, 2479, 2563, 2594, 2719, 2726, 2846, 2910, 2962, 2968, 3003, 3065, 3083, 3103]. Things-Integrated [3103]. Think [661]. Thinking [293, 1210]. Third [1757, 2181]. Third-Party [1757]. Thoughts [479, 2852]. Threaded [632]. Threat [3027]. Threatcasting [2079]. Threats [12, 505, 1925, 2301, 2416, 2524, 3033]. Three [397, 513, 567, 872, 980, 1472, 2620]. Three-Layer [980]. Three-Step [872]. Threshold [1755]. Thrive [1753]. Throughput [1579, 1723, 2355]. Thwarting [1184, 2387]. Tier [2804]. Time [43, 308, 319, 440, 886, 1159, 1165, 1177, 1199, 1215, 1400, 1441, 1488, 1655, 1662, 2039, 2188, 2215, 2414, 2941, 2959, 3019, 3087]. Time-Sharing [440]. Times [921, 1576]. Tips [605]. Tizen [1189]. Today [976, 2077, 2322]. Together [319, 1016, 2618, 2781]. Tohoku [308]. Told [1226]. Tolerance [2002]. Tolerant [1691].

Tom [257]. Tomorrow [55, 2083, 2322, 2448]. Too [289, 468, 1532, 1964]. Tool [843, 986]. Tools [48, 310, 349, 648, 659, 720, 934, 984, 1331, 1370, 1893, 1943, 1964, 2271, 3090]. Top [161, 577, 2268]. TOP500 [1718]. Topics [3016]. Topology [2329]. Topology-Aware [2329]. Tor [1839]. Touch [584]. Touchscreens [1330]. Toxic [941]. Traceability [1899, 3103]. Tracing [723]. Trackers [1840, 2206]. Tracking [460, 1122, 1472, 1561, 1997, 2998]. Trade [91, 974, 2221]. Trade-Off [91]. Tradeoff [1145]. Tradeoffs [502]. Trading [53, 437]. Traditional [288, 291]. Traffic [172, 2646]. Tragedy [2280]. Train [2721]. Trained [1059]. Training [1121, 3065]. Transaction [1655]. Transaction-Time [1655]. Transactions [1076, 2424, 2439, 2452, 2453, 2484, 2489, 2511, 2571, 2702, 2764, 2766, 2780, 2812, 2816, 2817, 2841, 2844, 2879, 2884, 2893, 2900, 2930, 2940, 2999, 3036, 3039, 3056, 3070, 3076, 3088, 3111]. Transcending [1211]. Transdisciplinary [742]. Transfer [291, 2621, 2741]. Transform [1466, 1783]. Transformation [907, 2590]. Transformative [2931]. Transforming [342, 863, 1418, 2382, 2929]. Transitioning [250, 3037]. Transitions [1424]. Translation [1217, 2991]. Translational [2991, 2994]. Transparency [197, 1069, 2618, 2727, 2757]. Trappings [995]. Traversal [955]. Treatment [482]. Tree [487]. Trenches [1385]. Trends [396, 607, 720, 2272, 2560, 2630, 2643]. Triangle [2318]. Trifecta [912]. Trigger [2436]. Trigger-Action [2436]. Trojan [355, 2892]. Trojans [176, 1997]. Troll [881]. Trolling [2547, 2670]. TrueNorth [2872]. Trump [989, 2495]. Trumps [109]. Trust [84, 183, 355, 571, 796, 894, 1484, 1757, 1784, 1811, 2019, 2159, 2392, 2563, 2646, 2878, 2910]. Trusted [1477]. Trusting [2827]. Trustworthiness [483, 3069]. Trustworthy [156, 176, 355, 1844]. Truth [2823]. Truths [2247]. TSA [1626]. Tsunami [308]. Tsunamis [1716]. TTTC [257]. Tuition [551, 1969]. Tumors [482]. Turing [581, 596, 2075, 2496]. Turning [429, 1316, 2502]. Tutorial [1068]. TV [70, 117]. Twins [358, 799, 2726, 3097]. Twitter [423, 1386, 1929, 2572]. Two [165, 455, 484, 1089, 1520, 2804]. Two-Tier [2804]. Types [3000]. Typical [2592]. Tyranny [1452, 2908].

U.S. [3023]. Ubiquitous [36, 332, 333, 340, 508, 822, 823, 868, 1778, 2499]. Ugly [2767]. Ultimate [334, 1366]. Ultrascale [935]. UMTS [1432]. Uncertainty [258]. Uncharted [451]. Unconstrained [29]. Uncontrollable [2357]. Undeciphered [69]. Under-Rubble [1896]. Undergraduate [109, 770, 1731, 2526, 2815]. Undergraduates [1074]. Undermining [453]. Underrepresented [686]. Understand [41, 524, 3094]. Understandable [2694, 2762]. Understanding [328, 664, 866, 974, 1014, 1356, 1416, 1551, 1687, 1732, 2043, 2616, 2621]. Underwater [2394]. Underwater-Vehicle [2394]. Unexpected [1395]. UniBwM [2466]. Unified [2469]. Unintended [864]. Uninterrupted [653]. unique [2765]. Universe [139]. Universität [2053]. Universities [1239]. University [146, 788, 1240, 2394, 2650]. Unknown [882]. Unlimited [2344, 2356, 2372, 2400, 2411]. Unlocking [2044]. Unrest [1017]. Unseen [735, 1268]. Update [1206]. Upend [2574]. ups [1406]. Upsilon [1700, 2363]. Upton [947]. Urban [331, 639, 2741]. Urgency [2908]. Usable [826, 1532]. Usage [172, 1556]. Use [18, 641, 1539, 1876, 2935]. Used [404]. User [12, 418, 430, 584, 708, 764, 1042, 1556, 1691, 1790, 1875, 1957, 1958, 1962, 2157, 2616, 2672]. User-Centered [764]. User-Independent [1691]. Users [66, 244, 289, 871, 937, 1964, 2206, 3000, 3035]. Using [55, 63, 97, 236, 249, 251, 328, 350, 372, 380, 464, 524, 594, 610, 657, 673, 680, 723, 931, 1074, 1269, 1347, 1685, 1710, 1719, 1783, 1786, 1831, 1840, 1871, 1894, 1955, 2068, 2073, 2097, 2210, 2235, 2266, 2325, 2408, 2477, 2544, 2595, 2597, 2620, 2621, 2667, 3067, 3103]. USNA [2818]. UTA [419]. Utopia [1049]. Utter [58, 166].

V [2014, 2677]. Validated [934]. Validation [2383]. Validity [648]. Valuable [2502]. Value [95, 122, 163, 862, 891, 3003]. Values [2592]. Vapors [1351]. Variability [469, 676, 982, 1075, 1759]. Variable [160]. Variables [1643]. Variance [1216]. Vehicle [1619, 2103, 2394, 2718, 2767]. Vehicles [2466, 2467]. Vehicular [545, 1177, 1783, 2740]. Vendor [2000]. Vendors [1563]. Venture [276, 1242]. Verge [879]. Verification [25, 49, 2134, 2383]. Verified [2215]. Versus [91, 268, 1126, 2521, 2727]. Vertex [1643]. Very [48, 246, 2003]. Vetting [1563]. via [45, 269, 1015, 1184, 1339, 2646]. Video [497, 532, 1157, 1160, 1316, 2414, 2876]. Videogame [262]. View [101, 271, 677, 961, 1298, 2783, 2964, 3040]. Viewer [2128]. Viewpoint [54]. VIII [2365, 2979]. Vint [715]. Viral [526]. Virginia [788]. Virtual [100, 359, 456, 729, 1210– 1213, 1476, 1579, 1892, 1934]. Virtualization [1345, 2101]. Virtualized [477, 724, 1588]. Visa [1972]. Vision [9, 100, 790, 1125, 1213, 1530, 2327, 2466, 2867]. Vision-Based [1125]. Visitation [554, 1536]. Visors [314]. Visual [107, 842, 882–884, 886, 887, 916, 2211, 2758]. Visual-Concept [107]. Visualization [352, 421, 457, 836– 841, 844, 885, 935, 1014, 2233, 2956]. Visually [1125, 2181]. VLSI [2291]. VMware [336]. Voice [266]. Volatile [2560]. Voltage [2355, 2987]. Volunteer [209, 2179]. Voracious [1703]. Voted [461, 730, 1023, 1391, 1761, 2138, 2480, 2747, 3112]. Voting [104, 571, 1123, 1283, 2077]. vs [1843, 2337]. VSSC [1283]. VSSC/1622 [1283]. Vu [751, 3042]. Vulnerabilities [391, 436, 506, 2651]. Vulnerability [3101].

W. [296]. Wager [1873]. Wall [218, 574]. Wallace [296]. Want [151, 948, 1819]. Wanted [2075]. War [2104, 2350]. Warehouse [656]. Warehouse-Scale [656]. Warning [1177]. Wars [1760, 2445]. Was [367, 2200]. Waste [1626]. Wasted [402]. Watch [1679, 1800, 1904, 2131]. Water [757]. Watson [1792]. Wave [149, 1010, 1335, 2934]. Waves [1716]. Way [190, 1027, 1269, 1415]. Ways [772, 825, 2648]. Weakest [1589]. Weakest-Link [1589]. Weaponization [2416]. Weaponizing [2572]. Weapons [1315]. Wearable [1553, 1555, 1644, 1691, 2207, 2210, 2211]. Wearables [1547, 1557, 1574, 1603, 1617, 1689, 2701, 2706]. Weather [1714]. Weave [1755]. Web [5, 15, 49, 56, 60, 92, 121, 127, 202, 204, 310, 312, 506, 727, 1066, 1096, 1098, 1464, 1522– 1524, 1526, 1528, 1593, 1653, 2245, 2301, 2335, 2389, 2436, 2493, 2615, 2624, 2964]. Web-Scale [1464]. Website [2887]. Websites [173, 1497, 1625]. We’d [1395]. Weight [3065]. Weird [2357]. Weiser [1778]. Welfare [3026]. We’ll [1122, 1974, 2094, 2337, 2653]. Well-Being [1122, 2094, 2337, 2653]. Wellth [1347]. Were [2725]. Weren’t [1395]. Whatever [153]. Wheels [379]. Where [286, 411, 605, 709, 927, 1193, 1769, 2087, 2114, 2145, 2175, 2196]. wherever [1293, 2568]. Whether [3106]. Which [671, 2335]. while [214, 1132, 2765]. White [1381]. Who [817, 832, 2543]. Whole [1443]. Wi [190, 430, 520, 2158]. Wi-Fi [190, 430, 520, 2158]. Wide [814, 824].

Wide-Area [814]. WikiLeaks [529]. Wikipedia [281, 531]. Wikis [1282]. Wild [546, 567, 658, 1353]. Wiley [1331, 1370]. Wiley-IEEE [1331, 1370]. Will [22, 41, 60, 70, 134, 297, 331, 562, 757, 908, 927, 1470, 1511, 1760, 1903, 2746]. Williams [257]. Win [278, 336, 484]. Windows [2277]. Winner [238, 317, 645]. Winners [106, 513, 731]. Wins [32, 68, 257, 296, 399, 419, 438, 439, 621]. Wireless [78, 104, 429, 653, 654, 681, 1154, 1260, 2804, 2935]. Wirelessly [61, 1561]. Wirth [604]. Wisdom [266, 1087, 2518, 3031]. Wise [2518]. Withers [487]. within [1414]. Without [229, 430, 2912]. Women [109, 790, 1314]. Won [596]. Work [214, 367, 599, 733, 1044, 1293, 1903, 2009, 2271, 2843, 3002]. Work/Play [733]. Workflows [159, 1655, 1723]. Workforce [1367, 2813, 2820]. Working [852, 925, 1193, 1480, 2218, 2299]. Workload [1518, 1684, 1812, 2330]. Workloads [1588]. Workplace [2080]. Works [2073]. Workshops [1251]. World [7, 75, 99, 101, 138, 158, 163, 319, 389, 435, 456, 550, 626, 729, 750, 1028, 1376, 1523, 1526, 1679, 1769, 1800, 1824, 1847, 1904, 2126, 2131, 2169, 2435, 2522, 2562, 2602, 2809, 2823, 2858, 2864, 2922, 3024]. Worlds [359, 1131]. Worm [191]. Worrying [1275]. Worse [240]. Wrappers [3102]. Write [1693, 1752, 1930, 1961, 2020, 2530, 2707]. Writing [1003].

X [1988]. Xbox [912]. XIX [1739]. XVIII [1342].

Yale [419]. Yau [2288]. Years [4, 21, 42, 77, 96, 115, 131, 152, 171, 189, 212, 228, 247, 267, 287, 307, 325, 363, 385, 408, 428, 452, 539, 561, 583, 606, 649, 672, 695, 704, 716, 739, 758, 779, 809, 833, 1037, 1091, 1117, 1148, 1179, 1207, 1235, 1262, 1300, 1333, 1372, 1408, 1435, 1461, 1493, 1521, 1552, 1611, 1646, 1681, 1712, 1746, 1777, 1806, 1833, 1889, 1921, 1953, 1986, 2029, 2061, 2091, 2119, 2140, 2202, 2229, 2257, 2259, 2288, 2349, 2378, 2406, 2434, 2464, 2492, 2516, 2538, 2561, 2613, 2641, 2872, 3018, 3053]. Years [59, 475, 496, 519, 567, 628, 857, 878, 901, 926, 949, 975, 1005, 1009, 1062, 1389, 1433, 1522, 1583, 1609, 1863, 2433, 2716, 2736, 2753, 2779, 2808, 2838, 2865, 2896, 2928, 2958, 2990, 3085]. Young [2200]. Youth [686, 1019].

Zero [2872]. ZigBee [553]. Zone [875]. Zooming [1469].

References

Anonymous:2010:CHSa

[1] Anonymous. Computer highlights society magazines. Computer, 43(1): 4–5, January 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Isaak:2010:CSI

[2] James D. Isaak. Computer Society 2010: Innovation, leadership, community, and careers. Computer, 43(1): 6–8, January 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:DF

[3] David Alan Grier. Designing the future. Computer, 43(1):9–11, January 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAa

[4] Anonymous. 32 and 16 years ago. Computer, 43(1):12–13, January 2010.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2010:FRW

[5] Sixto Ortiz Jr. Is 3D finally ready for the Web? Computer, 43(1): 14–16, January 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:NLA

[6] Linda Dailey Paulson. New lithography approach promises powerful chips. Computer, 43(1):17–20, January 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smarr:2010:PGO

[7] Larry Smarr. Project GreenLight: Optimizing cyber-infrastructure for a carbon-constrained world. Computer, 43(1):22–27, January 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parnas:2010:RRF

[8] David Lorge Parnas. Really rethinking ‘formal methods’. Computer, 43 (1):28–34, January 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dobson:2010:FVA

[9] Simon Dobson, Roy Sterritt, Paddy Nixon, and Mike Hinchey. Fulfilling the vision of autonomic computing. Computer, 43(1):35–41, January 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conti:2010:OOC

[10] Marco Conti and Mohan Kumar. Opportunities in opportunistic computing. Computer, 43(1):42–50, January 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schumaker:2010:DSP

[11] Robert P. Schumaker and Hsinchun Chen. A discrete stock price prediction engine based on financial news. Computer, 43(1):51–56, January 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stafford:2010:OST

[12] Thomas F. Stafford and Robin Poston. Online security threats and computer user intentions. Computer, 43 (1):58–64, January 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:CSP

[13] Bob Ward. Computer Society programs serve members. Computer, 43 (1):65–69, January 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:CCa

[14] Bob Ward. Call and calendar. Computer, 43(1):70–71, January 2010. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hendler:2010:WDS

[15] James Hendler. Web 3.0: The dawn of semantic search. Computer, 43(1): 77–80, January 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woh:2010:MSN

[16] Mark Woh, Scott Mahlke, Trevor Mudge, and Chaitali Chakrabarti. Mobile supercomputers for the nextgeneration cell phone. Computer, 43 (1):81–85, January 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2010:ECO

[17] Dirk Riehle. The economic case for open source foundations. Computer, 43 (1):86–90, January 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nordman:2010:PNS

[18] Bruce Nordman and Ken Christensen. Proxying: The next step in reducing IT energy use. Computer, 43(1): 91–93, January 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2010:P

[19] Neville Holmes. The profession. Computer, 43(2):96, 94–95, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:RB

[20] David Alan Grier. Raising barns. Computer, 43(2):7–9, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAb

[21] Anonymous. 32 and 16 years ago. Computer, 43(2):10–11, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2010:WND

[22] Neal Leavitt. Will NoSQL databases live up to their promise? Computer, 43 (2):12–14, February 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:OCK

[23] Linda Dailey Paulson. Onboard cooler keeps chips comfortable. Computer, 43 (2):15–18, February 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rashid:2010:AOS

[24] Awais Rashid, Thomas Cottenier, Phil Greenwood, Ruzanna Chitchyan, Regine Meunier, Roberta Coelho, Mario Sudholt, and Wouter Joosen. Aspect-oriented software development in practice: Tales from AOSD–Europe. Computer, 43(2):19–26, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:2010:ABV

[25] Ron Vetter. Authentication by biometric verification. Computer, 43(2): 28–29, February 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ross:2010:IRP

[26] Arun Ross. Iris recognition: The path forward. Computer, 43(2):30– 35, February 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jain:2010:FM

[27] Anil K. Jain, Jianjiang Feng, and Karthik Nandakumar. Fingerprint matching. Computer, 43(2):36–44, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chellappa:2010:FRC

[28] Rama Chellappa, Pawan Sinha, and P. Jonathon Phillips. Face recognition by computers and humans. Computer, 43(2):46–55, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Savvides:2010:UBI

[29] Marios Savvides, Karl Ricanek, Jr., Damon L. Woodard, and Gerry Dozier. Unconstrained biometric identification: Emerging technologies. Computer, 43 (2):56–62, February 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siponen:2010:CIS

[30] Mikko Siponen, Seppo Pahnila, and M. Adam Mahmood. Compliance with information security policies: An empirical investigation. Computer, 43(2): 64–71, February 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:Ba

[31] Anonymous. Bookshelf. Computer, 43 (2):78, February 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:MWK

[32] Bob Ward. Moore wins Karlsson Award. Computer, 43(2):79–84, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:CCa

[33] Anonymous. Call and calendar. Computer, 43(2):85–86, February 2010.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liebowitz:2010:SIS

[34] Jay Liebowitz. Strategic intelligence, social networking, and knowledge retention. Computer, 43(2):87–89, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

OHagan:2010:ACC

[35] Patricia O’Hagan, Edward Hanna, and Roy Sterritt. Addressing the corrections crisis with software technology. Computer, 43(2):90–93, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schmidt:2010:UCW

[36] Albrecht Schmidt. Ubiquitous computing: Are we there yet? Computer, 43 (2):95–97, February 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gates:2010:DIC

[37] Ann Q. Gates. Discovery, innovation, and creativity: The core of computing. Computer, 43(2):98–100, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Agresti:2010:FFS

[38] William W. Agresti. The four forces shaping cybersecurity. Computer, 43 (2):101–104, February 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kiesewetter-Kobinger:2010:PC

[39] Swen Kiesewetter-Köbinger. Programmers’ capital. Computer, 43(3):108, 106–107, March 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:LEa

[40] Anonymous. Letters to the Editor. Computer, 43(3):6–9, March 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:SYW

[41] David Alan Grier. Someday, you will understand. Computer, 43(3): 11–13, March 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:AYA

[42] Anonymous. 32 and16 years ago. Computer, 43(3):14–15, March 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2010:IRT

[43] David Geer. Is it really time for real-time search? Computer, 43(3): 16–19, March 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:SDC

[44] Linda Dailey Paulson. Scientists develop camera made of fabric. Computer, 43(3):20–22, March 2010. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sodan:2010:PMM

[45] Angela C. Sodan, Jacob Machina, Arash Deshmeh, Kevin Macnaughton, and Bryan Esbaugh. Parallelism via multithreaded and multicore CPUs.

Computer, 43(3):24–32, March 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bach:2010:APP

[46] Moshe (Maury) Bach, Mark Charney, Robert Cohn, Elena Demikhovsky, Tevi Devor, Kim Hazelwood, Aamer Jaleel, Chi-Keung Luk, Gail Lyons, Harish Patil, and Ady Tal. Analyzing parallel programs with pin. Computer, 43(3):34–41, March 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chamberlain:2010:APS

[47] Roger D. Chamberlain, Mark A. Franklin, Eric J. Tyson, James H. Buckley, Jeremy Buhler, Greg Galloway, Saurabh Gayen, Michael Hall, E. F. Berkley Shands, and Naveen Singla. Auto-pipe: Streaming applications on architecturally diverse systems. Computer, 43(3):42–49, March 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pasetto:2010:TVF

[48] Davide Pasetto, Fabrizio Petrini, and Virat Agarwal. Tools for very fast regular expression matching. Computer, 43 (3):50–58, March 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hall:2010:RVW

[49] Sylvain Hall, Tevfik Bultan, Graham Hughes, Muath Alkhalaf, and Roger Villemaire. Runtime verification of Web service interface contracts. Computer, 43(3):59–66, March 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mandal:2010:ITS

[50] Suman K. Mandal, Rabi N. Mahapatra, Praveen S. Bhojwani, and Saraju P. Mohanty. IntellBatt: Toward a smarter battery. Computer, 43 (3):67–71, March 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:CSN

[51] Bob Ward. Computer Society names award recipients. Computer, 43(3): 77–80, March 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:CCb

[52] Bob Ward. Call and calendar. Computer, 43(3):81–82, March 2010. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2010:TGI

[53] Kirk W. Cameron. Trading in green IT. Computer, 43(3):83–85, March 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowen:2010:VFC

[54] Nicholas Bowen and Jim Spohrer. Viewpoint: The future of computing practice and education. Computer, 43 (3):86–89, March 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanAalst:2010:AUP

[55] Wil M. P. van Aalst, Kees M. van Hee, Jan Martijn van Werf, and Marc Verdonk. Auditing 2.0: Using process mining to support tomorrow’s auditor. Computer, 43(3):90–93, March 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:2010:WFC

[56] Thomas M. Chen and Victoria Wang. Web filtering and censoring. Computer, 43(3):94–97, March 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2010:DCE

[57] Neville Holmes. The digital climate of English change. Computer, 43(4): 100, 98–99, April 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:UC

[58] David Alan Grier. Utter chaos. Computer, 43(4):6–9, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAc

[59] Anonymous. 32 and 16 years ago. Computer, 43(4):10–11, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughan-Nichols:2010:WHR

[60] Steven J. Vaughan-Nichols. Will HTML 5 restandardize the Web? Computer, 43(4):13–15, April 2010. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:NTP

[61] Linda Dailey Paulson. New technique provides energy wirelessly. Computer, 43(4):16–19, April 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Falkner:2010:PBL

[62] Nickolas Falkner, Raja Sooriamurthi, and Zbigniew Michalewicz. Puzzlebased learning for engineering and computer science. Computer, 43(4): 20–28, April 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Aiken:2010:UCC

[63] Peter Aiken, Robert M. Stanley, Juanita Billings, and Luke Anderson. Using codes of conduct to resolve legal disputes. Computer, 43(4):29–34, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Avetisyan:2010:OCG

[64] Arutyun I. Avetisyan, Roy Campbell, Indranil Gupta, Michael T. Heath, Steven Y. Ko, Gregory R. Ganger, Michael A. Kozuch, David O’Hallaron, Marcel Kunze, Thomas T. Kwan, Kevin Lai, Martha Lyons, Dejan S. Milojicic, Hing Yan Lee, Yeng Chai Soh, Ng Kwang Ming, Jing-Yuan Luke, and Han Namgoong. Open Cirrus: a global cloud computing testbed. Computer, 43(4):35–43, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walker:2010:LLS

[65] Edward Walker, Walter Brisken, and Jonathan Romney. To lease or not to lease from storage clouds. Computer, 43(4):44–50, April 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:2010:CCM

[66] Karthik Kumar and Yung-Hsiang Lu. Cloud computing for mobile users: Can offloading computation save energy? Computer, 43(4):51–56, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Basili:2010:LSD

[67] Victor R. Basili, Mikael Lindvall, Myrna Regardie, Carolyn Seaman, Jens Heidrich, Jürgen Münch, Dieter Rombach, and Adam Trendowicz. Linking software development and business strategy through measurement. Computer, 43(4):57–65, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:KBW

[68] Bob Ward. Ken Birman wins Kanai Award. Computer, 43(4):66–68, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rao:2010:PAA

[69] Rajesh P. N. Rao. Probabilistic analysis of an ancient undeciphered script. Computer, 43(4):76–80, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Breitman:2010:WTD

[70] Karin Breitman, Markus Endler, Rafael Pereira, and Marcello Azambuja. When TV dies, will it go to the cloud? Computer, 43(4):81–83, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buechley:2010:QI

[71] Leah Buechley. Questioning invisibility. Computer, 43(4):84–86, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vallino:2010:CAD

[72] Jim Vallino. Cutting across the disciplines. Computer, 43(4):87–89, April 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2010:ILP

[73] James Bret Michael, John F. Sarkesain, Thomas C. Wingfield, Georgios Dementis, and Gonzalo Nuno Baptista Sousa. Integrating legal and policy factors in cyberpreparedness. Computer, 43(4):90–92, April 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:Bb

[74] Anonymous. Bookshelf. Computer, 43(4):93, April 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Olsen:2010:CBW

[75] Kai A. Olsen. Computing a better world. Computer, 43(5):96, 94–95, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:AA

[76] David Alan Grier. The age of accountability. Computer, 43(5):6–8, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAd

[77] Anonymous. 32 and 16 years ago. Computer, 43(5):10–11, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2010:FIW

[78] George Lawton. Fighting intrusions into wireless networks. Computer, 43 (5):12–15, May 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:LFC

[79] Linda Dailey Paulson. Libraries face the challenge of archiving digital material. Computer, 43(5):16–19, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wan:2010:WNS

[80] Yun Wan and Gang Peng. What’s next for shopbots? Computer, 43 (5):20–26, May 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coyle:2010:GEI

[81] Lorcan Coyle, Mike Hinchey, Bashar Nuseibeh, and José Luiz Fiadeiro. Guest Editors’ introduction: Evolving critical systems. Computer, 43 (5):28–33, May 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karsai:2010:EES

[82] Gabor Karsai, Fabio Massacci, Leon Osterweil, and Ina Schieferdecker. Evolving embedded systems. Computer, 43(5):34–40, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mens:2010:ESA

[83] Tom Mens, Jeff Magee, and Bernhard Rumpe. Evolving software architecture descriptions of critical systems. Computer, 43(5):42–48, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lund:2010:ERR

[84] Mass Soldal Lund, Bjørnar Solhaug, and Ketil Stølen. Evolution in relation to risk and trust management. Computer, 43(5):49–55, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cohen:2010:WCS

[85] Bernard Cohen and Philip Boxer. Why critical systems need help to evolve. Computer, 43(5):56–63, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Robinson:2010:RCR

[86] William Robinson. A roadmap for comprehensive requirements modeling. Computer, 43(5):64–72, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:Bc

[87] Anonymous. Bookshelf. Computer, 43(5):73, May 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:BMR

[88] Bob Ward. Bertrand Meyer receives 2010 Mills Award. Computer, 43 (5):77–79, May 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2010:CEP

[89] Kirk W. Cameron. The challenges of energy-proportional computing. Computer, 43(5):82–83, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kushiki:2010:FES

[90] Yoshiaki Kushiki. The future of embedded software: Adapting to drastic change. Computer, 43(5):84–86, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vinekar:2010:AVM

[91] Vishnu Vinekar and Christopher L. Huntley. Agility versus maturity: Is there really a trade-off? Computer, 43(5):87–89, May 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeRoure:2010:SW

[92] David De Roure. e-Science and the Web. Computer, 43(5):90–93, May 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2010:RRD

[93] Neville Holmes. The rise and rise of digital gluttony. Computer, 43(5): 96, 94–95, May 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:CHSb

[94] Anonymous. Computer highlights society magazines. Computer, 43(6):4–5, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:VGN

[95] David Alan Grier. The value of a good name. Computer, 43(6):7–9, June 2010.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAe

[96] Anonymous. 32 and 16 years ago. Computer, 43(6):10–11, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2010:CAU

[97] David Geer. Are companies actually using secure development life cycles? Computer, 43(6):12–16, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:SDO

[98] Linda Dailey Paulson. Steganography development offers promise. Computer, 43(6):18–21, June 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bimber:2010:CW

[99] Oliver Bimber. Capturing the world. Computer, 43(6):22–23, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leberl:2010:ACV

[100] Franz Leberl, Horst Bischof, Thomas Pock, Arnold Irschara, and Stefan Kluckner. Aerial computer vision for a 3D virtual habitat. Computer, 43 (6):24–31, June 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anguelov:2010:GSV

[101] Dragomir Anguelov, Carole Dulong, Daniel Filip, Christian Frueh, Stéphane Lafon, Richard Lyon, Abhijit Ogale, Luc Vincent, and Josh Weaver. Google Street View: Capturing the world at street level. Computer, 43(6):32–38, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Agarwal:2010:RR

[102] Sameer Agarwal, Yasutaka Furukawa, Noah Snavely, Brian Curless, Steven M. Seitz, and Richard Szeliski. Reconstructing Rome. Computer, 43(6):40– 47, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goesele:2010:SRC

[103] Michael Goesele, Jens Ackermann, Simon Fuhrmann, Ronny Klowsky, Fabian Langguth, Patrick Mücke, and Martin Ritz. Scene reconstruction from community photo collections. Computer, 43(6):48–53, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tepandi:2010:WPS

[104] Jaak Tepandi, Ilja Tšahhirov, and Stanislav Vassiljev. Wireless PKI security and mobile voting. Computer, 43(6):54–60, June 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sarma:2010:CSC

[105] Anita Sarma, David Redmiles, and André van der Hoek. Categorizing the spectrum of coordination technology. Computer, 43(6):61–67, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:CSA

[106] Bob Ward. Computer Society announces award winners. Computer, 43 (6):72–73, June 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Snoek:2010:VCS

[107] Cees G. M. Snoek and Arnold W. M. Smeulders. Visual-concept search solved? Computer, 43(6):76–78, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhn:2010:AAR

[108] D. Richard Kuhn, Edward J. Coyne, and Timothy R. Weil. Adding attributes to role-based access control. Computer, 43(6):79–81, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barker:2010:STM

[109] Lecia Barker, Joanne McGrath Cohoon, and Lucy Sanders. Strategy trumps money: Recruiting undergraduate women into computing. Computer, 43(6):82–85, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harrison:2010:AIS

[110] Chris Harrison. Appropriated interaction surfaces. Computer, 43(6): 86–89, June 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Margaria:2010:SDA

[111] Tiziana Margaria and Bernhard Steffen. Simplicity as a driver for agile innovation. Computer, 43(6):90–92, June 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rutkowski:2010:GPI

[112] Anne-F. Rutkowski and Carol S. Saunders. Growing pains with information overload. Computer, 43(6):96, 94–95, June 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:CHSc

[113] Anonymous. Computer highlights Society magazines. Computer, 43(7):4–5, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:SC

[114] David Alan Grier. The spirit of combination. Computer, 43(7):6–9, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAf

[115] Anonymous. 32 and 16 years ago. Computer, 43(7):10–11, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2010:TBI

[116] Sixto Ortiz Jr. Taking business intelligence to the masses. Computer, 43 (7):12–15, July 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:NAC

[117] Linda Dailey Paulson. New approach combines TV and social networking. Computer, 43(7):16–19, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Domingo:2010:MHT

[118] Mari Carmen Domingo. Managing healthcare through social networks. Computer, 43(7):20–25, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Piniewski:2010:EHP

[119] Brigitte Piniewski, Johan Muskens, Leonardo Estevez, Randy Carroll, and Rick Cnossen. Empowering healthcare patients with smart technology. Computer, 43(7):27–34, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Palta:2010:LPT

[120] Jatinder R. Palta, Vincent A. Frouhar, and Robert A. Zlotecki. Leveraging pervasive technologies to improve collection of prostate cancer outcome data. Computer, 43(7):35–42, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weitzel:2010:WMP

[121] Mark Weitzel, Andy Smith, Scott de Deugd, and Robert Yates. A Web 2.0 model for patient-centered health informatics applications. Computer, 43 (7):43–50, July 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fleischmann:2010:VCC

[122] Kenneth Fleischmann and William Wallace. Value conflicts in computational modeling. Computer, 43 (7):57–63, July 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huedo:2010:GAM

[123] Eduardo Huedo, Ruben Montero, and Ignacio Llorente. Grid architecture from a metascheduling perspective. Computer, 43(7):51–56, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:Bd

[124] Anonymous. Bookshelf. Computer, 43(7):75, July 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carter:2010:DEE

[125] John Carter and Karthick Rajamani. Designing energy-efficient servers and data centers. Computer, 43(7):76–78, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gowan:2010:EIT

[126] Art Gowan and Han Reichgelt. Emergence of the information technology discipline. Computer, 43(7):79–81, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schall:2010:PHS

[127] Daniel Schall, Schahram Dustdar, and M. Brian Blake. Programming human and software-based Web services. Computer, 43(7):82–85, July 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2010:FCPa

[128] Neville Holmes. The future of the computing profession. Computer, 43(7): 88, 86–87, July 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:CHSd

[129] Anonymous. Computer highlights Society magazines. Computer, 43(8): 4–5, August 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:MD

[130] David Alan Grier. Mental discipline. Computer, 43(8):6–8, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAg

[131] Anonymous. 32 and 16 years ago. Computer, 43(8):9–10, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2010:RFK

[132] Neal Leavitt. Researchers fight to keep implanted medical devices safe from hackers. Computer, 43(8):11–14, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:PCA

[133] Linda Dailey Paulson. Project converts arms and hands into input devices. Computer, 43(8):15–18, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Izydorczyk:2010:MSW

[134] Jacek Izydorczyk and Michal Izydorczyk. Microprocessor scaling: What limits will hold? Computer, 43(8):

20–26, August 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tian:2010:SMC

[135] Yonghong Tian, Jaideep Srivastava, Tiejun Huang, and Noshir Contractor. Social multimedia computing. Computer, 43(8):27–36, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ko:2010:SNC

[136] Moo Nam Ko, Gorrell P. Cheek, Mohamed Shehab, and Ravi Sandhu. Social-networks connect services. Computer, 43(8):37–43, August 2010. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:2010:SHF

[137] Fei-Yue Wang, Daniel Zeng, James A. Hendler, Qingpeng Zhang, Zhuo Feng, Yanqing Gao, Hui Wang, and Guanpi Lai. A study of the human flesh search engine: Crowd-powered expansion of online knowledge. Computer, 43(8): 45–53, August 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bal:2010:RWD

[138] Henri E. Bal, Jason Maassen, Rob V. van Nieuwpoort, Niels Drost, Roelof Kemp, Nick Palmer, Thilo Kielmann, Frank Seinstra, and Ceriel Jacobs. Real-world distributed computer with Ibis. Computer, 43(8):54–62, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zwart:2010:SUI

[139] Simon Portegies Zwart, Derek Groen, Tomoaki Ishiyama, Keigo Nitadori, Junichiro Makino, Cees de Laat, Steve McMillan, Kei Hiraki, Stefan Harfst, and Paola Grosso. Simulating the universe on an intercontinental Grid. Computer, 43(8):63–70, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:ICS

[140] Anonymous. IEEE Computer Society election. Computer, 43(8):71–78, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:IPE

[141] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 43(8):79–83, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:Be

[142] Anonymous. Bookshelf. Computer, 43(8):90, August 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2010:CCD

[143] James Bret Michael, Eneken Tikk, Peter Wahlgren, and Thomas C. Wingfield. From chaos to collective defense. Computer, 43(8):91–94, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Iyengar:2010:PFO

[144] S. S. Iyengar, Supratik Mukhopadhyay, Christopher Steinmuller, and Xin Li. Preventing future oil spills with software-based event detection. Computer, 43(8):95–97, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Frezza:2010:CSI

[145] Stephen T. Frezza. Computer science: Is it really the scientific foundation for software engineering? Computer, 43 (8):98–101, August 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spohrer:2010:IUP

[146] James C. Spohrer. IBM’s university programs. Computer, 43(8):102– 104, August 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hodges:2010:HG

[147] Steve Hodges and Nicolas Villar. The hardware is not a given. Computer, 43(8):106–109, August 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cleve:2010:DIS

[148] Anthony Cleve, Tom Mens, and Jean-Luc Hainaut. Data-intensive system evolution. Computer, 43(8):110– 112, August 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sharkey:2010:CRC

[149] Noel Sharkey, Marc Goodman, and Nick Ross. The coming robot crime wave. Computer, 43(8):116, 114– 115, August 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:CHSe

[150] Anonymous. Computer highlights Society magazines. Computer, 43(9): 4–5, September 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:FWY

[151] David Alan Grier. Finding what you want. Computer, 43(9):8–10, September 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAh

[152] Anonymous. 32 and 16 years ago. Computer, 43(9):11–12, September 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2010:WHN

[153] David Geer. Whatever happened to network-access-control technology? Computer, 43(9):13–16, September 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:NSP

[154] Linda Dailey Paulson. New service promises to improve online search. Computer, 43(9):17–20, September 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mead:2010:ASA

[155] Nancy R. Mead and Joe Jarzombek. Advancing software assurance with public-private collaboration. Computer, 43(9):21–30, September 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sheldon:2010:MTT

[156] Frederick T. Sheldon and Claire Vishik. Moving toward trustworthy systems: R&D essentials. Computer, 43(9):31– 40, September 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Redwine:2010:FSA

[157] Samuel T. Redwine, Jr. Fitting software assurance into higher education. Computer, 43(9):41–46, September 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sobel:2010:ISS

[158] Ann E. K. Sobel and Gary Mc-Graw. Interview: Software security in the real world. Computer, 43(9): 47–53, September 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tan:2010:NAS

[159] Wei Tan, Jia Zhang, and Ian Foster. Network analysis of scientific workflows: a gateway to reuse. Computer, 43(9):54–61, September 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mytkowicz:2010:EOV

[160] Todd Mytkowicz, Amer Diwan, Matthias Hauswirth, and Peter Sweeney. The effect of omitted-variable bias on the evaluation of compiler optimizations. Computer, 43(9):62– 67, September 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:SHF

[161] Bob Ward. Society honors five top computing professionals. Computer, 43(9): 74–76, September 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:CCc

[162] Bob Ward. Call and calendar. Computer, 43(9):77–78, September 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schell:2010:IVM

[163] George P. Schell. IT value: From Moore’s Law to a flat world. Computer, 43(9):79–81, September 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kohavi:2010:OEP

[164] Ron Kohavi, Roger Longbotham, and Toby Walker. Online experiments: Practical lessons. Computer, 43(9): 82–85, September 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2010:TTG

[165] Kirk W. Cameron. A tale of two green lists. Computer, 43(9):86–88, September 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2010:SDS

[166] Neville Holmes. Seven digital steps to avoid utter hell. Computer, 43(9):92, 90–91, September 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:CHSf

[167] Anonymous. Computer highlights Society magazines. Computer, 43(10): 4–5, October 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:L

[168] Anonymous. Letters. Computer, 43 (10):6, October 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:MSH

[169] Anonymous. In memoriam: Scott Hamilton. Computer, 43(10):7, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:CC

[170] David Alan Grier. The curious and the confident. Computer, 43(10):8– 10, October 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAi

[171] Anonymous. 32 and 16 years ago. Computer, 43(10):11–12, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2010:NUC

[172] Neal Leavitt. Network-usage changes push Internet traffic to the edge. Computer, 43(10):13–15, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:HSS

[173] Linda Dailey Paulson. How safe are secure websites? Computer, 43(10): 17–20, October 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marin:2010:RIG

[174] Sergio Toral Marin, Maria del Rocio Martinez-Torres, and Federico Barrero. Reforming ICT graduate programs to meet professional needs. Computer, 43(10):22–29, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roy:2010:EPI

[175] Jarrod A. Roy, Farinaz Koushanfar, and Igor L. Markov. Ending piracy of integrated circuits. Computer, 43(10): 30–38, October 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karri:2010:THI

[176] Ramesh Karri, Jeyavijayan Rajendran, Kurt Rosenfeld, and Mohammad Tehranipoor. Trustworthy hardware: Identifying and classifying hardware trojans. Computer, 43(10):39– 46, October 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2010:CCD

[177] Nir Kshetri. Cloud computing in developing economies. Computer, 43(10): 47–55, October 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamers:2010:SBR

[178] Juan Hamers and Lieven Eeckhout. Scenario-based resource prediction for QoS-aware media processing. Computer, 43(10):56–63, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:SFN

[179] Bob Ward. Society forms new technical committee on semantic computing. Computer, 43(10):64–65, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:Bf

[180] Anonymous. Bookshelf. Computer, 43 (10):76, October 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramakrishnan:2010:MEH

[181] Naren Ramakrishnan, David Hanauer, and Benjamin Keller. Mining electronic health records. Computer, 43 (10):77–81, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

LeBlanc:2010:WSA

[182] Richard J. LeBlanc. What should all CS graduates know? Computer, 43 (10):82–84, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:2010:TWT

[183] Keith W. Miller, Jeffrey Voas, and Phil Laplante. In trust we trust. Computer, 43(10):85–87, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Broy:2010:ACP

[184] Manfred Broy and Ralf Reussner. Architectural concepts in programming languages. Computer, 43(10):88–91, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lukowicz:2010:BSG

[185] Paul Lukowicz, Oliver Amft, Daniel Roggen, and Jingyuan Cheng. Onbody sensing: From gesture-based input to activity-driven interaction. Computer, 43(10):92–96, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2010:FCPb

[186] Neville Holmes. The future of the computing profession: Readers’ E-mails. Computer, 43(10):100, 98–99, October 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:CHSg

[187] Anonymous. Computer highlights Society magazines. Computer, 43(11): 4–5, November 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:S

[188] David Alan Grier. Sabotage! Computer, 43(11):6–8, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAj

[189] Anonymous. 32 and 16 years ago. Computer, 43(11):9–10, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughan-Nichols:2010:GWF

[190] Steven J. Vaughan-Nichols. Gigabit Wi-Fi is on its way. Computer, 43(11): 11–14, November 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:WTI

[191] Anonymous. Worm targets industrialplant operations. Computer, 43(11): 15–18, November 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pirolli:2010:CSA

[192] Peter Pirolli, Jenny Preece, and Ben Shneiderman. Cyberinfrastructure for social action on national priorities. Computer, 43(11):20–21, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kraut:2010:SFC

[193] Robert Kraut, Mary Lou Maher, Judith Olson, Thomas W. Malone, Peter Pirolli, and John C. Thomas. Scientific foundations: a case for technologymediated social-participation theory. Computer, 43(11):22–28, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chi:2010:ADT

[194] Ed H. Chi, Sean Munson, Gerhard Fischer, Sarah Vieweg, and Cynthia Parr. Advancing the design of technology-mediated social participation systems. Computer, 43(11):29– 35, November 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Olson:2010:NMS

[195] Gary Olson, Gloria Mark, Elizabeth Churchill, and Dana Rotman. New missions for a sociotechnical infrastructure. Computer, 43(11):37–43, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hesse:2010:SPH

[196] Bradford W. Hesse, Derek Hansen, Thomas Finholt, Sean Munson, Wendy Kellogg, and John C. Thomas. Social participation in Health 2.0. Computer, 43(11):45–52, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bertot:2010:SMT

[197] John Carlo Bertot, Paul T. Jaeger, Sean Munson, and Tom Glaisyer. Social media technology and government transparency. Computer, 43(11):53– 59, November 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lampe:2010:EPT

[198] Cliff Lampe, Paul Resnick, Andrea Forte, Sarita Yardi, Dana Rotman, Todd Marshall, and Wayne Lutters. Educational priorities for technologymediated social participation. Computer, 43(11):60–67, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:BGA

[199] Bob Ward. Blue Gene architect Gara receives 2010 Cray Award. Computer, 43(11):78–80, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:CCb

[200] Anonymous. Call and calendar. Computer, 43(11):81–82, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:Bg

[201] Anonymous. Bookshelf. Computer, 43 (11):83, November 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hannay:2010:WCW

[202] Timo Hannay. What can the Web do for science? Computer, 43(11): 84–87, November 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Banerjee:2010:OIH

[203] Prith Banerjee, Rich Friedrich, and Lueny Morell. Open innovation at HP Labs. Computer, 43(11):88–90, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Umapathy:2010:SIW

[204] Karthikeyan Umapathy and Sandeep Purao. Systems integration and Web services. Computer, 43(11):91– 94, November 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2010:GE

[205] Kirk W. Cameron. Green with envy. Computer, 43(11):95–96, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2010:SSA

[206] Neville Holmes. Some social aspects of computers and numbers. Computer, 43(11):100, 98–99, November 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:CHSh

[207] Anonymous. Computer highlights Society magazines. Computer, 43(12): 4–5, December 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:LEb

[208] Anonymous. Letters to the Editor. Computer, 43(12):6–8, December 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:2010:IBL

[209] Carl K. Chang. It has been a long journey — from the heart of a loyal Computer Society volunteer. Computer, 43 (12):9–10, December 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Isaak:2010:LF

[210] James D. Isaak. 2010 to 2020 — looking forward. Computer, 43(12): 11–13, December 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2010:II

[211] David A. Grier. Investing in ignorance. Computer, 43(12):14–16, December 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:YAk

[212] Anonymous. 32 and 16 years ago. Computer, 43(12):17–18, December 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2010:PAM

[213] Neal Leavitt. Payment applications make E-commerce mobile. Computer, 43(12):19–22, December 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2010:ESP

[214] Linda Dailey Paulson. Energy-saving PCs work while sleeping. Computer, 43 (12):23–26, December 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huang:2010:MIM

[215] Tiejun Huang, Yonghong Tian, Wen Gao, and Jian Lu. Mediaprinting: Identifying multimedia content for digital rights management. Computer, 43 (12):28–35, December 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yingsaeree:2010:CF

[216] Chaiyakorn Yingsaeree, Giuseppe Nuti, and Philip Treleaven. Computational finance. Computer, 43(12): 36–43, December 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chandramouli:2010:DSM

[217] Badrish Chandramouli, Mohamed Ali, Jonathan Goldstein, Beysim Sezgin, and Balan Sethu Raman. Data stream management systems for computational finance. Computer, 43(12): 45–52, December 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spiers:2010:HPC

[218] Brad Spiers and Denis Wallez. Highperformance computing on Wall Street. Computer, 43(12):53–59, December 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coronato:2010:FDA

[219] Antonio Coronato and Giuseppe De Pietro. Formal design of ambient intelligence applications. Computer, 43(12): 60–68, December 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2010:Bh

[220] Anonymous. Bookshelf. Computer, 43 (12):85, December 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peck:2010:YBY

[221] Evan M. Peck, Erin Treacy Solovey, Krysta Chauncey, Angelo Sassaroli, Sergio Fantini, Robert J. K. Jacob, Audrey Girouard, and Leanne M. Hirshfield. Your brain, your computer, and you. Computer, 43(12): 86–89, December 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurlburt:2010:NPH

[222] George Hurlburt, Jeffrey Voas, Keith Miller, Phil Laplante, and Bret Michael. A nonlinear perspective on higher education. Computer, 43(12):

90–92, December 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vassev:2010:CDA

[223] Emil Vassev and Mike Hinchey. The challenge of developing autonomic systems. Computer, 43(12):93–96, December 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CHSa

[224] Anonymous. Computer highlights Society magazines. Computer, 44(1): 4–5, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2011:PIF

[225] Sorel Reisman. Planning for an inevitable future. Computer, 44(1): 6–8, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:2011:PMO

[226] Ron Vetter. Print, mobile, and online. Computer, 44(1):9–11, January 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:MM

[227] David Alan Grier. The migration to the middle. Computer, 44(1): 12–14, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAa

[228] Anonymous. 32 and 16 years ago. Computer, 44(1):15–16, January 2011.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2011:DGC

[229] George Lawton. 3D displays without glasses: Coming to a screen near you. Computer, 44(1):17–19, January 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2011:SFN

[230] Linda Dailey Paulson. Scanning the future with new barcodes. Computer, 44 (1):20–23, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thiagarajan:2011:ABC

[231] Geetha Thiagarajan, Venkatesh Sarangan, Ramasubramanian Suriyanarayanan, Pragathichitra Sethuraman, Anand Sivasubramaniam, and Avinash Yegyanarayanan. Automating a building’s carbon management. Computer, 44(1): 24–30, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuller:2011:CPG

[232] Samuel H. Fuller and Lynette I. Millett. Computing performance: Game over or next level? Computer, 44(1): 31–38, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ranganathan:2011:MNR

[233] Parthasarathy Ranganathan. From microprocessors to nanostores: Rethinking data-centric systems. Computer, 44 (1):39–48, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stevenson:2011:BID

[234] Daniel S. Stevenson and Robert O. Conn. Bridging the interconnection density gap for exascale computation. Computer, 44(1):49–57, January 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sterritt:2011:ACP

[235] Roy Sterritt. Apoptotic computing: Programmed death by default for computer-based systems. Computer, 44 (1):59–65, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Squicciarini:2011:UMS

[236] Anna Squicciarini, Sathya Dev Rajasekaran, and Marco Casassa Mont. Using modeling and simulation to evaluate enterprises’ risk exposure to social networks. Computer, 44(1):66– 73, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:Ba

[237] Anonymous. Bookshelf. Computer, 44 (1):81, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CSWa

[238] Anonymous. Computer Society winner nets $50,000 prize. Computer, 44 (1):82–84, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CCa

[239] Anonymous. Call and calendar. Computer, 44(1):85–86, January 2011. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tomlinson:2011:CME

[240] Bill Tomlinson, M. Six Silberman, and Jim White. Can more efficient IT be worse for the environment? Computer, 44(1):87–89, January 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferguson:2011:ASI

[241] Neil Ferguson. Achieving synergy in the industry-academia relationship. Computer, 44(1):90–92, January 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riedl:2011:PPS

[242] John Riedl. The promise and peril of social computing. Computer, 44(1): 93–95, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ricanek:2011:DHI

[243] Karl Ricanek, Jr. Dissecting the human identity. Computer, 44(1): 96–97, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2011:CTU

[244] Neville Holmes. Computers and their users. Computer, 44(1):100, 98– 99, January 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CHSb

[245] Anonymous. Computer highlights Society magazines. Computer, 44(2):

4–5, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:VBI

[246] David Alan Grier. A very bad idea. Computer, 44(2):6–8, February 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAb

[247] Anonymous. 32 and 16 years ago. Computer, 44(2):9–10, February 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2011:PPD

[248] Sixto Ortiz Jr. Is peer-to-peer on the decline? Computer, 44(2):11– 13, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2011:IPP

[249] Linda Dailey Paulson. IBM project proposes using light to make chips faster. Computer, 44(2):14–17, February 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Menon:2011:TMN

[250] Vivek Menon, Shamik Das, Bharat Jayaraman, and Venu Govindaraju. Transitioning from microelectronics to nanoelectronics. Computer, 44(2):18– 19, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Snider:2011:SCU

[251] Greg Snider, Rick Amerson, Dick Carter, Hisham Abdalla, Muhammad Shakeel Qureshi, Jasmin Leveille, Massimiliano Versace, Heather Ames, Sean Patrick, Benjamin Chandler, Anatoli Gorchetchnikov, and Ennio Mingolla. From synapses to circuitry: Using memristive memory to explore the electronic brain. Computer, 44(2):21–28, February 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schinke:2011:CNF

[252] Daniel Schinke, Neil Di Spigna, Mihir Shiveshwarkar, and Paul Franzon. Computing with novel floatinggate devices. Computer, 44(2):29– 36, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeHon:2011:CSB

[253] Andre DeHon and Benjamin Gojman. Crystals and snowflakes: Building computation from nanowire crossbars. Computer, 44(2):37–45, February 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rao:2011:TFS

[254] Wenjing Rao, Chengmo Yang, Ramesh Karri, and Alex Orailoglu. Toward future systems with nanoscale devices: Overcoming the reliability challenge. Computer, 44(2):46–53, February 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yan:2011:CRS

[255] Jeff Yan and Ahmad Salah El Ahmad. Captcha robustness: a security engineering perspective. Computer, 44(2): 54–60, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Welke:2011:SOA

[256] Richard Welke, Rudy Hirschheim, and Andrew Schwarz. Service-oriented architecture maturity. Computer, 44(2): 61–67, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2011:TWW

[257] Bob Ward. Tom Williams wins TTTC award. Computer, 44(2):74–76, February 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2011:CTU

[258] Phil Laplante, George Hurlburt, Keith Miller, and Jeffrey Voas. Certainty through uncertainty? Computer, 44 (2):79–81, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaefer:2011:FMS

[259] Ina Schaefer and Reiner Hahnle. Formal methods in software product line engineering. Computer, 44(2):82– 85, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schmidt:2011:PBH

[260] Albrecht Schmidt, Marc Langheinrich, and Kritian Kersting. Perception beyond the here and now. Computer, 44 (2):86–88, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sobel:2011:AME

[261] Ann E. K. Sobel. Academic misconduct in the early 21st Century. Computer, 44 (2):89–90, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sung:2011:RVC

[262] Kelvin Sung. Recent videogame console technologies. Computer, 44(2): 91–93, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fabian:2011:CP

[263] Robert Fabian. Computing professionals. Computer, 44(2):96, 94– 95, February 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CHSc

[264] Anonymous. Computer highlights Society magazines. Computer, 44(3): 4–5, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:LEa

[265] Anonymous. Letters to the editor. Computer, 44(3):6, March 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:VW

[266] David Alan Grier. The voice of wisdom. Computer, 44(3):8–10, March 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAc

[267] Anonymous. 32 and 16 years ago. Computer, 44(3):11–12, March 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2011:EVM

[268] George Lawton. 4G: Engineering versus marketing. Computer, 44(3): 14–16, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2011:AIO

[269] Linda Dailey Paulson. Accessing the Internet via overhead lights. Computer, 44(3):18–21, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Srinivasan:2011:NCC

[270] Savitha Srinivasan and Vladimir Getov. Navigating the cloud computing landscape — technologies, services, and adopters. Computer, 44 (3):22–23, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Narasimhan:2011:SCA

[271] Balakrishna Narasimhan and Ryan Nichols. State of cloud applications and platforms: The cloud adopters’ view. Computer, 44(3):24–28, March 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khalidi:2011:BCC

[272] Yousef Khalidi. Building a cloud computing platform for new possibilities. Computer, 44(3):29–34, March 2011.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Banerjee:2011:ESP

[273] Prith Banerjee, Richard Friedrich, Cullen Bash, Patrick Goldsack, Bernardo Huberman, John Manley, Chandrakant Patel, Parthasarathy Ranganathan, and Alistair Veitch. Everything as a service: Powering the new information economy. Computer, 44 (3):36–43, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rochwerger:2011:RWO

[274] Benny Rochwerger, David Breitgand, Amir Epstein, David Hadas, Irit Loy, Kenneth Nagin, Johan Tordsson, Carmelo Ragusa, Massimo Villari, Stuart Clayman, Eliezer Levy, Alessandro Maraschini, Philippe Massonet, Henar Munoz, and Giovanni Toffetti. Reservoir — when one cloud is not enough. Computer, 44(3):44–51, March 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Olden:2011:ACS

[275] Eric Olden. Architecting a cloudscale identity fabric. Computer, 44 (3):52–59, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vasan:2011:VPC

[276] Robin Vasan. A venture perspective on cloud computing. Computer, 44 (3):60–62, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sukwong:2011:CAS

[277] Orathai Sukwong, Hyong Kim, and James Hoe. Commercial antivirus software effectiveness: An empirical study. Computer, 44(3):63–70, March 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2011:CHW

[278] Bob Ward. Chen, Hescott win Society Education Awards. Computer, 44 (3):71–73, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:Bb

[279] Anonymous. Bookshelf. Computer, 44(3):83, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Phillips:2011:IFR

[280] P. Jonathon Phillips. Improving face recognition technology. Computer, 44 (3):84–86, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lam:2011:PPF

[281] Shyong (Tony) K. Lam and John Riedl. The past, present, and future of Wikipedia. Computer, 44(3): 87–90, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2011:CRR

[282] Kirk W. Cameron. Computing’s role in resource accounting. Computer, 44 (3):91–92, March 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Denning:2011:CSI

[283] Peter J. Denning and Neville Holmes. Computer science: An interview. Computer, 44(3):96, 94–95, March 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CHSd

[284] Anonymous. Computer highlights Society magazines. Computer, 44(4):4–5, April 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:LEb

[285] Anonymous. Letters to the Editor. Computer, 44(4):6–7, April 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:WY

[286] David Alan Grier and Erin Dian Dumbacher. Where are you from? Computer, 44(4):8–10, April 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAd

[287] Anonymous. 32 and 16 years ago. Computer, 44(4):11–12, April 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2011:CAB

[288] Sixto Ortiz Jr. Chipmakers ARM for battle in traditional computing market. Computer, 44(4):14–17, April 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2011:SUF

[289] Linda Dailey Paulson. Some users find the speed of light too slow for their networks. Computer, 44(4):18–21, April 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tarkoma:2011:AMC

[290] Sasu Tarkoma and Eemil Lagerspetz. Arching over the mobile computing chasm: Platforms and runtimes. Computer, 44(4):22–28, April 2011. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fontana:2011:TDA

[291] Steven Fontana. Technology development as an alternative to traditional technology transfer models. Computer, 44(4):30–36, April 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:2011:EII

[292] Thomas Miller III, Stephen Walsh, Seth Hollar, Elaine Rideout, and Beryl Pittman. Engineering and innovation: An immersive start-up experience. Computer, 44(4):38–46, April 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gates:2011:CET

[293] Ann Gates, Rodrigo Romero, Miguel Alonso, Jr., Fanny Klett, J. Fernando Naveda, and Dina Requena. Cultivating entrepreneurial thinking through IEEE-CS Student Chapters. Computer, 44(4):48–55, April 2011. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Savoia:2011:EIG

[294] Alberto Savoia and Patrick Copeland. Entrepreneurial innovation at Google. Computer, 44(4):56–61, April 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:2011:CRI

[295] Toby Myers, R. Geoff Dromey, and Peter Fritzson. Comodeling: From requirements to an integrated software/hardware model. Computer, 44 (4):62–70, April 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2011:AWW

[296] Bob Ward. Akyildiz wins W. Wallace McDowell Award. Computer, 44 (4):71–73, April 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kruger:2011:HCW

[297] Antonio Kruger, Johannes Schoning, and Patrick Olivier. How computing will change the face of retail. Computer, 44(4):84–87, April 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vandierendonck:2011:ANS

[298] Hans Vandierendonck and Tom Mens. Averting the next software crisis. Computer, 44(4):88–90, April 2011. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:2011:LS

[299] Thomas M. Chen and Saeed Abu-Nimeh. Lessons from Stuxnet. Computer, 44(4):91–93, April 2011. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leyvand:2011:KIT

[300] Tommer Leyvand, Casey Meekhof, Yi-Chen Wei, Jian Sun, and Baining Guo. Kinect Identity: Technology and experience. Computer, 44(4):94–96, April 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sluss:2011:EEA

[301] James J. Sluss, Jr. Engineering education activities in electric energy systems. Computer, 44(4):97–98, April 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ye:2011:MLA

[302] Jieping Ye, Teresa Wu, Jing Li, and Kewei Chen. Machine learning approaches for the neuroimaging study of Alzheimer’s disease. Computer, 44 (4):99–101, April 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2011:EDP

[303] Dejan Milojicic and Rich Wolski. Eucalyptus: Delivering a private cloud. Computer, 44(4):102–104, April 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morris:2011:DT

[304] Chris Morris and Kai A. Olsen. Democracy and technology. Computer, 44(4): 108, 106–107, April 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2011:CHSa

[305] Bob Ward. Computer highlights Society magazines. Computer, 44(5):4–5, May 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:AM

[306] David Alan Grier. Not for all markets. Computer, 44(5):6–8, May 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAe

[307] Anonymous. 32 and 16 years ago. Computer, 44(5):9–10, May 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

World:2011:FGT

[308] Linda World. Forecasting in geologic time: The Tohoku quake and tsunami. Computer, 44(5):11–13, May 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:FRD

[309] Anonymous. Flying robots designed to form emergency network. Computer, 44(5):14–16, May 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moreno:2011:TEO

[310] Lourdes Moreno, Paloma Martinez, Belen Ruiz, and Ana Iglesias. Toward an equal opportunity Web: Applications, standards, and tools that increase accessibility. Computer, 44 (5):18–26, May 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schilit:2011:MCL

[311] Bill N. Schilit. Mobile computing: Looking to the future. Computer, 44 (5):28–29, May 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mikkonen:2011:RWD

[312] Tommi Mikkonen and Antero Taivalsaari. Reports of the Web’s death are greatly exaggerated. Computer, 44 (5):30–36, May 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spagnolli:2011:EFG

[313] Anna Spagnolli, Nicola Corradi, Luciano Gamberini, Eve Hoggan, Giulio Jacucci, Cecilia Katzeff, Loove Broms, and Li Jonsson. Eco-feedback on the go: Motivating energy awareness. Computer, 44(5):38–45, May 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garcia-Macias:2011:BIT

[314] J. Antonio Garcia-Macias, Jorge Alvarez-Lozano, Paul Estrada-Martinez, and Edgardo Aviles-Lopez. Browsing the Internet of Things with sentient visors. Computer, 44(5):46–52, May 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bose:2011:MSA

[315] Raja Bose, Jorg Brakensiek, Keun-Young Park, and Jonathan Lester. Morphing Smartphones into automotive application platforms. Computer, 44(5):53–61, May 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Medvidovic:2011:EHR

[316] Nenad Medvidovic, Hossein Tajalli, Joshua Garcia, Ivo Krka, Yuriy Brun, and George Edwards. Engineering heterogeneous robotics systems: a software architecture-based approach. Computer, 44(5):62–71, May 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2011:GPF

[317] Bob Ward. Grid pioneer Foster named 2011 Kanai Award winner. Computer, 44(5):83–85, May 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huntley:2011:DP

[318] Christopher L. Huntley. The developer’s perspective. Computer, 44 (5):88–90, May 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ranganathan:2011:SWO

[319] Parthasarathy Ranganathan and Jichuan Chang. Saving the world, one server at a time, together. Computer, 44 (5):91–93, May 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jain:2011:MFS

[320] Anil K. Jain and Brendan Klare. Matching forensic sketches and mug shots to apprehend criminals. Computer, 44(5):94–96, May 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sen:2011:FF

[321] Shilad Sen and John Riedl. Folksonomy formation. Computer, 44(5): 97–101, May 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2011:ADO

[322] Neville Holmes. Aspects of data obesity. Computer, 44(5):104, 102–103, May 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2011:CHSb

[323] Bob Ward. Computer highlights Society magazines. Computer, 44(6):4–5, June 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:DGS

[324] David Alan Grier. Dumb grids and smart markets. Computer, 44(6):6–8, June 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAf

[325] Anonymous. 32 and 16 years ago. Computer, 44(6):9–10, June 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2011:MSF

[326] Neal Leavitt. Mobile security: Finally a serious problem? Computer, 44 (6):11–14, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:NBa

[327] Anonymous. News briefs. Computer, 44(6):18–20, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Honicky:2011:UUR

[328] R. J. Honicky. Understanding and using rendezvous to enhance mobile crowdsourcing applications. Computer, 44(6):22–28, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Helal:2011:IFG

[329] Sumi Helal. IT footprinting — groundwork for future smart cities. Computer, 44(6):30–31, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Naphade:2011:SCT

[330] Milind Naphade, Guruduth Banavar, Colin Harrison, Jurij Paraszczak, and Robert Morris. Smarter cities and their innovation challenges. Computer, 44 (6):32–39, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuikkaniemi:2011:SSH

[331] Kai Kuikkaniemi, Giulio Jacucci, Marko Turpeinen, Eve Hoggan, and Jorg Muller. From space to stage: How interactive screens will change urban life. Computer, 44(6):40–47, June 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gil-Castineira:2011:EIU

[332] Felipe Gil-Castineira, Enrique Costa-Montenegro, Francisco J. Gonzalez-Castano, Cristina Lopez-Bravo, Timo Ojala, and Raja Bose. Experiences inside the ubiquitous Oulu smart city. Computer, 44(6):48–55, June 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2011:BIS

[333] Jungwoo Lee, Songhoon Baik, and Choonhwa Lee. Building an integrated service management platform for ubiquitous cities. Computer, 44 (6):56–63, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mukherjee:2011:RGU

[334] Nilanjan Mukherjee, Janusz Rajski, Grzegorz Mrugalski, Artur Pogiel, and Jerzy Tyszer. Ring generator: An ultimate linear feedback shift register. Computer, 44(6):64–71, June 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2011:STC

[335] Dejan Milojicic and Phil Laplante. Special technical communities. Computer, 44(6):84–88, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2011:KRC

[336] Bob Ward. Kuck receives Computer Pioneer Award; VMware’s Greene and Rosenblum win Computer Entrepreneur Award; dozens of new CSDP and CSDA holders named. Computer, 44(6):89–91, June 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2011:CC

[337] Bob Ward. Call and calendar. Computer, 44(6):92–93, June 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pardo:2011:HQA

[338] Cesar Pardo, Francisco J. Pino, Felix Garcia, and Mario Piattini. Harmonizing quality assurance processes and product characteristics. Computer, 44 (6):94–96, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Solaro:2011:KDB

[339] John Solaro. The Kinect digital outof-box experience. Computer, 44(6): 97–99, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Breznitz:2011:UDC

[340] Dan Breznitz, Michael Murphree, and Seymour Goodman. Ubiquitous data collection: Rethinking privacy debates. Computer, 44(6):100–102, June 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rudin:2011:CDM

[341] Cynthia Rudin, Rebecca J. Passonneau, Axinia Radeva, Steve Ierome, and Delfina F. Isaac. 21st-Century data miners meet 19th-Century electrical cables. Computer, 44(6):103–105, June 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cuny:2011:TCS

[342] Jan Cuny. Transforming computer science education in high schools. Computer, 44(6):107–109, June 2011. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2011:CSF

[343] Neville Holmes. Computing and the step function. Computer, 44(6):112, 110–111, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:EH

[344] David Alan Grier. Experienced hands. Computer, 44(7):6–8, July 2011. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2011:TTKa

[345] Brian M. Gaff and Catherine J. Toppin. Ten things to know when applying for a patent. Computer, 44(7):11–12, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2011:PCC

[346] Sixto Ortiz Jr. The problem with cloud-computing standardization. Computer, 44(7):13–16, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2011:NB

[347] Lee Garber. News briefs. Computer, 44(7):17–19, July 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:2011:ESC

[348] Daqing Zhang, Bin Guo, and Zhiwen Yu. The emergence of social and community intelligence. Computer, 44 (7):21–28, July 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bimber:2011:CAR

[349] Oliver Bimber and Carl K. Chang. Computational archaeology: Reviving the past with present-day tools. Computer, 44(7):30–31, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Edmunds:2011:UCD

[350] Michael Edmunds and Tony Freeth. Using computation to decode the first known computer. Computer, 44(7): 32–39, July 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chng:2011:SLH

[351] Eugene Ch’ng, Henry Chapman, Vincent Gaffney, Phil Murgatroyd, Chris Gaffney, and Wolfgang Neubauer. From sites to landscapes: How computing technology is shaping archaeological practice. Computer, 44(7):40–46, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Scopigno:2011:MCH

[352] Roberto Scopigno, Marco Callieri, Paolo Cignoni, Massimiliano Corsini, Matteo Dellepiane, Federico Ponchio, and Guido Ranzuglia. 3D models for cultural heritage: Beyond plain visualization. Computer, 44(7):48–55, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Petrovic:2011:DAD

[353] Vid Petrovic, Aaron Gidding, Tom Wypych, Falko Kuester, Thomas A. DeFanti, and Thomas E. Levy. Dealing with archaeology’s data avalanche. Computer, 44(7):56–60, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2011:CAA

[354] Renju Li, Tao Luo, Hongbin Zha, and Wei Lu. Computer-assisted archaeological line drawing. Computer, 44 (7):62–65, July 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tehranipoor:2011:THT

[355] Mohammad Tehranipoor, Hassan Salmani, Xuehui Zhang, Xiaoxiao Wang, Ramesh Karri, Jeyavijayan Rajendran, and Kurt Rosenfeld. Trustworthy hardware: Trojan detection and design-for-trust challenges. Computer, 44(7):66–74, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2011:CSO

[356] Dirk Riehle. Controlling and steering open source projects. Computer, 44 (7):93–96, July 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Larrick:2011:CBM

[357] Richard P. Larrick and Kirk W. Cameron. Consumption-based metrics: From autos to IT. Computer, 44(7):97–99, July 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowyer:2011:WSD

[358] Kevin W. Bowyer. What surprises do identical twins have for identity science? Computer, 44(7):100–102, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lam:2011:EMI

[359] Shyong (Tony) K. Lam and John Riedl. Expressing my inner gnome: Appearance and behavior in virtual worlds. Computer, 44(7):103–105, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Olsen:2011:PP

[360] Kai A. Olsen. Programmed politeness. Computer, 44(7):108, 106–107, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CHSe

[361] Anonymous. Computer highlights Society magazines. Computer, 44(8): 4–5, August 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:CB

[362] David Alan Grier. The chicken bus. Computer, 44(8):6–8, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAg

[363] Anonymous. 32 and 16 years ago. Computer, 44(8):9–10, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2011:TTKb

[364] Brian M. Gaff and Catherine J. Toppin. Ten things to know when your patent application is pending. Computer, 44 (8):11–12, August 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughan-Nichols:2011:ONG

[365] Steven J. Vaughan-Nichols. OpenFlow: The next generation of the network? Computer, 44(8):13–15, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:NBb

[366] Anonymous. News briefs. Computer, 44(8):16–18, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goth:2011:IPR

[367] Greg Goth. IBM PC retrospective: There was enough right to make it work. Computer, 44(8):26–33, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bradley:2011:PHI

[368] David Bradley. A personal history of the IBM PC. Computer, 44(8): 19–25, August 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bride:2011:IPC

[369] Edward Bride. The IBM Personal Computer: a software-driven market. Computer, 44(8):34–39, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singh:2011:IPS

[370] Gurbir Singh. The IBM PC: The silicon story. Computer, 44(8):40–45, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Simoens:2011:RDS

[371] Pieter Simoens, Filip De Turck, Bart Dhoedt, and Piet Demeester. Remote display solutions for mobile cloud computing. Computer, 44(8):46–53, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yeo:2011:UMM

[372] Sungkap Yeo and Hsien-Hsin Lee. Using mathematical modeling in provisioning a heterogeneous cloud computing environment. Computer, 44(8): 55–62, August 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chesnais:2011:PIC

[373] Alain Chesnais and Sorel Reisman. From the Presidents of the IEEE Computer Society and ACM. Computer, 44 (8):63–64, August 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:ICS

[374] Anonymous. IEEE Computer Society election. Computer, 44(8):65–73, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:IPE

[375] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 44(8):75–80, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CCb

[376] Anonymous. Call and calendar. Computer, 44(8):81–82, August 2011. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bryant:2011:STE

[377] John Bryant and Jason Bay. Stereoscopic technologies and effects. Computer, 44(8):93–95, August 2011. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vassev:2011:KRR

[378] Emil Vassev and Mike Hinchey. Knowledge representation and reasoning for intelligent software systems. Computer, 44(8):96–99, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tscheligi:2011:ICW

[379] Manfred Tscheligi, Alexander Meschtscherjakov, and David Wilfinger. Interactive computing on wheels. Computer, 44(8): 100–102, August 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marwah:2011:UDM

[380] Manish Marwah, Amip Shah, Cullen Bash, Chandrakant Patel, and Naren Ramakrishnan. Using data mining to help design sustainable products. Computer, 44(8):103–106, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2011:TDE

[381] Charles Severance. Toward developing an education app store. Computer, 44(8):107–109, August 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2011:CPK

[382] Neville Holmes. The cultural potential of keyboards. Computer, 44(8):112, 110–111, August 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CHSf

[383] Anonymous. Computer highlights Society magazines. Computer, 44(9): 4–5, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:LS

[384] David Alan Grier. Leisure science. Computer, 44(9):6–8, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAh

[385] Anonymous. 32 and 16 years ago. Computer, 44(9):9–10, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2011:TTKc

[386] Brian M. Gaff and Catherine J. Toppin. Ten things to know when your patent application is allowed. Computer, 44 (9):11–13, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2011:ICA

[387] Neal Leavitt. IPv6: Any closer to adoption? Computer, 44(9):14– 16, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:NBc

[388] Anonymous. News briefs. Computer, 44(9):18–20, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oppliger:2011:SPO

[389] Rolf Oppliger. Security and privacy in an online world. Computer, 44(9): 21–22, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abu-Nimeh:2011:MSP

[390] Saeed Abu-Nimeh, Thomas M. Chen, and Omar Alzubi. Malicious and spam posts in online social networks. Computer, 44(9):23–28, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Saiedian:2011:SVS

[391] Hossein Saiedian and Dan S. Broyles. Security vulnerabilities in the sameorigin policy: Implications and alternatives. Computer, 44(9):29–36, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kerschbaum:2011:SCS

[392] Florian Kerschbaum, Axel Schropfer, Antonio Zilli, Richard Pibernik, Octavian Catrina, Sebastiaan de Hoogh, Berry Schoenmakers, Stelvio Cimato, and Ernesto Damiani. Secure collaborative supply-chain management. Computer, 44(9):38–43, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rocha:2011:FFC

[393] Francisco Rocha, Salvador Abreu, and Miguel Correia. The final frontier: Confidentiality and privacy in the cloud. Computer, 44(9):44–50, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roman:2011:SIT

[394] Rodrigo Roman, Pablo Najera, and Javier Lopez. Securing the Internet of Things. Computer, 44(9):51– 58, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pearson:2011:SPA

[395] Siani Pearson and Marco Casassa Mont. Sticky policies: An approach for managing privacy across multiple parties. Computer, 44(9):60–68, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ryckbosch:2011:TSE

[396] Frederick Ryckbosch, Stijn Polfliet, and Lieven Eeckhout. Trends in server energy proportionality. Computer, 44 (9):69–72, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Menon:2011:TRC

[397] Vivek Menon, Bharat Jayaraman, and Venu Govindaraju. The three Rs of cyberphysical spaces. Computer, 44(9): 73–79, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:Bc

[398] Anonymous. Bookshelf. Computer, 44 (9):87, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:BWP

[399] Anonymous. Barrett wins pre-college education award. Computer, 44(9): 88–90, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CCc

[400] Anonymous. Call and calendar. Computer, 44(9):91–92, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lieberman:2011:LGC

[401] Henry Lieberman, Karthik Dinakar, and Birago Jones. Let’s gang up on cyberbullying. Computer, 44(9):93– 96, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhattacharya:2011:SBW

[402] Suparna Bhattacharya, K. Gopinath, Karthick Rajamani, and Manish Gupta. Software bloat and wasted joules: Is modularity a hurdle to green software? Computer, 44(9):97– 101, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huntley:2011:OMA

[403] Christopher L. Huntley. Onshore mobile app development: Successes and challenges. Computer, 44(9):102– 104, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ricanek:2011:WSB

[404] Karl Ricanek, Jr. and Benjamin Barbour. What are soft biometrics and how can they be used? Computer, 44 (9):106–108, September 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Malizia:2011:EBI

[405] Alessio Malizia and Kai A. Olsen. Has everything been invented? on software development and the future of apps. Computer, 44(9):112, 110– 111, September 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CHSg

[406] Anonymous. Computer highlights Society magazines. Computer, 44(10): 4–5, October 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:HC

[407] David Alan Grier. The habit of change. Computer, 44(10):6–8, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAi

[408] Anonymous. 32 and 16 years ago. Computer, 44(10):9–10, October 2011.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2011:BSS

[409] Sixto Ortiz Jr. Bringing 3D to the small screen. Computer, 44(10):11– 13, October 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:NBd

[410] Anonymous. News briefs. Computer, 44(10):14–16, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:2011:WSE

[411] Carl K. Chang, David M. Weiss, and Mike Hinchey. Where software engineering meets …. Computer, 44(10): 17–18, October 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Broy:2011:CPN

[412] Manfred Broy. Can practitioners neglect theory and theoreticians neglect practice? Computer, 44(10): 19–24, October 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fitzgerald:2011:OSS

[413] Brian Fitzgerald. Open source software: Lessons from and for software engineering. Computer, 44(10): 25–30, October 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harman:2011:SEM

[414] Mark Harman. Software engineering meets evolutionary computation.

Computer, 44(10):31–39, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lutz:2011:SES

[415] Robyn Lutz. Software engineering for space exploration. Computer, 44(10): 41–46, October 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yau:2011:SEM

[416] Stephen Yau and Ho An. Software engineering meets services and cloud computing. Computer, 44(10): 47–53, October 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parnas:2011:SEM

[417] David Parnas. Software engineering — missing in action: a personal perspective. Computer, 44(10):54–58, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:2011:DPM

[418] Jae Min Kim, Minyong Kim, Joonho Kong, Hyung Beom Jang, and Sung Woo Chung. Display power management that detects user intent. Computer, 44(10):60–66, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:UYP

[419] Anonymous. UTA’s Yale Patt wins first IEEE CS Rau Award. Computer, 44(10):67–69, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clark:2011:ODE

[420] Justin Clark. Object digitization for everyone. Computer, 44(10):81–83, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:2011:DAV

[421] Min Chen, Anne Trefethen, Rene Banares-Alcantara, Marina Jirotka, Bob Coecke, Thomas Ertl, and Albrecht Schmidt. From data analysis and visualization to causality discovery. Computer, 44(10):84–87, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Basten:2011:GSD

[422] Dirk Basten and Ali Sunyaev. Guidelines for software development effort estimation. Computer, 44(10):88–90, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bollen:2011:TMS

[423] Johan Bollen and Huina Mao. Twitter mood as a stock market predictor. Computer, 44(10):91–94, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Guzdial:2011:LHP

[424] Mark Guzdial. Learning how to prepare computer science high school teachers. Computer, 44(10):95–97, October 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2011:DMA

[425] Neville Holmes. Digital machinery and analog brains. Computer, 44(10):100, 98–99, October 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CHSh

[426] Anonymous. Computer highlights Society magazines. Computer, 44(11): 4–5, November 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:HGT

[427] David Alan Grier. The honest give-and-take. Computer, 44(11):6– 8, November 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAj

[428] Anonymous. 32 and 16 years ago. Computer, 44(11):9–10, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2011:TLW

[429] Lee Garber. Turning on the lights for wireless communications. Computer, 44(11):11–14, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:SIU

[430] Anonymous. System identifies user location without GPS or Wi-Fi. Computer, 44(11):15–17, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Getov:2011:CSA

[431] Vladimir Getov, Adolfy Hoisie, and Harvey J. Wasserman. Codesign for systems and applications: Charting the path to exascale computing. Computer, 44(11):19–21, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shalf:2011:RHS

[432] John Shalf, Dan Quinlan, and Curtis Janssen. Rethinking hardware– software codesign for exascale systems. Computer, 44(11):22–30, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sur:2011:CIC

[433] Sayantan Sur, Sreeram Potluri, Krishna Kandalla, Hari Subramoni, Dhabaleswar K. Panda, and Karen Tomko. Codesign for InfiniBand clusters. Computer, 44(11):31–36, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kerbyson:2011:CCE

[434] Darren Kerbyson, Abhinav Vishnu, Kevin Barker, and Adolfy Hoisie. Codesign challenges for exascale systems: Performance, power, and reliability. Computer, 44(11):37–43, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stanzione:2011:ICC

[435] Dan Stanzione. The iPlant collaborative: Cyberinfrastructure to feed the world. Computer, 44(11):44– 52, November 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Padmanabhuni:2011:DAB

[436] Bindu Madhavi Padmanabhuni and Hee Beng Kuan Tan. Defending against buffer-overflow vulnerabilities. Computer, 44(11):53–60, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nuti:2011:AT

[437] Giuseppe Nuti, Mahnoosh Mirghaemi, Philip Treleaven, and Chaiyakorn Yingsaeree. Algorithmic trading. Computer, 44(11):61–69, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CSWb

[438] Anonymous. Chuck Seitz wins Cray Award. Computer, 44(11):70, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CMW

[439] Anonymous. Cleve Moler wins Fernbach Award. Computer, 44(11):71, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walden:2011:FAM

[440] David Walden. Fiftieth anniversary of MIT’s Compatible Time-Sharing System. Computer, 44(11):71, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:SGR

[441] Anonymous. Susan Graham receives Kennedy Award. Computer, 44(11): 72, November 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CCd

[442] Anonymous. Call and calendar. Computer, 44(11):73–74, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lu:2011:EEE

[443] Yung-Hsiang Lu, Qinru Qiu, Ali R. Butt, and Kirk W. Cameron. End-to-end energy management. Computer, 44 (11):75–77, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ross:2011:HER

[444] Arun Ross and Ayman Abaza. Human ear recognition. Computer, 44(11): 79–81, November 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gomez-Barroso:2011:OMS

[445] Jose Luis Gomez-Barroso, Claudio Feijoo, and Ramon Compano. Opportunities in the mobile search market. Computer, 44(11):83–85, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mukherjee:2011:IAL

[446] Shubu Mukherjee. If anything in this life is certain, it’s that you can kill any ISA. Computer, 44(11):87– 88, November 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Masli:2011:CM

[447] Mikhil Masli. Crowdsourcing maps. Computer, 44(11):90–93, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Olsen:2011:APA

[448] Kai A. Olsen and Alessio Malizia. Automated personal assistants. Computer, 44(11):112, 110–111, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CHSi

[449] Anonymous. Computer highlights Society magazines. Computer, 44(12): 4–5, December 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2011:LBL

[450] Sorel Reisman. Looking back, looking forward. Computer, 44(12):8– 10, December 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2011:UT

[451] David Alan Grier. Uncharted territory. Computer, 44(12):12–14, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:YAk

[452] Anonymous. 32 and 16 years ago. Computer, 44(12):15–16, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2011:ISU

[453] Neal Leavitt. Internet security under attack: The undermining of digital certificates. Computer, 44(12):

17–20, December 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:SSF

[454] Anonymous. Serious security flaws identified in cloud systems. Computer, 44(12):21–23, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Heines:2011:TAI

[455] Jesse M. Heines, Gena R. Greher, S. Alex Ruthmann, and Brendan L. Reilly. Two approaches to interdisciplinary computing + music courses. Computer, 44(12):25–32, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geigel:2011:AVW

[456] Joe Geigel, Marla Schweppe, David Huynh, and Brian Johnstone. Adapting a virtual world for theatrical performance. Computer, 44(12):33–38, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yamaoka:2011:CAL

[457] So Yamaoka, Lev Manovich, Jeremy Douglass, and Falko Kuester. Cultural analytics in large-scale visualization environments. Computer, 44(12): 39–48, December 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marowka:2011:BTC

[458] Ami Marowka. Back to thin-core massively parallel processors. Computer, 44 (12):49–54, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lacity:2011:ROD

[459] Mary Lacity, Erran Carmel, and Joseph Rottman. Rural outsourcing: Delivering ITO and BPO services from remote domestic locations. Computer, 44(12):55–62, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferrari:2011:DCD

[460] Laura Ferrari and Marco Mamei. Discovering city dynamics through sports tracking applications. Computer, 44 (12):63–68, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:GVC

[461] Anonymous. Grier voted 2012 Computer Society President-Elect. Computer, 44(12):69–70, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:NAN

[462] Anonymous. North America new PhD numbers hold steady. Computer, 44 (12):71–73, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:CCe

[463] Anonymous. Call and calendar. Computer, 44(12):74–76, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanderAalst:2011:UPM

[464] Wil van der Aalst. Using process mining to bridge the gap between BI and BPM. Computer, 44(12):77– 80, December 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baldwin:2011:CSR

[465] Douglas Baldwin. Is computer science a relevant academic discipline for the 21st Century? Computer, 44(12): 81–83, December 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tamaki:2011:PCH

[466] Emi Tamaki and Jun Rekimoto. PossessedHand: Controlling hand movements with computer output. Computer, 44(12):84–86, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hills:2011:MYF

[467] Alexander Hills. Museums at your fingertips. Computer, 44(12):87–89, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaffer:2011:PRT

[468] Kim Schaffer. Are password requirements too difficult? Computer, 44(12): 90–92, December 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Capilla:2011:PCR

[469] Rafael Capilla and Jan Bosch. The promise and challenge of runtime variability. Computer, 44(12):93–95, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2011:PDT

[470] Neville Holmes. The profession and digital technology. Computer, 44(12): 116, 114–115, December 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSa

[471] Anonymous. Computer highlights Society magazines. Computer, 45(1): 4–5, January 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walz:2012:COF

[472] John W. Walz. Creating our future. Computer, 45(1):6–7, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:2012:DCN

[473] Ron Vetter. Digital Computer: The next-generation flagship publication. Computer, 45(1):8–9, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:SOE

[474] Charles Severance. The second-order effects of Steve Jobs. Computer, 45 (1):10–11, January 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAa

[475] Anonymous. 32 & 16 years ago. Computer, 45(1):12–13, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:IPPa

[476] Brian M. Gaff, Ralph A. Loren, and Gabriel J. McCool. Intellectual property, Part I. Computer, 45(1):14– 16, January 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2012:CSV

[477] Lee Garber. The challenges of securing the virtualized environment. Computer, 45(1):17–20, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:NBa

[478] Anonymous. News briefs. Computer, 45(1):21–23, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reed:2012:IFT

[479] Daniel A. Reed, Dennis B. Gannon, and James R. Larus. Imagining the future: Thoughts on computing. Computer, 45(1):25–30, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Song:2012:CDP

[480] Dawn Song, Elaine Shi, Ian Fischer, and Umesh Shankar. Cloud data protection for the masses. Computer, 45 (1):39–45, January 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stroustrup:2012:SDI

[481] Bjarne Stroustrup. Software development for infrastructure. Computer, 45 (1):47–58, January 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Iyengar:2012:TMP

[482] S. S. Iyengar, Xin Li, Huanhuan Xu, Supratik Mukhopadhyay, N. Balakrishnan, Amit Sawant, and Puneeth Iyengar. Toward more precise radiotherapy treatment of lung tumors. Computer, 45(1):59–65, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bertolino:2012:ESF

[483] Antonia Bertolino, Guglielmo De Angelis, Sampo Kellomaki, and Andrea Polini. Enhancing service federation trustworthiness through online testing. Computer, 45(1):66–72, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:TCM

[484] Anonymous. Two CS members win US Presidential honors. Computer, 45 (1):73–76, January 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCa

[485] Anonymous. Call and calendar. Computer, 45(1):77–78, January 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jain:2012:ADA

[486] Anil K. Jain and Soweon Yoon. Automatic detection of altered fingerprints. Computer, 45(1):79–82, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lyon:2012:JTW

[487] Doug Lyon. The Java tree withers. Computer, 45(1):83–85, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2012:ITF

[488] Hal Berghel. Identity theft and financial fraud: Some strangeness in the proportions. Computer, 45(1): 86–89, January 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lam:2012:OOF

[489] Shyong (Tony) K. Lam and John Riedl. Are our online “friends” really friends? Computer, 45(1):91–93, January 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tolentino:2012:OPG

[490] Matthew Tolentino and Kirk W. Cameron. The optimist, the pessimist, and the global race to exascale in 20 megawatts. Computer, 45 (1):95–97, January 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:CF

[491] David Alan Grier and Erin Dian Dumbacher. The common future. Computer, 45(1):108, January 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSb

[492] Anonymous. Computer highlights Society magazines. Computer, 45(2): 4–5, February 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:La

[493] Anonymous. Letters. Computer, 45 (2):6, February 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:JDL

[494] Charles Severance. JavaScript: Designing a language in 10 days. Computer, 45(2):7–8, February 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:IPPb

[495] Brian M. Gaff, Ralph A. Loren, and Elizabeth A. Spinney. Intellectual property, Part II. Computer, 45(2): 9–11, February 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAb

[496] Anonymous. 32 & 16 years ago. Computer, 45(2):12–13, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2012:CSB

[497] George Lawton. Cloud streaming brings video to mobile devices. Computer, 45(2):14–16, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2012:NBa

[498] Lee Garber. News briefs. Computer, 45 (2):17–19, February 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shim:2012:GEI

[499] Simon S. Y. Shim. Guest Editor’s introduction: The CAP theorem’s growing impact. Computer, 45(2):21– 22, February 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brewer:2012:PCS

[500] Eric Brewer. Pushing the CAP: Strategies for consistency and availability. Computer, 45(2):23–29, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gilbert:2012:PCT

[501] Seth Gilbert and Nancy A. Lynch. Perspectives on the CAP theorem. Computer, 45(2):30–36, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abadi:2012:CTM

[502] Daniel J. Abadi. Consistency tradeoffs in modern distributed database system design: CAP is only part of the story. Computer, 45(2):37–42, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramakrishnan:2012:CCD

[503] Raghu Ramakrishnan. CAP and cloud data management. Computer, 45(2): 43–49, February 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Birman:2012:OCC

[504] Kenneth P. Birman, Daniel A. Freedman, Qi Huang, and Patrick Dowell. Overcoming CAP with consistent softstate replication. Computer, 45(2):

50–58, February 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kessler:2012:ISN

[505] Gary C. Kessler. Information security: New threats or familiar problems? Computer, 45(2):59–65, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Antunes:2012:DAW

[506] Nuno Antunes and Marco Vieira. Defending against Web application vulnerabilities. Computer, 45(2):66–72, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jee:2012:DEF

[507] Justin Jee, Lee Case Klippel, M. Shahriar Hossain, Naren Ramakrishnan, and Bud Mishra. Discovering the ebb and flow of ideas from text corpora. Computer, 45(2):73–77, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Muller-Tomfelde:2012:TIH

[508] Christian Muller-Tomfelde and Morten Fjeld. Tabletops: Interactive horizontal displays for ubiquitous computing. Computer, 45(2):78–81, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stavrou:2012:BSS

[509] Angelos Stavrou, Jeffrey Voas, Tom Karygiannis, and Steve Quirolgico. Building security into off-the-shelf smartphones. Computer, 45(2):82– 84, February 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gambier:2012:ISG

[510] Jimmy Gambier. An independent social game company’s perspective. Computer, 45(2):85–87, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brinkman:2012:NSE

[511] Bo Brinkman. The next step for ethics education is enforcing the code. Computer, 45(2):88–90, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hsu:2012:IP

[512] Francis Hsu. The impossible and the possible. Computer, 45(2):91– 92, February 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2012:CSN

[513] Bob Ward. Computer Society names winners of three key awards. Computer, 45(2):93–95, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCb

[514] Anonymous. Call and calendar. Computer, 45(2):97–98, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:NI

[515] David Alan Grier and Erin Dian Dumbacher. The new internationalists. Computer, 45(2):112, February 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSc

[516] Anonymous. Computer highlights Society magazines. Computer, 45(3): 4–5, March 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:MMC

[517] Charles Severance. Monash Museum of Computing History. Computer, 45 (3):6–7, March 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:PDS

[518] Brian M. Gaff, Thomas J. Smedinghoff, and Socheth Sor. Privacy and data security. Computer, 45(3):8–10, March 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAc

[519] Anonymous. 32 & 16 years ago. Computer, 45(3):11–12, March 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2012:WFR

[520] Lee Garber. Wi-Fi races into a faster future. Computer, 45(3):13–16, March 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2012:NBb

[521] Lee Garber. News briefs. Computer, 45 (3):18–20, March 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Murali:2012:GEI

[522] T. M. Murali. Guest Editor’s introduction: Computationally driven experimental biology. Computer, 45(3): 22–23, March 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mayhew:2012:CEC

[523] Michael B. Mayhew, Xin Guo, Steven B. Haase, and Alexander J. Hartemink. Close encounters of the collaborative kind. Computer, 45(3): 24–30, March 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koyuturk:2012:UPI

[524] Mehmet Koyuturk. Using protein interaction networks to understand complex diseases. Computer, 45(3):31–38, March 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vandin:2012:AGS

[525] Fabio Vandin, Eli Upfal, and Benjamin J. Raphael. Algorithms and genome sequencing: Identifying driver pathways in cancer. Computer, 45 (3):39–46, March 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Skiena:2012:RVG

[526] Steven S. Skiena. Redesigning viral genomes. Computer, 45(3):47–53, March 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shar:2012:DAC

[527] Lwin Khin Shar and Hee Beng Kuan Tan. Defending against cross-site scripting attacks. Computer, 45(3):

55–62, March 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Quan:2012:LGA

[528] Xiaojun Quan, Liu Wenyin, Wenyu Dou, Hui Xiong, and Yong Ge. Link graph analysis for business site selection. Computer, 45(3):64–69, March 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2012:WMP

[529] Hal Berghel. WikiLeaks and the matter of Private Manning. Computer, 45 (3):70–73, March 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2012:GAL

[530] Kirk W. Cameron and Rong Ge. Generalizing Amdahl’s Law for power and energy. Computer, 45(3):75–77, March 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Halfaker:2012:BCW

[531] Aaron Halfaker and John Riedl. Bots and cyborgs: Wikipedia’s immune system. Computer, 45(3):79–82, March 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Held:2012:IVS

[532] Cornelius Held, Julia Krumm, Petra Markel, and Ralf P. Schenke. Intelligent video surveillance. Computer, 45 (3):83–84, March 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CSCa

[533] Anonymous. Computer Society connection. Computer, 45(3):85–87, March 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2012:CC

[534] Bob Ward. Call and calendar. Computer, 45(3):88–89, March 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:FA

[535] David Alan Grier and Erin Dian Dumbacher. Free agency. Computer, 45 (3):104, March 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSd

[536] Anonymous. Computer highlights Society magazines. Computer, 45(4):4–5, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:DJO

[537] Charles Severance. Discovering JavaScript object notation. Computer, 45(4):6–8, April 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:DAP

[538] Brian M. Gaff, Brian P. Murphy, and Peter J. Cuomo. Defending against patent infringement. Computer, 45 (4):9–12, April 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAd

[539] Anonymous. 32 & 16 years ago. Computer, 45(4):13–14, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2012:NAK

[540] Sixto Ortiz Jr. New approach keeps security in context. Computer, 45 (4):15–17, April 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:NBb

[541] Anonymous. News briefs. Computer, 45(4):18–20, April 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schmidt:2012:IBK

[542] Albrecht Schmidt and Elizabeth Churchill. Interaction beyond the keyboard. Computer, 45(4):21–24, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanErp:2012:BCI

[543] Jan B. F. van Erp, Fabien Lotte, and Michael Tangermann. Brain-computer interfaces: Beyond medical applications. Computer, 45(4):26–34, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gellersen:2012:NIK

[544] Hans Gellersen and Florian Block. Novel interactions on the keyboard. Computer, 45(4):36–40, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riener:2012:GIV

[545] Andreas Riener. Gestural interaction in vehicular applications. Computer, 45(4):42–47, April 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Beaudouin-Lafon:2012:MIW

[546] Michel Beaudouin-Lafon, Stephane Huot, Mathieu Nancel, Wendy Mackay, Emmanuel Pietriga, Romain Primet, Julie Wagner, Olivier Chapuis, Clement Pillias, James R. Eagan, Tony Gjerlufsen, and Clemens Klokmose. Multisurface interaction in the WILD room. Computer, 45(4):48–56, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bianchi:2012:OSD

[547] Andrea Bianchi, Ian Oakley, and Dong-Soo Kwon. Open Sesame: Design guidelines for invisible passwords. Computer, 45(4):58–65, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:2012:PLS

[548] Kai hui Chang and Chris Browy. Parallel logic simulation: Myth or reality? Computer, 45(4):67–73, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yellin:2012:OIP

[549] Daniel Yellin, Yifat Yulevich, Alexander Pyasic, and Avshalom Houri. Optimized interactive processes. Computer, 45(4):74–81, April 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marsden:2012:MTI

[550] Gary Marsden, Edward Cutrell, Matt Jones, Amit A. Nanavati, and Nitendra Rajput. Making technology invisible in the developing world. Computer, 45(4):82–85, April 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hodge:2012:ITA

[551] David C. Hodge. Increasing tuition’s affordability. Computer, 45(4):86–88, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fitzgerald:2012:SC

[552] Brian Fitzgerald. Software crisis 2.0. Computer, 45(4):89–91, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Melaragno:2012:SZP

[553] Anthony Patrick Melaragno, Damindra Bandara, Duminda Wijesekera, and James Bret Michael. Securing the ZigBee protocol in the Smart Grid. Computer, 45(4):92–94, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tsai:2012:MAM

[554] Henry Tsai and Kelvin Sung. Mobile applications and museum visitation. Computer, 45(4):95–98, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CSR

[555] Anonymous. Computer Society recognizes outstanding professionals. Computer, 45(4):99–101, April 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCc

[556] Anonymous. Call and calendar. Computer, 45(4):102–103, April 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:LB

[557] David Alan Grier and Erin Dian Dumbacher. Learning from the best. Computer, 45(4):116, April 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSe

[558] Anonymous. Computer highlights Society magazines. Computer, 45(5):4–5, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:VJG

[559] Charles Severance. Van Jacobson: Getting NSFNet off the ground. Computer, 45(5):6–9, May 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:DNL

[560] Brian M. Gaff, Stephen G. Huggard, and Gregory W. Carey. Do I need A lawyer? If you have to ask, you probably do. Computer, 45(5):10–12, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAe

[561] Anonymous. 32 & 16 years ago. Computer, 45(5):13–14, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2012:WPP

[562] Neal Leavitt. Will power problems curtail processor progress? Computer, 45(5):15–17, May 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2012:NBc

[563] Lee Garber. News briefs. Computer, 45(5):18–20, May 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Want:2012:IDS

[564] Roy Want and Bill N. Schilit. Interactive digital signage. Computer, 45 (5):21–24, May 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Taylor:2012:SCA

[565] Nick Taylor and Keith Cheverst. Supporting community awareness with interactive displays. Computer, 45(5): 26–32, May 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Friday:2012:RLT

[566] Adrian Friday, Nigel Davies, and Christos Efstratiou. Reflections on long-term experiments with public displays. Computer, 45(5):34–41, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ojala:2012:MIP

[567] Timo Ojala, Vassilis Kostakos, Hannu Kukka, Tommi Heikkinen, Tomas Linden, Marko Jurmu, Simo Hosio, Fabio Kruger, and Daniele Zanni. Multipurpose interactive public displays in the wild: Three years later. Computer, 45(5):42–49, May 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alt:2012:APD

[568] Florian Alt, Jorg Muller, and Albrecht Schmidt. Advertising on public display networks. Computer, 45(5):50–56, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davies:2012:ODN

[569] Nigel Davies, Marc Langheinrich, Rui Jose, and Albrecht Schmidt. Open display networks: a communications medium for the 21st Century. Computer, 45(5):58–64, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Erbes:2012:FEI

[570] Jamie Erbes, Hamid R. Motahari-Nezhad, and Sven Graupner. The future of enterprise IT in the cloud. Computer, 45(5):66–72, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nestas:2012:BMT

[571] Lars Hopland Nestas and Kjell J. Hole. Building and maintaining trust in Internet voting. Computer, 45 (5):74–80, May 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hey:2012:FDI

[572] Tony Hey, Dennis Gannon, and Jim Pinkelman. The future of dataintensive science. Computer, 45(5): 81–82, May 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xu:2012:MGI

[573] Zhiwei Xu. Measuring green IT in society. Computer, 45(5):83–85, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2012:BFW

[574] Hal Berghel. Breaking the fourth wall of electronic crime: Blame it on the thespians. Computer, 45(5):86–88, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Malizia:2012:TNS

[575] Alessio Malizia and Kai A. Olsen. Toward a new search paradigm — can we learn from ants? Computer, 45 (5):89–91, May 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jain:2012:TIM

[576] Anil K. Jain, Rong Jin, and Jung-Eun Lee. Tattoo image matching and retrieval. Computer, 45(5):93–96, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CSH

[577] Anonymous. Computer Society honors top scientists. Computer, 45 (5):97–99, May 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCd

[578] Anonymous. Call and calendar. Computer, 45(5):100–102, May 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:TI

[579] David Alan Grier and Erin Dian Dumbacher. Teaching the intangible. Computer, 45(5):116, May 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSf

[580] Anonymous. Computer highlights Society magazines. Computer, 45(6):4–5, June 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:ATB

[581] Charles Severance. Alan Turing and Bletchley Park. Computer, 45(6):6–8, June 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:OSS

[582] Brian M. Gaff and Gregory J. Ploussios. Open source software. Computer, 45(6):9–11, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAf

[583] Anonymous. 32 & 16 years ago. Computer, 45(6):12–13, June 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2012:TUI

[584] Lee Garber. Tangible user interfaces: Technology you can touch. Computer, 45(6):15–18, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:NBc

[585] Anonymous. News briefs. Computer, 45(6):19–21, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:2012:CAS

[586] Carl Chang, Vladimir Getov, and Kelvin Sung. Computing in Asia: a sampling of recent success stories. Computer, 45(6):23–25, June 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mei:2012:ISP

[587] Hong Mei, Gang Huang, and Tao Xie. Internetware: a software paradigm for Internet computing. Computer, 45 (6):26–31, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Aoyama:2012:CNG

[588] Mikio Aoyama. Computing for the next-generation automobile. Computer, 45(6):32–37, June 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:2012:CAA

[589] Lijuan Wang, Yao Qian, Matthew Scott, Gang Chen, and Frank Soong. Computer-assisted audiovisual language learning. Computer, 45(6):38–47, June 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sung:2012:BAG

[590] Wing-Kin Sung. Bioinformatics applications in genomics. Computer, 45 (6):57–63, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:2012:FLB

[591] Zhenyu Zhang, W. K. Chan, and T. H. Tse. Fault localization based only on failed runs. Computer, 45 (6):64–71, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Israel:2012:WFC

[592] Jerome Israel. Why the FBI can’t build a case management system. Computer, 45(6):73–80, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Franqueira:2012:RBA

[593] Virginia N. L. Franqueira and Roel J. Wieringa. Role-based access control in retrospect. Computer, 45 (6):81–88, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Munson:2012:UTA

[594] Jeff Munson and Phillip Pasqual. Using technology in autism research: The promise and the perils. Computer, 45 (6):89–91, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Simmonds:2012:PPE

[595] Devon M. Simmonds. The programming paradigm evolution. Computer, 45(6):93–95, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lipton:2012:MTW

[596] Richard J. Lipton. Might Turing have won a Turing Award? Computer, 45 (6):96–97, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harle:2012:SSO

[597] Robert Harle and Andy Hopper. Sports sensing: An Olympic challenge for computing. Computer, 45(6):98–101, June 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oppliger:2012:CMC

[598] Rolf Oppliger and Bruno Wildhaber. Common misconceptions in computer and information security. Computer, 45 (6):102–104, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:SHF

[599] Anonymous. Society honors Fagin for work on databases. Computer, 45(6): 105–106, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCe

[600] Anonymous. Call and calendar. Computer, 45(6):107–109, June 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:CS

[601] David Alan Grier and Erin Dian Dumbacher. Controlling share. Computer, 45(6):128, June 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSg

[602] Anonymous. Computer highlights Society magazines. Computer, 45(7):4–5, July 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:Lb

[603] Anonymous. Letters. Computer, 45 (7):6–7, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:ATC

[604] Charles Severance. The art of teaching computer science: Niklaus Wirth. Computer, 45(7):8–10, July 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:RCW

[605] Brian M. Gaff, Richard N. Kimball, and Jill M. Hanson. Raising capital: Where to find it, how to secure it, and tips on what to avoid. Computer, 45 (7):11–13, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAg

[606] Anonymous. 32 & 16 years ago. Computer, 45(7):14–15, July 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2012:CTL

[607] Sixto Ortiz Jr. Computing trends lead to new programming languages. Computer, 45(7):17–20, July 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:NBd

[608] Anonymous. News briefs. Computer, 45(7):21–23, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bimber:2012:WRA

[609] Oliver Bimber. What’s real about augmented reality? Computer, 45 (7):24–25, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gervautz:2012:AIU

[610] Michael Gervautz and Dieter Schmalstieg. Anywhere interfaces using handheld augmented reality. Computer, 45 (7):26–31, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mine:2012:PBA

[611] Mark Mine, David Rose, Bei Yang, Jeroen van Baar, and Anselm Grundhöfer. Projection-based augmented reality in Disney theme parks. Computer, 45 (7):32–40, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barry:2012:ARP

[612] Ailsa Barry, Jonathan Trout, Paul Debenham, and Graham Thomas. Augmented reality in a public space: The Natural History Museum, London. Computer, 45(7):42–47, July 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Navab:2012:FDA

[613] Nassir Navab, Tobias Blum, Lejing Wang, Aslı, Okur, and Thomas Wendler. First deployments of augmented reality in operating rooms. Computer, 45(7):48–55, July 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Billinghurst:2012:ARC

[614] Mark Billinghurst and Andreas Dünser. Augmented reality in the classroom. Computer, 45(7):56–63, July 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

OLeary:2012:CBP

[615] Daniel E. O’Leary. Computer-based political action: The battle and Internet blackout over PIPA. Computer, 45(7):64–72, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:2012:DSA

[616] Jianyong Chen, Yang Wang, and Xiaomin Wang. On-demand security architecture for cloud computing. Computer, 45(7):73–78, July 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2012:ODM

[617] Hal Berghel. The OuijAmazon desktop metaphor. Computer, 45(7):80–82, July 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Frachtenberg:2012:HDD

[618] Eitan Frachtenberg. Holistic datacenter design in the Open Compute Project. Computer, 45(7):83–85, July 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campisi:2012:EAP

[619] Patrizio Campisi, Daria La Rocca, and Gaetano Scarano. EEG for automatic person recognition. Computer, 45(7):87–89, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Uduwage:2012:ATO

[620] Anuradha Uduwage and Joseph A. Konstan. The answer is out there: Online Q&A sites. Computer, 45 (7):90–92, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:AWE

[621] Anonymous. Avižienis wins Eckert– Mauchly Award. Computer, 45(7): 93–95, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCf

[622] Anonymous. Call and calendar. Computer, 45(7):96–98, July 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:R

[623] David Alan Grier and Erin Dian Dumbacher. Repurposers. Computer, 45 (7):116, July 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSh

[624] Anonymous. Computer highlights Society magazines. Computer, 45(8): 4–5, August 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:Lc

[625] Anonymous. Letters. Computer, 45 (8):6–7, August 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:TWD

[626] Charles Severance. Teaching the world: Daphne Koller and Coursera. Computer, 45(8):8–9, August 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:DBC

[627] Brian M. Gaff, Kenneth Choy, and Jamie Chan. Doing business in China. Computer, 45(8):11–13, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAh

[628] Anonymous. 32 & 16 years ago. Computer, 45(8):14–15, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2012:CIA

[629] Lee Garber. Converged infrastructure: Addressing the efficiency challenge. Computer, 45(8):17–20, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:NBe

[630] Anonymous. News briefs. Computer, 45(8):21–23, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keckler:2012:MMC

[631] Stephen W. Keckler and Steven K. Reinhardt. Massively multithreaded computing systems. Computer, 45(8): 24–25, August 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stratton:2012:ADO

[632] John A. Stratton, Christopher Rodrigues, I-Jui Sung, Li-Wen Chang, Nasser Anssari, Geng Liu, Wen mei W. Hwu, and Nady Obeid. Algorithm and data optimization techniques for scaling to massively threaded systems. Computer, 45(8):26–32, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tzeng:2012:GTP

[633] Stanley Tzeng, Brandon Lloyd, and John D. Owens. A GPU taskparallel model with dependency resolution. Computer, 45(8):34–41, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaster:2012:CGP

[634] Benedict R. Gaster and Lee Howes. Can GPGPU programming be liberated from the data-parallel bottleneck? Computer, 45(8):42–52, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tumeo:2012:DNG

[635] Antonino Tumeo, Simone Secchi, and Oreste Villa. Designing nextgeneration massively multithreaded architectures for irregular applications. Computer, 45(8):53–61, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davies:2012:CCB

[636] Stephen Davies, Stacey Aylor Seal, and Jesse Hatfield. Cinefile: a categorybased analytic browser. Computer, 45 (8):62–69, August 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Haricombe:2012:OAE

[637] Lorraine J. Haricombe, L. Ada Emmett, and Perry Alexander. Open access: An evolving alternative. Computer, 45(8):70–72, August 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ng-Thow-Hing:2012:PR

[638] Victor Ng-Thow-Hing and Sandra Okita. Playdates with robots. Computer, 45(8):73–75, August 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zambonelli:2012:TSU

[639] Franco Zambonelli. Toward sociotechnical urban superorganisms. Computer, 45(8):76–78, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mens:2012:CSS

[640] Tom Mens. On the complexity of software systems. Computer, 45(8): 79–81, August 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boothby:2012:WCU

[641] William H. Boothby, Wolff Heintschel von Heinegg, James Bret Michael, Michael N. Schmitt, and Thomas C. Wingfield. When is a cyberattack a use of force or an armed attack? Computer, 45(8):82–84, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:ICS

[642] Anonymous. IEEE Computer Society election. Computer, 45(8):85–93, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:IPE

[643] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 45(8):94–99, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCg

[644] Anonymous. Call and calendar. Computer, 45(8):100–103, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:WTI

[645] David Alan Grier and Erin Dian Dumbacher. The winner takes it all. Computer, 45(8):120, August 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSi

[646] Anonymous. Computer highlights Society magazines. Computer, 45(9): 4–5, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:BMS

[647] Charles Severance. Bertrand Meyer: Software engineering and the Eiffel programming language. Computer, 45(9): 6–8, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:PRN

[648] Brian M. Gaff, Brian P. Murphy, Ralph A. Loren, Christine A. Dudzik, and Lisa Swiszcz. Patent reform: New tools for challenging patent validity as of September 2012. Computer, 45(9): 9–11, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAi

[649] Anonymous. 32 & 16 years ago. Computer, 45(9):12–13, September 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2012:MPR

[650] Neal Leavitt. Are mobile payments ready to cash in yet? Computer, 45(9): 15–18, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:NBf

[651] Anonymous. News briefs. Computer, 45(9):19–20, September 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chowdhury:2012:MSS

[652] Kaushik R. Chowdhury, Dave Cavalcanti, Mostafa El-Said, Tommaso Mazza, and Chittabrata Ghosh. Modeling and simulation of smart and green computing systems. Computer, 45(9): 22–23, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

De:2012:TUO

[653] Swades De and Riya Singhal. Toward uninterrupted operation of wireless sensor networks. Computer, 45(9): 24–30, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Venkata:2012:BGM

[654] Prabhakar T. Venkata, S. N. Akshay Uttama Nambi, R. Venkatesha Prasad, and Ignas Niemegeers. Bond graph modeling for energy-harvesting wireless sensor networks. Computer, 45 (9):31–38, September 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kocaoglu:2012:FGC

[655] Murat Kocaoglu, Derya Malak, and Ozgur B. Akan. Fundamentals of green communications and computing: Modeling and simulation. Computer, 45(9): 40–46, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yeo:2012:SHW

[656] Sungkap Yeo and Hsien-Hsin S. Lee. SimWare: a holistic warehouse-scale computer simulator. Computer, 45(9): 48–55, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Aksanli:2012:UDS

[657] Baris Aksanli, Jagannathan Venkatesh, and Tajana Simunic Rosing. Using datacenter simulation to evaluate green energy integration. Computer, 45(9): 56–64, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chandramohan:2012:DMM

[658] Mahinthan Chandramohan and Hee Beng Kuan Tan. Detection of mobile malware in the wild. Computer, 45(9): 65–71, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ohrt:2012:CPD

[659] Julian Ohrt and Volker Turau. Crossplatform development tools for Smartphone applications. Computer, 45(9): 72–79, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yan:2012:MFD

[660] Xiaoqi Yan, Chi-Wing Fu, and Andrew J. Hanson. Multitouching the fourth dimension. Computer, 45(9): 80–88, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nordman:2012:TGD

[661] Bruce Nordman, Ken Christensen, and Alan Meier. Think globally, distribute power locally: The promise of nanogrids. Computer, 45(9):89– 91, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2012:DS

[662] Hal Berghel. The dumbing down of SportPhishing. Computer, 45(9):92– 94, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ricanek:2012:FAB

[663] Karl Ricanek, Jr. and Chris Boehnen. Facial analytics: From big data to law enforcement. Computer, 45(9):95– 97, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lampe:2012:UFS

[664] Cliff Lampe and Nicole Ellison. Understanding Facebook: Social computing isn’t ‘just’ social. Computer, 45(9): 98–100, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2012:IMA

[665] Dejan S. Milojicic, Martin Arlitt, Doree Duncan Seligmann, George K. Thiruvathukal, and Christian Timmerer. Innovation mashups: Academic rigor meets social networking buzz. Computer, 45(9):101–105, September 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCh

[666] Anonymous. Call and calendar. Computer, 45(9):106–108, September 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:C

[667] David Alan Grier and Erin Dian Dumbacher. It’s complicated. Computer, 45 (9):124, September 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSj

[668] Anonymous. Computer highlights Society magazines. Computer, 45(10):

4–5, October 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:Ld

[669] Anonymous. Letters. Computer, 45 (10):6–7, October 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:ASF

[670] Charles Severance. The Apache Software Foundation: Brian Behlendorf. Computer, 45(10):8–9, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:SBW

[671] Brian M. Gaff and Karl P. Fryzel. Starting a business: Which entity structure is best for you? Computer, 45(10):10–12, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAj

[672] Anonymous. 32 & 16 years ago. Computer, 45(10):14–15, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2012:UMA

[673] Lee Garber. Using in-memory analytics to quickly crunch big data. Computer, 45(10):16–18, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:NBg

[674] Anonymous. News briefs. Computer, 45(10):19–21, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hinchey:2012:BDS

[675] Mike Hinchey, Sooyong Park, and Klaus Schmid. Building dynamic software product lines. Computer, 45(10): 22–26, October 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bosch:2012:DVS

[676] Jan Bosch and Rafael Capilla. Dynamic variability in software-intensive embedded system families. Computer, 45(10):28–35, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bencomo:2012:VDS

[677] Nelly Bencomo, Svein Hallsteinsen, and Eduardo Santana de Almeida. A view of the dynamic software product line landscape. Computer, 45(10): 36–41, October 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baresi:2012:SOD

[678] Luciano Baresi, Sam Guinea, and Liliana Pasquale. Service-oriented dynamic software product lines. Computer, 45(10):42–48, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2012:ESB

[679] Jaejoon Lee, Gerald Kotonya, and Daniel Robinson. Engineering servicebased dynamic software product lines. Computer, 45(10):49–55, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sawyer:2012:UCP

[680] Pete Sawyer, Raúl Mazo, Daniel Diaz, Camille Salinesi, and Danny Hughes. Using constraint programming to manage configurations in self-adaptive systems. Computer, 45(10):56–63, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hedley:2012:AWL

[681] Mark Hedley and Jian Zhang. Accurate wireless localization in sports. Computer, 45(10):64–70, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

AbdelBaky:2012:EHP

[682] Moustafa AbdelBaky, Manish Parashar, Hyunjoo Kim, Kirk E. Jordan, Vipin Sachdeva, James Sexton, Hani Jamjoom, Zon-Yin Shae, Gergina Pencheva, Reza Tavakoli, and Mary F. Wheeler. Enabling high-performance computing as a service. Computer, 45 (10):72–80, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sobel:2012:SEG

[683] Ann E. K. Sobel. Should everyone go to college? Computer, 45(10): 82–83, October 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hollan:2012:AEC

[684] James D. Hollan. Activity-enriched computing: Capturing and mining activity histories. Computer, 45(10): 84–87, October 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boland:2012:JCC

[685] Tim Boland and Paul E. Black. Juliet 1.1 C/C++ and Java Test Suite. Computer, 45(10):88–90, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gruenbaum:2012:UYC

[686] Peter Gruenbaum. Underrepresented youth creating culturally relevant games. Computer, 45(10):91–93, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Madhavan:2012:PM

[687] Krishna P. C. Madhavan, Mihaela Vorvoreanu, Niklas Elmqvist, Aditya Johri, Naren Ramakrishnan, G. Alan Wang, and Ann McKenna. Portfolio mining. Computer, 45(10):95– 99, October 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reinicke:2012:WDB

[688] Bryan Reinicke. It’s 2012, so why am I driving a 1972 bug? Computer, 45(10): 100–103, October 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:NEN

[689] Anonymous. New EICs named for IEEE Computer Society publications. Computer, 45(10):104–106, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCi

[690] Anonymous. Call and calendar. Computer, 45(10):107–109, October 2012.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:CM

[691] David Alan Grier and Erin Dian Dumbacher. The computer in the mirror. Computer, 45(10):128, October 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSk

[692] Anonymous. Computer highlights Society magazines. Computer, 45(11): 4–5, November 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:IPR

[693] Charles Severance. Inventing PHP: Rasmus Lerdorf. Computer, 45(11): 6–7, November 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2012:BOA

[694] Brian M. Gaff, Ralph A. Loren, and Amy M. DeCloux. A brief overview of the America Invents Act. Computer, 45(11):8–10, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAk

[695] Anonymous. 32 & 16 years ago. Computer, 45(11):12–13, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2012:BIM

[696] Neal Leavitt. Big iron moves toward exascale computing. Computer, 45(11):

14–17, November 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:NBh

[697] Anonymous. News briefs. Computer, 45(11):18–20, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sobel:2012:MTE

[698] Ann E. K. Sobel. The move toward electronic health records. Computer, 45 (11):22–23, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fridsma:2012:EHR

[699] Doug Fridsma. Electronic health records: The HHS perspective. Computer, 45(11):24–26, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

CarrionSenor:2012:PHR

[700] Inmaculada Carrión Señor, José Luis Fernández Alemán, and Ambrosio Toval. Personal health records: New means to safely handle health data? Computer, 45(11):27–33, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhlisch:2012:ECR

[701] Raik Kuhlisch, Ben Kraufmann, and Hannes Restel. Electronic case records in a box: Integrating patient data in healthcare networks. Computer, 45 (11):34–40, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Basu:2012:FMH

[702] Sujoy Basu, Alan H. Karp, Jun Li, James Pruyne, Jerry Rolia, Sharad Singhal, Jaap Suermondt, and Ram Swaminathan. Fusion: Managing healthcare records at cloud scale. Computer, 45(11):42–49, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brady:2012:TNH

[703] Kevin Brady, Ram D. Sriram, Bettijoyce Lide, and Kathleen Roberts. Testing the nation’s healthcare information infrastructure: NIST perspective. Computer, 45(11):50–57, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fagin:2012:DSG

[704] Barry Fagin and Dale Skrien. Debugging on the shoulders of giants: von Neumann’s programs 65 years later. Computer, 45(11):59–68, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dedrick:2012:CII

[705] Jason Dedrick, Jian Tang, and Kenneth L. Kraemer. China’s indigenous innovation policy: Impact on multinational R&D. Computer, 45(11):70– 78, November 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Curry:2012:SIC

[706] Edward Curry, Bill Guyon, Charles Sheridan, and Brian Donnellan. Sustainable IT: Challenges, postures, and outcomes. Computer, 45(11):79–81, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2012:SS

[707] Hal Berghel. Stratfor or stratagainst. Computer, 45(11):82–85, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jain:2012:BAS

[708] Anil K. Jain and Karthik Nandakumar. Biometric authentication: System security and user privacy. Computer, 45(11):87–92, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bi:2012:CSW

[709] Frank Bi and Joseph A. Konstan. Customer service 2.0: Where social computing meets customer relations. Computer, 45(11):93–95, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CSCb

[710] Anonymous. Computer Society connection. Computer, 45(11):96–98, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCj

[711] Anonymous. Call and calendar. Computer, 45(11):99–101, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:LS

[712] David Alan Grier and Erin Dian Dumbacher. Love the show! Computer, 45 (11):120, November 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CHSl

[713] Anonymous. Computer highlights Society magazines. Computer, 45(12): 4–5, December 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walz:2012:ISR

[714] John W. Walz. Implementing SP7: a review of progress in 2012. Computer, 45(12):6–9, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2012:VCB

[715] Charles Severance. Vint Cerf: a brief history of packets. Computer, 45(12): 10–12, December 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:YAl

[716] Anonymous. 32 & 16 years ago. Computer, 45(12):16–17, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:NBi

[717] Anonymous. News briefs. Computer, 45(12):22–24, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kessler:2012:ASD

[718] Gary C. Kessler. Advancing the science of digital forensics. Computer, 45(12): 25–27, December 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Young:2012:DSH

[719] Joel Young, Kristina Foster, Simson Garfinkel, and Kevin Fairbanks. Distinct sector hashes for target file detection. Computer, 45(12):28–35, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hunt:2012:NFA

[720] Ray Hunt and Sherali Zeadally. Network forensics: An analysis of techniques, tools, and trends. Computer, 45 (12):36–43, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmed:2012:SSC

[721] Irfan Ahmed, Sebastian Obermeier, Martin Naedele, and Golden G. Richard III. SCADA systems: Challenges for forensic investigators. Computer, 45(12):44–51, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:2012:SSC

[722] Yong Wang, Kevin Streff, and Sonell Raman. Smartphone security challenges. Computer, 45(12):52–58, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spear:2012:UTS

[723] Aaron Spear, Markus Levy, and Mathieu Desnoyers. Using tracing to solve the multicore system debug problem. Computer, 45(12):60–64, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moreno-Vozmediano:2012:ICA

[724] Rafael Moreno-Vozmediano, Ruben S. Montero, and Ignacio M. Llorente. IaaS cloud architecture: From virtualized datacenters to federated cloud infrastructures. Computer, 45(12):65– 72, December 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sundaravaradan:2012:FES

[725] Naren Sundaravaradan, Naren Ramakrishnan, and David A. Hanauer. Factorizing event sequences. Computer, 45(12):73–75, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lau:2012:DF

[726] Manfred Lau, Jun Mitani, and Takeo Igarashi. Digital fabrication. Computer, 45(12):76–79, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:2012:ALS

[727] Arnold Brown, Benjamin Apple, James Bret Michael, and Michael Schumann. Atomic-level security for Web applications in a cloud environment. Computer, 45(12):80–83, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vassev:2012:ASI

[728] Emil Vassev and Mike Hinchey. Awareness in software-intensive systems. Computer, 45(12):84–87, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hallisey:2012:BVW

[729] Brad Hallisey. Building a virtual world: The pipeline and process. Computer, 45 (12):90–92, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:MVC

[730] Anonymous. Milojicic voted 2013 Computer Society President-Elect. Computer, 45(12):93–95, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CAW

[731] Anonymous. CS announces 2012 winners of Kennedy and Rau Awards. Computer, 45(12):96–98, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:CCk

[732] Anonymous. Call and calendar. Computer, 45(12):99–101, December 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2012:WP

[733] David Alan Grier and Erin Dian Dumbacher. Work/play. Computer, 45(12): 116, December 2012. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSa

[734] Anonymous. Computer highlights Society magazines. Computer, 46(1): 4–5, January 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:US

[735] David Alan Grier. The unseen and the seen. Computer, 46(1):6–7, January 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:2013:EDP

[736] Ron Vetter. Embracing digital publishing. Computer, 46(1):8–10, January 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:VJC

[737] Charles Severance. Van Jacobson: Content-centric networking. Computer, 46(1):11–13, January 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:PBI

[738] Brian M. Gaff, Ralph A. Loren, and Gareth Dickson. Protecting bioinformatics as intellectual property. Computer, 46(1):15–17, January 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:YAa

[739] Anonymous. 32 & 16 years ago. Computer, 46(1):18–19, January 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2013:BBA

[740] Neal Leavitt. Bringing big analytics to the masses. Computer, 46(1): 20–23, January 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:NBa

[741] Lee Garber. News briefs. Computer, 46 (1):24–26, January 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:2013:MHR

[742] Santosh Kumar, Wendy Nilsen, Misha Pavel, and Mani Srivastava. Mobile health: Revolutionizing healthcare through transdisciplinary research. Computer, 46(1):28–35, January 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schirner:2013:FHL

[743] Gunar Schirner, Deniz Erdogmus, Kaushik Chowdhury, and Taskin Padir. The future of human-in-the-loop cyberphysical systems. Computer, 46(1): 36–45, January 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rikakis:2013:EMD

[744] Thanassis Rikakis, Aisling Kelliher, and Nicole Lehrer. Experiential media and digital culture. Computer, 46 (1):46–54, January 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sharkey:2013:RSC

[745] Noel Sharkey and Amanda Sharkey. Robotic surgery: On the cutting edge of ethics. Computer, 46(1):56–64, January 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hole:2013:MHR

[746] Kjell Jørgen Hole. Management of hidden risks. Computer, 46(1):65– 70, January 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Varshney:2013:SAM

[747] Upkar Varshney. A smart approach to medication management. Computer, 46 (1):71–76, January 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Murugesan:2013:CCN

[748] San Murugesan. Cloud computing: The new normal? Computer, 46(1): 77–79, January 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2013:WSF

[749] Brian David Johnson. When science fiction and science fact meet. Computer, 46(1):80–82, January 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2013:EOPa

[750] Kirk W. Cameron. Energy oddities, part 1: Why the energy world is odd. Computer, 46(1):83–84, January 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2013:RDV

[751] Hal Berghel. RFIDiocy: It’s déjà vu all over again. Computer, 46(1): 85–88, January 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riedl:2013:CMR

[752] John Riedl and Eric Riedl. Crowdsourcing medical research. Computer, 46(1):89–92, January 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:HN

[753] David Alan Grier. The here and now. Computer, 46(1):108, January 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSb

[754] Anonymous. Computer highlights Society magazines. Computer, 46(2): 4–5, February 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:L

[755] Lee Garber. Letters. Computer, 46(2): 6, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:MBM

[756] Charles Severance. Mitchell Baker: The Mozilla Foundation. Computer, 46 (2):7–9, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:PBA

[757] Brian M. Gaff, Stephen G. Huggard, and Gregory W. Carey. Paying bribes abroad will get you in hot water in the US. Computer, 46(2):11–13, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:YAb

[758] Anonymous. 32 & 16 years ago. Computer, 46(2):14–15, 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:GGM

[759] Lee Garber. GPUs go mobile. Computer, 46(2):16–19, 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:NBb

[760] Lee Garber. News briefs. Computer, 46(2):20–23, 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feki:2013:ITN

[761] Mohamed Ali Feki, Fahim Kawsar, Mathieu Boussard, and Lieven Trappeniers. The Internet of Things: The next technological revolution. Computer, 46(2):24–25, 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hodges:2013:PCD

[762] Steve Hodges, Stuart Taylor, Nicolas Villar, James Scott, Dominik Bial, and Patrick Tobias Fischer. Prototyping connected devices for the Internet of Things. Computer, 46(2):26–34, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roggen:2013:OHA

[763] Daniel Roggen, Gerhard Tröster, Paul Lukowicz, Alois Ferscha, José del R. Millán, and Ricardo Chavarriaga. Opportunistic human activity and context recognition. Computer, 46(2):36– 45, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2013:AHL

[764] Sang Wan Lee, Oliver Prenzel, and Zeungnam Bien. Applying human learning principles to user-centered IoT systems. Computer, 46(2):46–52, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kortuem:2013:EIT

[765] Gerd Kortuem, Arosha K. Bandara, Neil Smith, Mike Richards, and Marian Petre. Educating the Internet-of-Things generation. Computer, 46(2): 53–61, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balasubramaniam:2013:RIN

[766] Sasitharan Balasubramaniam and Jussi Kangasharju. Realizing the Internet of Nano Things: Challenges, solutions, and applications. Computer, 46(2):62– 68, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2013:CED

[767] Jane W. S. Liu, Chi-Sheng Shih, and Edward T.-H. Chu. Cyberphysical elements of disaster-prepared smart environments. Computer, 46(2):69–75, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singhal:2013:CMC

[768] Mukesh Singhal, Santosh Chandrasekhar, Tingjian Ge, Ravi Sandhu, Ram Krishnan, Gail-Joon Ahn, and Elisa Bertino. Collaboration in multicloud computing environments: Framework and security issues. Computer, 46 (2):76–84, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hall:2013:BSE

[769] Jon G. Hall and Lucia Rapanotti. Beauty in software engineering. Computer, 46(2):85–87, 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lutz:2013:USE

[770] Michael Lutz. Undergraduate software engineering at RIT: Past, present, and future. Computer, 46(2):88–90, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dewri:2013:CPG

[771] Rinku Dewri and Ramakrishna Thurimella. Can a phone’s GPS ‘lie’ intelligently? Computer, 46(2):91–93, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pace:2013:WWPa

[772] Jason Pace. The ways we play, part 1: From PCs to consoles to mobile phones. Computer, 46(2):94–96, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:EAS

[773] Anonymous. EIC applicants sought for CS periodicals, 2014–2015 terms. Computer, 46(2):106–108, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCa

[774] Anonymous. Call and calendar. Computer, 46(2):109–111, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:ED

[775] David Alan Grier. The ethical dimension. Computer, 46(2):112, 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSc

[776] Anonymous. Computer highlights Society magazines. Computer, 46(3): 4–5, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:PSB

[777] Charles Severance. Pooja Sankar: Building the Piazza Collaboration System. Computer, 46(3):6–7, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:DP

[778] Brian M. Gaff and Peter J. Cuomo. Design patents. Computer, 46(3): 8–10, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2013:YAa

[779] Neville Holmes. 32 & 16 years ago. Computer, 46(3):12–13, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2013:TCH

[780] Neal Leavitt. A technology that comes highly recommended. Computer, 46 (3):14–17, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:NBc

[781] Lee Garber. News briefs. Computer, 46 (3):18–20, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Prey:2013:FGD

[782] Jane Chu Prey and Alfred C. Weaver. Fostering gender diversity in computing. Computer, 46(3):22–23, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DuBow:2013:DCW

[783] Wendy M. DuBow. Diversity in computing: Why it matters and how organizations can achieve it. Computer, 46 (3):24–29, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Whitney:2013:PPA

[784] Telle Whitney, Denise Gammal, Barbara Gee, Jody Mahoney, and Caroline Simard. Priming the pipeline: Addressing gender-based barriers in computing. Computer, 46(3):30–36, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fernandez:2013:PBM

[785] Mary Fernandez. A path between: Mentoring the next generation of computing professionals. Computer, 46 (3):38–41, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stephenson:2013:MTG

[786] Chris Stephenson and Rebecca Dovi. More than gender: Taking a systemic approach to improving K-12 computer science education. Computer, 46(3): 42–46, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eney:2013:BPW

[787] Crystal Eney, Ed Lazowska, Helene Martin, and Stuart Reges. Broadening participation: The why and the how. Computer, 46(3):48–51, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cohoon:2013:EDC

[788] James P. Cohoon, J. McGrath Cohoon, and Mary Lou Soffa. Educating diverse computing students at the University of Virginia. Computer, 46 (3):52–55, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Klawe:2013:IFP

[789] Maria Klawe. Increasing female participation in computing: The Harvey Mudd College story. Computer, 46 (3):56–58, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:2013:SWP

[790] Deepak Kumar. Serving women with a purposeful vision. Computer, 46 (3):60–61, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schnabel:2013:PRG

[791] Robert B. Schnabel. Personal reflections on gender diversity in computing. Computer, 46(3):62–63, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Starbird:2013:RMI

[792] Kate Starbird. Returning to my inner nerd: Following the ‘social’ disruption of computing. Computer, 46 (3):63–65, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anton:2013:FCH

[793] Annie I. Anton. Facing challenges and having fun. Computer, 46(3): 65–66, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bonhomme-Biais:2013:PPC

[794] Alice Bonhomme-Biais and Raquel Romano. Pursuing professional changes. Computer, 46(3):66–68, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shar:2013:DSI

[795] Lwin Khin Shar and Hee Beng Kuan Tan. Defeating SQL injection. Computer, 46(3):69–77, March 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hang:2013:SGT

[796] Chung-Wei Hang, Zhe Zhang, and Munindar P. Singh. Shin: Generalized trust propagation with limited evidence. Computer, 46(3):78–85, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2013:CCE

[797] Nir Kshetri and San Murugesan. Cloud computing and EU data privacy regulations. Computer, 46(3):86–89, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2013:EOPb

[798] Kirk W. Cameron. Energy oddities, part 2: Why green computing is odd. Computer, 46(3):90–93, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ricanek:2013:BHI

[799] Karl Ricanek, Jr. and Gayathri Mahalingam. Biometrically, how identical are identical twins? Computer, 46 (3):94–96, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2013:SHG

[800] Hal Berghel. The SCDOR hack: Great security theater in five stages. Computer, 46(3):97–99, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riedl:2013:CTI

[801] John Riedl. Crowdfunding technology innovation. Computer, 46(3): 100–103, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2013:JRC

[802] Brian David Johnson. Jimmy: The robot that changed it all. Computer, 46 (3):104–107, March 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CS

[803] Anonymous. Computer SocietyConnection. Computer, 46(3):120–121, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCb

[804] Anonymous. Call and calendar. Computer, 46(3):122–123, March 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:AC

[805] David Alan Grier. Attention to context. Computer, 46(3):124, March 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSd

[806] Anonymous. Computer highlights Society magazines. Computer, 46(4):4–5, April 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:LSB

[807] Charles Severance. Larry Smarr: Building Mosaic. Computer, 46(4):6–8, April 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:PM

[808] Brian M. Gaff and Ralph A. Loren. Patent monetization. Computer, 46 (4):9–11, April 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2013:YAb

[809] Neville Holmes. 32 & 16 years ago. Computer, 46(4):12–13, April 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:AGL

[810] Lee Garber. Analytics goes on location with new approaches. Computer, 46(4):14–17, April 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:NBd

[811] Lee Garber. News briefs. Computer, 46(4):18–20, April 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zeadally:2013:SCC

[812] Sherali Zeadally, Gregorio Martinez, and Han-Chieh Chao. Securing cyberspace in the 21st Century. Computer, 46(4):22–23, April 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Basaras:2013:DIS

[813] Pavlos Basaras, Dimitrios Katsaros, and Leandros Tassiulas. Detecting influential spreaders in complex, dynamic networks. Computer, 46(4): 24–29, April 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alcaraz:2013:WAS

[814] Cristina Alcaraz and Javier Lopez. Wide-area situational awareness for critical infrastructure protection. Computer, 46(4):30–37, April 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paredes-Oliva:2013:SDU

[815] Ignasi Paredes-Oliva, Pere Barlet-Ros, and Josep Sole-Pareta. Scan detection under sampling: A new perspective. Computer, 46(4):38–44, April 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ning:2013:CSI

[816] Huansheng Ning, Hong Liu, and Laurence T. Yang. Cyberentity security in the Internet of Things. Computer, 46(4):46–53, April 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rashid:2013:WAD

[817] Awais Rashid, Alistair Baron, Paul Rayson, Corinne May-Chahal, Phil Greenwood, and James Walkerdine. Who am I? Analyzing digital personas in cybercrime investigations. Computer, 46(4):54–61, April 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alhomoud:2013:NGA

[818] Adeeb Alhomoud, Irfan Awan, Jules Ferdinand Pagna Disso, and Muhammad Younas. A next-generation approach to combating botnets. Computer, 46(4):62–66, April 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rahwan:2013:GMP

[819] Iyad Rahwan, Sohan Dsouza, Alex Rutherford, Victor Naroditskiy, James McInerney, Matteo Venanzi, Nicholas R. Jennings, and Manuel Cebrian. Global manhunt pushes the limits of social mobilization. Computer, 46(4):68–75, April 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ma:2013:CPP

[820] Xiaohan Ma, Zhigang Deng, Mian Dong, and Lin Zhong. Characterizing the performance and power consumption of 3D mobile games. Computer, 46(4):76–82, April 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sanders:2013:WLF

[821] Ruth H. Sanders. Why learn a foreign language? Computer, 46(4): 83–85, April 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Elmqvist:2013:UAI

[822] Niklas Elmqvist and Pourang Irani. Ubiquitous analytics: Interacting with big data anywhere, anytime. Computer, 46(4):86–89, April 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hodges:2013:BIP

[823] Steve Hodges. Batteries not included: Powering the ubiquitous computing dream. Computer, 46(4):90–93, April 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferraiolo:2013:EEW

[824] David Ferraiolo, Serban Gavrila, and Wayne Jansen. Enabling an enterprisewide, data-centric operating environment. Computer, 46(4):94–96, April 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pace:2013:WWPb

[825] Jason Pace. The ways we play, part 2: Mobile game changers. Computer, 46(4):97–99, April 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moreno:2013:HPB

[826] Ana M. Moreno, Ahmed Seffah, Rafael Capilla, and Maria-Isabel Sanchez-Segura. HCI practices for building usable software. Computer, 46(4): 100–102, April 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CSCa

[827] Anonymous. Computer Society connection. Computer, 46(4):110–113, April 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCc

[828] Anonymous. Call and calendar. Computer, 46(4):114–115, April 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:EBB

[829] David Alan Grier. Equal before the Bar. Computer, 46(4):116, April 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSe

[830] Anonymous. Computer highlights Society magazines. Computer, 46(5):4–5, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:BME

[831] Charles Severance. Bob Metcalfe: Ethernet at forty. Computer, 46 (5):6–9, May 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:WIW

[832] Brian M. Gaff. Who invents what. Computer, 46(5):10–12, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2013:YAc

[833] Neville Holmes. 32 & 16 years ago. Computer, 46(5):13–14, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2013:HCM

[834] Neal Leavitt. Hybrid clouds move to the forefront. Computer, 46(5): 15–18, May 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:NBe

[835] Lee Garber. News briefs. Computer, 46(5):19–21, May 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rhyne:2013:CER

[836] Theresa-Marie Rhyne and Min Chen. Cutting-edge research in visualization. Computer, 46(5):22–24, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Majumder:2013:LAD

[837] Aditi Majumder and Behzad Sajadi. Large area displays: The changing face of visualization. Computer, 46 (5):26–33, May 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Childs:2013:RCV

[838] Hank Childs, Berk Geveci, Will Schroeder, Jeremy Meredith, Kenneth Moreland, Christopher Sewell, Torsten Kuhlen, and E. Wes Bethel. Research challenges for visualization software. Computer, 46(5):34–42, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kosara:2013:SNS

[839] Robert Kosara and Jock Mackinlay. Storytelling: The next step for visualization. Computer, 46(5):44–50, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keefe:2013:RSV

[840] Daniel F. Keefe and Tobias Isenberg. Reimagining the scientific visualization interaction paradigm. Computer, 46 (5):51–57, May 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shneiderman:2013:IHI

[841] Ben Shneiderman, Catherine Plaisant, and Bradford W. Hesse. Improving healthcare with interactive visualization. Computer, 46(5):58–66, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schreck:2013:VAS

[842] Tobias Schreck and Daniel Keim. Visual analysis of social media data. Computer, 46(5):68–75, May 2013.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2013:ICE

[843] Qing Li, Chris Larsen, and Tim van der Horst. IPv6: A catalyst and evasion tool for botnets and malware delivery networks. Computer, 46(5):76–82, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Szczepanski:2013:DAV

[844] A. F. Szczepanski, Jian Huang, T. Baer, Y. C. Mack, and S. Ahern. Data analysis and visualization in highperformance computing. Computer, 46 (5):84–92, May 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2013:CTS

[845] Hal Berghel. Computing technology and survivable journalism. Computer, 46(5):93–97, May 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Talia:2013:CSB

[846] Domenico Talia. Clouds for scalable big data analytics. Computer, 46 (5):98–101, May 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Poursaberi:2013:SAT

[847] Ahmad Poursaberi, Svetlana Yanushkevich, Marina L. Gavrilova, Vlad P. Shmerko, and Patrick S. P. Wang. Situational awareness through biometrics. Computer, 46(5):102–104, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2013:SSF

[848] Brian David Johnson. Secret science fiction. Computer, 46(5):105–107, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Turner:2013:PMC

[849] Joseph Turner. Power management as choice architecture. Computer, 46 (5):108–111, May 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CSCb

[850] Anonymous. Computer Society connection. Computer, 46(5):118–120, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCd

[851] Anonymous. Call and calendar. Computer, 46(5):121–123, May 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:WC

[852] David Alan Grier. Working in conference. Computer, 46(5):124, May 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSf

[853] Anonymous. Computer highlights Society magazines. Computer, 46(6):4–5, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:GBB

[854] Charles Severance. Gordon Bell: Building blocks of computing. Computer, 46(6):6–7, June 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:La

[855] Anonymous. Letters. Computer, 46(6):8, June 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:YEY

[856] Brian M. Gaff. Are you exhausted by your first sale? Computer, 46 (6):9–11, June 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2013:YAd

[857] Neville Holmes. 32 & 16 years ago. Computer, 46(6):12–13, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:GAE

[858] Lee Garber. Game accessibility: Enabling everyone to play. Computer, 46 (6):14–18, June 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:NBf

[859] Lee Garber. News briefs. Computer, 46(6):19–21, June 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2013:BDN

[860] Katina Michael and Keith W. Miller. Big data: New opportunities and new challenges [Guest Editors’ introduction]. Computer, 46(6):22–24, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hemerly:2013:PPC

[861] Jess Hemerly. Public policy considerations for data-driven innovation. Computer, 46(6):25–31, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tallon:2013:CGB

[862] Paul P. Tallon. Corporate governance of big data: Perspectives on value, risk, and cost. Computer, 46(6):32–38, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pitt:2013:TBD

[863] Jeremy Pitt, Aikaterini Bourazeri, Andrzej Nowak, Magda Roszczynska-Kurasinska, Agnieszka Rychwalska, Inmaculada Rodriguez Santiago, Maite Lopez Sanchez, Monica Florea, and Mihai Sanduleac. Transforming big data into collective awareness. Computer, 46 (6):40–45, June 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wigan:2013:BDB

[864] Marcus R. Wigan and Roger Clarke. Big data’s big unintended consequences. Computer, 46(6):46–53, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McGregor:2013:BDN

[865] Carolyn McGregor. Big data in neonatal intensive care. Computer, 46 (6):54–59, June 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anthonysamy:2013:SNP

[866] Pauline Anthonysamy, Phil Greenwood, and Awais Rashid. Social networking privacy: Understanding the disconnect from policy to controls. Computer, 46(6):60–67, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ewing:2013:GIE

[867] E. Thomas Ewing, Samah Gad, and Naren Ramakrishnan. Gaining insights into epidemics by mining historical newspapers. Computer, 46 (6):68–72, June 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Henze:2013:ERT

[868] Niels Henze, Alireza Sahami Shrazi, Albrecht Schmidt, Martin Pielot, and Florian Michahelles. Empirical research through ubiquitous data collection. Computer, 46(6):74–76, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Repenning:2013:MPA

[869] Alexander Repenning. Making programming accessible and exciting. Computer, 46(6):78–81, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sobel:2013:ME

[870] Ann E. K. Sobel. The modern educator. Computer, 46(6):82–83, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2013:EUT

[871] James Bret Michael. Empowering users through secure on-demand data provisioning. Computer, 46(6):84–85, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holzmann:2013:TSP

[872] Gerard J. Holzmann. A three-step program for recovering hackers. Computer, 46(6):86–88, June 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CSCc

[873] Anonymous. Computer Society connection. Computer, 46(6):90–91, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCe

[874] Anonymous. Call and calendar. Computer, 46(6):92–96, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:CZ

[875] David Alan Grier. The comfort zone. Computer, 46(6):104, June 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSg

[876] Anonymous. Computer highlights Society magazines. Computer, 46(7):4–5, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:KHO

[877] Charles Severance. Katie Hafner: The origins of the Internet. Computer, 46(7):6–7, July 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2013:YAe

[878] Neville Holmes. 32 & 16 years ago. Computer, 46(7):8–9, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2013:SDN

[879] Sixto Ortiz Jr. Software-defined networking: On the verge of a breakthrough? Computer, 46(7):10–12, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:NBg

[880] Lee Garber. News briefs. Computer, 46(7):13–15, July 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:TUB

[881] Brian M. Gaff. The troll under the bridge. Computer, 46(7):17–19, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Streit:2013:VAS

[882] Marc Streit and Oliver Bimber. Visual analytics: Seeking the unknown. Computer, 46(7):20–21, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fekete:2013:VAI

[883] Jean-Daniel Fekete. Visual analytics infrastructures: From data management to exploration. Computer, 46 (7):22–29, July 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gorg:2013:VAS

[884] Carsten Gorg, Youn ah Kang, Zhicheng Liu, and John Stasko. Visual analytics support for intelligence analysis. Computer, 46(7):30–38, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ma:2013:LSG

[885] Kwan-Liu Ma and Chris W. Muelder. Large-scale graph visualization and analytics. Computer, 46(7):39–46, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keim:2013:RTV

[886] Daniel A. Keim, Milos Krstajic, Christian Rohrdantz, and Tobias Schreck. Real-time visual analytics for text streams. Computer, 46(7):47–55, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanWijk:2013:ECV

[887] Jarke J. van Wijk. Evaluation: A challenge for visual analytics. Computer, 46(7):56–60, July 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cook:2013:CSH

[888] Diane J. Cook, Aaron S. Crandall, Brian L. Thomas, and Narayanan C. Krishnan. CASAS: A smart home in a box. Computer, 46(7):62–69, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zidenberg:2013:ORA

[889] Tsahee Zidenberg, Isaac Keslassy, and Uri Weiser. Optimal resource allocation with MultiAmdahl. Computer, 46 (7):70–77, July 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2013:N

[890] Brian David Johnson. From NASA. Computer, 46(7):78–81, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Witman:2013:SMS

[891] Paul Witman. Social media for social value. Computer, 46(7):82–85, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2013:TPD

[892] Hal Berghel. Through the PRISM darkly. Computer, 46(7):86–90, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stansberry:2013:FGI

[893] Matt Stansberry. The future of green IT: Solving the accountability issue. Computer, 46(7):91–93, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khan:2013:TCS

[894] Khaled M. Khan and Qutaibah Malluhi. Trust in cloud services: Providing more controls to clients. Computer, 46(7):94–96, July 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CSCd

[895] Anonymous. Computer Society connection. Computer, 46(7):97–98, July 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCf

[896] Anonymous. Call and calendar. Computer, 46(7):99–102, July 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:STL

[897] David Alan Grier. Short-term loan. Computer, 46(7):112, July 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSh

[898] Anonymous. Computer highlights Society magazines. Computer, 46(8): 4–5, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:LKT

[899] Charles Severance. Len Kleinrock: The theory of packets [computing conversations]. Computer, 46(8):6–8, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:TTK

[900] Brian M. Gaff and Joanna M. Dombrowski. Ten things to know about applying for non-US patents. Computer, 46(8):9–11, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2013:YAf

[901] Neville Holmes. 32 & 16 years ago. Computer, 46(8):12–13, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:LAR

[902] Lee Garber. The lowly API is ready to step front and center. Computer, 46 (8):14–17, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:NBh

[903] Lee Garber. News briefs. Computer, 46 (8):18–20, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atwood:2013:NGM

[904] Greg Atwood, Soo-Ik Chae, and Simon S. Y. Shim. Next-generation memory [Guest Editors’ introduction]. Computer, 46(8):21–22, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2013:NFM

[905] Yan Li and Khandker N. Quader. NAND flash memory: Challenges and opportunities. Computer, 46(8):23–29, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Song:2013:WLA

[906] Yoon-Jong Song, Gitae Jeong, In-Gyu Baek, and Jungdal Choi. What lies ahead for resistance-based memory technologies? Computer, 46(8): 30–36, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huffman:2013:NMT

[907] Amber Huffman and Dale Juenemann. The nonvolatile memory transformation of client storage. Computer, 46 (8):38–44, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Badam:2013:HPM

[908] Anirudh Badam. How persistent memory will change software systems. Computer, 46(8):45–51, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Swanson:2013:RRR

[909] Steven Swanson and Adrian M. Caulfield. Refactor, reduce, recycle: Restructuring the I/O stack for the future of storage. Computer, 46(8): 52–59, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2013:BCA

[910] Kaiping Liu, Hee Beng Kuan Tan, and Xu Chen. Binary code analysis. Computer, 46(8):60–68, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vandenHoven:2013:MTI

[911] Elise van den Hoven, Evelien van de Garde-Perik, Serge Offermans, Koen van Boerdonk, and Kars-Michiel H. Lenssen. Moving tangible interaction systems to the next level. Computer, 46 (8):70–76, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brechner:2013:GDX

[912] Eric Brechner. Getting diagonal on the Xbox One Trifecta. Computer, 46 (8):77–78, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

OGorman:2013:PKF

[913] Lawrence O’Gorman. Putting a kinder face on public cameras. Computer, 46 (8):79–81, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vassev:2013:ARE

[914] Emil Vassev and Mike Hinchey. Autonomy requirements engineering. Computer, 46(8):82–84, August 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hartmann:2013:DBI

[915] Bjorn Hartmann and Paul K. Wright. Designing bespoke interactive devices. Computer, 46(8):85–89, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fiaux:2013:BVA

[916] Patrick Fiaux, Maoyuan Sun, Lauren Bradel, Chris North, Naren Ramakrishnan, and Alex Endert. Bixplorer: Visual analytics with biclusters. Computer, 46(8):90–94, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilson:2013:WNC

[917] Cameron Wilson. What’s up next for Code.org? Computer, 46(8): 95–97, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:ICSa

[918] Anonymous. IEEE Computer Society election. Computer, 46(8):98– 106, August 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:IPE

[919] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 46(8):107–112, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCg

[920] Anonymous. Call and calendar. Computer, 46(8):113–117, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:TOT

[921] David Alan Grier. The tenor of our times. Computer, 46(8):128, August 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSi

[922] Anonymous. Computer highlights Society magazines. Computer, 46(9): 6–7, September 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:Lb

[923] Anonymous. Letters. Computer, 46(9): 10–12, September 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:JCH

[924] Charles Severance. John C. Hollar: History of computing. Computer, 46 (9):13–14, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:WYL

[925] Brian M. Gaff and Alyssa Harvey Dawson. Working with your legal department. Computer, 46(9):17–19, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2013:YAg

[926] Neville Holmes. 32 & 16 years ago. Computer, 46(9):20–21, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2013:SCW

[927] Neal Leavitt. Storage challenge: Where will all that big data go? Computer, 46(9):22–25, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:NBa

[928] Anonymous. News briefs. Computer, 46(9):26–28, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Guzdial:2013:CEB

[929] Mark Guzdial and Betsy DiSalvo. Computing education: Beyond the classroom. Computer, 46(9):30–31, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Knobelsdorf:2013:AFR

[930] Maria Knobelsdorf and Jan Vahrenhold. Addressing the full range of students: Challenges in K–12 computer science education. Computer, 46(9):

32–37, September 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peppler:2013:SPC

[931] Kylie Peppler. STEAM-powered computing education: Using e-textiles to integrate the arts and STEM. Computer, 46(9):38–43, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fincher:2013:PCP

[932] Sally Fincher and Daniel Knox. The porous classroom: Professional practices in the computing curriculum. Computer, 46(9):44–51, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brennan:2013:LCT

[933] Karen Brennan. Learning computing through creating and connecting. Computer, 46(9):52–59, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tew:2013:CVT

[934] Allison Elliott Tew and Brian Dorn. The case for validated tools in computer science education research. Computer, 46(9):60–66, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Williams:2013:UVC

[935] Dean N. Williams, Timo Bremer, Charles Doutriaux, John Patchett, Sean Williams, Galen Shipman, Ross Miller, David R. Pugmire, Brian Smith, Chad Steed, E. Wes Bethel, Hank Childs, Harinarayan Krishnan, Prabhat Prabhat, Michael Wehner, Claudio T. Silva, Emanuele Santos, David Koop, Tommy Ellqvist, Jorge Poco, Berk Geveci, Aashish Chaudhary, Andy Bauer, Alexander Pletzer, Dave Kindig, Gerald L. Potter, and Thomas P. Maxwell. Ultrascale visualization of climate data. Computer, 46(9):68–76, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Niu:2013:ESC

[936] Jianwei Niu, Jing Peng, Lei Shu, Chao Tong, and Wanjiun Liao. An empirical study of a Chinese online social network — Renren. Computer, 46(9): 78–84, September 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Costa:2013:PQE

[937] Rogerio Luis de Carvalho Costa and Pedro Furtado. Providing quality of experience for users: The next DBMS challenge. Computer, 46(9): 86–93, September 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ricanek:2013:NBC

[938] Karl Ricanek. The next biometric challenge: Medical alterations. Computer, 46(9):94–96, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cohen:2013:PNO

[939] Beth Cohen. PaaS: New opportunities for cloud application development. Computer, 46(9):97–100, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2013:CM

[940] Brian David Johnson. That Copernican moment. Computer, 46(9):101– 103, September 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2013:TC

[941] Hal Berghel. Toxic cookies. Computer, 46(9):104–107, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CSCe

[942] Anonymous. Computer Society connection. Computer, 46(9):108–109, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCh

[943] Anonymous. Call and calendar. Computer, 46(9):110–113, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:EHC

[944] David Alan Grier. The errant hashtag: For the common good. Computer, 46 (9):124, September 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSj

[945] Anonymous. Computer highlights Society magazines. Computer, 46(10): 6–7, October 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:Lc

[946] Anonymous. Letters. Computer, 46 (10):10–12, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:EUR

[947] Charles Severance. Eben Upton: Raspberry Pi. Computer, 46(10):14–16, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:YWS

[948] Brian M. Gaff and Greg Hendershott. So, you want to start a software company? Computer, 46(10): 17–19, October 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2013:YAh

[949] Neville Holmes. 32 & 16 years ago. Computer, 46(10):20–21, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:GTM

[950] Lee Garber. Gestural technology: Moving interfaces in a new direction [technology news]. Computer, 46(10): 22–25, October 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:NBb

[951] Anonymous. News briefs. Computer, 46(10):26–28, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Devadas:2013:TCM

[952] Srinivas Devadas. Toward a coherent multicore memory model. Computer, 46(10):30–31, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schuchhardt:2013:IDD

[953] Matthew Schuchhardt, Abhishek Das, Nikos Hardavellas, Gokhan Memik, and Alok Choudhary. The impact of dynamic directories on multicore interconnects. Computer, 46(10):32– 39, October 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shriraman:2013:ATA

[954] Arrvindh Shriraman, Hongzhou Zhao, and Sandhya Dwarkadas. An application-tailored approach to hardware cache coherence. Computer, 46 (10):40–47, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krishna:2013:SCM

[955] Tushar Krishna, Chia-Hsin Owen Chen, Sunghyun Park, Woo-Cheol Kwon, Suvinay Subramanian, Anantha P. Chandrakasan, and Li-Shiuan Peh. Single-cycle multihop asynchronous repeated traversal: A SMART future for reconfigurable onchip networks. Computer, 46(10): 48–55, October 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shi:2013:THS

[956] Qingchuan Shi and Omer Khan. Toward holistic soft-error-resilient sharedmemory multicores. Computer, 46(10):

56–64, October 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanGenuchten:2013:QSI

[957] Michiel van Genuchten and Les Hatton. Quantifying software’s impact. Computer, 46(10):66–72, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alcaraz:2013:CCS

[958] Cristina Alcaraz and Sherali Zeadally. Critical control system protection in the 21st Century. Computer, 46(10): 74–83, October 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2013:EUC

[959] Nir Kshetri and San Murugesan. EU and US cybersecurity strategies and their impact on businesses and consumers. Computer, 46(10):84–88, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Metri:2013:HSC

[960] Grace Metri, Manuj Sabharwal, Sundar Iyer, and Abhishek Agrawal. Hardware/software codesign to optimize SoC device battery life. Computer, 46(10):89–92, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:MIV

[961] Charles Severance. MOOCs: An insider’s view. Computer, 46(10):93– 96, October 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morgan:2013:KFI

[962] Lorraine Morgan and Kieran Conboy. Key factors impacting cloud computing adoption. Computer, 46(10): 97–99, October 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2013:KMT

[963] Keri Johnson, Jebediah Pavleas, and Jack Chang. Kinecting to mathematics through embodied interactions. Computer, 46(10):101–104, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kunze:2013:ARM

[964] Kai Kunze, Masakazu Iwamura, Koichi Kise, Seiichi Uchida, and Shinichiro Omachi. Activity recognition for the mind: Toward a cognitive “quantified self”. Computer, 46(10):105– 108, October 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CSCf

[965] Anonymous. Computer Society connection. Computer, 46(10):109–110, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCi

[966] Anonymous. Call and calendar. Computer, 46(10):114–119, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:PP

[967] David Alan Grier. Not part of the plan. Computer, 46(10):136, October 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:FCa

[968] Anonymous. [Front cover]. Computer, 46(11):c1, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:Ma

[969] Anonymous. [Masthead]. Computer, 46 (11):1, November 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:TCa

[970] Anonymous. Table of contents. Computer, 46(11):2–3, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:ICSb

[971] Anonymous. IEEE Computer Society membership [advertisement]. Computer, 46(11):4–5, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSk

[972] Anonymous. Computer highlights Society magazines. Computer, 46(11): 6–7, November 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:IHS

[973] Charles Severance. Ian Horrocks: Standardizing OWL. Computer, 46(11): 8–9, November 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:UUT

[974] Brian M. Gaff, David J. Levine, and Raymond Gabriel. Understanding US trade controls. Computer, 46(11):10– 13, November 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2013:YAi

[975] Neville Holmes. 32 & 16 years ago. Computer, 46(11):14–15, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2013:TMS

[976] Neal Leavitt. Today’s mobile security requires a new approach. Computer, 46(11):16–19, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:NBi

[977] Lee Garber. News briefs. Computer, 46 (11):20–22, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Margaria:2013:SIP

[978] Tiziana Margaria and Mike Hinchey. Simplicity in IT: The power of less. Computer, 46(11):23–25, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Floyd:2013:ISR

[979] Barry D. Floyd and Steve Bosselmann. ITSy — simplicity research in information and communication technology. Computer, 46(11):26–32, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bosch:2013:AST

[980] Jan Bosch. Achieving simplicity with the three-layer product model. Computer, 46(11):34–39, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Seceleanu:2013:CMR

[981] Cristina Seceleanu and Ivica Crnkovic. Component models for reasoning. Computer, 46(11):40–47, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lamprecht:2013:VMB

[982] Anna-Lena Lamprecht, Stefan Naujokat, and Ina Schaefer. Variability management beyond feature models. Computer, 46(11):48–54, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neubauer:2013:PPH

[983] Johannes Neubauer and Bernhard Steffen. Plug-and-play higher-order process integration. Computer, 46(11): 56–62, November 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Prasad:2013:TCY

[984] Chandra Prasad and Wolfram Schulte. Taking control of your engineering tools. Computer, 46(11):63–66, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:ICSc

[985] Anonymous. IEEE Computer Society information. Computer, 46(11):

67, November 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Seceleanu:2013:TIF

[986] Tiberiu Seceleanu and Gaetana Sapienza. A tool integration framework for sustainable embedded systems development. Computer, 46(11):68–71, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Herzlinger:2013:AKH

[987] Regina Herzlinger, Margo Seltzer, and Mark Gaynor. Applying KISS to healthcare information technology. Computer, 46(11):72–74, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bahga:2013:RPM

[988] Arshdeep Bahga and Vijay K. Madisetti. Rapid prototyping of multitier cloud-based services and systems. Computer, 46(11):76–83, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2013:DST

[989] James Bret Michael, Phillip A. Laplante, Jeffery Payne, Paul E. Black, and Jeffrey M. Voas. Does security trump reliability? Computer, 46(11): 84–86, November 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wei:2013:EDI

[990] David S. L. Wei, San Murugesan, Sy-Yen Kuo, Kshirasagar Naik, and Danny Krizanc. Enhancing data integrity and privacy in the cloud: An agenda. Computer, 46(11):87–90, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Klontz:2013:CSA

[991] Joshua C. Klontz and Anil K. Jain. A case study of automated face recognition: The Boston Marathon bombings suspects. Computer, 46(11):91– 94, November 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2013:JTM

[992] Brian David Johnson. Jimmy takes Manhattan. Computer, 46(11):95– 98, November 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CSCg

[993] Anonymous. Computer Society connection. Computer, 46(11):99–101, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCj

[994] Anonymous. Call and calendar. Computer, 46(11):102–105, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:TP

[995] David Alan Grier. The trappings of power. Computer, 46(11):120, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CA

[996] Anonymous. CES [advertisement]. Computer, 46(11):c4, November 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:FCb

[997] Anonymous. [Front cover]. Computer, 46(12):c1, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:ICC

[998] Anonymous. IEEE Cloud Computing. Computer, 46(12):c2, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:Mb

[999] Anonymous. [Masthead]. Computer, 46 (12):1, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:TCb

[1000] Anonymous. Table of contents. Computer, 46(12):2–3, December 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CHSl

[1001] Anonymous. Computer highlights Society magazines. Computer, 46(12): 4–5, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:CCI

[1002] David Alan Grier. The cycles of continuous improvement. Computer, 46 (12):6–8, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2013:ATW

[1003] Charles Severance. Andrew Tanenbaum: Writing the book on networks. Computer, 46(12):9–10, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2013:LCE

[1004] Brian M. Gaff and Judy St. John. Legal careers for engineers. Computer, 46 (12):11–13, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2013:YAj

[1005] Neville Holmes. 32 & 16 years ago. Computer, 46(12):14–15, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2013:RND

[1006] Lee Garber. Robot OS: A new day for robot design. Computer, 46(12): 16–20, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:NBc

[1007] Anonymous. News briefs. Computer, 46(12):21–23, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Getov:2013:CLO

[1008] Vladimir Getov. Computing laws: Origins, standing, and impact [guest editors’ introduction]. Computer, 46(12):

24–25, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Metcalfe:2013:MLA

[1009] Bob Metcalfe. Metcalfe’s Law after 40 years of Ethernet. Computer, 46(12): 26–31, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Makimoto:2013:IMW

[1010] Tsugio Makimoto. Implications of Makimoto’s Wave. Computer, 46(12): 32–37, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Amdahl:2013:CAA

[1011] Gene M. Amdahl. Computer architecture and Amdahl’s Law. Computer, 46 (12):38–46, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chien:2013:MLF

[1012] Andrew A. Chien and Vijay Karamcheti. Moore’s Law: The first ending and a new beginning. Computer, 46 (12):48–53, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ryan:2013:WCG

[1013] Patrick S. Ryan, Sarah Falvey, and Ronak Merchant. When the cloud goes local: The global problem with data localization. Computer, 46(12): 54–59, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanPelt:2013:UBF

[1014] Roy van Pelt and Anna Vilanova. Understanding blood-flow dynamics: New challenges for visualization. Computer, 46(12):60–67, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

ElDefrawy:2013:BDS

[1015] Karim El Defrawy and Sky Faber.

Blindfolded data search via secure pattern matching. Computer, 46(12):68– 75, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaffer:2013:PPP

[1016] Kim Schaffer. Passwords, privacy, and policies: Can they do business together? Computer, 46(12):76– 79, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hua:2013:ACU

[1017] Ting Hua, Chang-Tien Lu, Naren Ramakrishnan, Feng Chen, Jaime Arredondo, David Mares, and Kristen Summers. Analyzing civil unrest through social media. Computer, 46 (12):80–84, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sobel:2013:ECC

[1018] Ann E. K. Sobel. The escalating cost of college. Computer, 46(12): 85–87, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Resnick:2013:BDG

[1019] Hy Resnick. Busted: A digital group game to help at-risk youth stay out of jail. Computer, 46(12):88–90, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2013:IFH

[1020] Hal Berghel. The intimidation factor: How a surveillance state can affect what you read in professional publications. Computer, 46(12):91– 95, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Han:2013:MTB

[1021] Hao Han and Yinxing Xue. Mashup technology: Beyond open programming interfaces. Computer, 46(12): 96–99, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:ICSd

[1022] Anonymous. IEEE Computer Society information. Computer, 46(12): 100, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CVC

[1023] Anonymous. Conte voted 2014 Computer Society President-Elect. Computer, 46(12):101–103, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CSCh

[1024] Anonymous. Computer Society connection. Computer, 46(12):105–106, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:ICSe

[1025] Anonymous. IEEE Computer Society membership [advertisement]. Computer, 46(12):107, December 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2013:CCk

[1026] Anonymous. Call and calendar. Computer, 46(12):108–111, December 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2013:WWO

[1027] David Alan Grier. The way we organize ourselves. Computer, 46(12): 124, December 2013. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:IWC

[1028] Anonymous. IEEE 10th World Congress on Services (Services 2014) [house advertisement]. Computer, 47 (1):c2, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Ma

[1029] Anonymous. Masthead. Computer, 47(1):1, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:TCa

[1030] Anonymous. Table of contents. Computer, 47(1):2–3, January 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CHSa

[1031] Anonymous. Computer highlights Society magazines. Computer, 47(1): 4–5, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2014:RR

[1032] Dejan S. Milojicic. Reinventing relevance. Computer, 47(1):6–8, January 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Machiraju:2014:FMP

[1033] Raghu Machiraju. Fixing the Mercator projection for the Internet Age. Computer, 47(1):9, January 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:ICSa

[1034] Anonymous. IEEE Computer Society [house advertisement]. Computer, 47 (1):10, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2014:MBB

[1035] Charles Severance. Massimo Banzi: Building Arduino. Computer, 47(1): 11–12, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2014:CRS

[1036] Brian M. Gaff. Corporate risks from social media. Computer, 47(1):13– 15, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:YAa

[1037] Anonymous. 32 & 16 years ago.

Computer, 47(1):16–17, January 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:NBa

[1038] Anonymous. News briefs. Computer, 47(1):18–21, January 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:2014:DMF

[1039] Ron Vetter. Digital magazines: The future of publishing is here. Computer, 47 (1):22–23, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Strauss:2014:WFH

[1040] Karin Strauss and Doug Burger. What the future holds for solid-state memory. Computer, 47(1):24–31, January 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gollakota:2014:ERP

[1041] Shyamnath Gollakota, Matthew S.

Reynolds, Joshua R. Smith, and David J. Wetherall. The emergence of RF-powered computing. Computer, 47 (1):32–39, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:2014:ERD

[1042] Anuj Kumar, Florian Metze, and Matthew Kam. Enabling the rapid development and adoption of speechuser interfaces. Computer, 47(1): 40–47, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Murphy-Hill:2014:FSL

[1043] Emerson Murphy-Hill. The future of social learning in software engineering. Computer, 47(1):48–54, January 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kum:2014:SGP

[1044] Hye-Chung Kum, Ashok Krishnamurthy, Ashwin Machanavajjhala, and Stanley C. Ahalt. Social genome: Putting big data to work for population informatics. Computer, 47(1): 56–63, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Montuschi:2014:ARP

[1045] Paolo Montuschi and Alfredo Benso. Augmented reading: The present and future of electronic scientific publications. Computer, 47(1):64–74, January 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tilevich:2014:CBE

[1046] Eli Tilevich and Young-Woo Kwon. Cloud-based execution to improve mobile application energy efficiency. Computer, 47(1):75–77, January 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2014:PIP

[1047] Hal Berghel. Privacy informatics: A primer on defensive tactics for a society under siege. Computer, 47(1): 78–82, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2014:JCC

[1048] Nir Kshetri. Japan’s changing cybersecurity landscape. Computer, 47(1): 83–86, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2014:UR

[1049] Brian David Johnson. Utopia rising.

Computer, 47(1):87–89, January 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CSCa

[1050] Anonymous. Computer Society connection. Computer, 47(1):90, January 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CCa

[1051] Anonymous. Call and calendar. Computer, 47(1):91–93, January 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2014:JI

[1052] David Alan Grier. Just out of it.

Computer, 47(1):104, January 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:ICCa

[1053] Anonymous. IEEE Cloud Computing [house advertisement]. Computer, 47 (1):c4, January 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:FCa

[1054] Anonymous. Front cover. Computer, 47(2):c1, 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:ICSb

[1055] Anonymous. IEEE Computer Society membership house advertisement. Computer, 47(2):c2, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Mb

[1056] Anonymous. Masthead. Computer, 47 (2):1, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:TCb

[1057] Anonymous. Table of contents. Computer, 47(2):2–3, February 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CHSb

[1058] Anonymous. Computer highlights Society magazines. Computer, 47(2): 4–5, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Forsyth:2014:ODD

[1059] David Forsyth. Object detection with discriminatively trained partbased models. Computer, 47(2):6– 7, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2014:JHN

[1060] Charles Severance. Joseph Hardin: NCSA Mosaic. Computer, 47(2):8– 10, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2014:ILF

[1061] Brian M. Gaff. IP 2013–2014: Looking forward and looking back. Computer, 47(2):11–13, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2014:YA

[1062] Neville Holmes. 32 & 16 years ago.

Computer, 47(2):14–15, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2014:NBa

[1063] Lee Garber. News briefs. Computer, 47 (2):16–20, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bryce:2014:STG

[1064] Renee Bryce and Rick Kuhn. Software testing [guest editors’ introduction]. Computer, 47(2):21–22, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ali:2014:EFO

[1065] Midhat Ali, Francesco De Angelis, Daniele Fani, Antonia Bertolino, Guglielmo De Angelis, and Andrea Polini. An extensible framework for online testing of choreographed services. Computer, 47(2):23–29, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Antunes:2014:PTW

[1066] Nuno Antunes and Marco Vieira. Penetration testing for Web services. Computer, 47(2):30–36, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yilmaz:2014:MFC

[1067] Cemal Yilmaz, Sandro Fouche, Myra B. Cohen, Adam Porter, Gulsen Demiroz, and Ugur Koc. Moving forward with combinatorial interaction testing. Computer, 47(2):37–45, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gao:2014:MAT

[1068] Jerry Gao, Xiaoying Bai, Wei-Tek Tsai, and Tadahiro Uehara. Mobile application testing: A tutorial. Computer, 47 (2):46–55, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dahn:2014:ATM

[1069] Andreas Dahn and Clemens H. Cap. Application transparency and manipulation. Computer, 47(2):56–61, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chimiak:2014:EII

[1070] William J. Chimiak, Samuel T. Patton, and Stephen Janansky. Enhanced IP: IPv4 with 64-bit addresses. Computer, 47(2):62–69, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Broy:2014:CEC

[1071] Manfred Broy and Albrecht Schmidt. Challenges in engineering cyberphysical systems. Computer, 47(2): 70–72, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2014:MHN

[1072] Hal Berghel. Moral hazards, negative externalities, and the surveillance economy. Computer, 47(2):73–77, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vassilev:2014:IEI

[1073] Apostol Vassilev and Timothy A. Hall.

The importance of entropy to information security. Computer, 47(2): 78–81, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gruenbaum:2014:UTG

[1074] Peter Gruenbaum. Undergraduates teach game programming using scratch. Computer, 47(2):82–84, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Capilla:2014:CVC

[1075] Rafael Capilla, Oscar Ortiz, and Mike Hinchey. Context variability for context-aware systems. Computer, 47 (2):85–87, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Froyd:2014:NSC

[1076] Jeffrey E. Froyd. The nature of scholarship in computing education: An approach from IEEE Transactions on Education. Computer, 47(2):88– 90, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CSCb

[1077] Anonymous. Computer Society connection. Computer, 47(2):91, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CCb

[1078] Anonymous. Call and calendar. Computer, 47(2):92–93, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2014:PT

[1079] David Alan Grier. Practice and theory. Computer, 47(2):104, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:ICCb

[1080] Anonymous. IEEE Cloud Computing Magazine [house advertisement]. Computer, 47(2):c3, February 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:RSMa

[1081] Anonymous. Rock stars of mobile cloud [house advertisement]. Computer, 47 (2):c4, February 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:FCb

[1082] Anonymous. Front cover. Computer, 47(3):c1, 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:SESa

[1083] Anonymous. Software experts summit [house advertisement]. Computer, 47(3):c2, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Mc

[1084] Anonymous. Masthead. Computer, 47(3):1, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:TCc

[1085] Anonymous. Table of contents. Computer, 47(3):2–3, March 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CHSc

[1086] Anonymous. Computer highlights Society magazines. Computer, 47(3): 4–5, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brusilovsky:2014:CCW

[1087] Peter Brusilovsky. Collective class wisdom: Collaborative lecture annotation. Computer, 47(3):6, March 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:La

[1088] Anonymous. Letters. Computer, 47 (3):7–9, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2014:CCL

[1089] Charles Severance. Computing conversations: Len Kleinrock: The first two packets on the Internet. Computer, 47 (3):10–11, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2014:C

[1090] Brian M. Gaff. Contracts 101. Computer, 47(3):12–15, March 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:YAb

[1091] Anonymous. 32 & 16 years ago. Computer, 47(3):16–17, March 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2014:NBb

[1092] Lee Garber. News briefs. Computer, 47 (3):18–21, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shah:2014:CIS

[1093] Chirag Shah, Robert Capra, and Preben Hansen. Collaborative information seeking [guest editors’ introduction]. Computer, 47(3):22–25, March 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xu:2014:IKB

[1094] Zhe Xu, Jay Ramanathan, and Rajiv Ramnath. Identifying knowledge brokers and their role in enterprise research through social media. Computer, 47(3):26–31, March 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tao:2014:ICS

[1095] Yihan Tao and Anastasios Tombros. Investigating collaborative sensemaking behavior in collaborative information seeking. Computer, 47(3): 38–45, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yue:2014:IQR

[1096] Zhen Yue, Shuguang Han, Daqing He, and Jiepu Jiang. Influences on query reformulation in collaborative Web search. Computer, 47(3): 46–53, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hearst:2014:WMC

[1097] Marti A. Hearst. What’s missing from collaborative search? Computer, 47 (3):58–61, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:2014:WIY

[1098] Zheng Chen, Jian-Tao Sun, and Xuedong Huang. Web information at your fingertips: Paper as an interaction metaphor. Computer, 47(3): 62–66, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Timmerer:2014:SME

[1099] Christian Timmerer and Benjamin Rainer. The social multimedia experience. Computer, 47(3):67–69, March 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2014:SR

[1100] Hal Berghel. STEM, revisited. Computer, 47(3):70–73, March 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

daSilva:2014:HPB

[1101] Hugo Placido da Silva and Ana Fred. Harnessing the power of biosignals. Computer, 47(3):74–77, March 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2014:EPM

[1102] Kirk Cameron. The evolution of power measurement. Computer, 47 (3):78–80, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McGraw:2014:SFS

[1103] Gary McGraw. Security fatigue? shift your paradigm. Computer, 47(3): 81–83, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2014:DA

[1104] Brian David Johnson. The Darwin App. Computer, 47(3):84–86, March 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CSCc

[1105] Anonymous. Computer Society connection. Computer, 47(3):87, March 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CCc

[1106] Anonymous. Call and calendar. Computer, 47(3):88–89, March 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2014:AT

[1107] David Alan Grier. Assets to the table. Computer, 47(3):100, March 2014.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:RSMb

[1108] Anonymous. Rock stars of mobile cloud [house advertisement]. Computer, 47 (3):c4, March 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:FCc

[1109] Anonymous. Front cover. Computer, 47(4):c1, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:FYIa

[1110] Anonymous. Fuel your imagination [house advertisement]. Computer, 47 (4):c2, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Md

[1111] Anonymous. Masthead. Computer, 47(4):1, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:TCd

[1112] Anonymous. Table of contents. Computer, 47(4):2–3, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CHSd

[1113] Anonymous. Computer highlights Society magazines. Computer, 47(4):4–5, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2014:SSR

[1114] Ling Liu. Service selection and recommendation through collective intelligence. Computer, 47(4):6, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2014:DVH

[1115] Charles Severance. Doug Van Houweling: Building the NSFNet. Computer, 47(4):7–9, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2014:RCI

[1116] Brian M. Gaff. Readers’ choice [intellectual property laws]. Computer, 47 (4):10–12, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:YAc

[1117] Anonymous. 32 & 16 years ago. Computer, 47(4):13–14, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:NBb

[1118] Anonymous. News briefs. Computer, 47(4):15–19, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:2014:ACG

[1119] Carl Chang and Bill N. Schilit. Aware computing [Guest Editors’ introduction]. Computer, 47(4):20–21, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vieweg:2014:RCL

[1120] Sarah Vieweg and Adam Hodges. Rethinking context: Leveraging human and machine computation in disaster response. Computer, 47(4):22–27, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoque:2014:RNS

[1121] Mohammed Ehsan Hoque and Rosalind W. Picard. Rich nonverbal sensing technology for automated social skills training. Computer, 47 (4):28–35, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Matthews:2014:TMW

[1122] Mark Matthews, Saeed Abdullah, Geri Gay, and Tanzeem Choudhury. Tracking mental well-being: Balancing rich sensing and patient needs. Computer, 47(4):36–43, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Beroggi:2014:IVE

[1123] Giampiero E. G. Beroggi. Internet voting: An empirical evaluation. Computer, 47(4):44–50, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:SESb

[1124] Anonymous. Software experts summit [house advertisement]. Computer, 47(4):51, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Terven:2014:NOC

[1125] Juan R. Terven, Joaquin Salas, and Bogdan Raducanu. New opportunities for computer vision-based assistive technology systems for the visually impaired. Computer, 47(4):52–58, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2014:AVP

[1126] Ying-Dar Lin, Chia-Yin Lee, Yu-Sung Wu, Pei-Hsiu Ho, Fu-Yu Wang, and Yi-Lang Tsai. Active versus passive malware collection. Computer, 47 (4):59–65, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2014:MSL

[1127] Hal Berghel. Mr. Snowden’s legacy. Computer, 47(4):66–70, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vega:2014:BTB

[1128] Katia Vega and Hugo Fuks. Beauty technology: Body surface computing. Computer, 47(4):71–75, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:2014:SMI

[1129] Jooyoung Kim, Dong Sun Shin, Dae Ok Kim, Dae Hyun Lee, Jaekeun Kwak, Antonio Di Carlo, Alberto Paoluzzi, Kazuomi Sugamoto, and Young Lae Moon. Standardizing 3D medical imaging. Computer, 47(4):76–79, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Beecham:2014:MSE

[1130] Sarah Beecham, Padraig O’Leary, Sean Baker, Ita Richardson, and John Noll. Making software engineering research relevant. Computer, 47(4):80–83, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurlburt:2014:BDN

[1131] George F. Hurlburt and Jeffrey Voas.

Big data, networked worlds. Computer, 47(4):84–87, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arnez:2014:LWB

[1132] Fernando Arnez, Jason Pace, and Kelvin Sung. Learning while building games for teaching. Computer, 47 (4):88–91, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2014:FOIa

[1133] Sorel Reisman. The future of online instruction, Part 1. Computer, 47 (4):92–93, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Butler:2014:SIA

[1134] Patrick Butler, Naren Ramakrishnan, Elaine O. Nsoesie, and John S. Brownstein. Satellite imagery analysis: What can hospital parking lots tell us about a disease outbreak? Computer, 47 (4):94–97, April 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pingali:2014:CAE

[1135] Gopal Pingali and T. S. Mohan. The cloud ascends in emerging markets: Highlights of IEEE CCEM 2013. Computer, 47(4):98–101, April 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CSCd

[1136] Anonymous. Computer Society connection. Computer, 47(4):102, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CCd

[1137] Anonymous. Call and calendar. Computer, 47(4):103–105, April 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2014:AGG

[1138] David Alan Grier. All that glitters is not gold. Computer, 47(4):116, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:IAIa

[1139] Anonymous. Instant access to IEEE Publications [house advertisement]. Computer, 47(4):c3, April 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:RSMc

[1140] Anonymous. Rock stars of mobile cloud [house advertisement]. Computer, 47(4):c4, April 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:FCd

[1141] Anonymous. Front cover. Computer, 47(5):c1, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Me

[1142] Anonymous. Masthead. Computer, 47(5):1, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:TCe

[1143] Anonymous. Table of contents. Computer, 47(5):2–3, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CHSe

[1144] Anonymous. Computer highlights Society magazines. Computer, 47(5):4–5, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zomaya:2014:PPT

[1145] Albert Zomaya. Power-performance tradeoff in a federated cloud. Computer, 47(5):6–7, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CPY

[1146] Anonymous. Conferences in the palm of your hand [house advertisement]. Computer, 47(5):8, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2014:NBM

[1147] Charles Severance. Nathaniel Borenstein: Multipurpose Internet Mail Extensions (MIME). Computer, 47 (5):9–11, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:YAd

[1148] Anonymous. 32 & 16 years ago. Computer, 47(5):12–13, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:SCH

[1149] Anonymous. Stay connected [house advertisement]. Computer, 47(5):14, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Lb

[1150] Anonymous. Letters. Computer, 47 (5):15–17, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:NBc

[1151] Anonymous. News briefs. Computer, 47(5):18–22, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisslein:2014:SCN

[1152] Martin Reisslein, Bernhard Rinner, and Amit Roy-Chowdhury. Smart camera networks [guest editors’ introduction]. Computer, 47(5):23–25, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:2014:CNH

[1153] Ching-Hui Chen, Julien Favre, Gregorij Kurillo, Thomas P. Andriacchi, Ruzena Bajcsy, and Rama Chellappa. Camera networks for healthcare, teleimmersion, and surveillance. Computer, 47 (5):26–36, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abas:2014:WSC

[1154] Kevin Abas, Caio Porto, and Katia Obraczka. Wireless smart camera networks for the surveillance of public spaces. Computer, 47(5):37–44, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Prati:2014:ICS

[1155] Andrea Prati and Faisal Z. Qureshi. Integrating consumer smart cameras into camera networks: Opportunities and obstacles. Computer, 47(5):45–51, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mitra:2014:TCC

[1156] Kaushik Mitra, Ashok Veeraraghavan, Aswin C. Sankaranarayanan, and Richard G. Baraniuk. Toward compressive camera networks. Computer, 47(5):52–59, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Guan:2014:CAS

[1157] Zhangyu Guan and Tommaso Melodia.

Cloud-assisted smart camera networks for energy-efficient 3D video streaming. Computer, 47(5):60–66, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

SanMiguel:2014:SRS

[1158] Juan C. SanMiguel, Christian Micheloni, Karen Shoop, Gian Luca Foresti, and Andrea Cavallaro. Selfreconfigurable smart camera networks. Computer, 47(5):67–73, May 2014.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sun:2014:CAM

[1159] Xian-He Sun and Dawei Wang. Concurrent average memory access time. Computer, 47(5):74–80, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gao:2014:ISA

[1160] Wen Gao, Tiejun Huang, Cliff Reader, Weibei Dou, and Xilin Chen. IEEE standards for advanced audio and video coding in emerging applications. Computer, 47(5):81–83, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Papadopoulos:2014:SMC

[1161] Symeon Papadopoulos and Yiannis Kompatsiaris. Social multimedia crawling for mining and search. Computer, 47(5):84–87, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cho:2014:DGA

[1162] Michael Cheng Yi Cho, Pokai Chen, and Shiuhpyng Winston Shieh. Dmail: A globally authenticated email service. Computer, 47(5):88–91, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Preist:2014:AEE

[1163] Chris Preist, Dan Schien, Paul Shabajee, Stephen Wood, and Christopher Hodgson. Analyzing end-to-end energy consumption for digital services. Computer, 47(5):92–95, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowyer:2014:CCL

[1164] Kevin W. Bowyer and James S. Doyle.

Cosmetic contact lenses and iris recognition spoofing. Computer, 47(5): 96–98, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2014:NDT

[1165] Brian David Johnson. A new doorway into time. Computer, 47(5): 99–101, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2014:PFC

[1166] Hal Berghel. PII, the FTC, car dealers, and you. Computer, 47(5): 102–106, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CSCe

[1167] Anonymous. Computer Society connection. Computer, 47(5):107, May 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CCe

[1168] Anonymous. Call and calendar. Computer, 47(5):108–111, May 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:FYJ

[1169] Anonymous. Focus on your job search [house advertisement]. Computer, 47 (5):119, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2014:RCE

[1170] David Alan Grier. The real change [the errant hashtag]. Computer, 47 (5):120, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:MMHa

[1171] Anonymous. Membership matters [house advertisement]. Computer, 47 (5):c3, May 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:FCe

[1172] Anonymous. Front cover. Computer, 47(6):c1, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:RSCa

[1173] Anonymous. Rock stars of cybersecurity [house advertisement]. Computer, 47(6):c2, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Mf

[1174] Anonymous. Masthead. Computer, 47(6):1, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:TCf

[1175] Anonymous. Table of contents. Computer, 47(6):2–3, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CHSf

[1176] Anonymous. Computer highlights Society magazines. Computer, 47(6):4–5, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Guo:2014:TWD

[1177] Song Guo. On-time warning delivery for vehicular ad hoc networks. Computer, 47(6):6, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2014:PBD

[1178] Brian M. Gaff, Heather Egan Sussman, and Jennifer Geetter. Privacy and big data. Computer, 47(6):7–9, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:YAe

[1179] Anonymous. 32 & 16 years ago. Computer, 47(6):10–11, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:2014:NBc

[1180] Lee Garber. News briefs. Computer, 47(6):12–17, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Lc

[1181] Anonymous. Letters. Computer, 47 (6):18–19, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:ISP

[1182] Anonymous. IEEE Security & Privacy [house advertisement]. Computer, 47(6):20, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2014:MAS

[1183] Ying-Dar Lin, Chun-Ying Huang, Matthew Wright, and Georgios Kambourakis. Mobile application security. Computer, 47(6):21–23, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Suarez-Tangil:2014:TOM

[1184] Guillermo Suarez-Tangil, Juan E. Tapiador, Flavio Lombardi, and Roberto Di Pietro. Thwarting obfuscated malware via differential fault analysis. Computer, 47(6):24–31, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shao:2014:RPR

[1185] Yuru Shao, Xiapu Luo, and Chenxiong Qian. RootGuard: Protecting rooted Android phones. Computer, 47 (6):32–40, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:MMHb

[1186] Anonymous. Membership matters [house advertisement]. Computer, 47 (6):41, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marmol:2014:SAE

[1187] Felix Gomez Marmol, Gregor Rozinaj, Sebastian Schumann, Ondrej Labaj, and Juraj Kacur. Smart AppStore: Expanding the frontiers of Smartphone ecosystems. Computer, 47(6):42–47, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Armando:2014:SBY

[1188] Alessandro Armando, Gabriele Costa, Luca Verderame, and Alessio Merlo. Securing the “bring your own device” paradigm. Computer, 47(6): 48–56, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gadyatskaya:2014:SFT

[1189] Olga Gadyatskaya, Fabio Massacci, and Yury Zhauniarovich. Security in the Firefox OS and Tizen mobile platforms. Computer, 47(6):57–63, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2014:LFN

[1190] Hal Berghel. Leadership failures in the national security complex. Computer, 47(6):64–67, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Combemale:2014:GML

[1191] Benoit Combemale, Julien DeAntoni, Benoit Baudry, Robert B. France, Jean-Marc Jezequel, and Jeff Gray. Globalizing modeling languages. Computer, 47(6):68–71, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bojanova:2014:IIA

[1192] Irena Bojanova, George Hurlburt, and Jeffrey Voas. Imagineering an Internet of Anything. Computer, 47 (6):72–77, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pace:2014:WSW

[1193] Jason Pace. Working with students where theory and practice intersect. Computer, 47(6):78–80, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:ICSc

[1194] Anonymous. IEEE Computer Society information [house advertisement]. Computer, 47(6):81, June 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2014:FOIb

[1195] Sorel Reisman. The future of online instruction, part 2. Computer, 47 (6):82–84, June 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sturrus:2014:OHC

[1196] Edwin Sturrus and Olga Kulikova. Orchestrating hybrid cloud deployment: An overview. Computer, 47(6):85–87, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CSCf

[1197] Anonymous. Computer Society connection. Computer, 47(6):88–89, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CCf

[1198] Anonymous. Call and calendar. Computer, 47(6):90–93, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2014:POT

[1199] David Alan Grier. Peace in our time. Computer, 47(6):104, June 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:FCf

[1200] Anonymous. Front cover. Computer, 47(7):c1, July 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Mg

[1201] Anonymous. Masthead. Computer, 47(7):1, July 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:TCg

[1202] Anonymous. Table of contents. Computer, 47(7):2–3, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CHSg

[1203] Anonymous. Computer highlights Society magazines. Computer, 47(7):4–5, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pei:2014:MDI

[1204] Jian Pei. Managing data-intensive applications in the cloud. Computer, 47(7):6, July 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2014:ATI

[1205] Charles Severance. Andrew S. Tanenbaum: The impact of MINIX. Computer, 47(7):7–8, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2014:SCI

[1206] Brian M. Gaff. Supreme Court IP update. Computer, 47(7):9–12, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:YAf

[1207] Anonymous. 32 & 16 years ago. Computer, 47(7):13–14, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Ld

[1208] Anonymous. Letters. Computer, 47(7):15, July 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:NBd

[1209] Anonymous. News briefs. Computer, 47(7):16–21, July 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bimber:2014:TV

[1210] Oliver Bimber. Thinking virtual. Computer, 47(7):22–23, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Slater:2014:TSI

[1211] Mel Slater and Maria V. Sanchez-Vives. Transcending the self in immersive virtual reality. Computer, 47 (7):24–30, July 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rizzo:2014:VRE

[1212] Albert Rizzo, Arno Hartholt, Mario Grimani, Andrew Leeds, and Matt Liewer. Virtual reality exposure therapy for combat-related posttraumatic stress disorder. Computer, 47(7):31–37, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stowers:2014:REA

[1213] John R. Stowers, Anton Fuhrmann, Maximilian Hofbauer, Martin Streinzer, Axel Schmid, Michael H. Dickinson, and Andrew D. Straw. Reverse engineering animal vision with virtual reality and genetics. Computer, 47 (7):38–45, July 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuchs:2014:IT

[1214] Henry Fuchs, Andrei State, and Jean-Charles Bazin. Immersive 3D telepresence. Computer, 47(7):46–52, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yang:2014:SBT

[1215] Xia Yang, Nan Sang, and Jim Alves-Foss. Shortening the boot time of Android OS. Computer, 47(7):53–58, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:2014:MVB

[1216] Minyong Kim, Young Geun Kim, Sung Woo Chung, and Cheol Hong Kim. Measuring variance between Smartphone energy consumption and battery life. Computer, 47(7):59–65, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahn:2014:MTL

[1217] Youngjung Ahn, Yongsuk Lee, Jin-Young Choi, Gyungho Lee, and Dongkyun Ahn. Monitoring translation lookahead buffers to detect code injection attacks. Computer, 47(7):66–72, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2014:TF

[1218] Brian David Johnson. Talking to the future. Computer, 47(7):73–75, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2014:RDS

[1219] Hal Berghel. Robert David Steele on OSINT. Computer, 47(7):76–81, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nikolopoulos:2014:EET

[1220] Dimitrios S. Nikolopoulos, Hans Vandierendonck, Nikolaos Bellas, Christos D. Antonopoulos, Spyros Lalis, Georgios Karakonstantis, Andreas Burg, and Uwe Naumann. Energy efficiency through significancebased computing. Computer, 47(7):82– 85, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cesar:2014:TCS

[1221] Pablo Cesar, Rene Kaiser, and Marian F. Ursu. Toward connected shared experiences. Computer, 47(7):86–89, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:2014:CPO

[1222] Chi-Wei Wang, Chia-Wei Wang, Chu-An Hsieh, and Shiuhpyng Winston Shieh. Cloudebug: A programmable online malware testbed. Computer, 47 (7):90–92, July 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nikolich:2014:LMS

[1223] Paul Nikolich. 802 LAN/MAN standards: High quality and market relevant. Computer, 47(7):93–95, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CSCg

[1224] Anonymous. Computer Society connection. Computer, 47(7):96–97, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CCg

[1225] Anonymous. Call and calendar. Computer, 47(7):98–101, July 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2014:HTL

[1226] David Alan Grier. The half-told lie. Computer, 47(7):108, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:IIC

[1227] Anonymous. IEEE International Conference on Cloud Engineering (IC2E 2015). Computer, 47(7):c4, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:FYIb

[1228] Anonymous. Fuel your imagination [house advertisement]. Computer, 47 (8):c2, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Mh

[1229] Anonymous. Masthead. Computer, 47(8):1, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:TCh

[1230] Anonymous. Table of contents. Computer, 47(8):2–3, August 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CHSh

[1231] Anonymous. Computer highlights Society magazines. Computer, 47(8): 4–5, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grundy:2014:DCD

[1232] John Grundy. Determining the cause of design model inconsistencies. Computer, 47(8):6, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:2014:EFC

[1233] Charles Severance. Elizabeth Fong: Creating the SQL database standards. Computer, 47(8):7–8, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaff:2014:SPA

[1234] Brian M. Gaff and Bruce Rubinger.

The significance of prior art. Computer, 47(8):9–11, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:YAg

[1235] Anonymous. 32 & 16 years ago.

Computer, 47(8):12–13, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:NBe

[1236] Anonymous. News briefs. Computer, 47(8):14–18, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:RMDa

[1237] Anonymous. 2015 Richard E. Merwin Distinguished Service Award [call for award nominations]. Computer, 47 (8):19, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:2014:TCI

[1238] Ron Vetter. Technology commercialization in the innovation economy [Guest Editor’s introduction]. Computer, 47 (8):20–21, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:2014:RPC

[1239] Gary L. Miller and Robert R. Hoon.

The role of public comprehensive universities in closing the innovation deficit. Computer, 47(8):22–27, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ball:2014:UCM

[1240] Elyse N. Ball and Kenneth G. Preston. University commercialization models: The University of Akron. Computer, 47 (8):28–36, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Speicher:2014:TCD

[1241] Audrey Speicher. tekMountain: A community-driven corporate incubator. Computer, 47(8):37–39, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Srinivasan:2014:VCF

[1242] Savitha Srinivasan, Isaac Barchas, Mark Gorenberg, and Evangelos Simoudis. Venture capital: Fueling the innovation economy. Computer, 47 (8):40–47, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gorton:2014:CBD

[1243] Ian Gorton. Cyberinfrastructures: Bridging the divide between scientific research and software engineering. Computer, 47(8):48–55, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boring:2014:DPP

[1244] Sebastian Boring, Saul Greenberg, Jo Vermeulen, Jakub Dostal, and Nicolai Marquardt. The dark patterns of proxemic sensing. Computer, 47(8): 56–60, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:RSCb

[1245] Anonymous. Rock stars of cybersecurity [house advertisement]. Computer, 47(8):61, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mendonca:2014:AOC

[1246] Nabor C. Mendonca. Architectural options for cloud migration. Computer, 47 (8):62–66, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McDuffie:2014:FCE

[1247] Ernest L. McDuffie and Victor P. Piotrowski. The future of cybersecurity education. Computer, 47(8): 67–69, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2014:FDM

[1248] Hal Berghel. The future of digital money laundering. Computer, 47(8): 70–75, August 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Basten:2014:OEE

[1249] Dirk Basten and Thomas Hoerstrup.

Organizational effort estimation. Computer, 47(8):76–79, August 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wheeler:2014:PH

[1250] David A. Wheeler. Preventing Heartbleed. Computer, 47(8):80–83, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vonWangenheim:2014:TGP

[1251] Christiane Gresse von Wangenheim and Aldo von Wangenheim. Teaching game programming in family workshops. Computer, 47(8):84–87, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:ICSd

[1252] Anonymous. IEEE Computer Society election. Computer, 47(8):88–97, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:IPE

[1253] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 47(8):98–102, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CCh

[1254] Anonymous. Call and calendar. Computer, 47(8):103–107, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2014:GF

[1255] David Alan Grier. The guest at the feast. Computer, 47(8):116, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:CPHa

[1256] Anonymous. Call for papers [house advertisement]. Computer, 47(8):c3, August 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2014:Mi

[1257] Anonymous. [Masthead]. Computer, 47(9):1, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/09/mco2014090001.pdf.

Anonymous:2014:TCi

[1258] Anonymous. Table of contents. Computer, 47(9):2–3, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2014/09/mco2014090002.
pdf.

Anonymous:2014:CHSi

[1259] Anonymous. Computer highlights Society magazines. Computer, 47(9): 4–5, September 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090004.html.

Mohapatra:2014:CBE

[1260] Prasant Mohapatra. Channel bonding in evolving wireless LANs. Computer, 47(9):6, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/09/mco2014090006.html.

Severance:2014:NQB

[1261] Charles Severance. Nii Quaynor:

Bringing the Internet to Africa. Computer, 47(9):7–9, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/

                                        csdl/mags/co/2014/09/mco2014090007.
                                        html.

Anonymous:2014:YAh

[1262] Anonymous. 32 & 16 years ago. Computer, 47(9):10–11, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        09/mco2014090010.html.

Anonymous:2014:RMDb

[1263] Anonymous. Richard E. Merwin Distinguished Service Award. Computer, 47(9):12, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/09/mco2014090012.pdf.

Anonymous:2014:IAIb

[1264] Anonymous. Instant access to IEEE publications [advertisement]. Computer, 47(9):13, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2014/09/mco2014090013.
                                        pdf.

Anonymous:2014:Le

[1265] Anonymous. Letters. Computer, 47 (9):14–16, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/09/mco2014090014.html.

Anonymous:2014:NBf

[1266] Anonymous. News briefs. Computer, 47(9):17–21, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/09/mco2014090017.html.

Anonymous:2014:RSBa

[1267] Anonymous. Rock stars of big data analytics [advertisement]. Computer, 47(9):22, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/09/mco2014090022.pdf.

Pankratius:2014:CAS

[1268] Victor Pankratius and Chris Mattmann.

Computing in astronomy: To see the unseen. Computer, 47(9):23– 25, September 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090023.html.

Jenne:2014:SMW

[1269] Mark Jenne, Owen Boberg, Hasan Kurban, and Mehmet Dalkilic. Studying the Milky Way galaxy using ParaHeap-k. Computer, 47(9):26– 33, September 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090026-abs.html.

Frascarelli:2014:HPC

[1270] Daniel Frascarelli, Sergio Nesmachnow, and Gonzalo Tancredi. Highperformance computing of self-gravity for small solar system bodies. Computer, 47(9):34–39, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/

csdl/mags/co/2014/09/mco2014090034-
abs.html.

Anonymous:2014:MMA

[1271] Anonymous. Membership matters [advertisement]. Computer, 47(9): 40, September 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090040.pdf.

Muller:2014:SAP

[1272] Stefan C. Muller, Gustavo Alonso, and Andre Csillaghy. Scaling astroinformatics: Python + automatic parallelization. Computer, 47(9):41– 47, September 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090041-abs.html.

Jongerius:2014:EEC

[1273] Rik Jongerius, Stefan Wijnholds, Ronald Nijboer, and Henk Corporaal. An end-to-end computing model for the square kilometre array. Computer, 47 (9):48–54, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/09/mco2014090048-abs.html.

Anonymous:2014:ICSe

[1274] Anonymous. IEEE Computer Society Harlan D. Mills Award: Call for software engineering award nominations. Computer, 47(9):55, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090055.pdf.

Gerhardt:2014:AAC

[1275] Lisa Gerhardt, Juan Carlos Diaz Velez, and Spencer R. Klein. Adventures in Antarctic computing, or how i learned to stop worrying and love the neutrino. Computer, 47(9):56– 61, September 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090056-abs.html.

Bogert:2014:CAA

[1276] F. Alexander Bogert, Nicholas Smith, John Holdener, Eric M. De Jong, Andrew F. Hart, Luca Cinquini, Shakeh E. Khudikyan, David R. Thompson, Chris A. Mattmann, Kiri Wagstaff, Joseph Lazio, Dayton L. Jones, Alice Allen, Lior Shamir, and Peter Teuben. Computing in astronomy: Applications and examples. Computer, 47(9):62–69, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090062-abs.html.

Collier:2014:CSM

[1277] Zachary A. Collier, Daniel DiMase, Steve Walters, Mark Mohammad Tehranipoor, James H. Lambert, and Igor Linkov. Cybersecurity standards: Managing risk and creating resilience. Computer, 47(9):70– 76, September 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090070-abs.html.

Anonymous:2014:CP

[1278] Anonymous. Call for papers. Computer, 47(9):77, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2014/09/mco2014090077.
                                        pdf.

Hunter:2014:ASS

[1279] Hillery Hunter, Luis A. Lastras-Montano, and Bishwaranjan Bhattacharjee. Adapting server systems for new memory technologies. Computer, 47(9):78–84, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/09/mco2014090078-abs.html.

Hubschman:2014:MSR

[1280] Julie Hubschman. My summer with a robot. Computer, 47(9):85–86, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        09/mco2014090085.html.

Ricanek:2014:BRP

[1281] Karl Ricanek. Beyond recognition: The promise of biometric analytics. Computer, 47(9):87–89, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        09/mco2014090087-abs.html.

Berghel:2014:SW

[1282] Hal Berghel. Sticky wikis. Computer, 47(9):90–93, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/09/mco2014090090-abs.html.

Wack:2014:IVV

[1283] John Wack. IEEE VSSC/1622: Voting System Standards. Computer, 47 (9):94–97, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/09/mco2014090094.html.

Hossfeld:2014:CQE

[1284] Tobias Hossfeld, Christian Keimel, and Christian Timmerer. Crowdsourcing quality-of-experience assessments. Computer, 47(9):98–102, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090098-abs.html.

Gandhi:2014:GIF

[1285] Robin A. Gandhi, Keesha Crosby, Harvey Siy, and Sayonnha Mandal. Gauging the impact of FISMA on software security. Computer, 47(9): 103–107, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/09/mco2014090103-abs.html.

Anonymous:2014:CSCh

[1286] Anonymous. Computer Society connection. Computer, 47(9):108, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
09/mco2014090108.html.

Anonymous:2014:CCi

[1287] Anonymous. Call and calendar. Computer, 47(9):109–112, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        09/mco2014090109.html.

Anonymous:2014:SHA

[1288] Anonymous. Symantec house advertisement. Computer, 47(9):113, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        09/mco2014090113.pdf.

Anonymous:2014:COa

[1289] Anonymous. Career opportunities. Computer, 47(9):114–119, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        09/mco2014090114.pdf.

Grier:2014:PR

[1290] David Alan Grier. The point of recognition. Computer, 47(9):120, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        09/mco2014090120.html.

Anonymous:2014:IAa

[1291] Anonymous. IEEE advertisement. Computer, 47(9):c3, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/

csdl/mags/co/2014/09/mco20140900c3.
pdf.

Anonymous:2014:ICH

[1292] Anonymous. 2015 International CES House advertisement. Computer, 47 (9):c2, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/09/mco20140900c2.pdf.

Anonymous:2014:KOS

[1293] Anonymous. Keep the office safe wherever you work. Computer, 47 (9):c4, September 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/09/mco20140900c4.pdf.

Anonymous:2014:Mj

[1294] Anonymous. Masthead. Computer, 47(10):1, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/10/mco2014100001.pdf.

Anonymous:2014:CPHb

[1295] Anonymous. Call for papers house advertisement. Computer, 47(10): 3, October 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
10/mco2014100003.pdf.

Anonymous:2014:ISA

[1296] Anonymous. IEEE SOSE 2015 advertisement. Computer, 47(10):5, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        10/mco2014100005.pdf.

Anonymous:2014:CHSj

[1297] Anonymous. Computer highlights Society magazines. Computer, 47(10): 6–7, October 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        10/mco2014100006.html.

Heilig:2014:VCC

[1298] Leonard Heilig. A view on cloud computing from above the clouds. Computer, 47(10):8, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/10/mco2014100008.html.

Gaff:2014:I

[1299] Brian M. Gaff. 3D IP. Computer, 47(10):9–11, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/10/mco2014100009-abs.html.

Anonymous:2014:YAi

[1300] Anonymous. 32 & 16 years ago. Computer, 47(10):12–13, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        10/mco2014100012.html.

Anonymous:2014:NBg

[1301] Anonymous. News briefs. Computer, 47(10):14–20, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/10/mco2014100014.html.

Anonymous:2014:ICAa

[1302] Anonymous. 2015 International CES advertisement. Computer, 47(10): 21, October 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
10/mco2014100021.pdf.

Jimenez:2014:GIL

[1303] Carlos E. Jimenez, Agusti Solanas, and Francisco Falcone. E-government interoperability: Linking open and smart government. Computer, 47 (10):22–24, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/10/mco2014100022.html.

Bovalis:2014:PIE

[1304] Konstantinos Bovalis, Vassilios Peristeras, Margarida Abecasis, Raul-Mario Abril-Jimenez, Miguel Alvarez Rodriguez, Corinne Gattegno, Athanasios Karalopoulos, Ioannis Sagias, Szabolcs Szekacs, and Suzanne Wigard. Promoting interoperability in Europe’s e-government. Computer, 47(10):25– 33, October 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
10/mco2014100025-abs.html.

Kovac:2014:HDG

[1305] Mario Kovac. E-health demystified:

An e-government showcase. Computer, 47(10):34–42, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/10/mco2014100034-abs.html.

Anonymous:2014:IAb

[1306] Anonymous. IEEE advertisement. Computer, 47(10):43, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2014/10/mco2014100043.
pdf.

Janssen:2014:IBO

[1307] Marijn Janssen, Elsa Estevez, and Tomasz Janowski. Interoperability in big, open, and linked data — organizational maturity, capabilities, and data portfolios. Computer, 47(10): 44–49, October 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
10/mco2014100044-abs.html.

Colpaert:2014:QIO

[1308] Pieter Colpaert, Mathias Van Compernolle, Laurens De Vocht, Anastasia Dimou, Miel Vander Sande, Ruben Verborgh, Peter Mechant, and Erik Mannens. Quantifying the interoperability of open government datasets. Computer, 47(10):50–56, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
10/mco2014100050-abs.html.

leVrang:2014:EBJ

[1309] Martin le Vrang, Agis Papantoniou, Erika Pauwels, Pieter Fannes, Dominique Vandensteen, and Johan De Smedt. ESCO: Boosting job matching in Europe with semantic interoperability. Computer, 47(10):57– 64, October 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
10/mco2014100057-abs.html.

Cano:2014:DFE

[1310] Jesus Cano, Roberto Hernandez, and Salvador Ros. Distributed framework for electronic democracy in smart cities. Computer, 47(10):65– 71, October 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
10/mco2014100065-abs.html.

Liu:2014:CSC

[1311] Pu Liu and Zhenghong Peng. China’s smart city pilots: A progress report. Computer, 47(10):72–81, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
10/mco2014100072-abs.html.

Secci:2014:CNE

[1312] Stefano Secci and San Murugesan.

Cloud networks: Enhancing performance and resiliency. Computer, 47 (10):82–85, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/10/mco2014100082-abs.html.

Bowen:2014:FAS

[1313] Jonathan P. Bowen, Mike Hinchey, Helge Janicke, Martin Ward, and Hussein Zedan. Formality, agility, security, and evolution in software development. Computer, 47(10):86– 89, October 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        10/mco2014100086-abs.html.

DuBow:2014:ARW

[1314] Wendy DuBow. Attracting and retaining women in computing. Computer, 47(10):90–93, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/10/mco2014100090-abs.html.

Berghel:2014:DAB

[1315] Hal Berghel. Defending against Big Dada: Defensive tactics for weapons of mass deception. Computer, 47 (10):94–98, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/10/mco2014100094-abs.html.

Bay:2014:TVG

[1316] Jason W. Bay. Turning video gamers into software developers. Computer, 47 (10):99–101, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/10/mco2014100099-abs.html.

Vassilev:2014:ACC

[1317] Apostol Vassilev and Christopher Celi. Avoiding cyberspace catastrophes through smarter testing. Computer, 47 (10):102–106, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/10/mco2014100102-abs.html.

Anonymous:2014:CCj

[1318] Anonymous. Call and calendar.

Computer, 47(10):107–109, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
10/mco2014100107.html.

Anonymous:2014:COb

[1319] Anonymous. Career opportunities. Computer, 47(10):110–125, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
10/mco2014100110.pdf.

Anonymous:2014:RSBb

[1320] Anonymous. Rock stars of big data analytics house advertisement. Computer, 47(10):126, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/10/mco2014100126.pdf.

Anonymous:2014:ICCc

[1321] Anonymous. IEEE cloud computing house advertisement. Computer, 47(10):127, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/10/mco2014100127.pdf.

Grier:2014:LF

[1322] David Alan Grier. Lying to the future. Computer, 47(10):128, October 2014. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        10/mco2014100128.html.

Anonymous:2014:IAc

[1323] Anonymous. IEEE advertisement. Computer, 47(10):c3, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2014/10/mco20141000c3.
                                        pdf.

Anonymous:2014:SAN

[1324] Anonymous. Samsung advertisement: The next big thing for business is here. Computer, 47(10): c2, October 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        10/mco20141000c2.pdf.

Anonymous:2014:SAO

[1325] Anonymous. Samsung advertisement: One giant leap forward. Computer, 47(10):c4, October 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/10/mco20141000c4.pdf.

Anonymous:2014:Mk

[1326] Anonymous. Masthead. Computer, 47(11):1, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/11/mco2014110001.pdf.

Anonymous:2014:BILa

[1327] Anonymous. Bring IT to life advertisement. Computer, 47(11):3, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110003.pdf.

Anonymous:2014:BILb

[1328] Anonymous. Bring IT to life advertisement. Computer, 47(11):5, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110005.pdf.

Anonymous:2014:CHSk

[1329] Anonymous. Computer highlights Society magazines. Computer, 47(11): 6–7, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110006.html.

Klatzky:2014:RTF

[1330] Roberta L. Klatzky and J. Edward Colgate. Rendering tactile features on touchscreens. Computer, 47(11): 8, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110008.html.

Anonymous:2014:DBTa

[1331] Anonymous. Discover the best tools and solutions from Wiley and Wiley-IEEE Computer Press house advertisement. Computer, 47(11):9, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110009.pdf.

Severance:2014:IFG

[1332] Charles Severance. Ian Foster and the Globus project. Computer, 47 (11):10–11, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/11/mco2014110010.html.

Anonymous:2014:YAj

[1333] Anonymous. 32 & 16 years ago. Computer, 47(11):12–13, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110012.html.

Garber:2014:NBd

[1334] Lee Garber. News briefs. Computer, 47 (11):14–18, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/11/mco2014110014.html.

Lin:2014:SDN

[1335] Ying-Dar Lin, Dan Pitt, David Hausheer, Erica Johnson, and Yi-Bing Lin. Software-defined networking: Standardization for cloud computing’s second wave. Computer, 47 (11):19–21, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/11/mco2014110019.html.

Tourrilhes:2014:SOE

[1336] Jean Tourrilhes, Puneet Sharma, Sujata Banerjee, and Justin Pettit. SDN and OpenFlow evolution: A standards perspective. Computer, 47(11): 22–29, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110022-abs.html.

Bauer:2014:ATM

[1337] Rick Bauer, Ron Milford, and Li Zhen.

Aligning technology and market drivers in an open source standards testing program. Computer, 47(11):30– 36, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110030-abs.html.

Anonymous:2014:ICAb

[1338] Anonymous. 2015 International CES advertisement. Computer, 47(11): 37, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110037.pdf.

Quinn:2014:SFC

[1339] Paul Quinn and Jim Guichard. Service function chaining: Creating a service plane via network service headers. Computer, 47(11):38–44, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110038-abs.html.

Anonymous:2014:IBH

[1340] Anonymous. IEEE BigDataService 2015 house advertisement. Computer, 47(11):45, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/11/mco2014110045.pdf.

Rothenberg:2014:WOS

[1341] Christian Esteve Rothenberg, Roy Chua, Josh Bailey, Martin Winter, Carlos N. A. Correa, Sidney C. de Lucena, Marcos Rogerio Salvador, and Thomas D. Nadeau. When open source meets network control planes. Computer, 47(11):46–54, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110046-abs.html.

Anonymous:2014:CCX

[1342] Anonymous. COOL Chips XVIII house advertisement. Computer, 47 (11):55, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/11/mco2014110055.pdf.

Canini:2014:SDN

[1343] Marco Canini, Anja Feldmann, Dan Levin, Fabian Schaffert, and Stefan Schmid. Software-defined networks: Incremental deployment with Panopticon. Computer, 47(11):56– 60, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110056-abs.html.

Anonymous:2014:IAIc

[1344] Anonymous. Instant access to IEEE publications house advertisement. Computer, 47(11):61, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110061.pdf.

Dillon:2014:VHN

[1345] Marion Dillon and Timothy Winters. Virtualization of home network gateways. Computer, 47(11): 62–65, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110062-abs.html.

Lowther:2014:DB

[1346] Lindsey Lowther. Defensive biometrics. Computer, 47(11):66–69, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110066-abs.html.

mcschraefel:2014:WCU

[1347] m.c. schraefel and Elizabeth F.

Churchill. Wellth creation: Using computer science to support proactive health. Computer, 47(11):70– 72, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110070-abs.html.

Anonymous:2014:SA

[1348] Anonymous. Symantec advertisement. Computer, 47(11):73, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110073.pdf.

Faghmous:2014:TGD

[1349] James H. Faghmous, Arindam Banerjee, Shashi Shekhar, Michael Steinbach, Vipin Kumar, Auroop R. Ganguly, and Nagiza Samatova. Theory-guided data science for climate change. Computer, 47(11):74–78, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/11/mco2014110074-abs.html.

Johnson:2014:IBR

[1350] Brian David Johnson. Imagining and building robots (in capes!). Computer, 47(11):79–81, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/11/mco2014110079.html.

Berghel:2014:WCG

[1351] Hal Berghel. Why clouds give me a case of the vapors. Computer, 47 (11):82–85, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/11/mco2014110082.html.

Jain:2014:TSL

[1352] Ramesh Jain. Toward social life networks. Computer, 47(11):86–88, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110086-abs.html.

Cameron:2014:EEW

[1353] Kirk W. Cameron. Energy efficiency in the wild: Why datacenters fear power management. Computer, 47 (11):89–92, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/11/mco2014110089-abs.html.

Anonymous:2014:FYIc

[1354] Anonymous. Fuel your imagination advertisement. Computer, 47(11): 93, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110093.pdf.

Hurlburt:2014:SPM

[1355] George Hurlburt, Irena Bojanova, Ann Sobel, and Keesha Crosby. Security or privacy? A matter of perspective. Computer, 47(11):94– 98, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110094-abs.html.

Joannou:2014:UIL

[1356] Paul Joannou and Alan Wassyng.

Understanding integrity level concepts. Computer, 47(11):99–101, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110099-abs.html.

Anonymous:2014:CSCi

[1357] Anonymous. Computer society connection. Computer, 47(11):102, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110102.html.

Anonymous:2014:ISCa

[1358] Anonymous. 2015 IEEE Services Congress IEEE BigData Congress house advertisement. Computer, 47 (11):103, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/11/mco2014110103.pdf.

Anonymous:2014:CCk

[1359] Anonymous. Call and calendar. Computer, 47(11):104–107, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110104.html.

Anonymous:2014:COc

[1360] Anonymous. Career opportunities. Computer, 47(11):108–123, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110108.pdf.

Grier:2014:EE

[1361] David Alan Grier. The external environment. Computer, 47(11):124, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco2014110124.html.

Anonymous:2014:IIP

[1362] Anonymous. 29th IEEE international parallel & distributed processing symposium house advertisement. Computer, 47(11):c3, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco20141100c3.pdf.

Anonymous:2014:NBTa

[1363] Anonymous. The next big thing for business is here. Computer, 47(11): c2, November 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco20141100c2.pdf.

Anonymous:2014:OGLa

[1364] Anonymous. One giant leap forward. Computer, 47(11):c4, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
11/mco20141100c4.pdf.

Anonymous:2014:Ml

[1365] Anonymous. Masthead. Computer, 47(12):1, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/12/mco2014120001.pdf.

Anonymous:2014:UCA

[1366] Anonymous. The ultimate competitive advantage bring IT to life advertisement. Computer, 47(12):

3, December 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120003.pdf.

Anonymous:2014:TEW

[1367] Anonymous. Technology to empower the workforce bring IT to life advertisement. Computer, 47(12): 5, December 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120005.pdf.

Anonymous:2014:CHSl

[1368] Anonymous. Computer highlights Society magazines. Computer, 47(12): 6–7, December 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120006.html.

Jin:2014:WDG

[1369] Hai Jin. When data grows big. Computer, 47(12):8, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2014/12/mco2014120008.
html.

Anonymous:2014:DBTb

[1370] Anonymous. Discover the best tools and solutions from Wiley and Wiley-IEEE Computer Press house advertisement. Computer, 47(12):9, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120009.pdf.

Milojicic:2014:R

[1371] Dejan S. Milojicic. Revitalization. Computer, 47(12):10–13, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120010.html.

Anonymous:2014:YAk

[1372] Anonymous. 32 & 16 years ago. Computer, 47(12):14–15, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120014.html.

Anonymous:2014:NBh

[1373] Anonymous. News briefs. Computer, 47(12):16–20, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/12/mco2014120016.html.

Anonymous:2014:FYId

[1374] Anonymous. Fuel your imagination house advertisement. Computer, 47 (12):21, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/12/mco2014120021.pdf.

Johnson:2014:MCG

[1375] Brian David Johnson. Making in the 21st Century [Guest Editor’s introduction]. Computer, 47(12):22– 23, December 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120022.html.

Axup:2014:WM

[1376] Jun Axup, AnnMarie Thomas, Ariel Waldman, Susan Faulkner, Carolina Odman-Govender, Jim St. Leger, Jessica Polka, Melissa Gregg, and Brian David Johnson. The world of making. Computer, 47(12):24– 40, December 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120024-abs.html.

Anonymous:2014:ISCb

[1377] Anonymous. 2015 IEEE Service Congress and IEEE BigData Congress house advertisement. Computer, 47 (12):41, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/12/mco2014120041.pdf.

Beasley:2014:PPW

[1378] Crystal Beasley, Emmeline Parsons, Melissa Pierce, Sam Baumgarten, Andrius Neviera, Shubham Banerjee, Tyesha Snow, Carla Diana, Nicholas Nakadate, and Shruti Satsangi. Passionate projects: This is what I made. Computer, 47(12):42–54, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120042-abs.html.

Anonymous:2014:LCA

[1379] Anonymous. Linkedin Corp. advertisement. Computer, 47(12):55, December 2014. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120055.pdf.

Sheri:2014:FM

[1380] Yasaman Sheri, Enrique Gabriel Legaspi, Cynthia Erenas, Brian Kessler, Eri Gentry, and Ruben Estrada. The future of making. Computer, 47(12):56–63, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120056-abs.html.

Edwards:2014:WSN

[1381] William Edwards, Judith Gebauer, and Bryan Reinicke. White space networks: Architecture, application, and opportunity. Computer, 47(12):64– 71, December 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120064-abs.html.

Bartoli:2014:ASR

[1382] Alberto Bartoli, Giorgio Davanzo, Andrea De Lorenzo, Eric Medvet, and Enrico Sorio. Automatic synthesis of regular expressions from examples. Computer, 47(12):72–80, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120072-abs.html.

Anonymous:2014:HA

[1383] Anonymous. House advertisement.

Computer, 47(12):81, December 2014.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2014/12/mco2014120081.
pdf.

Jung:2014:COD

[1384] Eun-Sung Jung and Rajkumar Kettimuthu. Challenges and opportunities for data-intensive computing in the cloud. Computer, 47(12):82– 85, December 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120082-abs.html.

Ardito:2014:HCD

[1385] Carmelo Ardito, Paolo Buono, Danilo Caivano, Maria Francesca Costabile, Rosa Lanzilotti, and Yvonne Dittrich. Human-centered design in industry: Lessons from the trenches. Computer, 47(12):86–89, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/12/mco2014120086-abs.html.

Jin:2014:MPA

[1386] Fang Jin, Wei Wang, Liang Zhao, Edward Dougherty, Yang Cao, Chang-Tien Lu, and Naren Ramakrishnan. Misinformation propagation in the age of Twitter. Computer, 47(12):90– 94, December 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120090-abs.html.

Berghel:2014:JRC

[1387] Hal Berghel. Judson Rosebush on computer graphics and animation. Computer, 47(12):95–99, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120095.html.

Chau:2014:CGT

[1388] Brian Chau, Adam Robinson, Jason Pace, Rob Nash, and Kelvin Sung. Corrupted: A game to teach programming concepts. Computer, 47(12): 100–103, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/12/mco2014120100-abs.html.

Kott:2014:SOS

[1389] Alexander Kott, Ananthram Swami, and Patrick McDaniel. Security outlook: Six cyber game changers for the next 15 years. Computer, 47(12): 104–106, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2014/12/mco2014120104-abs.html.

Anonymous:2014:CCl

[1390] Anonymous. Call and calendar.

Computer, 47(12):107–110, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco2014120107.html.

Anonymous:2014:FVC

[1391] Anonymous. Fujii voted 2015 Computer Society President-Elect. Computer, 47(12):111–113, December 2014. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        12/mco2014120111.html.

Anonymous:2014:COd

[1392] Anonymous. Career opportunities. Computer, 47(12):114–127, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        12/mco2014120114.pdf.

Grier:2014:SS

[1393] David Alan Grier. State secrets. Computer, 47(12):128, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        12/mco2014120128.html.

Anonymous:2014:FCg

[1394] Anonymous. Front cover. Computer, 47(12):c1, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2014/12/mco20141200c1.pdf.

Anonymous:2014:IEU

[1395] Anonymous. If “expect the unexpected” weren’t so expected, we’d say that’s exactly what to expect. Computer, 47(12):c3, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2014/
                                        12/mco20141200c3.pdf.

Anonymous:2014:NBTb

[1396] Anonymous. The next big thing for business is here. Computer, 47(12): c2, December 2014. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco20141200c2.pdf.

Anonymous:2014:OGLb

[1397] Anonymous. One giant leap forward. Computer, 47(12):c4, December 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2014/
12/mco20141200c4.pdf.

Anonymous:2015:TCa

[1398] Anonymous. Table of contents.

Computer, 48(1):1–2, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/01/mco2015010001.
pdf.

Anonymous:2015:Ma

[1399] Anonymous. [Masthead]. Computer, 48(1):3, January 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010003.pdf.

Conte:2015:TC

[1400] Tom Conte. A time of change.

Computer, 48(1):4–6, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/01/mco2015010004.
html.

Helal:2015:SCI

[1401] Sumi Helal. Sustaining Computer’s impact and adapting to change. Computer, 48(1):7–9, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        01/mco2015010007.html.

Anonymous:2015:CHSa

[1402] Anonymous. Computer highlights Society magazines. Computer, 48(1): 10–11, January 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        01/mco2015010010.html.

Xu:2015:CPA

[1403] Ying Xu. Carrying out phylogenetic analyses through computational model checking. Computer, 48(1): 12, January 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        01/mco2015010012.html.

Anonymous:2015:SRSa

[1404] Anonymous. Startup rock stars [advertisement]. Computer, 48(1):13, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        01/mco2015010013.pdf.

Severance:2015:KAC

[1405] Charles Severance. Khan Academy and computer science. Computer, 48 (1):14–15, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/01/mco2015010014.html.

Gaff:2015:IIS

[1406] Brian M. Gaff. IP issues for startups. Computer, 48(1):16–18, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010016-abs.html.

Anonymous:2015:RSPa

[1407] Anonymous. Rock stars of 3D printing [advertisement]. Computer, 48 (1):19, January 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010019.pdf.

Anonymous:2015:YAa

[1408] Anonymous. 32 & 16 years ago. Computer, 48(1):20–21, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010020.html.

Anonymous:2015:Na

[1409] Anonymous. News. Computer, 48 (1):22–27, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/01/mco2015010022.html.

Want:2015:EIT

[1410] Roy Want, Bill N. Schilit, and Scott Jenson. Enabling the Internet of Things. Computer, 48(1):

28–35, January 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010028-abs.html.

Gebara:2015:SGB

[1411] Fadi H. Gebara, H. Peter Hofstee, and Kevin J. Nowka. Second-generation big data systems. Computer, 48(1): 36–41, January 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010036-abs.html.

Fonash:2015:CMM

[1412] Peter Fonash and Phyllis Schneck. Cybersecurity: From months to milliseconds. Computer, 48(1):42–50, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010042-abs.html.

Anonymous:2015:II

[1413] Anonymous. IEEE IRI-2015. Computer, 48(1):51, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/01/mco2015010051.pdf.

Patsakis:2015:IPA

[1414] Constantinos Patsakis, Paul Laird, Michael Clear, Melanie Bouroche, and Agusti Solanas. Interoperable privacy-aware e-participation within smart cities. Computer, 48(1):52– 58, January 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010052-abs.html.

Crago:2015:HCC

[1415] Stephen P. Crago and John Paul Walters. Heterogeneous cloud computing: The way forward. Computer, 48(1):59–61, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/01/mco2015010059-abs.html.

Caviglione:2015:UIH

[1416] Luca Caviglione and Wojciech Mazurczyk. Understanding information hiding in iOS. Computer, 48(1):62– 65, January 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010062-abs.html.

Daniels:2015:MGJ

[1417] Mitchell E. Daniels. Measuring great jobs and great lives: The Gallup– Purdue Index. Computer, 48(1):66– 69, January 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010066-abs.html.

Popper:2015:TBT

[1418] Ari Popper. Transforming business through science fiction prototyping. Computer, 48(1):70–72, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010070.html.

Costley:2015:LSL

[1419] Jennifer Costley. Leveraging the standards life cycle for green IT. Computer, 48(1):73–75, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/01/mco2015010073-abs.html.

Berghel:2015:PG

[1420] Hal Berghel. A paperless gamble. Computer, 48(1):76–78, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        01/mco2015010076.html.

Anonymous:2015:CCa

[1421] Anonymous. CS connection. Computer, 48(1):79–80, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2015/01/mco2015010079.
                                        html.

Anonymous:2015:CCb

[1422] Anonymous. Call and calendar. Computer, 48(1):81–83, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2015/01/mco2015010081.
                                        html.

Anonymous:2015:COa

[1423] Anonymous. Career opportunities. Computer, 48(1):84–91, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010084.pdf.

Grier:2015:ST

[1424] David Alan Grier. State transitions. Computer, 48(1):92, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco2015010092.html.

Anonymous:2015:CPa

[1425] Anonymous. Call for papers. Computer, 48(1):c3, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/01/mco20150100c3.pdf.

Anonymous:2015:FCa

[1426] Anonymous. [Front cover]. Computer, 48(1):c1, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/01/mco20150100c1.pdf.

Anonymous:2015:ISC

[1427] Anonymous. 2015 IEEE Services Congress and IEEE Big Data Congress. Computer, 48(1):c2, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
01/mco20150100c2.pdf.

Anonymous:2015:OGL

[1428] Anonymous. One giant leap forward. Computer, 48(1):c4, January 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        01/mco20150100c4.pdf.

Anonymous:2015:TCb

[1429] Anonymous. Table of contents. Computer, 48(2):1–2, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2015/02/mco2015020001.
                                        pdf.

Anonymous:2015:Mb

[1430] Anonymous. [Masthead]. Computer, 48(2):3, February 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        02/mco2015020003.pdf.

Anonymous:2015:CHSb

[1431] Anonymous. Computer highlights Society magazines. Computer, 48(2): 4–5, February 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        02/mco2015020004.html.

Bertino:2015:DSA

[1432] Elisa Bertino. Denial-of-service attacks to UMTS. Computer, 48(2): 6, February 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        02/mco2015020006.html.

Severance:2015:GVRa

[1433] Charles Severance. Guido van Rossum: The early years of Python. Computer, 48(2):7–9, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/02/mco2015020007.html.

Gaff:2015:BO

[1434] Brian M. Gaff. BYOD? OMG! Computer, 48(2):10–11, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/02/mco2015020010.
html.

Anonymous:2015:YAb

[1435] Anonymous. 32 & 16 years ago. Computer, 48(2):12–13, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
02/mco2015020012.html.

Anonymous:2015:Nb

[1436] Anonymous. News. Computer, 48 (2):14–20, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/02/mco2015020014.html.

Weaver:2015:TAM

[1437] Alf Weaver and Renee Bryce. Technological advances in medicine: It’s personal [Guest Editors’ introduction]. Computer, 48(2):21–23, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
02/mco2015020021.html.

Li:2015:EPP

[1438] Jingquan Li. Ensuring privacy in a personal health record system. Computer, 48(2):24–31, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/02/mco2015020024-abs.html.

Velikova:2015:IDS

[1439] Marina Velikova, Peter J. F. Lucas, and Maarten van der Heijden. Intelligent disease self-management with mobile technology. Computer, 48 (2):32–39, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/02/mco2015020032-abs.html.

Anonymous:2015:IIO

[1440] Anonymous. Intuit, Inc. has openings for the following positions (by location) house advertisement. Computer, 48(2):40, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/02/mco2015020040.pdf.

Kang:2015:MGQ

[1441] Kyungtae Kang, Qixin Wang, Junbeom Hur, Kyung-Joon Park, and Lui Sha. Medical-grade quality of service for real-time mobile healthcare. Computer, 48(2):41–49, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
02/mco2015020041-abs.html.

Bahga:2015:HDI

[1442] Arshdeep Bahga and Vijay K. Madisetti. Healthcare data integration and informatics in the cloud. Computer, 48(2):50–57, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/02/mco2015020050-abs.html.

Ayday:2015:WGS

[1443] Erman Ayday, Emiliano De Cristofaro, Jean-Pierre Hubaux, and Gene Tsudik. Whole genome sequencing: Revolutionary medicine or privacy nightmare? Computer, 48(2): 58–66, February 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
02/mco2015020058-abs.html.

Spring:2015:CSC

[1444] Adam P. Spring. Creating substance from a cloud: Low-cost product generation. Computer, 48(2):67– 74, February 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
02/mco2015020067-abs.html.

Byrd:2015:SSI

[1445] Greg Byrd. Spotlighting student innovation. Computer, 48(2):75–76, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
02/mco2015020075.html.

Berghel:2015:CCS

[1446] Hal Berghel. Cyber chutzpah: The Sony hack and the celebration of hyperbole. Computer, 48(2):77–80, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        02/mco2015020077.html.

Vassev:2015:KKR

[1447] Emil Vassev and Mike Hinchey. KnowLang: Knowledge representation for self-adaptive systems. Computer, 48(2):81–84, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/02/mco2015020081-abs.html.

Hu:2015:ABA

[1448] Vincent C. Hu, D. Richard Kuhn, and David F. Ferraiolo. Attributebased access control. Computer, 48 (2):85–88, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/02/mco2015020085-abs.html.

Anonymous:2015:CCc

[1449] Anonymous. CS connection. Computer, 48(2):89–90, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2015/02/mco2015020089.
                                        html.

Anonymous:2015:CCd

[1450] Anonymous. Call and calendar. Computer, 48(2):91–92, February 2015.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/02/mco2015020091.
html.

Anonymous:2015:COb

[1451] Anonymous. Career opportunities. Computer, 48(2):93–99, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
02/mco2015020093.pdf.

Grier:2015:TG

[1452] David Alan Grier. The tyranny of geography. Computer, 48(2):100, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
02/mco2015020100.html.

Anonymous:2015:CPb

[1453] Anonymous. Call for papers.

Computer, 48(2):c2, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/02/mco20150200c2.
pdf.

Anonymous:2015:FCb

[1454] Anonymous. [Front cover]. Computer, 48(2):c1, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/02/mco20150200c1.pdf.

Anonymous:2015:IAIa

[1455] Anonymous. Instant access to IEEE publications house advertisement. Computer, 48(2):c3, February 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        02/mco20150200c3.pdf.

Anonymous:2015:SRSb

[1456] Anonymous. Startup rock stars [Advertisement]. Computer, 48(2): c4, February 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        02/mco20150200c4.pdf.

Anonymous:2015:Mc

[1457] Anonymous. Masthead. Computer, 48(3):3, March 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        03/mco2015030003.pdf.

Anonymous:2015:CHSc

[1458] Anonymous. Computer highlights Society magazines. Computer, 48 (3):4–5, March 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        03/mco2015030004.html.

Schuller:2015:DCP

[1459] Bjorn Schuller. Do computers have personality? Computer, 48(3):6– 7, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        03/mco2015030006.html.

Severance:2015:GVRb

[1460] Charles Severance. Guido van Rossum:

The modern era of Python. Computer, 48(3):8–10, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/03/mco2015030008.html.

Anonymous:2015:YAc

[1461] Anonymous. 32 & 16 years ago.

Computer, 48(3):11–12, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/03/mco2015030011.
html.

Garber:2015:N

[1462] Lee Garber. News. Computer, 48 (3):13–19, March 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
03/mco2015030013.html.

Gudivada:2015:BDP

[1463] Venkat N. Gudivada, Ricardo Baeza-Yates, and Vijay V. Raghavan. Big data: Promises and problems. Computer, 48(3):20–23, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/03/mco2015030020.
html.

Castellana:2015:MGD

[1464] Vito Giovanni Castellana, Alessandro Morari, Jesse Weaver, Antonino Tumeo, David Haglin, Oreste Villa, and John Feo. In-memory graph databases for Web-scale data. Computer, 48(3):24–35, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/03/mco2015030024-
abs.html.

Zhuang:2015:TRL

[1465] Zhenyun Zhuang, Haricharan Ramachandra, and Chaoyue Xiong. Taming replication latency of big data events with capacity planning. Computer, 48(3):36–41, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/03/mco2015030036-
abs.html.

Bansal:2015:IBD

[1466] Srividya K. Bansal and Sebastian Kagemann. Integrating big data: A semantic extract-transform-load framework. Computer, 48(3):42–50, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
03/mco2015030042-abs.html.

Anonymous:2015:FYJa

[1467] Anonymous. Focus on your job search house advertisement. Computer, 48 (3):51, March 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
03/mco2015030051.pdf.

Kaur:2015:MDH

[1468] Karamjit Kaur and Rinkle Rani. Managing data in healthcare information systems: Many models, one solution. Computer, 48(3):52–59, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
03/mco2015030052-abs.html.

Giese:2015:OZB

[1469] Martin Giese, Ahmet Soylu, Guillermo Vega-Gorgojo, Arild Waaler, Peter Haase, Ernesto Jimenez-Ruiz, Davide Lanti, Martin Rezk, Guohui Xiao, Ozgur Ozcep, and Riccardo Rosati. Optique: Zooming in on big data. Computer, 48(3):60–67, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
03/mco2015030060-abs.html.

Alkhatib:2015:WWL

[1470] Hasan Alkhatib, Paolo Faraboschi, Eitan Frachtenberg, Hironori Kasahara, Danny Lange, Phil Laplante, Arif Merchant, Dejan Milojicic, and Karsten Schwan. What will 2022 look like? the IEEE CS 2022 report. Computer, 48(3):68–76, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/03/mco2015030068.html.

Anonymous:2015:SCH

[1471] Anonymous. Stay connected house advertisement. Computer, 48(3):77, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
03/mco2015030077.pdf.

Thakoor:2015:PTC

[1472] Ninad S. Thakoor, Le An, Bir Bhanu, Santhoshkumar Sunderrajan, and B. S. Manjunath. People tracking in camera networks: Three open questions. Computer, 48(3):78–86, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        03/mco2015030078-abs.html.

Truong:2015:PEC

[1473] Hong-Linh Truong and Schahram Dustdar. Programming elasticity in the cloud. Computer, 48(3): 87–90, March 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        03/mco2015030087-abs.html.

Anonymous:2015:CPc

[1474] Anonymous. Call for papers. Computer, 48(3):91, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/03/mco2015030091.pdf.

Hornbaek:2015:FDR

[1475] Kasper Hornbaek. Flexible displays, rigid designs? Computer, 48(3): 92–96, March 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        03/mco2015030092-abs.html.

Schlieski:2015:VD

[1476] Tawny Schlieski. Virtual dystopia. Computer, 48(3):97–98, March 2015.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/03/mco2015030097.
html.

Michael:2015:TCE

[1477] James Bret Michael. Trusted computing: An elusive goal. Computer, 48(3):99–101, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/03/mco2015030099-abs.html.

Berghel:2015:N

[1478] Hal Berghel. Noirware. Computer, 48(3):102–107, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/03/mco2015030102.html.

Timmerer:2015:TNA

[1479] Christian Timmerer, Touradj Ebrahimi, and Fernando Pereira. Toward a new assessment of quality. Computer, 48 (3):108–110, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/03/mco2015030108-abs.html.

Aiello:2015:MHP

[1480] Bob Aiello. Managing the High-Performance Computer Society Standards Working Group. Computer, 48 (3):111–113, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/03/mco2015030111-abs.html.

Roach:2015:CCS

[1481] Steven Roach and Mehran Sahami. CS2013: Computer science curricula 2013. Computer, 48(3):114– 116, March 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        03/mco2015030114.html.

Anonymous:2015:CCe

[1482] Anonymous. Call and calendar. Computer, 48(3):117–119, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2015/03/mco2015030117.
                                        html.

Anonymous:2015:COc

[1483] Anonymous. Career opportunities. Computer, 48(3):120–135, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        03/mco2015030120.pdf.

Grier:2015:BT

[1484] David Alan Grier. Building trust. Computer, 48(3):136, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2015/03/mco2015030136.
                                        html.

Anonymous:2015:FCc

[1485] Anonymous. Front cover. Computer, 48(3):c1, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/03/mco20150300c1.pdf.

Anonymous:2015:IAIb

[1486] Anonymous. Instant access to IEEE publications advertisement. Computer, 48(3):c3, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/03/mco20150300c3.pdf.

Anonymous:2015:RSPb

[1487] Anonymous. Rock stars of 3D printing house advertisement. Computer, 48(3):c2, March 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
03/mco20150300c2.pdf.

Anonymous:2015:YDS

[1488] Anonymous. Are your developers spending time advertisement. Computer, 48(3):c4, March 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/03/mco20150300c4.pdf.

Anonymous:2015:Md

[1489] Anonymous. Masthead. Computer, 48(4):3, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040003.pdf.

Anonymous:2015:CHSd

[1490] Anonymous. Computer highlights society magazines. Computer, 48 (4):4–5, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040004.html.

Forsyth:2015:RL

[1491] David Forsyth. Representation learning. Computer, 48(4):6, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/04/mco2015040006.
html.

Gaff:2015:IICa

[1492] Brian M. Gaff. Important IP cases in 2015, part one. Computer, 48 (4):7–9, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040007.html.

Anonymous:2015:YAd

[1493] Anonymous. 32 & 16 years ago. Computer, 48(4):10–11, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/04/mco2015040010.
html.

Hsu:2015:CTS

[1494] D. Frank Hsu, Dorothy Marinucci, and Jeffrey M. Voas. Cybersecurity: Toward a secure and sustainable cyber ecosystem. Computer, 48 (4):12–14, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040012.html.

Anonymous:2015:GMLa

[1495] Anonymous. Get more, for less!

House advertisement. Computer, 48 (4):15, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040015.pdf.

Lee:2015:RCC

[1496] Ruby B. Lee. Rethinking computers for cybersecurity. Computer, 48 (4):16–25, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040016-abs.html.

Gillman:2015:PWA

[1497] David Gillman, Yin Lin, Bruce Maggs, and Ramesh K. Sitaraman. Protecting Websites from attack with secure delivery networks. Computer, 48(4):26–34, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/04/mco2015040026-abs.html.

Anonymous:2015:CPAa

[1498] Anonymous. Call for papers advertisement. Computer, 48(4):35, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040035.pdf.

Heckman:2015:DDC

[1499] Kristin E. Heckman, Frank J. Stech, Ben S. Schmoker, and Roshan K. Thomas. Denial and deception in cyber defense. Computer, 48(4): 36–44, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040036-abs.html.

Vagoun:2015:IFC

[1500] Tomas Vagoun and George O. Strawn. Implementing the Federal cybersecurity R&D strategy. Computer, 48 (4):45–55, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040045-abs.html.

Chai:2015:ESC

[1501] Sek Chai. Enabling smart camera networks with Smartphone processors. Computer, 48(4):56–63, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040056-abs.html.

Hagar:2015:ICT

[1502] Jon D. Hagar, Thomas L. Wissink, D. Richard Kuhn, and Raghu N. Kacker. Introducing combinatorial testing in a large organization. Computer, 48(4):64–72, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040064-abs.html.

Anonymous:2015:ICSa

[1503] Anonymous. IEEE Computer Society house advertisement. Computer, 48(4):73, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040073.pdf.

Byrd:2015:LIS

[1504] Greg Byrd. A little ingenuity solves an elephant-sized problem. Computer, 48(4):74–77, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/04/mco2015040074.html.

Shen:2015:CSC

[1505] Emily Shen, Mayank Varia, Robert K.

Cunningham, and W. Konrad Vesey. Cryptographically secure computation. Computer, 48(4):78–81, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/04/mco2015040078-
abs.html.

Berghel:2015:BED

[1506] Hal Berghel. Borderline executive disorder. Computer, 48(4):82– 86, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco2015040082.html.

Anonymous:2015:CCf

[1507] Anonymous. CS connection. Computer, 48(4):87–89, April 2015. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/csdl/mags/
co/2015/04/mco2015040087.html.

Anonymous:2015:CCg

[1508] Anonymous. Call and calendar. Computer, 48(4):90–93, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2015/04/mco2015040090.
                                        html.

Anonymous:2015:FYJb

[1509] Anonymous. Focus on your job search house advertisement. Computer, 48 (4):94, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        04/mco2015040094.pdf.

Anonymous:2015:COd

[1510] Anonymous. Career opportunities. Computer, 48(4):95–103, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2015/04/mco2015040095.
                                        pdf.

Grier:2015:AWH

[1511] David Alan Grier. Accidents will happen [The errant hashtag]. Computer, 48(4):104, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/04/mco2015040104.html.

Anonymous:2015:FCd

[1512] Anonymous. Front cover. Computer, 48(4):c1, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/04/mco20150400c1.pdf.

Anonymous:2015:IAIc

[1513] Anonymous. Instant access to IEEE publications advertisement. Computer, 48(4):c3, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/04/mco20150400c3.pdf.

Anonymous:2015:ISPa

[1514] Anonymous. IEEE Security & Privacy house advertisement. Computer, 48(4):c2, April 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
04/mco20150400c2.pdf.

Anonymous:2015:RSCa

[1515] Anonymous. Rock stars of cybersecurity house advertisement. Computer, 48(4):c4, April 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/04/mco20150400c4.pdf.

Anonymous:2015:Me

[1516] Anonymous. Masthead. Computer, 48(5):3, May 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050003.pdf.

Anonymous:2015:CHSe

[1517] Anonymous. Computer highlights Society magazines. Computer, 48 (5):4–5, May 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050004.html.

Liu:2015:ECS

[1518] Ling Liu. Enhancing cloud services through multitier workload analysis. Computer, 48(5):6, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/05/mco2015050006.html.

Severance:2015:JRB

[1519] Charles Severance. John Resig: Building JQuery. Computer, 48(5):7– 8, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        05/mco2015050007.html.

Gaff:2015:IICb

[1520] Brian M. Gaff. Important IP cases in 2015, part two. Computer, 48 (5):9–11, May 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        05/mco2015050009-abs.html.

Anonymous:2015:YAe

[1521] Anonymous. 32 & 16 years ago. Computer, 48(5):12–13, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/05/mco2015050012.html.

Vetter:2015:WNY

[1522] Ron Vetter and San Murugesan. The Web: The next 25 years. Computer, 48(5):14–17, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/05/mco2015050014.html.

Pendyala:2015:WEB

[1523] Vishnu S. Pendyala, Simon S. Y. Shim, and Christoph Bussler. The Web that extends beyond the world. Computer, 48(5):18–25, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/05/mco2015050018-abs.html.

Raggett:2015:WTC

[1524] Dave Raggett. The Web of things:

Challenges and opportunities. Computer, 48(5):26–32, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/05/mco2015050026-abs.html.

Anonymous:2015:CCP

[1525] Anonymous. Compsac call for papers house advertisement. Computer, 48(5):33, May 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050033.pdf.

Heuer:2015:TWT

[1526] Jorg Heuer, Johannes Hund, and Oliver Pfaff. Toward the Web of Things: Applying Web technologies to the physical world. Computer, 48 (5):34–42, May 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050034-abs.html.

Anonymous:2015:SCPa

[1527] Anonymous. STC 2015: Call for presentations house advertisement. Computer, 48(5):43, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/05/mco2015050043.pdf.

Vetter:2015:RDW

[1528] Ron Vetter and San Murugesan. Roundtable discussion: The Web we look forward to. Computer, 48(5): 44–50, May 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050044.html.

Riehle:2015:HOS

[1529] Dirk Riehle. How open source is changing the software developer’s career. Computer, 48(5):51–57, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050051-abs.html.

Schlessman:2015:TDE

[1530] Jason Schlessman and Marilyn Wolf. Tailoring design for embedded computer vision applications. Computer, 48(5):58–62, May 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050058-abs.html.

Anonymous:2015:FYJc

[1531] Anonymous. Focus on your job search house advertisement. Computer, 48 (5):63, May 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050063.pdf.

DeLuca:2015:SUS

[1532] Alexander De Luca and Janne Lindqvist. Is secure and usable Smartphone authentication asking too much? Computer, 48(5):64– 68, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050064-abs.html.

Draudt:2015:SIA

[1533] Alida Draudt, Jonjozuf Hadley, Ryan Hogan, Leticia Murray, Gregory Stock, and Julia Rose West. Six insights about science fiction prototyping. Computer, 48(5):69–71, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/05/mco2015050069.html.

Zhou:2015:OPR

[1534] Wenqi Zhou and Yuhong Liu. Online product rating manipulation and market performance. Computer, 48 (5):72–75, May 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050072-abs.html.

Hambrusch:2015:PP

[1535] Susanne Hambrusch and Ran Libeskind-Hadas. The PhD pipeline. Computer, 48(5):76–79, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/05/mco2015050076-abs.html.

Zhanikeev:2015:CVP

[1536] Marat Zhanikeev. A cloud visitation platform to facilitate cloud federation and fog computing. Computer, 48(5):80–83, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/05/mco2015050080-abs.html.

Berghel:2015:WMG

[1537] Hal Berghel. What makes a good column. Computer, 48(5):84–87, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        05/mco2015050084.html.

Hurlburt:2015:CS

[1538] George Hurlburt. Cyberhuman security. Computer, 48(5):88–91, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        05/mco2015050088-abs.html.

Gundy-Burlet:2015:USL

[1539] Karen Gundy-Burlet. The use of standards on the LADEE mission. Computer, 48(5):92–95, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/05/mco2015050092-abs.html.

Broy:2015:RNS

[1540] Manfred Broy. Rethinking nonfunctional software requirements. Computer, 48(5):96–99, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/05/mco2015050096-abs.html.

Anonymous:2015:CCh

[1541] Anonymous. Call and calendar.

Computer, 48(5):100–105, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/05/mco2015050100.
html.

Anonymous:2015:COe

[1542] Anonymous. Career opportunities. Computer, 48(5):106–115, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco2015050106.pdf.

Grier:2015:GE

[1543] David Alan Grier. The GitHub effect. Computer, 48(5):116, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/05/mco2015050116.
html.

Anonymous:2015:FCe

[1544] Anonymous. Front cover. Computer, 48(5):c1, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/05/mco20150500c1.pdf.

Anonymous:2015:GMLb

[1545] Anonymous. Get more, for less!

House advertisement. Computer, 48 (5):c2, May 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco20150500c2.pdf.

Anonymous:2015:LMHa

[1546] Anonymous. Learn more house advertisement. Computer, 48(5): c3, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
05/mco20150500c3.pdf.

Anonymous:2015:RSWa

[1547] Anonymous. Rock stars of wearables house advertisement. Computer, 48(5):c4, May 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/05/mco20150500c4.pdf.

Anonymous:2015:Mf

[1548] Anonymous. Masthead. Computer, 48(6):3, June 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
06/mco2015060003.pdf.

Anonymous:2015:CHSf

[1549] Anonymous. Computer highlights Society magazines. Computer, 48 (6):4–5, June 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
06/mco2015060004.html.

Mohapatra:2015:DOC

[1550] Prasant Mohapatra. Distributed optimal channel access in cognitive radio networks. Computer, 48(6): 6, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
06/mco2015060006.html.

Severance:2015:RFU

[1551] Charles Severance. Roy T. Fielding:

Understanding the REST style. Computer, 48(6):7–9, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/06/mco2015060007.html.

Anonymous:2015:YAf

[1552] Anonymous. 32 & 16 years ago. Computer, 48(6):10–11, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/06/mco2015060010.html.

Starner:2015:WCN

[1553] Thad Starner and Tom Martin. Wearable computing: The new dress code [Guest Editors’ introduction]. Computer, 48(6):12–15, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/06/mco2015060012.html.

Guo:2015:OPH

[1554] Anhong Guo, Xiaolong Wu, Zhengyang Shen, Thad Starner, Hannes Baumann, and Scott Gilliland. Order picking with head-up displays. Computer, 48(6):16–24, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/06/mco2015060016-abs.html.

Kirkham:2015:CDD

[1555] Reuben Kirkham. Can disability discrimination law expand the availability of wearable computers? Computer, 48(6):25–33, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/06/mco2015060025-abs.html.

Hammer:2015:IMU

[1556] Jon C. Hammer and Tingxin Yan. Inferring mobile user status with usage cues. Computer, 48(6):34– 40, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
06/mco2015060034-abs.html.

Rosales:2015:BSC

[1557] Andrea Rosales, Sergio Sayago, and Josep Blat. Beeping socks and chirping arm bands: Wearables that foster free play. Computer, 48(6): 41–48, June 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
06/mco2015060041-abs.html.

Anonymous:2015:SCPb

[1558] Anonymous. STC 2015: Call for presentations house advertisement. Computer, 48(6):49, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/06/mco2015060049.pdf.

Ullah:2015:SBR

[1559] Najeeb Ullah, Maurizio Morisio, and Antonio Vetro. Selecting the best reliability model to predict residual defects in open source software. Computer, 48(6):50–58, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/06/mco2015060050-abs.html.

Anonymous:2015:FYJd

[1560] Anonymous. Focus on your job search house advertisement. Computer, 48 (6):59, June 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
06/mco2015060059.pdf.

Byrd:2015:TCW

[1561] Greg Byrd. Tracking cows wirelessly. Computer, 48(6):60–63, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
06/mco2015060060.html.

Berghel:2015:FAGa

[1562] Hal Berghel. A farewell to air gaps, Part 1. Computer, 48(6): 64–68, June 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
06/mco2015060064.html.

Voas:2015:VMA

[1563] Jeffrey Voas. Vetting mobile app vendors. Computer, 48(6):69–71, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        06/mco2015060069-abs.html.

Hahnle:2015:DRA

[1564] Reiner Hahnle and Einar Broch Johnsen. Designing resource-aware cloud applications. Computer, 48 (6):72–75, June 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        06/mco2015060072-abs.html.

Anonymous:2015:CCi

[1565] Anonymous. CS connection. Computer, 48(6):76–77, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/06/mco2015060076.html.

Anonymous:2015:CCj

[1566] Anonymous. Call and calendar. Computer, 48(6):78–82, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/06/mco2015060078.html.

Anonymous:2015:ICSb

[1567] Anonymous. IEEE Computer Society house advertisement. Computer, 48(6):83, June 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        06/mco2015060083.pdf.

Anonymous:2015:COf

[1568] Anonymous. Career opportunities. Computer, 48(6):84–98, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/06/mco2015060084.
pdf.

Anonymous:2015:AIC

[1569] Anonymous. 39th Annual International Computers, Software & Applications Conference advertisement. Computer, 48(6):99, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/06/mco2015060099.pdf.

Grier:2015:DC

[1570] David Alan Grier. Disruptive computing. Computer, 48(6):100, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
06/mco2015060100.html.

Anonymous:2015:FCf

[1571] Anonymous. Front cover. Computer, 48(6):c1, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/06/mco20150600c1.pdf.

Anonymous:2015:GMLc

[1572] Anonymous. Get more, for less!

House advertisement. Computer, 48 (6):c2, June 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
06/mco20150600c2.pdf.

Anonymous:2015:LMHb

[1573] Anonymous. Learn more house advertisement. Computer, 48(6): c3, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        06/mco20150600c3.pdf.

Anonymous:2015:RSWb

[1574] Anonymous. Rock stars of wearables house advertisement. Computer, 48(6):c4, June 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/06/mco20150600c4.pdf.

Anonymous:2015:Mg

[1575] Anonymous. Masthead. Computer, 48(7):3, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        07/mco2015070003.pdf.

Gaudiot:2015:TTC

[1576] Jean-Luc Gaudiot. The times they are a-changin’. Computer, 48(7):4– 6, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        07/mco2015070004.html.

Anonymous:2015:CPCa

[1577] Anonymous. Call for participation and celebration house advertisement. Computer, 48(7):7, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/07/mco2015070007.pdf.

Anonymous:2015:CHSg

[1578] Anonymous. Computer highlights Society magazines. Computer, 48 (7):8–9, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
07/mco2015070008.html.

Keckler:2015:IIN

[1579] Stephen W. Keckler. Increasing interconnection network throughput with virtual channels. Computer, 48(7): 10, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
07/mco2015070010.html.

Anonymous:2015:SA

[1580] Anonymous. Symantec advertisement. Computer, 48(7):11, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/07/mco2015070011.
pdf.

Gaff:2015:CCF

[1581] Brian M. Gaff. “Curiouser and curiouser!” — the fallout from Alice. Computer, 48(7):12–14, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/07/mco2015070012-abs.html.

Anonymous:2015:ICSc

[1582] Anonymous. IEEE Computer Society house advertisement. Computer, 48(7):15, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
07/mco2015070015.pdf.

Anonymous:2015:YAg

[1583] Anonymous. 32 & 16 years ago. Computer, 48(7):16–17, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/07/mco2015070016.html.

Torresen:2015:SAS

[1584] Jim Torresen, Christian Plessl, and Xin Yao. Self-aware and selfexpressive systems. Computer, 48 (7):18–20, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
07/mco2015070018.html.

Rinner:2015:SAS

[1585] Bernhard Rinner, Lukas Esterle, Jennifer Simonjan, Georg Nebehay, Roman Pflugfelder, Gustavo Fernandez Dominguez, and Peter R. Lewis. Self-aware and self-expressive camera networks. Computer, 48(7):21– 28, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
07/mco2015070021-abs.html.

Amoretti:2015:TCS

[1586] Michele Amoretti and Stefano Cagnoni. Toward collective self-awareness and self-expression in distributed systems. Computer, 48(7):29–36, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
07/mco2015070029-abs.html.

Preden:2015:BSA

[1587] Jurgo S. Preden, Kalle Tammemae, Axel Jantsch, Mairo Leier, Andri Riid, and Emine Calis. The benefits of self-awareness and attention in fog and mist computing. Computer, 48 (7):37–45, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
07/mco2015070037-abs.html.

vanBeek:2015:SEM

[1588] Vincent van Beek, Jesse Donkervliet, Tim Hegeman, Stefan Hugtenburg, and Alexandru Iosup. Self-expressive management of business-critical workloads in virtualized datacenters. Computer, 48(7):46–54, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/07/mco2015070046-abs.html.

Parhami:2015:ARL

[1589] Behrooz Parhami. Amdahl’s Reliability Law: A simple quantification of the weakest-link phenomenon. Computer, 48(7):55–58, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/07/mco2015070055-abs.html.

Berghel:2015:FAGb

[1590] Hal Berghel. A farewell to air gaps, Part 2. Computer, 48(7): 59–63, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
07/mco2015070059.html.

Kshetri:2015:RUC

[1591] Nir Kshetri. Recent US cybersecurity policy initiatives: Challenges and implications. Computer, 48(7): 64–69, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        07/mco2015070064-abs.html.

Johnson:2015:SSR

[1592] Brian David Johnson. Steve the Steampunk Robot. Computer, 48 (7):70–72, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        07/mco2015070070.html.

Montalvo:2015:MIA

[1593] Soto Montalvo, Raquel Martinez, Victor Fresno, and Rafael Capilla. Multilingual information access on the Web. Computer, 48(7):73– 75, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        07/mco2015070073-abs.html.

Rao:2015:CEE

[1594] Dhananjai M. Rao. Computing efficiency in education. Computer, 48 (7):76–79, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        07/mco2015070076-abs.html.

Xu:2015:BDB

[1595] Xiaofei Xu, Quan Z. Sheng, Liang-Jie Zhang, Yushun Fan, and Schahram Dustdar. From big data to big service. Computer, 48(7):80–83, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
07/mco2015070080-abs.html.

Kristensson:2015:NGT

[1596] Per Ola Kristensson. Next-generation text entry. Computer, 48(7):84– 87, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
07/mco2015070084-abs.html.

Anonymous:2015:CCk

[1597] Anonymous. CS connection. Computer, 48(7):88–89, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/07/mco2015070088.html.

Anonymous:2015:CCl

[1598] Anonymous. Call and calendar. Computer, 48(7):90–93, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/07/mco2015070090.html.

Anonymous:2015:COg

[1599] Anonymous. Career opportunities.

Computer, 48(7):94–103, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/07/mco2015070094.
pdf.

Grier:2015:DWN

[1600] David Alan Grier. Do we need a software czar? Computer, 48 (7):104, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        07/mco2015070104.html.

Anonymous:2015:FCg

[1601] Anonymous. Front cover. Computer, 48(7):c1, July 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/07/mco20150700c1.pdf.

Anonymous:2015:MYCa

[1602] Anonymous. Move your career forward [Advertisement]. Computer, 48 (7):c4, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        07/mco20150700c4.pdf.

Anonymous:2015:RSWc

[1603] Anonymous. Rock stars of wearables [Advertisement]. Computer, 48 (7):c2, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        07/mco20150700c2.pdf.

Anonymous:2015:SCPc

[1604] Anonymous. STC 2015: Call for presentations learn more. Computer, 48(7):c3, July 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        07/mco20150700c3.pdf.

Anonymous:2015:Mh

[1605] Anonymous. Masthead. Computer, 48(8):3, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080003.pdf.

Anonymous:2015:CHSh

[1606] Anonymous. Computer highlights Society magazines. Computer, 48(8): 4–5, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080004.html.

Liu:2015:MNE

[1607] Shixia Liu and Leila De Floriani.

Multivariate network exploration and presentations. Computer, 48(8):6, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080006.html.

Anonymous:2015:MYCb

[1608] Anonymous. Move your career forward house advertisement. Computer, 48(8):7, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/08/mco2015080007.pdf.

Severance:2015:AJY

[1609] Charles Severance. Anil Jain: 25 years of biometric recognition. Computer, 48(8):8–10, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/08/mco2015080008.html.

Anonymous:2015:ICSd

[1610] Anonymous. IEEE Computer Society [Advertisement]. Computer, 48 (8):11, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080011.pdf.

Anonymous:2015:YAh

[1611] Anonymous. 32 & 16 years ago. Computer, 48(8):12–13, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080012.html.

Tumeo:2015:IAA

[1612] Antonino Tumeo and John Feo. Irregular applications: From architectures to algorithms [Guest Editors’ introduction]. Computer, 48(8):14– 16, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080014.html.

Anonymous:2015:RMD

[1613] Anonymous. 2016 Richard E. Merwin Distinguished Service Award house advertisement. Computer, 48(8):17, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080017.pdf.

Lloyd:2015:MDR

[1614] Scott Lloyd and Maya Gokhale.

In-memory data rearrangement for irregular, data-intensive computing. Computer, 48(8):18–25, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080018-abs.html.

Buono:2015:OSL

[1615] Daniele Buono, John A. Gunnels, Xinyu Que, Fabio Checconi, Fabrizio Petrini, Tai-Ching Tuan, and Chris Long. Optimizing sparse linear algebra for large-scale graph analytics. Computer, 48(8):26–34, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080026-abs.html.

Lenharth:2015:SRI

[1616] Andrew Lenharth and Keshav Pingali. Scaling runtimes for irregular algorithms to large-scale NUMA systems. Computer, 48(8):35–44, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080035-abs.html.

Anonymous:2015:RSWd

[1617] Anonymous. Rock stars of wearables house advertisement. Computer, 48 (8):45, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080045.pdf.

Halappanavar:2015:CLL

[1618] Mahantesh Halappanavar, Alex Pothen, Ariful Azad, Fredrik Manne, Johannes Langguth, and Arif Khan. Codesign lessons learned from implementing graph matching on multithreaded architectures. Computer, 48(8):46– 55, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        08/mco2015080046-abs.html.

Rajput:2015:AVL

[1619] Hitesh Rajput, Tanmoy Som, and Soumitra Kar. An automated vehicle license plate recognition system. Computer, 48(8):56–61, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        08/mco2015080056-abs.html.

Lewis:2015:AAS

[1620] Peter R. Lewis, Arjun Chandra, Funmilade Faniyi, Kyrre Glette, Tao Chen, Rami Bahsoon, Jim Torresen, and Xin Yao. Architectural aspects of self-aware and self-expressive computing systems: From psychology to engineering. Computer, 48 (8):62–70, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/08/mco2015080062-abs.html.

Anonymous:2015:RMS

[1621] Anonymous. Richard E. Merwin Scholarship house advertisement. Computer, 48(8):71, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/08/mco2015080071.pdf.

Byrd:2015:CTC

[1622] Greg Byrd. Cycling through cyberspace. Computer, 48(8):72–75, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080072.html.

Bruneliere:2015:SMR

[1623] Hugo Bruneliere, Jordi Cabot, Javier Luis Canovas Izquierdo, Leire Orue-Echevarria Arrieta, Oliver Strauss, and Manuel Wimmer. Software modernization revisited: Challenges and prospects. Computer, 48(8):76–80, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080076-abs.html.

Zhang:2015:IAI

[1624] Zhi-Kai Zhang, Michael Cheng Yi Cho, Zong-Yu Wu, and Shiuhpyng Winston Shieh. Identifying and authenticating IoT objects in a natural context. Computer, 48(8):81– 83, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080081-abs.html.

Davis:2015:TEI

[1625] Darold Davis. Taming the engineering of information services Websites with standards. Computer, 48(8): 84–89, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        08/mco2015080084.html.

Berghel:2015:TMC

[1626] Hal Berghel. TSA: Mission creep meets waste. Computer, 48(8):90– 94, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        08/mco2015080090.html.

Anonymous:2015:PPC

[1627] Anonymous. PerCom Pervasive Computing 2016 house advertisement. Computer, 48(8):95, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        08/mco2015080095.pdf.

Anonymous:2015:ICSe

[1628] Anonymous. IEEE Computer Society election. Computer, 48(8):96– 106, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        08/mco2015080096.html.

Anonymous:2015:CHAa

[1629] Anonymous. COMPSAC 2016 house advertisement. Computer, 48(8): 107, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        08/mco2015080107.pdf.

Anonymous:2015:IPE

[1630] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 48(8):108– 111, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080108.html.

Anonymous:2015:CCm

[1631] Anonymous. Call and calendar. Computer, 48(8):112–115, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/08/mco2015080112.
html.

Anonymous:2015:COh

[1632] Anonymous. Career opportunities. Computer, 48(8):116–126, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080116.pdf.

Anonymous:2015:DOH

[1633] Anonymous. Deadline: 31 October house advertisement. Computer, 48 (8):127, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
08/mco2015080127.pdf.

Grier:2015:SDD

[1634] David Alan Grier. In search of the diabolical disconnect. Computer, 48 (8):128, August 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        08/mco2015080128.html.

Anonymous:2015:CPCb

[1635] Anonymous. Call for participation and celebration house advertisement. Computer, 48(8):c3, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/08/mco20150800c3.pdf.

Anonymous:2015:CPH

[1636] Anonymous. Call for papers house advertisement. Computer, 48(8):c2, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        08/mco20150800c2.pdf.

Anonymous:2015:FCh

[1637] Anonymous. Front cover. Computer, 48(8):c1, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/08/mco20150800c1.pdf.

Anonymous:2015:IIP

[1638] Anonymous. 30th IEEE International Parallel and Distributed Processing Symposium house advertisement. Computer, 48(8):c4, August 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        08/mco20150800c4.pdf.

Anonymous:2015:Mi

[1639] Anonymous. Masthead. Computer, 48(9):3, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/09/mco2015090003.pdf.

Anonymous:2015:CHSi

[1640] Anonymous. Computer highlights Society magazines. Computer, 48(9): 4–5, September 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
09/mco2015090004.html.

Helal:2015:HYL

[1641] Sumi Helal. However you like it:

Computer’s enhanced digital format or affordable print option. Computer, 48(9):6, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/09/mco2015090006.html.

Anonymous:2015:FAH

[1642] Anonymous. The Future’s algorithm is here advertisement. Computer, 48(9):7, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/09/mco2015090007.pdf.

Bandeira:2015:EAE

[1643] Afonso S. Bandeira. An efficient algorithm for exact recovery of vertex variables from edge measurements. Computer, 48(9):8–9, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
09/mco2015090008.html.

Gaff:2015:LIW

[1644] Brian M. Gaff. Legal issues with wearable technology. Computer, 48 (9):10–12, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/09/mco2015090010-abs.html.

Anonymous:2015:RSCb

[1645] Anonymous. Rock stars of cybersecurity [Advertisement]. Computer, 48(9):13, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/09/mco2015090013.pdf.

Anonymous:2015:YAi

[1646] Anonymous. 32 & 16 years ago. Computer, 48(9):14–15, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco2015090014.html.

Want:2015:AIT

[1647] Roy Want and Schahram Dustdar. Activating the Internet of Things [Guest Editors’ introduction]. Computer, 48 (9):16–20, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/09/mco2015090016.html.

Anonymous:2015:FYJe

[1648] Anonymous. Focus on your job search [Advertisement]. Computer, 48(9): 21, September 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
09/mco2015090021.pdf.

Beal:2015:API

[1649] Jacob Beal, Danilo Pianini, and Mirko Viroli. Aggregate programming for the Internet of Things. Computer, 48 (9):22–30, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/09/mco2015090022-abs.html.

Anonymous:2015:ICSf

[1650] Anonymous. IEEE Computer Society 2015 call for major award nominations. Computer, 48(9):31, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
09/mco2015090031.pdf.

Belli:2015:DDI

[1651] Laura Belli, Simone Cirani, Luca Davoli, Andrea Gorrieri, Mirko Mancin, Marco Picone, and Gianluigi Ferrari. Design and deployment of an IoT application-oriented testbed. Computer, 48(9):32–40, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
09/mco2015090032-abs.html.

Anonymous:2015:ICSg

[1652] Anonymous. IEEE Computer Society Richard E. Merwin Student Leadership Scholarship. Computer, 48(9): 41, September 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco2015090041.pdf.

Mikusz:2015:RWA

[1653] Mateusz Mikusz, Sarah Clinch, Rachel Jones, Mike Harding, Christopher Winstanley, and Nigel Davies. Repurposing Web analytics to support the IoT. Computer, 48(9):42– 49, September 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco2015090042-abs.html.

Shim:2015:EMM

[1654] Keun Sup Shim, Mieszko Lis, Omer Khan, and Srinivas Devadas. The execution migration machine: Directoryless shared-memory architecture. Computer, 48(9):50–59, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco2015090050-abs.html.

Gessler:2015:SET

[1655] Damian D. G. Gessler and Evren Sirin. SSWAP: Enabling transactiontime reasoning for semantic workflows. Computer, 48(9):60–68, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco2015090060-abs.html.

Parssian:2015:QBS

[1656] Amir Parssian, William Yeoh, and Mong Shan Ee. Quality-based SQL: Specifying information quality in relational database queries. Computer, 48 (9):69–74, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/09/mco2015090069-abs.html.

Berghel:2015:SC

[1657] Hal Berghel. STEM crazy. Computer, 48(9):75–80, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/09/mco2015090075.html.

Anonymous:2015:SE

[1658] Anonymous. Software experts. Computer, 48(9):81, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/09/mco2015090081.
pdf.

Brown:2015:PS

[1659] Steve Brown. The power of story. Computer, 48(9):82–83, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
09/mco2015090082.html.

Kolias:2015:SMT

[1660] Constantinos Kolias, Angelos Stavrou, and Jeffrey Voas. Securely making “things” right. Computer, 48(9):84– 88, September 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
09/mco2015090084-abs.html.

Anonymous:2015:ICC

[1661] Anonymous. IEEE Cloud Computing call for papers. Computer, 48 (9):89, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/09/mco2015090089.pdf.

Klamma:2015:NRT

[1662] Ralf Klamma. Near-real-time social computing. Computer, 48(9):90– 92, September 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco2015090090-abs.html.

Chen:2015:TSC

[1663] Tao Chen and Rami Bahsoon. Toward a smarter cloud: Self-aware autoscaling of cloud configurations and resources. Computer, 48(9):93– 96, September 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco2015090093-abs.html.

Anonymous:2015:ISPb

[1664] Anonymous. IEEE Security & Privacy [Advertisement]. Computer, 48 (9):97, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/09/mco2015090097.pdf.

Gary:2015:PBL

[1665] Kevin Gary. Project-based learning. Computer, 48(9):98–100, September 2015. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
09/mco2015090098-abs.html.

Anonymous:2015:LCA

[1666] Anonymous. Linkedin Corp. advertisement. Computer, 48(9):101, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
09/mco2015090101.pdf.

Durant:2015:CGF

[1667] Eric Durant. CE2016: Guidelines for forward-looking computer engineering curricula. Computer, 48(9): 102–104, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/09/mco2015090102-abs.html.

Anonymous:2015:CCn

[1668] Anonymous. CS connection. Computer, 48(9):105–106, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/09/mco2015090105.
html.

Anonymous:2015:CCo

[1669] Anonymous. Call and calendar.

Computer, 48(9):107–111, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
09/mco2015090107.html.

Anonymous:2015:COi

[1670] Anonymous. Career opportunities. Computer, 48(9):112–119, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco2015090112.pdf.

Grier:2015:IL

[1671] David Alan Grier. Industrial leadership. Computer, 48(9):120, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco2015090120.html.

Anonymous:2015:CPCc

[1672] Anonymous. Call for participation and celebration. Computer, 48(9): c3, September 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco20150900c3.pdf.

Anonymous:2015:FCi

[1673] Anonymous. Front cover. Computer, 48(9):c1, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/09/mco20150900c1.pdf.

Anonymous:2015:LMAa

[1674] Anonymous. Learn more [Advertisement]. Computer, 48(9):c4, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        09/mco20150900c4.pdf.

Anonymous:2015:MYCc

[1675] Anonymous. Move your career forward IEEE Computer Society membership [Advertisement]. Computer, 48(9):c2, September 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/09/mco20150900c2.pdf.

Anonymous:2015:Mj

[1676] Anonymous. Masthead. Computer, 48(10):3, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/10/mco2015100003.pdf.

Anonymous:2015:CHSj

[1677] Anonymous. Computer highlights Society magazines. Computer, 48(10): 4–5, October 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100004.html.

Montuschi:2015:SNN

[1678] Paolo Montuschi. Spiking neural network architecture. Computer, 48 (10):6, October 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100006.html.

Anonymous:2015:WWL

[1679] Anonymous. Watch the world’s leading experts take multi-core strategies to new heights house advertisement. Computer, 48(10):7, October 2015. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100007.pdf.

Severance:2015:AAI

[1680] Charles Severance. Anant Agarwal: Inside edX. Computer, 48(10): 8–9, October 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100008.html.

Anonymous:2015:YAj

[1681] Anonymous. 32 & 16 years ago. Computer, 48(10):10–11, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100010.html.

Jacucci:2015:PC

[1682] Giulio Jacucci, Stephen Fairclough, and Erin T. Solovey. Physiological computing. Computer, 48(10): 12–16, October 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100012.html.

Anonymous:2015:ICSh

[1683] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. Computer, 48 (10):17, October 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100017.pdf.

Rozado:2015:CEP

[1684] David Rozado and Andreas Dunser.

Combining EEG with pupillometry to improve cognitive workload detection. Computer, 48(10):18–25, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100018-abs.html.

Sioni:2015:SDU

[1685] Riccardo Sioni and Luca Chittaro.

Stress detection using physiological sensors. Computer, 48(10):26–33, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100026-abs.html.

Peck:2015:DBC

[1686] Evan M. Peck, Emily Carlin, and Robert Jacob. Designing braincomputer interfaces for attentionaware systems. Computer, 48(10): 34–42, October 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100034-abs.html.

Gudivada:2015:USE

[1687] Venkat N. Gudivada, Dhana Rao, and Jordan Paris. Understanding searchengine optimization. Computer, 48 (10):43–52, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/10/mco2015100043-abs.html.

Anonymous:2015:CSIa

[1688] Anonymous. Call for special issue proposals house advertisement. Computer, 48(10):53, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/10/mco2015100053.pdf.

Bedri:2015:TSS

[1689] Abdelkareem Bedri, Himanshu Sahni, Pavleen Thukral, Thad Starner, David Byrd, Peter Presti, Gabriel Reyes, Maysam Ghovanloo, and Zehua Guo. Toward silent-speech control of consumer wearables. Computer, 48(10): 54–62, October 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100054-abs.html.

Anonymous:2015:CSA

[1690] Anonymous. Call for standards award nominations house advertisement. Computer, 48(10):63, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100063.pdf.

Blake:2015:UIS

[1691] Madison Blake, Rabih Younes, Jacob Dennis, Thomas L. Martin, and Mark Jones. A user-independent and sensor-tolerant wearable activity classifier. Computer, 48(10):64– 71, October 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100064-abs.html.

Berghel:2015:LTB

[1692] Hal Berghel. Legislating technology (Badly). Computer, 48(10):72– 78, October 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100072.html.

Anonymous:2015:CWUa

[1693] Anonymous. Computer write for us house advertisement. Computer, 48 (10):79, October 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100079.pdf.

Byrd:2015:CP

[1694] Greg Byrd. 21st century Pong. Computer, 48(10):80–84, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/10/mco2015100080.
html.

Mens:2015:ESE

[1695] Tom Mens and Philippe Grosjean.

The ecology of software ecosystems. Computer, 48(10):85–87, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100085-abs.html.

Wang:2015:CSP

[1696] Chia-Wei Wang, Michael Cheng Yi Cho, Chi-Wei Wang, and Shiuhpyng Winston Shieh. Combating software piracy in public clouds. Computer, 48(10):88–91, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100088-abs.html.

Anonymous:2015:CCq

[1697] Anonymous. CS connection. Computer, 48(10):92–93, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/10/mco2015100092.
html.

Anonymous:2015:CCCa

[1698] Anonymous. Cloud computing call for papers house advertisement. Computer, 48(10):94, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100094.pdf.

Anonymous:2015:CCr

[1699] Anonymous. Call and calendar. Computer, 48(10):95–98, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2015/10/mco2015100095.
html.

Anonymous:2015:UPE

[1700] Anonymous. Upsilon Pi Epsilon Student Excellence Award house advertisement. Computer, 48(10):99, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100099.pdf.

Anonymous:2015:COj

[1701] Anonymous. Career opportunities. Computer, 48(10):100–118, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100100.pdf.

Anonymous:2015:ICSi

[1702] Anonymous. IEEE Computer Society house advertisement. Computer, 48(10):119, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2015/10/mco2015100119.pdf.

Grier:2015:VD

[1703] David Alan Grier. The voracious discipline. Computer, 48(10):120, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco2015100120.html.

Anonymous:2015:ELB

[1704] Anonymous. Explore levels for beyond the next house advertisement. Computer, 48(10):c3, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2015/
10/mco20151000c3.pdf.

Anonymous:2015:EP

[1705] Anonymous. Endless possibilities. Computer, 48(10):c2, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        10/mco20151000c2.pdf.

Anonymous:2015:FCj

[1706] Anonymous. Front cover. Computer, 48(10):c1, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2015/10/mco20151000c1.pdf.

Anonymous:2015:LMHc

[1707] Anonymous. Learn more house advertisement. Computer, 48(10):c4, October 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2015/
                                        10/mco20151000c4.pdf.

Anonymous:2015:Mk

[1708] Anonymous. Masthead. Computer, 48(11):3, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/11/mco2015110003.pdf.

Anonymous:2015:CHSk

[1709] Anonymous. Computer highlights Society magazines. Computer, 48(11): 4–5, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        11/mco2015110004.html.

Sharples:2015:SLU

[1710] Mike Sharples. Seamless learning: Using location-aware technology to support art education. Computer, 48 (11):6, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110006.html.

Gaff:2015:PID

[1711] Brian M. Gaff. Patent infringement damages. Computer, 48(11): 7–9, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110007-abs.html.

Anonymous:2015:YAk

[1712] Anonymous. 32 & 16 years ago.

Computer, 48(11):10–11, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110010.html.

Getov:2015:SGC

[1713] Vladimir Getov. Scientific grand challenges: Toward exascale supercomputing and beyond. Computer, 48 (11):12–14, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2015/11/mco2015110012.html.

Miyoshi:2015:BED

[1714] Takemasa Miyoshi, Keiichi Kondo, and Koji Terasaki. Big ensemble data assimilation in numerical weather prediction. Computer, 48(11):15– 21, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110015-abs.html.

Anonymous:2015:CHAb

[1715] Anonymous. COMPSAC 2016 house advertisement. Computer, 48(11): 22, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        11/mco2015110022.pdf.

Keon:2015:POS

[1716] Dylan Keon, Cherri M. Pancake, and Harry Yeh. Protecting our shorelines: Modeling the effects of tsunamis and storm waves. Computer, 48(11):23– 32, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        11/mco2015110023-abs.html.

Romero:2015:QMD

[1717] Nichols A. Romero, Aiichiro Nakano, Katherine M. Riley, Fuyuki Shimojo, Rajiv K. Kalia, Priya Vashishta, and Paul C. Messina. Quantum molecular dynamics in the postpetaflops era. Computer, 48(11):33– 41, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        11/mco2015110033-abs.html.

Strohmaier:2015:TLP

[1718] Erich Strohmaier, Hans W. Meuer, Jack Dongarra, and Horst D. Simon. The TOP500 list and progress in highperformance computing. Computer, 48 (11):42–49, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/11/mco2015110042-abs.html.

Zwart:2015:UGE

[1719] Simon Portegies Zwart and Jeroen Bedorf. Using GPUs to enable simulation with computational gravitational dynamics in astrophysics. Computer, 48(11):50–58, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2015/11/mco2015110050-abs.html.

Anonymous:2015:GRY

[1720] Anonymous. Get the recognition you deserve house advertisement. Computer, 48(11):59, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110059.pdf.

Valdi:2015:STM

[1721] Andrea Valdi, Eros Lever, Simone Benefico, Davide Quarta, Stefano Zanero, and Federico Maggi. Scalable testing of mobile antivirus applications. Computer, 48(11):60– 68, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110060-abs.html.

Anonymous:2015:MYCd

[1722] Anonymous. Move your career forward house advertisement. Computer, 48(11):69, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2015/11/mco2015110069.pdf.

Liu:2015:HTG

[1723] Jia Liu, Yong Xue, Dominic Palmer-Brown, Ziqiang Chen, and Xingwei He. High-throughput geocomputational workflows in a grid environment. Computer, 48(11):70– 80, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        11/mco2015110070-abs.html.

Anonymous:2015:EA

[1724] Anonymous. ECSEE advertisement. Computer, 48(11):81, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        11/mco2015110081.pdf.

Berghel:2015:DAC

[1725] Hal Berghel. The dystory of Ahmed’s clock. Computer, 48(11): 82–85, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        11/mco2015110082.html.

Netto:2015:DWH

[1726] Marco A. S. Netto, Renato L. F. Cunha, and Nicole Sultanum. Deciding when and how to move HPC jobs to the cloud. Computer, 48 (11):86–89, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/11/mco2015110086-abs.html.

Finn:2015:WCB

[1727] Ed Finn. We can build the future. Computer, 48(11):90–91, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110090.html.

Schaffer:2015:ECA

[1728] Kim B. Schaffer. Expanding continuous authentication with mobile devices. Computer, 48(11):92–95, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110092-abs.html.

Tockey:2015:IHS

[1729] Steve Tockey. Insanity, hiring, and the software industry. Computer, 48 (11):96–101, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2015/11/mco2015110096-abs.html.

Saini:2015:APL

[1730] Mukesh Saini and Abdulmotaleb El Saddik. Absence privacy loss. Computer, 48(11):102–105, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110102-abs.html.

Ardis:2015:SCG

[1731] Mark Ardis, David Budgen, Gregory W. Hislop, Jeff Offutt, Mark Sebern, and Willem Visser. SE 2014:

Curriculum guidelines for undergraduate degree programs in software engineering. Computer, 48(11):106– 109, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110106-abs.html.

Eck:2015:UHC

[1732] Adam Eck, Leen-Kiat Soh, Kristen Olson, Allan L. McCutcheon, Jolene Smyth, and Robert F. Belli. Understanding the human condition through survey informatics. Computer, 48(11): 110–114, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2015/11/mco2015110110-abs.html.

Anonymous:2015:CCs

[1733] Anonymous. Call and calendar. Computer, 48(11):115–119, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110115.html.

Anonymous:2015:CCt

[1734] Anonymous. CS connection. Computer, 48(11):120, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://www.computer.org/csdl/mags/
co/2015/11/mco2015110120.html.

Anonymous:2015:CSIb

[1735] Anonymous. Call for special issue proposals house advertisement. Computer, 48(11):121, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2015/11/mco2015110121.pdf.

Anonymous:2015:COk

[1736] Anonymous. Career opportunities. Computer, 48(11):122–134, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110122.pdf.

Anonymous:2015:FYJf

[1737] Anonymous. Focus on your job search house advertisement. Computer, 48 (11):135, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2015/11/mco2015110135.pdf.

Grier:2015:HP

[1738] David Alan Grier. Hidden persuaders. Computer, 48(11):136, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco2015110136.html.

Anonymous:2015:CCX

[1739] Anonymous. Cool Chips XIX house advertisement. Computer, 48(11): c3, November 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
11/mco20151100c3.pdf.

Anonymous:2015:FCk

[1740] Anonymous. Front cover. Computer, 48(11):c1, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/11/mco20151100c1.pdf.

Anonymous:2015:LMAb

[1741] Anonymous. Learn more advertisement. Computer, 48(11):c2, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        11/mco20151100c2.pdf.

Anonymous:2015:SMF

[1742] Anonymous. Science minus the fiction advertisement. Computer, 48 (11):c4, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/11/mco20151100c4.pdf.

Anonymous:2015:Ml

[1743] Anonymous. Masthead. Computer, 48(12):3, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/12/mco2015120003.pdf.

Anonymous:2015:CHSl

[1744] Anonymous. Computer highlights Society magazines. Computer, 48(12): 4–5, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        12/mco2015120004.html.

Menzies:2015:CPD

[1745] Tim Menzies. Cross-project data for software engineering. Computer, 48 (12):6, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120006-abs.html.

Anonymous:2015:YAl

[1746] Anonymous. 32 & 16 years ago. Computer, 48(12):7–8, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120007.html.

Conte:2015:CSM

[1747] Thomas M. Conte. The Computer Society must change. Computer, 48 (12):9, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120009.html.

Conte:2015:RCN

[1748] Thomas M. Conte, Elie Track, and Erik DeBenedictis. Rebooting computing: New strategies for technology scaling. Computer, 48(12):10– 13, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120010.html.

Shalf:2015:CBM

[1749] John M. Shalf and Robert Leland. Computing beyond Moore’s Law. Computer, 48(12):14–23, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120014-abs.html.

Aly:2015:EEA

[1750] Mohamed M. Sabry Aly, Mingyu Gao, Gage Hills, Chi-Shuen Lee, Greg Pitner, Max M. Shulaker, Tony F. Wu, Mehdi Asheghi, Jeff Bokor, Franz Franchetti, Kenneth E. Goodson, Christos Kozyrakis, Igor Markov, Kunle Olukotun, Larry Pileggi, Eric Pop, Jan Rabaey, Christopher Re, H.- S. Philip Wong, and Subhasish Mitra. Energy-efficient abundant-data computing: The N3XT 1,000×. Computer, 48(12):24–33, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/12/mco2015120024-abs.html.

Holmes:2015:SCL

[1751] D. Scott Holmes, Alan M. Kadin, and Mark W. Johnson. Superconducting computing in large-scale hybrid systems. Computer, 48(12):34– 42, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        12/mco2015120034-abs.html.

Anonymous:2015:CWUb

[1752] Anonymous. Computer write for us house advertisement. Computer, 48 (12):43, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/12/mco2015120043.pdf.

Bresniker:2015:ATN

[1753] Kirk M. Bresniker, Sharad Singhal, and R. Stanley Williams. Adapting to thrive in a new economy of memory abundance. Computer, 48 (12):44–53, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2015/12/mco2015120044-abs.html.

Khasanvis:2015:ACI

[1754] Santosh Khasanvis, Mingyu Li, Mostafizur Rahman, Ayan K. Biswas, Mohammad Salehi-Fashami, Jayasimha Atulasimha, Supriyo Bandyopadhyay, and Csaba Andras Moritz. Architecting for causal intelligence at nanoscale. Computer, 48(12):54– 64, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120054-abs.html.

Mountain:2015:OWM

[1755] David J. Mountain, Mark McLean, Douglas Palmer, Joshua D. Prucnal, and Christopher D. Krieger. Ohmic weave: Memristor-based threshold gate networks. Computer, 48(12):65– 71, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120065-abs.html.

Papamichael:2015:CNC

[1756] Michael K. Papamichael and James C.

Hoe. The CONNECT network-on-chip generator. Computer, 48(12): 72–79, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120072-abs.html.

Voas:2015:TPS

[1757] Jeffrey Voas and George Hurlburt. Third-party software’s trust quagmire. Computer, 48(12):80–87, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        12/mco2015120080-abs.html.

Berghel:2015:WPG

[1758] Hal Berghel. What price Gonzo ethics? Computer, 48(12):88– 93, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        12/mco2015120088.html.

Brugali:2015:DVM

[1759] Davide Brugali, Rafael Capilla, and Mike Hinchey. Dynamic variability meets robotics. Computer, 48(12): 94–97, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        12/mco2015120094-abs.html.

Kott:2015:WCD

[1760] Alexander Kott, David S. Alberts, and Cliff Wang. Will cybersecurity dictate the outcome of future wars? Computer, 48(12):98– 101, December 2015. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2015/
                                        12/mco2015120098-abs.html.

Anonymous:2015:GVC

[1761] Anonymous. Gaudiot voted 2016 Computer Society President-Elect.

Computer, 48(12):102–104, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120102.html.

Anonymous:2015:CCu

[1762] Anonymous. Call and calendar.

Computer, 48(12):105–109, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120105.html.

Anonymous:2015:CCv

[1763] Anonymous. CS connection. Computer, 48(12):110–111, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/csdl/mags/co/2015/12/mco2015120110.
html.

Anonymous:2015:COl

[1764] Anonymous. Career opportunities. Computer, 48(12):112–127, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120112.pdf.

Grier:2015:QI

[1765] David Alan Grier. Quantum imagination. Computer, 48(12):128, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2015/
12/mco2015120128.html.

Anonymous:2015:FCl

[1766] Anonymous. Front cover. Computer, 48(12):c1, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/12/mco20151200c1.pdf.

Anonymous:2015:ICSj

[1767] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. Computer, 48 (12):c2, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/12/mco20151200c2.pdf.

Anonymous:2015:MYCe

[1768] Anonymous. Move your career forward house advertisement. Computer, 48(12):c3, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/12/mco20151200c3.pdf.

Anonymous:2015:WBW

[1769] Anonymous. Where a better world shows advertisement. Computer, 48 (12):c4, December 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2015/12/mco20151200c4.pdf.

Anonymous:2016:Ma

[1770] Anonymous. Masthead. Computer, 49(1):3, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        01/mco2016010003.pdf.

Fujii:2016:RBE

[1771] Roger U. Fujii. Raising the bar:

Ensuring quality products and services for our members. Computer, 49 (1):4, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
01/mco2016010004.html.

Helal:2016:OHT

[1772] Sumi Helal. 2016 outlook: Humans in the technology loop. Computer, 49(1):7–9, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/01/mco2016010007.html.

Nikolich:2016:S

[1773] Paul Nikolich, Kenneth Christensen, and Lori Cameron. 802 standards. Computer, 49(1):10, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
01/mco2016010010.html.

Anonymous:2016:FYJa

[1774] Anonymous. Focus on your job search house advertisement. Computer, 49 (1):11, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
01/mco2016010011.pdf.

Anonymous:2016:CHSa

[1775] Anonymous. Computer highlights Society magazines. Computer, 49(1): 12–13, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        01/mco2016010012.html.

Jones:2016:DER

[1776] Lynette A. Jones. Designing effective refreshable Braille displays. Computer, 49(1):14, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2016/01/mco2016010014.html.

Anonymous:2016:YAa

[1777] Anonymous. 32 & 16 years ago. Computer, 49(1):15–16, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        01/mco2016010015.html.

Abowd:2016:BWU

[1778] Gregory D. Abowd. Beyond Weiser: From ubiquitous to collective computing. Computer, 49(1):17–23, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        01/mco2016010017-abs.html.

Chang:2016:SAF

[1779] Carl K. Chang. Situation analytics: A foundation for a new software engineering paradigm. Computer, 49 (1):24–33, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2016/01/mco2016010024-abs.html.

Jiang:2016:CHC

[1780] Zhihao Jiang, Houssam Abbas, Kuk Jin Jang, and Rahul Mangharam. The challenges of high-confidence medical device software. Computer, 49 (1):34–42, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/01/mco2016010034-abs.html.

Milojicic:2016:OOS

[1781] Dejan Milojicic and Timothy Roscoe.

Outlook on operating systems. Computer, 49(1):43–51, January 2016. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://www.computer.org/csdl/mags/
co/2016/01/mco2016010043-abs.html.

Singh:2016:EPS

[1782] Lalit Kumar Singh, Gopika Vinod, and A. K. Tripathi. Early prediction of software reliability: A case study with a nuclear power plant system. Computer, 49(1):52–58, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/01/mco2016010052-abs.html.

Rajput:2016:URT

[1783] Hitesh Rajput, Tanmoy Som, and Soumitra Kar. Using Radon transform to recognize skewed images of vehicular license plates. Computer, 49(1):59–65, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/01/mco2016010059-abs.html.

Hole:2016:BTG

[1784] Kjell Jorgen Hole. Building trust in e-government services. Computer, 49 (1):66–74, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2016/01/mco2016010066-abs.html.

Berghel:2016:DP

[1785] Hal Berghel. Digital politics 2016. Computer, 49(1):75–79, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://www.computer.org/csdl/mags/
                                        co/2016/01/mco2016010075.html.

Petkovic:2016:ULA

[1786] Dragutin Petkovic. Using learning analytics to assess Capstone Project teams. Computer, 49(1):80– 83, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        01/mco2016010080-abs.html.

McGraw:2016:FSS

[1787] Gary McGraw. Four software security findings. Computer, 49(1): 84–87, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        01/mco2016010084.html.

Wolf:2016:ESC

[1788] Marilyn Wolf. Embedded software in crisis. Computer, 49(1):88– 90, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
01/mco2016010088.html.

Johnson:2016:BSF

[1789] Brian David Johnson. Beyond science fiction: The American dream. Computer, 49(1):91–93, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://www.computer.org/csdl/mags/
co/2016/01/mco2016010091.html.

Bulling:2016:PAU

[1790] Andreas Bulling. Pervasive attentive user interfaces. Computer, 49 (1):94–98, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/01/mco2016010094.html.

Spring:2016:FSW

[1791] Michael B. Spring. The future of standardization: Are we destined to repeat history? Computer, 49(1): 99–101, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
01/mco2016010099-abs.html.

Byrd:2016:IIW

[1792] Greg Byrd. IEEE/IBM Watson student showcase. Computer, 49(1):102– 104, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
01/mco2016010102.html.

Anonymous:2016:CCa

[1793] Anonymous. CS connection. Computer, 49(1):105, January 2016. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://www.computer.org/csdl/mags/
                                        co/2016/01/mco2016010105.html.

Anonymous:2016:CCb

[1794] Anonymous. Call and calendar. Computer, 49(1):106–109, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://www.computer.
                                        org/csdl/mags/co/2016/01/mco2016010106.
                                        html.

Anonymous:2016:COa

[1795] Anonymous. Career opportunities. Computer, 49(1):110–119, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        01/mco2016010110.pdf.

Grier:2016:NCT

[1796] David Alan Grier. New cycles of technology. Computer, 49(1): 120, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        01/mco2016010120.html.

Anonymous:2016:FCa

[1797] Anonymous. Front cover. Computer, 49(1):c1, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2016/01/mco20160100c1.pdf.

Anonymous:2016:ICSa

[1798] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. Computer, 49 (1):c3, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
01/mco20160100c3.pdf.

Anonymous:2016:MYCa

[1799] Anonymous. Move your career forward house advertisement. Computer, 49(1):c2, January 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
01/mco20160100c2.pdf.

Anonymous:2016:WWLa

[1800] Anonymous. Watch the world’s leading experts take multi-core strategies to new heights house advertisement. Computer, 49(1):c4, January 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
01/mco20160100c4.pdf.

Anonymous:2016:CSIa

[1801] Anonymous. Computer Society IEEE Fellows. Computer, 49(2): 3, February 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
02/mco2016020003.html.

Anonymous:2016:Mb

[1802] Anonymous. Masthead. Computer, 49 (2):4, February 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
02/mco2016020004.pdf.

Anonymous:2016:CHSb

[1803] Anonymous. Computer highlights Society magazines. Computer, 49(2): 5–6, February 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        02/mco2016020005.html.

Severance:2016:BSS

[1804] Charles Severance. Bruce Schneier: The security mindset. Computer, 49(2):7–8, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/02/mco2016020007.html.

Bertino:2016:SMA

[1805] Elisa Bertino. Securing mobile applications. Computer, 49(2):9, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        02/mco2016020009.html.

Anonymous:2016:YAb

[1806] Anonymous. 32 & 16 years ago. Computer, 49(2):10–11, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        02/mco2016020010.html.

Dustdar:2016:CC

[1807] Schahram Dustdar. Cloud computing. Computer, 49(2):12–13, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
02/mco2016020012.html.

Diaz-del-Rio:2016:EAL

[1808] Fernando Diaz del Rio, Javier Salmeron-Garcia, and Jose Luis Sevillano. Extending Amdahl’s Law for the cloud computing era. Computer, 49(2):14–22, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/02/mco2016020014-abs.html.

Anonymous:2016:ICSb

[1809] Anonymous. IEEE Computer Society:

Be at the center of it all. Computer, 49(2):23, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/02/mco2016020023.pdf.

Khoshkbarforoushha:2016:DEC

[1810] Alireza Khoshkbarforoushha, Meisong Wang, Rajiv Ranjan, Lizhe Wang, Leila Alem, Samee U. Khan, and Boualem Benatallah. Dimensions for evaluating cloud resource orchestration frameworks. Computer, 49(2): 24–33, February 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
02/mco2016020024-abs.html.

Noor:2016:MTC

[1811] Talal H. Noor, Quan Z. Sheng, Zakaria Maamar, and Sherali Zeadally. Managing trust in the cloud: State of the art and research challenges. Computer, 49(2):34–45, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/02/mco2016020034-abs.html.

Zhang:2016:EEW

[1812] Quan Zhang and Weisong Shi. Energyefficient workload placement in enterprise datacenters. Computer, 49 (2):46–52, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/02/mco2016020046-abs.html.

Anonymous:2016:FYJb

[1813] Anonymous. Focus on your job search house advertisement. Computer, 49 (2):53, February 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        02/mco2016020053.pdf.

Xu:2016:FCA

[1814] Lei Xu, Chunxiao Jiang, Yan Chen, Jian Wang, and Yong Ren. A framework for categorizing and applying privacy-preservation techniques in big data mining. Computer, 49(2):54– 62, February 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        02/mco2016020054-abs.html.

Berghel:2016:S

[1815] Hal Berghel. Secretocracy. Computer, 49(2):63–67, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/02/mco2016020063.html.

Barhamgi:2016:PPC

[1816] Mahmoud Barhamgi, Arosha K. Bandara, Yijun Yu, Khalid Belhajjame, and Bashar Nuseibeh. Protecting privacy in the cloud: Current practices, future directions. Computer, 49 (2):68–72, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/02/mco2016020068-abs.html.

Anonymous:2016:ICSc

[1817] Anonymous. IEEE Computer Society house advertisement. Computer, 49(2):73, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/02/mco2016020073.pdf.

Kshetri:2016:GSC

[1818] Nir Kshetri. Gambling with source code. Computer, 49(2):74–77, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
02/mco2016020074-abs.html.

Johnson:2016:FWW

[1819] Brian David Johnson and David Guston. Futures we want to inhabit. Computer, 49(2):78–79, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
02/mco2016020078.html.

Cerf:2016:TIN

[1820] Vint Cerf and Max Senges. Taking the Internet to the next physical level. Computer, 49(2):80–86, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
02/mco2016020080-abs.html.

Anonymous:2016:CCc

[1821] Anonymous. CS connection. Computer, 49(2):87–88, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2016/02/mco2016020087.
html.

Anonymous:2016:CCd

[1822] Anonymous. Call and calendar. Computer, 49(2):89–93, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2016/02/mco2016020089.
html.

Anonymous:2016:COb

[1823] Anonymous. Career opportunities. Computer, 49(2):94–103, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
02/mco2016020094.pdf.

Grier:2016:PMW

[1824] David Alan Grier. The post-MOOC world. Computer, 49(2): 104, February 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
02/mco2016020104.html.

Anonymous:2016:CPHa

[1825] Anonymous. Call for participation house advertisement. Computer, 49 (2):c4, February 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
02/mco20160200c4.pdf.

Anonymous:2016:FCb

[1826] Anonymous. Front cover. Computer, 49(2):c1, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/02/mco20160200c1.pdf.

Anonymous:2016:MYCb

[1827] Anonymous. Move your career forward house advertisement. Computer, 49(2):c2, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/02/mco20160200c2.pdf.

Anonymous:2016:RSRa

[1828] Anonymous. Rock stars of riskbased security house advertisement. Computer, 49(2):c3, February 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2016/02/mco20160200c3.
pdf.

Anonymous:2016:BK

[1829] Anonymous. Bodies of knowledge. Computer, 49(3):3, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.

                                        org/csdl/mags/co/2016/03/mco2016030003.
                                        html.

Anonymous:2016:Mc

[1830] Anonymous. Masthead. Computer, 49(3):4, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        03/mco2016030004.pdf.

DeFloriani:2016:UML

[1831] Leila De Floriani. Using 2D maps for 3D localization. Computer, 49 (3):5, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        03/mco2016030005.html.

Anonymous:2016:CHSc

[1832] Anonymous. Computer highlights Society magazines. Computer, 49 (3):6–7, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        03/mco2016030006.html.

Anonymous:2016:YAc

[1833] Anonymous. 32 & 16 years ago. Computer, 49(3):8–9, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2016/03/mco2016030008.html.

Lin:2016:CPU

[1834] Ying-Dar Lin, Jeffrey Voas, Antonio Pescape, and Peter Mueller. Communications and privacy under surveillance. Computer, 49(3):10– 13, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco2016030010.html.

Bellovin:2016:IST

[1835] Steven M. Bellovin, Matt Blaze, and Susan Landau. Insecure surveillance: Technical issues with remote computer searches. Computer, 49(3): 14–24, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco2016030014-abs.html.

Murdoch:2016:IDP

[1836] Steven J. Murdoch. Insecure by design: Protocols for encrypted phone calls. Computer, 49(3):25– 33, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco2016030025-abs.html.

Farrell:2016:RAR

[1837] Stephen Farrell. Requirements analysis required — otherwise targeted monitoring enables pervasive monitoring. Computer, 49(3):34–40, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco2016030034-abs.html.

Anonymous:2016:CNHa

[1838] Anonymous. Call for nominees house advertisement. Computer, 49 (3):41, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        03/mco2016030041.pdf.

Koch:2016:HAT

[1839] Robert Koch, Mario Golling, and Gabi Dreo Rodosek. How anonymous is the Tor network? A long-term black-box investigation. Computer, 49(3):42–49, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2016/03/mco2016030042-abs.html.

Metwalley:2016:UPM

[1840] Hassan Metwalley, Stefano Traverso, and Marco Mellia. Using passive measurements to demystify online trackers. Computer, 49(3):50– 55, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        03/mco2016030050-abs.html.

Yusof:2016:MCS

[1841] Imran Yusof and Al-Sakib Khan Pathan. Mitigating cross-site scripting attacks with a content security policy. Computer, 49(3):56– 63, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        03/mco2016030056-abs.html.

Sheth:2016:SCP

[1842] Amit Sheth, Pramod Anantharam, and Cory Henson. Semantic, cognitive, and perceptual computing: Paradigms that shape human experience. Computer, 49(3):64–72, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/03/mco2016030064-abs.html.

Berghel:2016:NNV

[1843] Hal Berghel. Net neutrality vs.

net neutering. Computer, 49(3): 73–77, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco2016030073.html.

Zawoad:2016:TDF

[1844] Shams Zawoad and Ragib Hasan.

Trustworthy digital forensics in the cloud. Computer, 49(3):78–81, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco2016030078-abs.html.

Charette:2016:SAB

[1845] Robert N. Charette. The STEM anxiety business. Computer, 49(3): 82–87, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco2016030082.html.

Walrad:2016:ICS

[1846] Charlene Chuck Walrad. The IEEE Computer Society and ACM’s Collaboration on Computing Education. Computer, 49(3):88–91, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco2016030088.html.

Bartleson:2016:PIP

[1847] Karen Bartleson. Practical insights from the peculiar world of standards. Computer, 49(3):92– 95, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        03/mco2016030092.html.

Misra:2016:HSA

[1848] Gaurav Misra and Jose M. Such. How socially aware are social media privacy controls? Computer, 49(3): 96–99, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        03/mco2016030096-abs.html.

Byrd:2016:IBM

[1849] Greg Byrd. Immortal bits: Managing our digital legacies. Computer, 49(3):100–103, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2016/03/mco2016030100.html.

Stavrou:2016:MED

[1850] Angelos Stavrou, Daniel Fleck, and Constantinos Kolias. On the move: Evading distributed denial-of-service attacks. Computer, 49(3):104–107, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        03/mco2016030104-abs.html.

Anonymous:2016:CCe

[1851] Anonymous. Call and calendar. Computer, 49(3):108–112, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/csdl/mags/co/2016/03/mco2016030108.
html.

Anonymous:2016:CPY

[1852] Anonymous. Conferences in the palm of your hand house advertisement. Computer, 49(3):113, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco2016030113.pdf.

Grier:2016:GSG

[1853] David Alan Grier. Go, stop, go, stop. Computer, 49(3):128, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco2016030128.html.

Anonymous:2016:FCc

[1854] Anonymous. Front cover. Computer, 49(3):c1, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/03/mco20160300c1.pdf.

Anonymous:2016:MYCc

[1855] Anonymous. Move your career forward house advertisement. Computer, 49(3):c3, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
03/mco20160300c3.pdf.

Anonymous:2016:RSBa

[1856] Anonymous. Rock stars of big data house advertisement. Computer, 49 (3):c4, March 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        03/mco20160300c4.pdf.

Anonymous:2016:RSRb

[1857] Anonymous. Rock stars of risk-based security house advertisement. Computer, 49(3):c2, March 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2016/03/mco20160300c2.pdf.

Anonymous:2016:Md

[1858] Anonymous. Masthead. Computer, 49(4):3, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        04/mco2016040003.pdf.

Anonymous:2016:CHSd

[1859] Anonymous. Computer highlights Society magazines. Computer, 49 (4):4–5, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        04/mco2016040004.html.

Anonymous:2016:CSP

[1860] Anonymous. Computer Society publications. Computer, 49(4):8– 9, April 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        04/mco2016040008.html.

Gaff:2016:NS

[1861] Brian M. Gaff. The need for speed. Computer, 49(4):10–12, April 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
04/mco2016040010-abs.html.

Severance:2016:BSB

[1862] Charles Severance. Bruce Schneier:

Building cryptographic systems. Computer, 49(4):13–15, April 2016. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://www.computer.org/csdl/mags/
co/2016/04/mco2016040013.html.

Holmes:2016:YA

[1863] Neville Holmes. 32 & 16 years ago. Computer, 49(4):16–17, April 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/csdl/mags/co/2016/04/mco2016040016.
html.

Yang:2016:CDD

[1864] Qiang Yang. Cross-domain data fusion. Computer, 49(4):18, April 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/csdl/mags/co/2016/04/mco2016040018.
html.

Anonymous:2016:CPHb

[1865] Anonymous. Call for papers: House advertisement. Computer, 49(4): 19, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
04/mco2016040019.pdf.

Ramakrishnan:2016:BD

[1866] Naren Ramakrishnan and Ravi Kumar. Big data. Computer, 49(4): 20–22, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
04/mco2016040020.html.

Patel:2016:ONS

[1867] Jignesh M. Patel. Operational NoSQL systems: What’s new and what’s next? Computer, 49(4): 23–30, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
04/mco2016040023-abs.html.

Gudivada:2016:RDM

[1868] Venkat N. Gudivada, Dhana Rao, and Vijay V. Raghavan. Renaissance in database management: Navigating the landscape of candidate systems. Computer, 49(4): 31–42, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
04/mco2016040031-abs.html.

Cherubini:2016:CSB

[1869] Giovanni Cherubini, Jens Jelitto, and Vinodh Venkatesan. Cognitive storage for big data. Computer, 49(4): 43–51, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
04/mco2016040043-abs.html.

Yu:2016:NCB

[1870] Hsiang-Fu Yu, Cho-Jui Hsieh, Hyokun Yun, S. V. N. Vishwanathan, and Inderjit Dhillon. Nomadic computing for big data analytics. Computer, 49(4):52–60, April 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/04/mco2016040052-abs.html.

Elbadrawy:2016:PSP

[1871] Asmaa Elbadrawy, Agoritsa Polyzou, Zhiyun Ren, Mackenzie Sweeney, George Karypis, and Huzefa Rangwala. Predicting student performance using personalized analytics. Computer, 49(4):61–69, April 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/04/mco2016040061-abs.html.

Parhami:2016:LAR

[1872] Behrooz Parhami. Low acceptance rates of conference papers considered harmful. Computer, 49(4): 70–73, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
04/mco2016040070-abs.html.

Berghel:2016:BST

[1873] Hal Berghel. The bogus science theorem (and the junk science wager). Computer, 49(4): 74–78, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
04/mco2016040074.html.

DeBenedictis:2016:BLT

[1874] Erik P. DeBenedictis. The Boolean logic tax. Computer, 49(4):79– 82, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        04/mco2016040079-abs.html.

Steimle:2016:SNU

[1875] Jurgen Steimle. Skin — the next user interface. Computer, 49(4): 83–87, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        04/mco2016040083-abs.html.

Agarwal:2016:TBS

[1876] Yuvraj Agarwal and Anind K. Dey. Toward building a safe, secure, and easy-to-use Internet of Things infrastructure. Computer, 49(4):88– 91, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        04/mco2016040088-abs.html.

Bennett:2016:A

[1877] Michael Bennett. Afrofuturism. Computer, 49(4):92–93, April 2016. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://www.computer.org/csdl/mags/
                                        co/2016/04/mco2016040092.html.

Knight:2016:DC

[1878] John Knight, Jack Davidson, Anh Nguyen-Tuong, Jason Hiser, and Michele Co. Diversity in cybersecurity. Computer, 49(4):94–98, April 2016. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
04/mco2016040094-abs.html.

Anonymous:2016:CCf

[1879] Anonymous. Call and calendar. Computer, 49(4):99–103, April 2016. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://www.computer.org/csdl/mags/
co/2016/04/mco2016040099.html.

Grier:2016:GP

[1880] David Alan Grier. The global profession. Computer, 49(4):120, April 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/csdl/mags/co/2016/
04/mco2016040120.html.

Anonymous:2016:CEAa

[1881] Anonymous. Computer Entrepreneur Award: House advertisement. Computer, 49(4):c3, April 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/04/mco20160400c3.pdf.

Anonymous:2016:FCd

[1882] Anonymous. Front cover. Computer, 49(4):c1, April 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2016/04/mco20160400c1.pdf.

Anonymous:2016:MYCd

[1883] Anonymous. Move your career forward: IEEE Computer Society membership: House advertisement. Computer, 49(4):c2, April 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //www.computer.org/csdl/mags/co/
                                        2016/04/mco20160400c2.pdf.

Anonymous:2016:RSBb

[1884] Anonymous. Rock stars of big data: House advertisement. Computer, 49 (4):c4, April 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

                                        computer.org/csdl/mags/co/2016/
                                        04/mco20160400c4.pdf.

Anonymous:2016:RSBc

[1885] Anonymous. Rock stars of big data: House advertisement. Computer, 49 (5):2, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        05/mco2016050002.pdf.

Anonymous:2016:Me

[1886] Anonymous. Masthead. Computer, 49(5):5, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        05/mco2016050005.pdf.

Anonymous:2016:H

[1887] Anonymous. Hiscox. Computer, 49 (5):6–7, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        05/mco2016050006.pdf.

Anonymous:2016:CHSe

[1888] Anonymous. Computer highlights Society magazines. Computer, 49 (5):8–9, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
05/mco2016050008.html.

Anonymous:2016:YAd

[1889] Anonymous. 32 & 16 years ago. Computer, 49(5):10–11, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/05/mco2016050010.html.

Anonymous:2016:MYCe

[1890] Anonymous. Move your career forward IEEE Computer Society membership: House advertisement. Computer, 49(5):12, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/05/mco2016050012.pdf.

Anonymous:2016:ICSd

[1891] Anonymous. IEEE Computer Society conferences. Computer, 49 (5):13, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
05/mco2016050013.html.

Lane:2016:VLO

[1892] H. Chad Lane and David Forsyth.

Virtual learning and object reconstruction. Computer, 49(5):14–15, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
05/mco2016050014.html.

Weaver:2016:CTT

[1893] Alfred C. Weaver, Greg Byrd, and Renee Bryce. Computing tools and techniques for emergency response. Computer, 49(5):16–18, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/05/mco2016050016.html.

Murphy:2016:EIU

[1894] Robin R. Murphy. Emergency informatics: Using computing to improve disaster management. Computer, 49(5):19–27, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/05/mco2016050019-abs.html.

Avvenuti:2016:ICM

[1895] Marco Avvenuti, Stefano Cresci, Fabio Del Vigna, and Maurizio Tesconi. Impromptu crisis mapping to prioritize emergency response. Computer, 49 (5):28–37, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        05/mco2016050028-abs.html.

Bozkurt:2016:BDS

[1896] Alper Bozkurt, Edgar Lobaton, and Mihail Sichitiu. A biobotic distributed sensor network for underrubble search and rescue. Computer, 49(5):38–46, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        05/mco2016050038-abs.html.

Besaleva:2016:ASN

[1897] Liliya I. Besaleva and Alfred C.

Weaver. Applications of social networks and crowdsourcing for disaster management improvement. Computer, 49(5):47–53, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/05/mco2016050047-abs.html.

Bal:2016:MSD

[1898] Henri Bal, Dick Epema, Cees de Laat, Rob van Nieuwpoort, John Romein, Frank Seinstra, Cees Snoek, and Harry Wijshoff. A medium-scale distributed system for computer science research: Infrastructure for the long term. Computer, 49(5):54–63, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/05/mco2016050054-abs.html.

Gayer:2016:LTA

[1899] Sebastian Gayer, Andrea Herrmann, Thorsten Keuler, Matthias Riebisch, and Pablo Oliveira Antonino. Lightweight traceability for the agile architect. Computer, 49(5):64–71, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2016/05/mco2016050064-
abs.html.

Anonymous:2016:ICSe

[1900] Anonymous. IEEE Computer Society information. Computer, 49 (5):72, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        05/mco2016050072.pdf.

Berghel:2016:AF

[1901] Hal Berghel. Attenuated FAQs. Computer, 49(5):73–77, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/05/mco2016050073.html.

Shi:2016:PEC

[1902] Weisong Shi and Schahram Dustdar. The promise of edge computing. Computer, 49(5):78–81, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/05/mco2016050078.html.

King:2016:WCP

[1903] John Leslie King and Jonathan Grudin. Will computers put us out of work? Computer, 49(5): 82–85, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        05/mco2016050082.html.

Anonymous:2016:WWLb

[1904] Anonymous. Watch the World’s leading experts take multi-core strategies to new heights: House advertisement. Computer, 49(5):86, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        05/mco2016050086.pdf.

Laplante:2016:SIT

[1905] Phillip A. Laplante, Jeffrey Voas, and Nancy Laplante. Standards for the Internet of Things: A case study in disaster response. Computer, 49 (5):87–90, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
05/mco2016050087.html.

Anonymous:2016:IMD

[1906] Anonymous. IEEE member digital library: House advertisement. Computer, 49(5):91, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/05/mco2016050091.pdf.

Niamut:2016:STR

[1907] Omar Niamut, Mu Mu, Spyros Denazis, and Nicholas Race. Social telemedia: The relationship between social information and networked media. Computer, 49(5): 92–97, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
05/mco2016050092-abs.html.

Byrd:2016:HSM

[1908] Greg Byrd. Home sweet mindcontrolled home. Computer, 49(5): 98–101, May 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
05/mco2016050098.html.

Voas:2016:IFM

[1909] Jeffrey Voas and Kim Schaffer. Insights on formal methods in cybersecurity. Computer, 49(5):102– 105, May 2016. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        05/mco2016050102.html.

Anonymous:2016:CCg

[1910] Anonymous. CS connection. Computer, 49(5):106–107, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2016/05/mco2016050106.
                                        html.

Anonymous:2016:CCh

[1911] Anonymous. Call and calendar. Computer, 49(5):108–111, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2016/05/mco2016050108.
                                        html.

Anonymous:2016:ICSf

[1912] Anonymous. IEEE Computer Society 2016 call for major award nominations: House advertisement. Computer, 49(5):112, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/05/mco2016050112.pdf.

Anonymous:2016:SCA

[1913] Anonymous. SanDisk corporation advertisement. Computer, 49(5): 113, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        05/mco2016050113.pdf.

Anonymous:2016:COc

[1914] Anonymous. Career opportunities. Computer, 49(5):114–127, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
05/mco2016050114.pdf.

Grier:2016:TI

[1915] David Alan Grier. Technological improvisation. Computer, 49(5): 128, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
05/mco2016050128.html.

Anonymous:2016:SGS

[1916] Anonymous. Samsung Galaxy S7 edge. Computer, 49(5):c2–1, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
05/mco20160500c2.pdf.

Anonymous:2016:CEAb

[1917] Anonymous. Computer entrepreneur award: House advertisement. Computer, 49(5):c3, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/05/mco20160500c3.pdf.

Anonymous:2016:FCe

[1918] Anonymous. Front cover. Computer, 49(5):c1, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/05/mco20160500c1.pdf.

Anonymous:2016:RSCa

[1919] Anonymous. Rock stars of cybersecurity: House advertisement. Computer, 49(5):c4, May 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/05/mco20160500c4.pdf.

Anonymous:2016:Mf

[1920] Anonymous. Masthead. Computer, 49(6):3, June 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        06/mco2016060003.pdf.

Anonymous:2016:YAe

[1921] Anonymous. 32 & 16 years ago. Computer, 49(6):4–5, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/06/mco2016060004.html.

Anonymous:2016:CHSf

[1922] Anonymous. Computer highlights Society magazines. Computer, 49 (6):6–7, June 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        06/mco2016060006.html.

Ibrahim:2016:DMB

[1923] Mohamed Ibrahim and Krishnendu Chakrabarty. Digital-microfluidic biochips. Computer, 49(6):8–9, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        06/mco2016060008.html.

Paulk:2016:IFP

[1924] Mark Paulk and Lori Cameron. IEEE floating point standard. Computer, 49(6):10, June 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco2016060010.html.

Bertino:2016:STP

[1925] Elisa Bertino. Security threats: Protecting the new cyberfrontier. Computer, 49(6):11–14, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/06/mco2016060011.html.

Vecchiato:2016:PAS

[1926] Daniel Vecchiato, Marco Vieira, and Eliane Martins. The perils of Android security configuration. Computer, 49(6):15–21, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/06/mco2016060015-abs.html.

Kotz:2016:PSM

[1927] David Kotz, Carl A. Gunter, Santosh Kumar, and Jonathan P. Weiner. Privacy and security in mobile health: A research agenda. Computer, 49 (6):22–30, June 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco2016060022-abs.html.

Loui:2016:HSC

[1928] Ronald P. Loui and Terrence D.

Loui. How to survive a cyber Pearl Harbor. Computer, 49(6):

31–37, June 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco2016060031.html.

Subrahmanian:2016:DTB

[1929] V. S. Subrahmanian, Amos Azaria, Skylar Durst, Vadim Kagan, Aram Galstyan, Kristina Lerman, Linhong Zhu, Emilio Ferrara, Alessandro Flammini, and Filippo Menczer. The DARPA Twitter bot challenge. Computer, 49(6):38–46, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/06/mco2016060038-abs.html.

Anonymous:2016:CWUa

[1930] Anonymous. Computer write for us: House advertisement. Computer, 49 (6):47, June 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco2016060047.pdf.

Chen:2016:MTC

[1931] Tsong Yueh Chen, Fei-Ching Kuo, Wenjuan Ma, Willy Susilo, Dave Towey, Jeffrey Voas, and Zhi Quan Zhou. Metamorphic testing for cybersecurity. Computer, 49(6):48– 55, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco2016060048-abs.html.

Zhang:2016:BBE

[1932] Qingpeng Zhang, Daniel Dajun Zeng, Fei-Yue Wang, Ronald Breiger, and James A. Hendler. Brokers or bridges? Exploring structural holes in a crowdsourcing system. Computer, 49(6): 56–64, June 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco2016060056-abs.html.

Anonymous:2016:MYCf

[1933] Anonymous. Move your career forward IEEE Computer Society membership: House advertisement. Computer, 49(6):65, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/06/mco2016060065.pdf.

Luo:2016:OMM

[1934] Qiuming Luo, Feng Xiao, Zhong Ming, Hao Li, Jianyong Chen, and Jianhua Zhang. Optimizing the memory management of a virtual machine monitor on a NUMA system. Computer, 49(6):66–74, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/06/mco2016060066-abs.html.

Anonymous:2016:RSBd

[1935] Anonymous. Rock stars of big data:

House advertisement. Computer, 49 (6):75, June 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco2016060075.pdf.

Anonymous:2016:CNHb

[1936] Anonymous. Call for nominees:

House advertisement. Computer, 49 (6):76, June 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        06/mco2016060076.pdf.

Reisman:2016:OID

[1937] Sorel Reisman and Eli Shmueli. Opening the instructional doors to computer science and IT — finally. Computer, 49(6):77–79, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/06/mco2016060077-abs.html.

Voas:2016:DIT

[1938] Jeffrey Voas. Demystifying the Internet of Things. Computer, 49(6): 80–83, June 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        06/mco2016060080-abs.html.

DeBenedictis:2016:RCL

[1939] Erik P. DeBenedictis. Rebooting computers as learning machines. Computer, 49(6):84–87, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/06/mco2016060084-abs.html.

Ullrich:2016:SRC

[1940] Daniel Ullrich and Andreas Butz. Social robots for the 21st Century. Computer, 49(6):88–92, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/06/mco2016060088-abs.html.

Anonymous:2016:IBA

[1941] Anonymous. Impact a broader audience: House advertisement. Computer, 49(6):93, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/06/mco2016060093.pdf.

Anonymous:2016:ICSg

[1942] Anonymous. IEEE Computer Society 2016 call for major award nominations: House advertisement. Computer, 49(6):94, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/06/mco2016060094.pdf.

Johnson:2016:SFC

[1943] Brian David Johnson. Science fiction and the coming age of sentient tools. Computer, 49(6):95– 97, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco2016060095.html.

Netkachova:2016:SIS

[1944] Kate Netkachova and Robin E.

Bloomfield. Security-informed safety. Computer, 49(6):98–102, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2016/06/mco2016060098-
abs.html.

Anonymous:2016:CCi

[1945] Anonymous. Call and calendar.

Computer, 49(6):103–107, June 2016.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2016/06/mco2016060103.
html.

Anonymous:2016:COd

[1946] Anonymous. Career opportunities. Computer, 49(6):108–119, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco2016060108.pdf.

Grier:2016:SC

[1947] David Alan Grier. Scientific conversations. Computer, 49(6):120, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco2016060120.html.

Anonymous:2016:EPA

[1948] Anonymous. Endless possibilities advertisement. Computer, 49(6): c2, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
06/mco20160600c2.pdf.

Anonymous:2016:FCf

[1949] Anonymous. Front cover. Computer, 49(6):c1, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/06/mco20160600c1.pdf.

Anonymous:2016:GSE

[1950] Anonymous. Galaxy S7 edge. Computer, 49(6):c4, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/06/mco20160600c4.pdf.

Anonymous:2016:RSCb

[1951] Anonymous. Rock stars of cybersecurity: House advertisement. Computer, 49(6):c3, June 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/06/mco20160600c3.pdf.

Anonymous:2016:Mg

[1952] Anonymous. Masthead. Computer, 49(7):3, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070003.pdf.

Anonymous:2016:YAf

[1953] Anonymous. 32 & 16 years ago. Computer, 49(7):4–5, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/07/mco2016070004.html.

Anonymous:2016:CHSg

[1954] Anonymous. Computer highlights Society magazines. Computer, 49 (7):6–7, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070006.html.

Schuller:2016:UCI

[1955] Bjorn Schuller and Jian Pei. Using computer intelligence for depression diagnosis and crowdsourcing. Computer, 49(7):8–9, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/07/mco2016070008.html.

Anonymous:2016:ICSh

[1956] Anonymous. IEEE Computer Society awards. Computer, 49(7): 10, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070010.html.

Oulasvirta:2016:UID

[1957] Antti Oulasvirta and Gregory D. Abowd. User interface design in the 21st Century. Computer, 49 (7):11–13, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070011.html.

Sugiura:2016:CUI

[1958] Yuta Sugiura, Takeo Igarashi, and Masahiko Inami. Cuddly user interfaces. Computer, 49(7):14–19, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070014-abs.html.

Sakamoto:2016:GIH

[1959] Daisuke Sakamoto, Yuta Sugiura, Masahiko Inami, and Takeo Igarashi. Graphical instruction for home robots. Computer, 49(7):20–25, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2016/07/mco2016070020-
abs.html.

Carroll:2016:SCS

[1960] John M. Carroll, Jiawei Chen, Chien Wen Tina Yuan, and Benjamin V. Hanrahan. In search of coproduction: Smart services as reciprocal activities. Computer, 49(7): 26–32, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070026-abs.html.

Anonymous:2016:CWUb

[1961] Anonymous. Computer write for us house advertisement. Computer, 49 (7):33, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070033.pdf.

Kato:2016:PED

[1962] Jun Kato, Takeo Igarashi, and Masataka Goto. Programming with examples to develop data-intensive user interfaces. Computer, 49(7):34– 42, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070034-abs.html.

Anonymous:2016:MYCg

[1963] Anonymous. Move your career forward IEEE Computer Society membership house advertisement. Computer, 49(7):43, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/07/mco2016070043.pdf.

Myers:2016:PUT

[1964] Brad A. Myers, Andrew J. Ko, Thomas D. LaToza, and YoungSeok Yoon. Programmers are users too: Human-centered methods for improving programming tools. Computer, 49 (7):44–52, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        07/mco2016070044-abs.html.

Kounev:2016:MBA

[1965] Samuel Kounev, Nikolaus Huber, Fabian Brosig, and Xiaoyun Zhu. A model-based approach to designing self-aware IT systems and infrastructures. Computer, 49(7):53– 61, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        07/mco2016070053-abs.html.

Yang:2016:FBF

[1966] Xia Yang, Peng Shi, Haiyong Sun, Wenxuan Zheng, and Jim Alves-Foss. A fast boot, fast shutdown technique for Android OS devices. Computer, 49(7):62–68, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        07/mco2016070062-abs.html.

Anonymous:2016:ICSi

[1967] Anonymous. IEEE Computer Society 2016 call for major award nominations house advertisement. Computer, 49(7):69, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/07/mco2016070069.pdf.

Bajcsy:2016:ESC

[1968] Peter Bajcsy, Antoine Vandecreme, Julien Amelot, Joe Chalfoun, Michael Majurski, and Mary Brady. Enabling stem cell characterization from large microscopy images. Computer, 49 (7):70–79, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070070-abs.html.

Grier:2016:FTC

[1969] David Alan Grier. Free tuition for community college: A role for the tech community? Computer, 49 (7):80–82, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070080-abs.html.

Anonymous:2016:CNEa

[1970] Anonymous. Call for nominees education awards nominations house advertisement. Computer, 49(7): 83, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070083.pdf.

Berghel:2016:DFD

[1971] Hal Berghel. The Dictator’s (false)

dilemma. Computer, 49(7):84–87, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        07/mco2016070084.html.

Matloff:2016:HVC

[1972] Norman Matloff. The H-1B visa controversy. Computer, 49(7):88– 93, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        07/mco2016070088.html.

Byrd:2016:LSS

[1973] Greg Byrd. Let the sun shine. Computer, 49(7):94–97, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/07/mco2016070094.html.

Hurlburt:2016:GES

[1974] George Hurlburt. “Good Enough” security: The best we’ll ever have. Computer, 49(7):98–101, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2016/07/mco2016070098-
                                        abs.html.

Anonymous:2016:CCj

[1975] Anonymous. CS connection. Computer, 49(7):102–103, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2016/07/mco2016070102.
                                        html.

Anonymous:2016:CCk

[1976] Anonymous. Call and calendar. Computer, 49(7):104–107, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
csdl/mags/co/2016/07/mco2016070104.
html.

Anonymous:2016:COe

[1977] Anonymous. Career opportunities. Computer, 49(7):108–119, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070108.pdf.

Grier:2016:AS

[1978] David Alan Grier. Aware of the situation. Computer, 49(7):120, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco2016070120.html.

Anonymous:2016:FCg

[1979] Anonymous. Front cover. Computer, 49(7):c1, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/07/mco20160700c1.pdf.

Anonymous:2016:FYJc

[1980] Anonymous. Focus on your job house advertisement. Computer, 49 (7):c3, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco20160700c3.pdf.

Anonymous:2016:RSBe

[1981] Anonymous. Rock stars of big data house advertisement. Computer, 49 (7):c4, July 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco20160700c4.pdf.

Anonymous:2016:RSPa

[1982] Anonymous. Rock stars of pervasive, predictive analytics house advertisement. Computer, 49(7): c2, July 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
07/mco20160700c2.pdf.

Anonymous:2016:Mh

[1983] Anonymous. Masthead. Computer, 49(8):3, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080003.pdf.

Anonymous:2016:CNEb

[1984] Anonymous. Call for nominees education awards nominations. Computer, 49(8):4, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/08/mco2016080004.pdf.

Anonymous:2016:ICSj

[1985] Anonymous. IEEE Computer Society digital library: IEEE’s first digital library. Computer, 49(8):5, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080005.html.

Anonymous:2016:YAg

[1986] Anonymous. 32 & 16 years ago. Computer, 49(8):6–7, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080006.html.

Anonymous:2016:CHSh

[1987] Anonymous. Computer highlights society magazines. Computer, 49(8): 8–9, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080008.html.

Bader:2016:EMT

[1988] David A. Bader. Evolving MPI+X toward exascale. Computer, 49(8): 10, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080010.html.

Anonymous:2016:RSPb

[1989] Anonymous. Rock stars of pervasive, predictive analytics. Computer, 49 (8):11, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080011.pdf.

Forte:2016:SCS

[1990] Domenic Forte, Ron Perez, Yongdae Kim, and Swarup Bhunia. Supply-chain security for cyberinfrastructure [Guest editors’ introduction]. Computer, 49(8):12–16, August 2016. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080012.html.

Anonymous:2016:Sa

[1991] Anonymous. Not so secure. Computer, 49(8):17, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/08/mco2016080017.pdf.

Ralston:2016:DCM

[1992] Parrish Ralston, David Fry, Scott Suko, Bryce Winters, Matthew King, and Robert Kober. Defeating counterfeiters with microscopic dielets embedded in electronic components. Computer, 49(8):18–26, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/08/mco2016080018-abs.html.

Anonymous:2016:RMA

[1993] Anonymous. 2017 Richard E. Merwin Award for Distinguished Service. Computer, 49(8):27, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/08/mco2016080027.pdf.

Skudlarek:2016:PSS

[1994] Joseph P. Skudlarek, Tom Katsioulas, and Michael Chen. A platform solution for secure supply-chain and chip life-cycle management. Computer, 49 (8):28–34, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/08/mco2016080028-abs.html.

Anonymous:2016:ISP

[1995] Anonymous. IEEE Security & Privacy. Computer, 49(8):35, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080035.pdf.

Ali:2016:SCS

[1996] Sk Subidh Ali, Mohamed Ibrahim, Jeyavijayan Rajendran, Ozgur Sinanoglu, and Krishnendu Chakrabarty. Supplychain security of digital microfluidic biochips. Computer, 49(8):36– 43, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080036-abs.html.

Hu:2016:DHT

[1997] Wei Hu, Baolei Mao, Jason Oberg, and Ryan Kastner. Detecting hardware Trojans with gate-level informationflow tracking. Computer, 49(8):44– 52, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080044-abs.html.

Anonymous:2016:ICSk

[1998] Anonymous. IEEE Computer Society 2016 call for major award nominations. Computer, 49(8):53, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080053.pdf.

Xiao:2016:SRC

[1999] Kan Xiao, Adib Nahiyan, and Mark Tehranipoor. Security rule checking in IC design. Computer, 49(8): 54–61, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        08/mco2016080054-abs.html.

Lysne:2016:VMD

[2000] Olav Lysne, Kjell J. Hole, Christian Otterstad, Oyvind Ytrehus, Raymond Aarseth, and Jorgen Tellnes. Vendor malware: Detection limits and mitigation. Computer, 49(8): 62–69, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        08/mco2016080062-abs.html.

Schaffer:2016:WHF

[2001] Kim Schaffer and Jeffrey Voas. What happened to formal methods for security? Computer, 49(8):70–79, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        08/mco2016080070.html.

Terry:2016:TNA

[2002] Doug Terry. Toward a new approach to IoT fault tolerance. Computer, 49(8):80–83, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/08/mco2016080080-abs.html.

Laporte:2016:SSE

[2003] Claude Y. Laporte and Rory V.

O’Connor. Systems and software engineering standards for very small entities: Accomplishments and overview. Computer, 49(8):84–87, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080084-abs.html.

DeBenedictis:2016:SCS

[2004] Erik P. DeBenedictis. The search for computing’s secretariat. Computer, 49(8):88–91, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/08/mco2016080088-abs.html.

Paulsen:2016:CSB

[2005] Celia Paulsen. Cybersecuring small businesses. Computer, 49(8):92–97, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080092-abs.html.

Berghel:2016:PD

[2006] Hal Berghel. The Pollyanna delusion. Computer, 49(8):98–102, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080098.html.

Anonymous:2016:ICSl

[2007] Anonymous. IEEE Computer Society information. Computer, 49(8):

103, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080103.pdf.

Ren:2016:RRB

[2008] Xiangshi Ren. Rethinking the relationship between humans and computers. Computer, 49(8):104–108, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080104-abs.html.

Atherton:2016:SFP

[2009] Evan Atherton. Science fiction prototyping at work. Computer, 49(8): 109–111, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080109.html.

Dastjerdi:2016:FCH

[2010] Amir Vahid Dastjerdi and Rajkumar Buyya. Fog computing: Helping the Internet of Things realize its potential. Computer, 49(8):112– 116, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080112-abs.html.

Anonymous:2016:FI

[2011] Anonymous. Facebook, Inc. Computer, 49(8):117, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/08/mco2016080117.pdf.

Anonymous:2016:ICSm

[2012] Anonymous. IEEE Computer Society election. Computer, 49(8):118– 128, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080118.html.

Anonymous:2016:MYCh

[2013] Anonymous. Move your career forward IEEE Computer Society membership. Computer, 49(8):129, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco2016080129.pdf.

Anonymous:2016:IDV

[2014] Anonymous. IEEE Division V delegate/director candidates. Computer, 49(8):130–131, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/08/mco2016080130.html.

Anonymous:2016:PAI

[2015] Anonymous. Proposed amendment to IEEE’s constitution. Computer, 49 (8):132–133, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/08/mco2016080132.html.

Anonymous:2016:IPE

[2016] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 49(8):134– 136, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        08/mco2016080134.html.

Anonymous:2016:II

[2017] Anonymous. Intuit Inc. Computer, 49(8):137, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/csdl/mags/co/
                                        2016/08/mco2016080137.pdf.

Anonymous:2016:CCl

[2018] Anonymous. Call and calendar. Computer, 49(8):138–141, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        csdl/mags/co/2016/08/mco2016080138.
                                        html.

Grier:2016:CT

[2019] David Alan Grier. Changing trust. Computer, 49(8):156, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        08/mco2016080156.html.

Anonymous:2016:CWU

[2020] Anonymous. Computer write for us. Computer, 49(8):c3, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/co/2016/
                                        08/mco20160800c3.pdf.

Anonymous:2016:FCh

[2021] Anonymous. Front cover. Computer, 49(8):c1, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/csdl/mags/co/
2016/08/mco20160800c1.pdf.

Anonymous:2016:IIP

[2022] Anonymous. 31st IEEE International Parallel and Distributed Processing Symposium. Computer, 49 (8):c2, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco20160800c2.pdf.

Anonymous:2016:RSCc

[2023] Anonymous. Rock stars of cybersecurity. Computer, 49(8):c4, August 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/csdl/mags/co/2016/
08/mco20160800c4.pdf.

Anonymous:2016:NMOa

[2024] Anonymous. New membership options for a better fit [Advertisement]. Computer, 49(9):2, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
09/mco2016090002.pdf.

Anonymous:2016:AYCa

[2025] Anonymous. Achieve your career goals with the fit that’s right for you [Advertisement]. Computer, 49(9): 3, September 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
09/mco2016090003.pdf.

Anonymous:2016:Mi

[2026] Anonymous. Masthead. Computer, 49(9):4, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/09/mco2016090004.pdf.

Anonymous:2016:TCa

[2027] Anonymous. Table of contents. Computer, 49(9):5–6, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2016/09/mco2016090005.
                                        pdf.

Anonymous:2016:ICSn

[2028] Anonymous. IEEE Computer Society membership. Computer, 49(9): 7, September 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        09/mco2016090007.html.

Anonymous:2016:YAh

[2029] Anonymous. 32 & 16 years ago. Computer, 49(9):8–9, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        09/mco2016090008.html.

Anonymous:2016:CHSi

[2030] Anonymous. Computer highlights Society magazines. Computer, 49(9): 10–11, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/09/mco2016090010.html.

Montuschi:2016:MCA

[2031] Paolo Montuschi and Jean-Michel Muller. Modern computer arithmetic. Computer, 49(9):12, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
09/mco2016090012.html.

Anonymous:2016:CNEc

[2032] Anonymous. Call for nominees education awards nominations. Computer, 49(9):13, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/09/mco2016090013.pdf.

Murugesan:2016:NGC

[2033] San Murugesan and Bob Colwell.

Next-generation computing paradigms. Computer, 49(9):14–20, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
09/mco2016090014.html.

Svore:2016:QFC

[2034] Krysta M. Svore and Matthias Troyer. The quantum future of computation. Computer, 49(9):21– 30, September 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
09/mco2016090021-abs.html.

VanMeter:2016:PSD

[2035] Rodney Van Meter and Simon J. Devitt. The path to scalable distributed quantum computing. Computer, 49 (9):31–42, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/09/mco2016090031-abs.html.

Coleman:2016:EMC

[2036] Victoria Coleman. Embodied molecular computation: Potential and challenges. Computer, 49(9):43–51, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
09/mco2016090043-abs.html.

Yang:2016:SIM

[2037] Xin-She Yang, Suash Deb, Simon Fong, Xingshi He, and Yu-Xin Zhao. From swarm intelligence to metaheuristics: Nature-inspired optimization algorithms. Computer, 49(9): 52–59, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/09/mco2016090052-abs.html.

Saproo:2016:CCC

[2038] Sameer Saproo, Josef Faller, Victor Shih, Paul Sajda, Nicholas R. Waytowich, Addison Bohannon, Vernon J. Lawhern, Brent J. Lance, and David Jangraw. Cortically coupled computing: A new paradigm for synergistic human-machine interaction. Computer, 49(9):60–68, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/09/mco2016090060-abs.html.

Sha:2016:RTC

[2039] Lui Sha, Marco Caccamo, Renato Mancuso, Jung-Eun Kim, Man-Ki Yoon, Rodolfo Pellizzoni, Heechul Yun, Russell B. Kegley, Dennis R. Perlman, Greg Arundale, and Richard Bradford. Real-time computing on multicore processors. Computer, 49(9): 69–77, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/09/mco2016090069-abs.html.

Maresca:2016:SSE

[2040] Massimo Maresca. The spreadsheet space: Eliminating the boundaries of data cross-referencing. Computer, 49 (9):78–85, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/09/mco2016090078-abs.html.

Zhang:2016:SCM

[2041] Xing Zhang, Zhenglei Yi, Zhi Yan, Geyong Min, Wenbo Wang, Ahmed Elmokashfi, Sabita Maharjan, and Yan Zhang. Social computing for mobile big data. Computer, 49(9):86– 90, September 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
09/mco2016090086-abs.html.

Campbell:2016:TCC

[2042] Scott Campbell. Teaching cloud computing. Computer, 49(9):91– 93, September 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        09/mco2016090091.html.

Byrd:2016:SU

[2043] Greg Byrd. Seeing is understanding. Computer, 49(9):94–97, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        09/mco2016090094.html.

Vassilev:2016:ESU

[2044] Apostol Vassilev and Robert Staples. Entropy as a service: Unlocking cryptography’s full potential. Computer, 49 (9):98–102, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/09/mco2016090098-abs.html.

Anonymous:2016:FIH

[2045] Anonymous. Facebook, Inc. house advertisement. Computer, 49(9): 103, September 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        09/mco2016090103.pdf.

Berghel:2016:CKF

[2046] Hal Berghel. Coda in the key of F2654hD4. Computer, 49(9):104– 109, September 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        09/mco2016090104.html.

Madria:2016:SRA

[2047] Sanjay K. Madria. Security and risk assessment in the cloud. Computer, 49 (9):110–113, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/09/mco2016090110-abs.html.

Charette:2016:DGL

[2048] Robert N. Charette. Dragging government legacy systems out of the shadows. Computer, 49(9):114–119, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
09/mco2016090114.html.

Anonymous:2016:CCm

[2049] Anonymous. CS connection. Computer, 49(9):120–121, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2016/09/mco2016090120.
html.

Anonymous:2016:CCn

[2050] Anonymous. Call and calendar.

Computer, 49(9):122–125, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
09/mco2016090122.html.

Anonymous:2016:COf

[2051] Anonymous. Career opportunities.

Computer, 49(9):126–143, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
09/mco2016090126.pdf.

Grier:2016:MP

[2052] David Alan Grier. The means of production. Computer, 49(9):144, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        09/mco2016090144.html.

Anonymous:2016:UBM

[2053] Anonymous. The Universität der Bundeswehr München [advertisement]. Computer, 49(9):c2–1, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        09/mco20160900c2.pdf.

Anonymous:2016:FCi

[2054] Anonymous. Front cover. Computer, 49(9):c1, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/09/mco20160900c1.pdf.

Anonymous:2016:MYCi

[2055] Anonymous. Move your career forward IEEE Computer Society membership [Advertisement]. Computer, 49(9):c3, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/09/mco20160900c3.pdf.

Anonymous:2016:Sb

[2056] Anonymous. SolidThinking. Computer, 49(9):c4, September 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2016/09/mco20160900c4.
pdf.

Anonymous:2016:AYCb

[2057] Anonymous. Achieve your career goals with the fit that’s right for you. Computer, 49(10):1, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/10/mco2016100001.pdf.

Anonymous:2016:Mj

[2058] Anonymous. Masthead. Computer, 49(10):2, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/10/mco2016100002.pdf.

Anonymous:2016:TCb

[2059] Anonymous. Table of contents.

Computer, 49(10):3–4, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2016/10/mco2016100003.
pdf.

Anonymous:2016:ICSo

[2060] Anonymous. IEEE Computer Society education. Computer, 49(10): 5, October 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100005.html.

Anonymous:2016:YAi

[2061] Anonymous. 32 & 16 years ago. Computer, 49(10):6–7, October 2016. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100006.html.

Anonymous:2016:CHSj

[2062] Anonymous. Computer highlights Society magazines. Computer, 49(10): 8–9, October 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100008.html.

Sethumadhavan:2016:HEP

[2063] Simha Sethumadhavan. Hardwareenforced privacy. Computer, 49(10): 10, October 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100010.html.

Anonymous:2016:CIS

[2064] Anonymous. COMPSAC 2017 — ITiP symposium. Computer, 49(10): 11, October 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100011.pdf.

Severance:2016:LCM

[2065] Charles Severance. The Living Computer Museum. Computer, 49(10): 12–13, October 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100012.html.

Getov:2016:NFE

[2066] Vladimir Getov, Adolfy Hoisie, and Pradip Bose. New frontiers in energyefficient computing [Guest Editors’ introduction]. Computer, 49(10):14– 18, October 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100014.html.

Anonymous:2016:ICSp

[2067] Anonymous. IEEE Computer Society. Computer, 49(10):19, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100019.pdf.

Wu:2016:UPP

[2068] Xingfu Wu, Valerie Taylor, Jeanine Cook, and Philip J. Mucci. Using performance-power modeling to improve energy efficiency of HPC applications. Computer, 49(10):20– 29, October 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100020-abs.html.

Acun:2016:PRP

[2069] Bilge Acun, Akhil Langer, Esteban Meneses, Harshitha Menon, Osman Sarood, Ehsan Totoni, and Laxmikant V. Kale. Power, reliability, and performance: One system to rule them all. Computer, 49 (10):30–37, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/10/mco2016100030-abs.html.

Grant:2016:SPM

[2070] Ryan E. Grant, Michael Levenhagen, Stephen L. Olivier, David DeBonis, Kevin T. Pedretti, and James H. Laros III. Standardizing power monitoring and control at exascale. Computer, 49(10):38–46, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/10/mco2016100038-abs.html.

Rotem:2016:HEH

[2071] Efraim Rotem, Uri C. Weiser, Avi Mendelson, Ran Ginosar, Eliezer Weissmann, and Yoni Aizik. H-EARtH: Heterogeneous multicore platform energy management. Computer, 49(10): 47–55, October 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        10/mco2016100047-abs.html.

Merkel:2016:NSB

[2072] Cory Merkel, Raqibul Hasan, Nicholas Soures, Dhireesha Kudithipudi, Tarek Taha, Sapan Agarwal, and Matthew Marinella. Neuromemristive systems: Boosting efficiency through braininspired computing. Computer, 49 (10):56–64, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/10/mco2016100056-abs.html.

Atluri:2016:BNP

[2073] Gowtham Atluri, Angus MacDonald III, Kelvin O. Lim, and Vipin Kumar. The brain-network paradigm: Using functional imaging data to study how the brain works. Computer, 49 (10):65–71, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/10/mco2016100065-abs.html.

Linthicum:2016:RDA

[2074] David Linthicum. Responsive data architecture for the Internet of Things. Computer, 49(10):72–75, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100072-abs.html.

DeBenedictis:2016:HWM

[2075] Erik P. DeBenedictis and R. Stanley Williams. Help wanted: A modernday Turing. Computer, 49(10):76– 79, October 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100076-abs.html.

Simos:2016:CMS

[2076] Dimitris E. Simos, Rick Kuhn, Artemios G. Voyiatzis, and Raghu Kacker. Combinatorial methods in security testing. Computer, 49(10): 80–83, October 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100080-abs.html.

Berghel:2016:DJT

[2077] Hal Berghel. Douglas Jones on today’s voting machines. Computer, 49 (10):84–89, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/10/mco2016100084.html.

Jacucci:2016:BSQ

[2078] Giulio Jacucci. Beyond search queries. Computer, 49(10):90–93, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        10/mco2016100090-abs.html.

Lee:2016:T

[2079] Ryan Lee. Threatcasting. Computer, 49(10):94–95, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/10/mco2016100094.html.

Wild:2016:FLW

[2080] Fridolin Wild. The future of learning at the workplace is augmented reality. Computer, 49(10):96–98, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        10/mco2016100096-abs.html.

Anonymous:2016:CCo

[2081] Anonymous. Call and calendar. Computer, 49(10):99–102, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2016/10/mco2016100099.
                                        html.

Anonymous:2016:COg

[2082] Anonymous. Career opportunities. Computer, 49(10):103–111, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100103.pdf.

Grier:2016:IT

[2083] David Alan Grier. Innovating for tomorrow. Computer, 49(10):112, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco2016100112.html.

Anonymous:2016:FCj

[2084] Anonymous. Front cover. Computer, 49(10):c1, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/10/mco20161000c1.pdf.

Anonymous:2016:MYCj

[2085] Anonymous. Move your career forward IEEE Computer Society membership. Computer, 49(10):c3, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
10/mco20161000c3.pdf.

Anonymous:2016:NMOb

[2086] Anonymous. New membership options for a better fit. Computer, 49(10):c2, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/10/mco20161000c2.pdf.

Anonymous:2016:ICSq

[2087] Anonymous. IEEE Computer Society is where you choose the resources that fit your career house advertisement. Computer, 49(11): 1, November 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
11/mco2016110001.pdf.

Anonymous:2016:Mk

[2088] Anonymous. Masthead. Computer, 49(11):2, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/11/mco2016110002.pdf.

Anonymous:2016:TCc

[2089] Anonymous. Table of contents. Computer, 49(11):3–4, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2016/11/mco2016110003.
pdf.

Anonymous:2016:ICSr

[2090] Anonymous. IEEE Computer Society technical committees. Computer, 49(11):5, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/11/mco2016110005.html.

Anonymous:2016:YAj

[2091] Anonymous. 32 & 16 years ago. Computer, 49(11):6–7, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
11/mco2016110006.html.

Anonymous:2016:CHSk

[2092] Anonymous. Computer highlights Society magazines. Computer, 49(11): 8–9, November 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
11/mco2016110008.html.

Dwyer:2016:FFN

[2093] Matthew B. Dwyer. Finding flaws in natural language requirements. Computer, 49(11):10, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2016/11/mco2016110010.
html.

Varshney:2016:SHW

[2094] Upkar Varshney and Carl K. Chang.

Smart health and well-being. Computer, 49(11):11–13, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2016/11/mco2016110011.
html.

Olshansky:2016:FSH

[2095] S. Jay Olshansky, Bruce A. Carnes, Yang Claire Yang, Norvell Miller, Janet Anderson, Hiram Beltran-Sanchez, and Karl Ricanek, Jr. The future of smart health. Computer, 49(11):14– 21, November 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
11/mco2016110014-abs.html.

Carroll:2016:KSF

[2096] Noel Carroll. Key success factors for smart and connected health software solutions. Computer, 49(11):22– 28, November 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
11/mco2016110022-abs.html.

Sprint:2016:USH

[2097] Gina Sprint, Diane J. Cook, Roschelle Fritz, and Maureen Schmitter-Edgecombe. Using smart homes to detect and analyze health events. Computer, 49 (11):29–37, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/11/mco2016110029-abs.html.

Hijazi:2016:MLC

[2098] Shurouq Hijazi, Alex Page, Burak Kantarci, and Tolga Soyata. Machine learning in cardiac health monitoring and decision support. Computer, 49 (11):38–48, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/11/mco2016110038-abs.html.

Su:2016:PSP

[2099] Xiang Su, Jarkko Hyysalo, Mika Rautiainen, Jukka Riekki, Jaakko Sauvola, Altti Ilari Maarala, Harri Hirvonsalo, Pingjiang Li, and Harri Honko. Privacy as a service: Protecting the individual in healthcare data processing. Computer, 49(11): 49–59, November 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
11/mco2016110049-abs.html.

Cervantes:2016:AAS

[2100] Humberto Cervantes, Rick Kazman, Jungwoo Ryoo, Duyoung Choi, and Duksung Jang. Architectural approaches to security: Four case studies. Computer, 49(11):60–67, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
11/mco2016110060-abs.html.

Lin:2016:BSC

[2101] Po-Ching Lin, Ying-Dar Lin, Cheng-Ying Wu, Yuan-Cheng Lai, and Yi-Chih Kao. Balanced service chaining in software-defined networks with network function virtualization. Computer, 49(11):68– 76, November 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
11/mco2016110068-abs.html.

Banifatemi:2016:ENB

[2102] Amir Banifatemi and Jean-Luc Gaudiot. Engineering the new boundaries of AI. Computer, 49(11):77– 79, November 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
11/mco2016110077.html.

Dhar:2016:ESP

[2103] Vasant Dhar. Equity, safety, and privacy in the autonomous vehicle era. Computer, 49(11):80–83, November 2016. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        11/mco2016110080.html.

Kott:2016:FWC

[2104] Alexander Kott, Ananthram Swami, and Bruce J. West. The fog of war in cyberspace. Computer, 49(11):84– 87, November 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        11/mco2016110084-abs.html.

Byrd:2016:TDB

[2105] Greg Byrd. Tactile digital Braille display. Computer, 49(11):88–90, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        11/mco2016110088.html.

Berghel:2016:CEG

[2106] Hal Berghel. Chasing Elbridge’s ghost: The digital gerrymander. Computer, 49 (11):91–95, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/11/mco2016110091.html.

Kogias:2016:CFE

[2107] Dimitrios G. Kogias, Michael G. Xevgenis, and Charalampos Z. Patrikakis. Cloud federation and the evolution of cloud computing. Computer, 49 (11):96–99, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/11/mco2016110096-abs.html.

Anonymous:2016:CCp

[2108] Anonymous. Call and calendar.

Computer, 49(11):100–103, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
11/mco2016110100.html.

Grier:2016:ICS

[2109] David Alan Grier. “I’m Not a Computer Scientist, but …. Computer, 49(11):104, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/11/mco2016110104.html.

Anonymous:2016:CSIb

[2110] Anonymous. Call for special issue proposals house advertisement. Computer, 49(11):c3, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/11/mco20161100c3.pdf.

Anonymous:2016:FCk

[2111] Anonymous. Front cover. Computer, 49(11):c1, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/11/mco20161100c1.pdf.

Anonymous:2016:MYCk

[2112] Anonymous. Move your career forward IEEE Computer Society membership house advertisement. Computer, 49(11):c4, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/11/mco20161100c4.pdf.

Anonymous:2016:NMOc

[2113] Anonymous. New membership options for a better fit house advertisement. Computer, 49(11):c2, November 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        11/mco20161100c2.pdf.

Anonymous:2016:ICSs

[2114] Anonymous. IEEE Computer Society is where you choose the resources that fit your career. Computer, 49 (12):1, December 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        12/mco2016120001.pdf.

Anonymous:2016:Ml

[2115] Anonymous. Masthead. Computer, 49(12):2, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/12/mco2016120002.pdf.

Anonymous:2016:TCd

[2116] Anonymous. Table of contents. Computer, 49(12):3–4, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2016/12/mco2016120003.
                                        pdf.

Anonymous:2016:ICSt

[2117] Anonymous. IEEE Computer Society’s Computer magazine. Computer, 49(12):5, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/12/mco2016120005.html.

Fujii:2016:OJF

[2118] Roger U. Fujii. Our journey to the future. Computer, 49(12):6– 9, December 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
12/mco2016120006.html.

Anonymous:2016:YAk

[2119] Anonymous. 32 & 16 years ago. Computer, 49(12):10–11, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
12/mco2016120010.html.

Anonymous:2016:CHSl

[2120] Anonymous. Computer highlights Society magazines. Computer, 49(12): 12–13, December 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
12/mco2016120012.html.

Anonymous:2016:AHL

[2121] Anonymous. Art of hacking elearning courses IEEE Computer Society. Computer, 49(12):14, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
12/mco2016120014.pdf.

Jadbabaie:2016:DDC

[2122] Ali Jadbabaie. Defining data clusters. Computer, 49(12):15, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        12/mco2016120015.html.

Want:2016:MIG

[2123] Roy Want, Trevor Pering, and Yuvraj Agarwal. Multidevice interaction [Guest Editors’ introduction]. Computer, 49(12):16–20, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        12/mco2016120016.html.

Anonymous:2016:CSI

[2124] Anonymous. Call for special issue proposals. Computer, 49(12): 21, December 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        12/mco2016120021.pdf.

Leppanen:2016:CEE

[2125] Jarno Leppanen, Mikko Pelkonen, Haipeng Guo, Samuli Hemminki, Petteri Nurmi, and Sasu Tarkoma. Collaborative and energy-efficient speech monitoring on smart devices. Computer, 49(12):22–30, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        12/mco2016120022-abs.html.

Anonymous:2016:CYI

[2126] Anonymous. Can you invent a better world through technology? Computer, 49(12):31, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/12/mco2016120031.pdf.

Pinciroli:2016:SOP

[2127] Carlo Pinciroli and Giovanni Beltrame.

Swarm-oriented programming of distributed robot networks. Computer, 49 (12):32–41, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/12/mco2016120032-abs.html.

Neate:2016:IMC

[2128] Timothy Neate, Matt Jones, and Michael Evans. Interdevice media: Choreographing content to maximize viewer engagement. Computer, 49 (12):42–49, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/12/mco2016120042-abs.html.

Bures:2016:SAC

[2129] Tomas Bures, Frantisek Plasil, Michal Kit, Petr Tuma, and Nicklas Hoch. Software abstractions for component interaction in the Internet of Things. Computer, 49(12):50–59, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
12/mco2016120050-abs.html.

Perera:2016:PKM

[2130] Charith Perera, Chang Liu, Rajiv Ranjan, Lizhe Wang, and Albert Y. Zomaya. Privacy-knowledge modeling for the Internet of Things: A look back. Computer, 49(12):60– 68, December 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        12/mco2016120060-abs.html.

Anonymous:2016:WWL

[2131] Anonymous. Watch the world’s leading experts take multi-core strategies to new heights. Computer, 49 (12):69, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2016/12/mco2016120069.pdf.

Kott:2016:IBT

[2132] Alexander Kott, Ananthram Swami, and Bruce J. West. The Internet of Battle Things. Computer, 49(12):70– 75, December 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        12/mco2016120070-abs.html.

DeBenedictis:2016:CCN

[2133] Erik P. DeBenedictis. Computational complexity and new computing approaches. Computer, 49(12): 76–79, December 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2016/
                                        12/mco2016120076-abs.html.

Hu:2016:ACP

[2134] Vincent C. Hu and Rick Kuhn. Access control policy verification. Computer, 49(12):80–83, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/12/mco2016120080-abs.html.

Johnson:2016:LTA

[2135] Brian David Johnson. So long, and thanks for all the robots. Computer, 49 (12):84–85, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/12/mco2016120084.html.

Zemrowski:2016:NBF

[2136] Kenneth M. Zemrowski. NIST bases flagship security engineering publication on ISO/IEC/IEEE 15288:2015. Computer, 49(12):86– 88, December 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
12/mco2016120086.html.

Anonymous:2016:NSS

[2137] Anonymous. Nominations are solicited for the Seymour Cray, Sidney Fernbach & Ken Kennedy awards. Computer, 49(12):89, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/12/mco2016120089.pdf.

Anonymous:2016:KVC

[2138] Anonymous. Kasahara voted 2017 Computer Society President-Elect.

Computer, 49(12):90–92, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
12/mco2016120090.html.

Anonymous:2016:CCq

[2139] Anonymous. Call and calendar. Computer, 49(12):93–95, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
12/mco2016120093.html.

Grier:2016:FYO

[2140] David Alan Grier. Fifteen years to open source. Computer, 49(12): 96, December 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
12/mco2016120096.html.

Anonymous:2016:FCl

[2141] Anonymous. Front cover. Computer, 49(12):c1, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/12/mco20161200c1.pdf.

Anonymous:2016:MYCl

[2142] Anonymous. Move your career forward IEEE Computer Society membership. Computer, 49(12):c3, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2016/
12/mco20161200c3.pdf.

Anonymous:2016:NMOd

[2143] Anonymous. New membership options for a better fit. Computer, 49 (12):c2, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/12/mco20161200c2.pdf.

Anonymous:2016:T

[2144] Anonymous. TechIgnite. Computer, 49(12):c4, December 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2016/12/mco20161200c4.pdf.

Anonymous:2017:ICSa

[2145] Anonymous. IEEE Computer Society is where you choose the resources that fit your career house advertisement. Computer, 50(1):1, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010001.pdf.

Anonymous:2017:Ma

[2146] Anonymous. Masthead. Computer, 50(1):2, January 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010002.pdf.

Anonymous:2017:TCa

[2147] Anonymous. Table of contents.

Computer, 50(1):3–4, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/01/mco2017010003.
pdf.

Gaudiot:2017:NCS

[2148] Jean-Luc Gaudiot. 2017: The new Computer Society. Computer, 50(1): 5–7, January 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010005.html.

Liu:2017:CDM

[2149] Shaoshan Liu, James Peng, and Jean-Luc Gaudiot. Computer, drive my car! Computer, 50(1):8, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010008.html.

Anonymous:2017:AICa

[2150] Anonymous. ACM–IEEE CS Eckert– Mauchly Award house advertisement. Computer, 50(1):9, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010009.pdf.

Neuhold:2017:CJa

[2151] Erich Neuhold. Computer, January 1967 and 1992. Computer, 50(1): 10–11, January 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010010.html.

Anonymous:2017:CHSa

[2152] Anonymous. Computer highlights Society magazines. Computer, 50(1): 12–13, January 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010012.html.

Helal:2017:CA

[2153] Sumi Helal. Computer’s 50th anniversary. Computer, 50(1):14–18, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010014.html.

Anonymous:2017:FYJa

[2154] Anonymous. Focus on your job search house advertisement. Computer, 50 (1):19, January 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010019.pdf.

Conte:2017:RCR

[2155] Thomas M. Conte, Erik P. DeBenedictis, Paolo A. Gargini, and Elie Track. Rebooting computing: The road ahead. Computer, 50(1):20– 29, January 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010020-abs.html.

Satyanarayanan:2017:EEC

[2156] Mahadev Satyanarayanan. The emergence of edge computing. Computer, 50(1):30–39, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/01/mco2017010030-abs.html.

Oulasvirta:2017:UID

[2157] Antti Oulasvirta. User interface design with combinatorial optimization. Computer, 50(1):40–47, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010040-abs.html.

Zhang:2017:TCS

[2158] Daqing Zhang, Hao Wang, and Dan Wu. Toward centimeter-scale human activity sensing with wi-fi signals. Computer, 50(1):48–57, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010048-abs.html.

Atkinson:2017:RBT

[2159] John Atkinson and Alvaro Maurelia. Redundancy-based trust in questionanswering systems. Computer, 50 (1):58–65, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/01/mco2017010058-abs.html.

Gangadharan:2017:OSS

[2160] G. R. Gangadharan. Open source solutions for cloud computing. Computer, 50(1):66–70, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/01/mco2017010066-abs.html.

Johnson:2017:IT

[2161] Brian David Johnson. The impossible thing. Computer, 50(1): 71, January 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010071.html.

Voas:2017:CCI

[2162] Jeffrey Voas and Philip Laplante. Curriculum considerations for the Internet of Things. Computer, 50(1): 72–75, January 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010072-abs.html.

Kshetri:2017:BA

[2163] Nir Kshetri and Jeffrey Voas. Banking on availability. Computer, 50 (1):76–80, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/01/mco2017010076-abs.html.

Anonymous:2017:MCH

[2164] Anonymous. My CS house advertisement. Computer, 50(1):81, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010081.pdf.

Berghel:2017:SSP

[2165] Hal Berghel. Software sophistry and political sleight of hand. Computer, 50(1):82–87, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/01/mco2017010082.html.

Anonymous:2017:CCa

[2166] Anonymous. CS connection. Computer, 50(1):88–90, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/01/mco2017010088.
html.

Anonymous:2017:SCS

[2167] Anonymous. Seymour Cray, Sidney Fernbach, & Ken Kennedy Awards house advertisement. Computer, 50 (1):91, January 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010091.pdf.

Anonymous:2017:CCb

[2168] Anonymous. Call and calendar. Computer, 50(1):92–94, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/01/mco2017010092.
html.

Anonymous:2017:CYI

[2169] Anonymous. Can you invent a better world through technologies? House advertisement. Computer, 50(1): 95, January 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010095.pdf.

Grier:2017:LCG

[2170] David Alan Grier. Local culture, global habits. Computer, 50(1):96, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco2017010096.html.

Anonymous:2017:FCa

[2171] Anonymous. Front cover. Computer, 50(1):c1, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/01/mco20170100c1.pdf.

Anonymous:2017:MYCa

[2172] Anonymous. Move your career forward IEEE Computer Society membership house advertisement. Computer, 50(1):c3, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/01/mco20170100c3.pdf.

Anonymous:2017:NMOa

[2173] Anonymous. New membership options for a better fit house advertisement. Computer, 50(1):c2, January 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
01/mco20170100c2.pdf.

Anonymous:2017:TRS

[2174] Anonymous. TechIgnite: A rock star technology event house advertisement. Computer, 50(1):c4, January 2017. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        01/mco20170100c4.pdf.

Anonymous:2017:ICSb

[2175] Anonymous. IEEE Computer Society is where you choose the resources that fit your career. Computer, 50 (2):1, February 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        02/mco2017020001.pdf.

Anonymous:2017:Mb

[2176] Anonymous. Masthead. Computer, 50 (2):2, February 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        02/mco2017020002.pdf.

Anonymous:2017:TCb

[2177] Anonymous. Table of contents. Computer, 50(2):3–4, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2017/02/mco2017020003.
                                        pdf.

McMillin:2017:FCS

[2178] Bruce McMillin and Tao Zhang. Fog computing for smart living. Computer, 50(2):5, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/02/mco2017020005.html.

Grier:2017:CFI

[2179] David Alan Grier. Computer’s first issues and the first volunteer. Computer, 50(2):6–7, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/02/mco2017020006.html.

Anonymous:2017:CHSb

[2180] Anonymous. Computer highlights Society magazines. Computer, 50(2): 8–9, February 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
02/mco2017020008.html.

Zientara:2017:TES

[2181] Peter A. Zientara, Sooyeon Lee, Gus H. Smith, Rorry Brenner, Laurent Itti, Mary B. Rosson, John M. Carroll, Kevin M. Irick, and Vijaykrishnan Narayanan. Third eye: A shopping assistant for the visually impaired. Computer, 50(2):16– 24, February 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
02/mco2017020016-abs.html.

Anonymous:2017:MH

[2182] Anonymous. myCS house. Computer, 50(2):25, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/02/mco2017020025.pdf.

Wahl:2017:PPS

[2183] Florian Wahl, Rui Zhang, Martin Freund, and Oliver Amft. Personalizing 3D-printed smart eyeglasses to augment daily life. Computer, 50 (2):26–35, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/02/mco2017020026-abs.html.

Zhang:2017:BBH

[2184] Cheng Zhang, Sinan Hersek, Yiming Pu, Danrui Sun, Qiuyue Xue, Thad E. Starner, Gregory D. Abowd, and Omer T. Inan. Bioacousticsbased human-body-mediated communication. Computer, 50(2):36–46, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
02/mco2017020036-abs.html.

Anonymous:2017:AICb

[2185] Anonymous. ACM–IEEE CS Eckert– Mauchly Award. Computer, 50(2): 47, February 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
02/mco2017020047.pdf.

Bhattacharyya:2017:DNP

[2186] Rajan Bhattacharyya, Brian A. Coffman, Jaehoon Choe, and Matthew E. Phillips. Does neurotechnology produce a better brain? Computer, 50 (2):48–58, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/02/mco2017020048-abs.html.

Jenson:2017:FIB

[2187] Scott Jenson. The future IoT: Building better Legos. Computer, 50(2):

68–71, February 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
02/mco2017020068-abs.html.

DeBenedictis:2017:TRM

[2188] Erik P. DeBenedictis. It’s time to redefine Moore’s Law again. Computer, 50(2):72–75, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/02/mco2017020072-abs.html.

Bertino:2017:BIT

[2189] Elisa Bertino and Nayeem Islam.

Botnets and Internet of Things security. Computer, 50(2):76–79, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
02/mco2017020076-abs.html.

Brown:2017:SLA

[2190] Barry Brown. The social life of autonomous cars. Computer, 50(2):92– 96, February 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
02/mco2017020092-abs.html.

Anonymous:2017:ICSc

[2191] Anonymous. IEEE Computer Society 2017 call for major award nominations. Computer, 50(2):98, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
02/mco2017020098.pdf.

Anonymous:2017:CNEa

[2192] Anonymous. Call for nominees education awards nominations house advertisement. Computer, 50(2):103, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        02/mco2017020103.pdf.

Anonymous:2017:MYCb

[2193] Anonymous. Move your career forward IEEE Computer Society membership. Computer, 50(2):c3, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        02/mco20170200c3.pdf.

Anonymous:2017:NMOb

[2194] Anonymous. New membership options for a better fit. Computer, 50(2):c2, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/02/mco20170200c2.pdf.

Anonymous:2017:Ta

[2195] Anonymous. TechIgnite. Computer, 50(2):c4, February 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/02/mco20170200c4.pdf.

Anonymous:2017:ICSd

[2196] Anonymous. IEEE Computer Society is where you choose the resources that fit your career. Computer, 50(3):1, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/03/mco2017030001.pdf.

Anonymous:2017:Mc

[2197] Anonymous. Masthead. Computer, 50(3):2, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030002.pdf.

Anonymous:2017:TCc

[2198] Anonymous. Table of contents. Computer, 50(3):3–4, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/03/mco2017030003.pdf.

Anonymous:2017:ICSe

[2199] Anonymous. IEEE Computer Society Richard E. Merwin Student Leadership Scholarship. Computer, 50 (3):5, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030005.pdf.

Grier:2017:WFW

[2200] David Alan Grier. When the field was young. Computer, 50(3):6–7, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030006.html.

Anonymous:2017:CHSc

[2201] Anonymous. Computer highlights Society magazines. Computer, 50 (3):8–9, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030008.html.

Anonymous:2017:YAa

[2202] Anonymous. 50 & 25 years ago. Computer, 50(3):10–11, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/03/mco2017030010.
html.

Sorin:2017:PM

[2203] Daniel J. Sorin. Persistent memory. Computer, 50(3):12, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/03/mco2017030012.
html.

Anonymous:2017:FYJb

[2204] Anonymous. Focus on your job search. Computer, 50(3):13, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/03/mco2017030013.
pdf.

Wac:2017:QLT

[2205] Katarzyna Wac, Homero Rivas, and Maddalena Fiordelli. Quality-of-life technologies. Computer, 50(3): 14–19, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030014.html.

Randriambelonoro:2017:CFT

[2206] Mirana Randriambelonoro, Yu Chen, and Pearl Pu. Can fitness trackers help diabetic and obese users make and sustain lifestyle changes? Computer, 50(3):20–29, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/03/mco2017030020-abs.html.

Sathyanarayana:2017:SSD

[2207] Aarti Sathyanarayana, Jaideep Srivastava, and Luis Fernandez-Luque. The science of sweet dreams: Predicting sleep efficiency from wearable device data. Computer, 50(3): 30–38, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030030-abs.html.

Anonymous:2017:CNEb

[2208] Anonymous. Call for nominees education awards nominations. Computer, 50(3):39, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/03/mco2017030039.pdf.

Zhang:2017:TPA

[2209] Tongda Zhang, Renate Fruchter, and Maria Frank. Are they paying attention? A model-based method to identify individuals’ mental states. Computer, 50(3):40–49, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030040.html.

Timmons:2017:UMW

[2210] Adela C. Timmons, Theodora Chaspari, Sohyun C. Han, Laura Perrone, Shrikanth S. Narayanan, and Gayla Margolin. Using multimodal wearable technology to detect conflict among couples. Computer, 50 (3):50–59, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        03/mco2017030050-abs.html.

Zarepour:2017:ACW

[2211] Eisa Zarepour, Mohammadreza Hosseini, Salil S. Kanhere, Arcot Sowmya, and Hamid R. Rabiee. Applications and challenges of wearable visual lifeloggers. Computer, 50(3): 60–69, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        03/mco2017030060-abs.html.

Walrad:2017:SEI

[2212] Chuck Walrad. Standards for the enterprise IT profession. Computer, 50(3):70–73, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/03/mco2017030070-abs.html.

Willis:2017:YNF

[2213] Scooter Willis. You never forget your first project. Computer, 50(3): 74–76, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        03/mco2017030074.html.

Anonymous:2017:NSS

[2214] Anonymous. Nominations are solicited for the Seymour Cray Sidney Fernbach & Ken Kennedy Awards house advertisement. Computer, 50 (3):77, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030077.pdf.

Stavrou:2017:VT

[2215] Angelos Stavrou and Jeffrey Voas.

Verified time. Computer, 50(3):78– 82, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bitcoin.

org/bitcoin.pdf;   https://www.
computer.org/csdl/mags/co/2017/
03/mco2017030078.html.

Anonymous:2017:CPY

[2216] Anonymous. Conferences in the palm of your hand. Computer, 50 (3):83, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030083.pdf.

Berghel:2017:PCA

[2217] Hal Berghel. On the problem of (cyber) attribution. Computer, 50(3): 84–89, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030084.html.

Brown:2017:NDW

[2218] Steve Brown. The new diversity:

Working with nonhumans. Computer, 50(3):90–91, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/03/mco2017030090.html.

Anonymous:2017:CCc

[2219] Anonymous. Call and calendar. Computer, 50(3):92–94, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/03/mco2017030092.
html.

Anonymous:2017:ICSf

[2220] Anonymous. IEEE Computer Society 2017 call for major award nominations. Computer, 50(3):95, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco2017030095.pdf.

Grier:2017:DT

[2221] David Alan Grier. Digital trade. Computer, 50(3):96, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/03/mco2017030096.html.

Anonymous:2017:FCb

[2222] Anonymous. Front cover. Computer, 50(3):c1, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/03/mco20170300c1.pdf.

Anonymous:2017:MYCc

[2223] Anonymous. Move your career forward IEEE Computer Society membership. Computer, 50(3):c3, March 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco20170300c3.pdf.

Anonymous:2017:NMOc

[2224] Anonymous. New membership options for a better fit. Computer, 50 (3):c2, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco20170300c2.pdf.

Anonymous:2017:Tb

[2225] Anonymous. TechIgnite. Computer, 50(3):c4, March 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
03/mco20170300c4.pdf.

Anonymous:2017:TCd

[2226] Anonymous. Table of contents. Computer, 50(4):1–2, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/04/mco2017040001.pdf.

Anonymous:2017:Md

[2227] Anonymous. Masthead. Computer, 50(4):3, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040003.pdf.

Anonymous:2017:EC

[2228] Anonymous. Elsewhere in the CS.

Computer, 50(4):4–5, April 2017.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/04/mco2017040004.
html.

Anonymous:2017:YAb

[2229] Anonymous. 50 & 25 years ago. Computer, 50(4):6, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/04/mco2017040006.html.

Dietrich:2017:CF

[2230] Sven Dietrich. Cybersecurity and the future. Computer, 50(4):7, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040007.html.

Morse:2017:ICF

[2231] Stephen P. Morse. The Intel 8086 chip and the future of microprocessor design. Computer, 50(4):8– 9, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040008.html.

Severance:2017:RCS

[2232] Charles Severance. Resurrecting the CDC 6500 supercomputer. Computer, 50(4):10–12, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/04/mco2017040010.html.

DeFloriani:2017:HLL

[2233] Leila De Floriani. A high-level language for interactive data visualization. Computer, 50(4):13, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040013.html.

Zanero:2017:CPS

[2234] Stefano Zanero. Cyber-physical systems. Computer, 50(4):14–16, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040014.html.

Howser:2017:UIF

[2235] Gerry Howser and Bruce McMillin. Using information-flow methods to analyze the security of cyber-physical systems. Computer, 50(4):17–26, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040017-abs.html.

Sanislav:2017:CIM

[2236] Teodora Sanislav, Sherali Zeadally, and George Dan Mois. A cloudintegrated, multilayered, agent-based cyber-physical system architecture. Computer, 50(4):27–37, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/04/mco2017040027-
abs.html.

Alippi:2017:FPS

[2237] Cesare Alippi and Manuel Roveri. The (not) far-away path to smart cyberphysical systems: An informationcentric framework. Computer, 50(4): 38–47, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040038-abs.html.

Oppliger:2017:NFA

[2238] Rolf Oppliger, Gunther Pernul, and Sokratis Katsikas. New frontiers: Assessing and managing security risks. Computer, 50(4):48– 51, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040048.html.

Wangen:2017:ISR

[2239] Gaute Wangen. Information security risk assessment: A method comparison. Computer, 50(4):52– 61, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040052-abs.html.

Rossebo:2017:ERA

[2240] Judith E. Y. Rossebo, Reinder Wolthuis, Frank Fransen, Gunnar Bjorkman, and Nuno Medeiros. An enhanced risk-assessment methodology for smart grids. Computer, 50 (4):62–71, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040062-abs.html.

Burnap:2017:DSR

[2241] Pete Burnap, Yulia Cherdantseva, Andrew Blyth, Peter Eden, Kevin Jones, Hugh Soulsby, and Kristan Stoddart. Determining and sharing risk data in distributed interdependent systems. Computer, 50(4): 72–79, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040072-abs.html.

Akhtar:2017:FFN

[2242] Zahid Akhtar and Ajita Rattani. A face in any form: New challenges and opportunities for face recognition technology. Computer, 50(4): 80–90, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040080-abs.html.

Montero:2017:ECN

[2243] Ruben S. Montero, Elisa Rojas, Alfonso A. Carrillo, and Ignacio M. Llorente. Extending the cloud to the network edge. Computer, 50(4): 91–95, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040091-abs.html.

DeBenedictis:2017:CAC

[2244] Erik P. DeBenedictis. Computer architecture’s changing role in rebooting computing. Computer, 50(4): 96–99, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        04/mco2017040096-abs.html.

Hurlburt:2017:SLD

[2245] George Hurlburt. Shining light on the Dark Web. Computer, 50(4): 100–105, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        04/mco2017040100-abs.html.

Cecchinato:2017:SDH

[2246] Marta E. Cecchinato and Anna L. Cox. Smartwatches: Digital handcuffs or magic bracelets? Computer, 50(4):106–109, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/04/mco2017040106-abs.html.

Berghel:2017:ANP

[2247] Hal Berghel. Alt-news and post-truths in the ‘fake news’ era. Computer, 50(4):110–114, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/04/mco2017040110.html.

Anonymous:2017:CCd

[2248] Anonymous. Call and calendar. Computer, 50(4):115–119, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2017/04/mco2017040115.
                                        html.

Grier:2017:RTI

[2249] David Alan Grier. The radical technology of Industrie 4.0. Computer, 50(4):120, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco2017040120.html.

Anonymous:2017:FCc

[2250] Anonymous. Front cover. Computer, 50(4):c1, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/04/mco20170400c1.pdf.

Anonymous:2017:Me

[2251] Anonymous. myCS. Computer, 50 (4):c4, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco20170400c4.pdf.

Anonymous:2017:MYCd

[2252] Anonymous. Move your career forward IEEE Computer Society membership. Computer, 50(4):c3, April 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco20170400c3.pdf.

Anonymous:2017:NMOd

[2253] Anonymous. New membership options for a better fit. Computer, 50(4):c2, April 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
04/mco20170400c2.pdf.

Anonymous:2017:TCe

[2254] Anonymous. Table of contents. Computer, 50(5):1–2, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/05/mco2017050001.pdf.

Anonymous:2017:Mf

[2255] Anonymous. Masthead. Computer, 50(5):3, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        05/mco2017050003.pdf.

Anonymous:2017:CHSd

[2256] Anonymous. Computer highlights Society magazines. Computer, 50 (5):4–5, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        05/mco2017050004.html.

Anonymous:2017:YAc

[2257] Anonymous. 50 & 25 years ago. Computer, 50(5):6–7, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/05/mco2017050006.html.

Lewis:2017:CC

[2258] Grace A. Lewis. Cloud computing. Computer, 50(5):8–9, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2017/05/mco2017050008.
                                        html.

DeMicheli:2017:NCY

[2259] Giovanni De Micheli and Luca Benini. Networks on chips: 15 years later. Computer, 50(5):10– 11, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050010.html.

Schuller:2017:CAC

[2260] Bjorn Schuller. Can affective computing save lives? Meet mobile health. Computer, 50(5):13, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050013.html.

Chang:2017:ALT

[2261] Carl K. Chang, Sorel Reisman, and Edmundo Tovar. Advances in learning technologies. Computer, 50(5): 14–17, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050014.html.

Bart:2017:BOA

[2262] Austin Cory Bart, Javier Tibau, Eli Tilevich, Clifford A. Shaffer, and Dennis Kafura. BlockPy: An open access data-science environment for introductory programmers. Computer, 50(5):18–26, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050018-abs.html.

Ghorashi:2017:ICL

[2263] Soroush Ghorashi and Carlos Jensen. Integrating collaborative and live coding for distance education. Computer, 50(5):27–35, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/05/mco2017050027-abs.html.

Rani:2017:SAS

[2264] Sujata Rani and Parteek Kumar. A sentiment analysis system to improve teaching and learning. Computer, 50 (5):36–43, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        05/mco2017050036-abs.html.

Ishida:2017:IED

[2265] Toru Ishida, Tetsuo Sawaragi, Kumiyo Nakakoji, and Takushi Sogo. Interdisciplinary education for design innovation. Computer, 50(5): 44–52, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        05/mco2017050044-abs.html.

Chan:2017:IER

[2266] Henry C. B. Chan. Internet of education resources using a chemistryinspired framework. Computer, 50 (5):54–60, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        05/mco2017050054-abs.html.

Xie:2017:SAI

[2267] Yang Xie, Chongxi Bao, and Ankur Srivastava. Security-aware 2.5D integrated circuit design flow against hardware IP piracy. Computer, 50 (5):62–71, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050062-abs.html.

Mann:2017:TEM

[2268] Zoltan Adam Mann. The top eight misconceptions about NP-hardness. Computer, 50(5):72–79, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/05/mco2017050072-abs.html.

Fairley:2017:BNT

[2269] Richard E. Fairley and Mary Jane Willshire. Better now than later: Managing technical debt in systems development. Computer, 50(5):80– 87, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050080-abs.html.

Voas:2017:WHS

[2270] Jeffrey Voas and Rick Kuhn. What happened to software metrics? Computer, 50(5):88–98, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/05/mco2017050088.html.

Johnson:2017:STF

[2271] Brian David Johnson. Sentient tools and the future of work. Computer, 50(5):99, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050099.html.

McGraw:2017:STT

[2272] Gary McGraw. Six tech trends impacting software security. Computer, 50(5):100–102, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/05/mco2017050100-abs.html.

Adrion:2017:HCS

[2273] W. Richards Adrion. How computer science departments and faculty can contribute to the CS for all initiative. Computer, 50(5):103– 105, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050103-abs.html.

Gray:2017:SAR

[2274] Andrew C. Gray and Eric M. Schwartz. A successful approach to the 2016 RobotX Challenge. Computer, 50(5): 106–109, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050106.html.

Valentino:2017:PPP

[2275] Nicholas A. Valentino, John Leslie King, and Walter W. Hill. Polling and prediction in the 2016 Presidential Election. Computer, 50(5): 110–115, May 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050110.html.

Bryson:2017:SED

[2276] Joanna Bryson and Alan Winfield.

Standardizing ethical design for artificial intelligence and autonomous systems. Computer, 50(5):116–119, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
05/mco2017050116.html.

Berghel:2017:QTW

[2277] Hal Berghel. A quick take on Windows security evolution. Computer, 50(5):120–124, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/05/mco2017050120.html.

Anonymous:2017:CCe

[2278] Anonymous. Call and calendar.

Computer, 50(5):125–129, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/05/mco2017050125.
html.

Anonymous:2017:CCf

[2279] Anonymous. CS connection. Computer, 50(5):130–131, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/05/mco2017050130.
html.

Grier:2017:RT

[2280] David Alan Grier. A Russian tragedy. Computer, 50(5):132, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        05/mco2017050132.html.

Anonymous:2017:FCd

[2281] Anonymous. Front cover. Computer, 50(5):c1, May 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/05/mco20170500c1.pdf.

Anonymous:2017:TCf

[2282] Anonymous. Table of contents. Computer, 50(6):1–2, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/06/mco2017060001.pdf.

Anonymous:2017:Mg

[2283] Anonymous. Masthead. Computer, 50(6):3, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        06/mco2017060003.pdf.

Garber:2017:CHS

[2284] Lee Garber. Computer highlights Society magazines. Computer, 50 (6):4–5, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        06/mco2017060004.html.

Jones:2017:HCN

[2285] Lynette A. Jones. A haptic compass for navigation. Computer, 50 (6):6, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060006.html.

Neuhold:2017:CJb

[2286] Erich Neuhold. Computer, June 1967 and 1992. Computer, 50(6): 7, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060007.html.

Snyder:2017:IE

[2287] Tom Snyder and Greg Byrd. The Internet of Everything. Computer, 50 (6):8–9, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060008.html.

Grier:2017:YY

[2288] David Alan Grier. The Yau years:

1981–1984. Computer, 50(6):10– 11, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060010.html.

Hill:2017:RAL

[2289] Mark D. Hill and Michael R. Marty.

Retrospective on Amdahl’s Law in the multicore era. Computer, 50 (6):12–14, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060012.html.

Anonymous:2017:Mi

[2290] Anonymous. myCS. Computer, 50 (6):15, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060015.pdf.

Wolf:2017:VIT

[2291] Marilyn Wolf and Saibal Mukhopadhyay. VLSI for the Internet of Things. Computer, 50(6):16–18, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060016.html.

Shen:2017:EHH

[2292] Chenguang Shen and Mani Srivastava. Exploring hardware heterogeneity to improve pervasive context inferences. Computer, 50(6): 19–26, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060019-abs.html.

Gao:2017:ACL

[2293] Mingze Gao, Qian Wang, Md Tanvir Arafin, Yongqiang Lyu, and Gang Qu. Approximate computing for low power and security in the Internet of Things. Computer, 50(6): 27–34, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060027-abs.html.

Soeken:2017:PCI

[2294] Mathias Soeken, Pierre-Emmanuel Gaillardon, Saeideh Shirinzadeh, Rolf Drechsler, and Giovanni De Micheli. A PLiM computer for the Internet of Things. Computer, 50(6): 35–40, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060035-abs.html.

Romani:2017:DLL

[2295] Aldo Romani, Marco Tartagni, and Enrico Sangiorgi. Doing a lot with a little: Micropower conversion and management for ambientpowered electronics. Computer, 50 (6):41–49, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060041-abs.html.

Gupta:2017:KSI

[2296] Gagan Gupta, Tony Nowatzki, Vinay Gangadhar, and Karthikeyan Sankaralingam. Kickstarting semiconductor innovation with open source hardware. Computer, 50(6):50–59, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060050-abs.html.

Eugster:2017:BDA

[2297] Patrick Eugster, Chamikara Jayalath, Kirill Kogan, and Julian Stephen. Big data analytics beyond the single datacenter. Computer, 50(6): 60–68, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        06/mco2017060060-abs.html.

Voas:2017:IBG

[2298] Jeffrey Voas and Phillip A. Laplante. The IoT blame game. Computer, 50 (6):69–73, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        06/mco2017060069-abs.html.

Davis:2017:MEW

[2299] Jennifer Davis. Marketing and engineering — working across the aisle. Computer, 50(6):74–75, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2017/06/mco2017060074.
                                        html.

DeBenedictis:2017:OCR

[2300] Erik P. DeBenedictis, Jesse K. Mee, and Michael P. Frank. The opportunities and controversies of reversible computing. Computer, 50 (6):76–80, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        06/mco2017060076-abs.html.

Huang:2017:WAS

[2301] Hsiu-Chuan Huang, Zhi-Kai Zhang, Hao-Wen Cheng, and Shiuhpyng Winston Shieh. Web application security: Threats, countermeasures, and pitfalls. Computer, 50(6):81– 85, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060081-abs.html.

Thimbleby:2017:III

[2302] Harold Thimbleby. Improve IT, improve healthcare. Computer, 50(6): 86–91, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
06/mco2017060086-abs.html.

Riley:2017:HAC

[2303] Justin Riley, John Noss, Wes Dillingham, James Cuff, and Ignacio M. Llorente. A high-availability cloud for research computing. Computer, 50(6):92–95, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/06/mco2017060092-abs.html.

Anonymous:2017:CCg

[2304] Anonymous. CS connection. Computer, 50(6):96–98, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/06/mco2017060096.html.

Anonymous:2017:CCh

[2305] Anonymous. Call and calendar.

Computer, 50(6):99–103, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/06/mco2017060099.
html.

Grier:2017:GC

[2306] David Alan Grier. Global connections. Computer, 50(6):104, June 2017. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        06/mco2017060104.html.

Anonymous:2017:FCe

[2307] Anonymous. Front cover. Computer, 50(6):c1, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/06/mco20170600c1.pdf.

Anonymous:2017:MYCe

[2308] Anonymous. Move your career forward IEEE Computer Society membership. Computer, 50(6):c3, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        06/mco20170600c3.pdf.

Anonymous:2017:NMOe

[2309] Anonymous. New membership options for a better fit. Computer, 50(6):c2, June 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        06/mco20170600c2.pdf.

Anonymous:2017:PCa

[2310] Anonymous. Prepare to connect. Computer, 50(6):c4, June 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/06/mco20170600c4.pdf.

Anonymous:2017:TCg

[2311] Anonymous. Table of contents. Computer, 50(7):1–2, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/07/mco2017070001.pdf.

Anonymous:2017:Mj

[2312] Anonymous. Masthead. Computer, 50(7):3, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070003.pdf.

Anonymous:2017:CHSe

[2313] Anonymous. Computer highlights Society magazines. Computer, 50 (7):4–5, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070004.html.

Anonymous:2017:NMO

[2314] Anonymous. New membership options for a better fit. Computer, 50(7):6, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070006.pdf.

Anonymous:2017:AYCa

[2315] Anonymous. Achieve your career goals with the fit that’s right for you. Computer, 50(7):7, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/07/mco2017070007.pdf.

Anonymous:2017:CJ

[2316] Anonymous. Computer, July 1967 and 1992. Computer, 50(7):8– 9, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070008.html.

Kazman:2017:SE

[2317] Rick Kazman. Software engineering. Computer, 50(7):10–11, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070010.html.

Henkel:2017:CEC

[2318] Jorg Henkel and Paolo Montuschi. Computer engineers’ challenges for the next decade: The triangle of power density, circuit degradation, and reliability. Computer, 50(7): 12, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070012.html.

Willis:2017:CBL

[2319] Scooter Willis, Greg Byrd, and Brian David Johnson. Challengebased learning. Computer, 50(7): 13–16, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070013.html.

Boyer:2017:ING

[2320] Nancy Boyer. Inspiring the next generation of scientists and engineers: K–12 and beyond. Computer, 50 (7):17–19, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070017.html.

Hitchcock:2017:GRS

[2321] Lindsey Hitchcock. Greenpower: Racing to a STEM finish. Computer, 50 (7):20–22, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070020.html.

Swearer:2017:IBC

[2322] Randy Swearer. Industry-backed competitions: Helping today’s students prepare for tomorrow’s careers. Computer, 50(7):23–25, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/07/mco2017070023.html.

Anonymous:2017:LCH

[2323] Anonymous. Low-cost high-altitude ballooning. Computer, 50(7):26– 28, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070026.html.

Scheid:2017:CBG

[2324] Randy Scheid. Code4GoodPBC:

Bridging the gap between innovation and philanthropy. Computer, 50 (7):29–31, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070029.html.

Thomas:2017:UDS

[2325] Richard S. Thomas. Using design slam to foster lifelong learning solutions. Computer, 50(7):32– 33, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        07/mco2017070032.html.

Anonymous:2017:CN

[2326] Anonymous. Call for nominees. Computer, 50(7):34, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/07/mco2017070034.pdf.

SanMiguel:2017:NCV

[2327] Juan C. SanMiguel and Andrea Cavallaro. Networked computer vision: The importance of a holistic simulator. Computer, 50(7):35– 43, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        07/mco2017070035-abs.html.

Liebig:2017:HSC

[2328] Jorg Liebig, Sven Apel, Andreas Janker, Florian Garbe, and Sebastian Oster. Handling static configurability in refactoring engines. Computer, 50(7):44–53, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/07/mco2017070044-abs.html.

Pasquale:2017:TAA

[2329] Liliana Pasquale, Carlo Ghezzi, Edoardo Pasi, Christos Tsigkanos, Menouer Boubekeur, Blanca Florentino-Liano, Tarik Hadzic, and Bashar Nuseibeh. Topology-aware access control of smart spaces. Computer, 50 (7):54–63, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070054-abs.html.

Allu:2017:BFH

[2330] Yamini Allu, Fred Douglis, Mahesh Kamat, Philip Shilane, Hugo Patterson, and Ben Zhu. Backup to the future: How workload and hardware changes continually redefine data domain file systems. Computer, 50 (7):64–72, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070064-abs.html.

Anonymous:2017:MYCf

[2331] Anonymous. Move your career forward. Computer, 50(7):73, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/07/mco2017070073.
pdf.

Salahuddin:2017:SIT

[2332] Mohammad A. Salahuddin, Ala Al-Fuqaha, Mohsen Guizani, Khaled Shuaib, and Farag Sallabi. Softwarization of Internet of Things infrastructure for secure and smart healthcare. Computer, 50(7):74–79, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/07/mco2017070074-abs.html.

Kolias:2017:DIM

[2333] Constantinos Kolias, Georgios Kambourakis, Angelos Stavrou, and Jeffrey Voas. DDoS in the IoT: Mirai and other botnets. Computer, 50 (7):80–84, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        07/mco2017070080-abs.html.

Vanatta:2017:EFE

[2334] Natalie Vanatta. Envisioning the future to empower action. Computer, 50(7):85, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        07/mco2017070085.html.

Berghel:2017:WMD

[2335] Hal Berghel. Which is more dangerous — the Dark Web or the Deep State? Computer, 50(7): 86–91, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        07/mco2017070086.html.

Crouch:2017:GAI

[2336] Al Crouch, Michael Laisne, and Martin Keim. Generalizing access to instrumentation embedded in a semiconductor device. Computer, 50(7): 92–95, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        07/mco2017070092.html.

Reisman:2017:TVR

[2337] Sorel Reisman. Teaching vs. research– optimizing your contribution for society’s well-being. Computer, 50(7): 96–98, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070096-abs.html.

Willis:2017:SSG

[2338] Scooter Willis. Stand on the shoulders of giants. Computer, 50(7): 99–102, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
07/mco2017070099.html.

Anonymous:2017:CCi

[2339] Anonymous. CS connection. Computer, 50(7):103–105, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/07/mco2017070103.
html.

Anonymous:2017:CCj

[2340] Anonymous. Call and calendar.

Computer, 50(7):106–111, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/07/mco2017070106.
html.

Grier:2017:CN

[2341] David Alan Grier. Connecting to Nigeria. Computer, 50(7):112, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        07/mco2017070112.html.

Anonymous:2017:FCf

[2342] Anonymous. Front cover. Computer, 50(7):c1, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/07/mco20170700c1.pdf.

Anonymous:2017:ICSg

[2343] Anonymous. IEEE Computer Society 2017 call for major award nominations. Computer, 50(7):c3, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        07/mco20170700c3.pdf.

Anonymous:2017:OMUa

[2344] Anonymous. One membership. Unlimited knowledge. Computer, 50 (7):c4, July 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        07/mco20170700c4.pdf.

Anonymous:2017:PCb

[2345] Anonymous. Prepare to connect. Computer, 50(7):c2, July 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/07/mco20170700c2.pdf.

Anonymous:2017:TCh

[2346] Anonymous. Table of contents. Computer, 50(8):1–2, August 2017.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/08/mco2017080001.
pdf.

Anonymous:2017:Mk

[2347] Anonymous. Masthead. Computer, 50(8):3, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco2017080003.pdf.

Grier:2017:BM

[2348] David Alan Grier. The better mix:

1985–1990. Computer, 50(8):4–5, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco2017080004.html.

Anonymous:2017:YAd

[2349] Anonymous. 50 & 25 years ago. Computer, 50(8):6–7, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco2017080006.html.

Guitton:2017:ICW

[2350] Pascal Guitton, Philippe Nain, and François Sillion. Inria: From Cold-War computing to digital sciences. Computer, 50(8):8–12, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco2017080008.html.

Anonymous:2017:ICSh

[2351] Anonymous. IEEE Computer Society information. Computer, 50 (8):13, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        08/mco2017080013.pdf.

DeBenedictis:2017:CDS

[2352] Erik P. DeBenedictis. Computer design starts over. Computer, 50(8): 14–17, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        08/mco2017080014.html.

Liu:2017:CAA

[2353] Shaoshan Liu, Jie Tang, Zhe Zhang, and Jean-Luc Gaudiot. Computer architectures for autonomous driving. Computer, 50(8):18–25, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        08/mco2017080018-abs.html.

Sen:2017:EPC

[2354] Rathijit Sen and David A. Wood. Energy-proportional computing: A new definition. Computer, 50(8): 26–33, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        08/mco2017080026-abs.html.

Xia:2017:VTP

[2355] Fei Xia, Ashur Rafiev, Ali Aalsaud, Mohammed Al-Hayanni, James Davis, Joshua Levine, Andrey Mokhov, Alexander Romanovsky, Rishad Shafik, Alex Yakovlev, and Sheng Yang. Voltage, throughput, power, reliability, and multicore scaling. Computer, 50 (8):34–45, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/08/mco2017080034-abs.html.

Nassi:2017:SCP

[2356] Ike Nassi. Scaling the computer to the problem: Application programming with unlimited memory. Computer, 50(8):46–51, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/08/mco2017080046-abs.html.

Teuscher:2017:WSU

[2357] Christof Teuscher. The weird, the small, and the uncontrollable: Redefining the frontiers of computing. Computer, 50(8):52–58, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco2017080052.html.

Anonymous:2017:ICSi

[2358] Anonymous. IEEE Computer Society Richard E. Merwin Student Leadership Scholarship. Computer, 50 (8):59, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco2017080059.pdf.

Stefik:2017:MIP

[2359] Andreas Stefik and Stefan Hanenberg. Methodological irregularities in programming-language research. Computer, 50(8):60–63, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        08/mco2017080060.html.

Hagel:2017:SSI

[2360] John Hagel and John Seely Brown. Shaping strategies for the IoT. Computer, 50(8):64–68, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2017/08/mco2017080064-
                                        abs.html.

Anonymous:2017:SML

[2361] Anonymous. Sustaining Moore’s law with 3D chips. Computer, 50(8): 69–73, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        08/mco2017080069-abs.html.

Netkachova:2017:CGY

[2362] Kate Netkachova and Robin Bloomfield. Is chocolate good for you — or, is the cloud secure? Computer, 50(8):74–78, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/08/mco2017080074-abs.html.

Anonymous:2017:UPE

[2363] Anonymous. Upsilon Pi Epsilon Student Excellence Award: Up to four $1,000 awards. Computer, 50 (8):79, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco2017080079.pdf.

Anonymous:2017:ICSj

[2364] Anonymous. IEEE Computer Society election. Computer, 50(8):80– 93, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco2017080080.html.

Anonymous:2017:IDV

[2365] Anonymous. IEEE Division VIII delegate/director candidates. Computer, 50(8):94–95, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/08/mco2017080094.html.

Anonymous:2017:IPE

[2366] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 50(8):96– 100, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco2017080096.html.

Anonymous:2017:CCk

[2367] Anonymous. CS connection. Computer, 50(8):101–103, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/08/mco2017080101.
html.

Grier:2017:LFT

[2368] David Alan Grier. The lingua franca of technology. Computer, 50(8):

104, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco2017080104.html.

Anonymous:2017:FCg

[2369] Anonymous. Front cover. Computer, 50(8):c1, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/08/mco20170800c1.pdf.

Anonymous:2017:LBTa

[2370] Anonymous. Looking for the BEST tech job for you? Computer, 50 (8):c2, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco20170800c2.pdf.

Anonymous:2017:MYCg

[2371] Anonymous. Move your career forward IEEE Computer Society membership. Computer, 50(8):c3, August 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco20170800c3.pdf.

Anonymous:2017:OMUb

[2372] Anonymous. One membership. Unlimited knowledge. Computer, 50 (8):c4, August 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
08/mco20170800c4.pdf.

Anonymous:2017:TCi

[2373] Anonymous. Table of contents. Computer, 50(9):1–2, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/09/mco2017090001.
pdf.

Anonymous:2017:Ml

[2374] Anonymous. Masthead. Computer, 50(9):3, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco2017090003.pdf.

Anonymous:2017:CHSf

[2375] Anonymous. Computer highlights Society magazines. Computer, 50(9): 4–7, September 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco2017090004.html.

DeMarle:2017:LUE

[2376] Ann DeMarle. Let us entertain you. Computer, 50(9):8–9, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco2017090008.html.

Magel:2017:RIA

[2377] Kenneth Magel. Revisiting the impact of the Ada programming language. Computer, 50(9):10–11, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco2017090010.html.

Anonymous:2017:YAe

[2378] Anonymous. 50 & 25 years ago. Computer, 50(9):12–13, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco2017090012.html.

Treleaven:2017:BTF

[2379] Philip Treleaven, Richard Gendal Brown, and Danny Yang. Blockchain technology in finance. Computer, 50 (9):14–17, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco2017090014.html.

Aste:2017:BTF

[2380] Tomaso Aste, Paolo Tasca, and Tiziana Di Matteo. Blockchain technologies: The foreseeable impact on society and industry. Computer, 50 (9):18–28, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco2017090018-abs.html.

Khan:2017:DLC

[2381] Chris Khan, Antony Lewis, Emily Rutland, Clemens Wan, Kevin Rutter, and Clark Thompson. A distributedledger consortium model for collaborative innovation. Computer, 50 (9):29–37, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco2017090029-abs.html.

Eyal:2017:BTT

[2382] Ittay Eyal. Blockchain technology:

Transforming libertarian cryptocurrency dreams to finance and banking realities. Computer, 50(9):38– 49, September 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco2017090038-abs.html.

Magazzeni:2017:VVS

[2383] Daniele Magazzeni, Peter McBurney, and William Nash. Validation and verification of smart contracts: A research agenda. Computer, 50(9):50– 57, September 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco2017090050-abs.html.

Taylor:2017:EBH

[2384] Michael Bedford Taylor. The evolution of Bitcoin hardware. Computer, 50(9):58–66, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco2017090058-abs.html.

Farahmand:2017:IHI

[2385] Fariborz Farahmand. The importance of human information processing: A behavioral economics model for predicting domain name choice. Computer, 50(9):67–74, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco2017090067-abs.html.

Anonymous:2017:SCa

[2386] Anonymous. Stay connected. Computer, 50(9):75, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2017/09/mco2017090075.
                                        pdf.

Ahmed:2017:TAF

[2387] Mohiuddin Ahmed. Thwarting DoS attacks: A framework for detection based on collective anomalies and clustering. Computer, 50(9):76– 82, September 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        09/mco2017090076-abs.html.

Embury:2017:SCC

[2388] Mark Embury and Jan Clayton. Standardization and certification challenges for biopharmaceutical plants. Computer, 50(9):83–86, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        09/mco2017090083.html.

Berghel:2017:OWT

[2389] Hal Berghel. Oh, what a tangled Web: Russian hacking, fake news, and the 2016 US Presidential election. Computer, 50(9):87– 91, September 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        09/mco2017090087.html.

Chang:2017:IFE

[2390] Chii Chang, Satish Narayana Srirama, and Rajkumar Buyya. Indie fog: An efficient fog-computing infrastructure for the Internet of Things. Computer, 50 (9):92–98, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco2017090092-abs.html.

Anonymous:2017:ICSk

[2391] Anonymous. IEEE Computer Society information. Computer, 50(9): 99, September 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco2017090099.pdf.

Chung:2017:ACT

[2392] Hyunji Chung, Michaela Iorga, Jeffrey Voas, and Sangjin Lee. “Alexa, can I trust you?”. Computer, 50(9): 100–104, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco2017090100-abs.html.

Grier:2017:BBH

[2393] David Alan Grier. Bringing the benefits home. Computer, 50(9): 105, September 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco2017090105.html.

Rakos:2017:PUT

[2394] Kyle Rakos, Alex Ruffino, Charles Li, Katherine Mao, Luke McBee, Ryan McBee, Matt Molo, Sanay Shah, Ian Sibley, and Tyler Stagge. Purdue University team tackles global underwatervehicle competition. Computer, 50(9): 106–110, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco2017090106.html.

Anonymous:2017:LBTb

[2395] Anonymous. Looking for the BEST tech job for you? Computer, 50 (9):111, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco2017090111.pdf.

Bennett:2017:MM

[2396] Julie Jenson Bennett. The meanings in the machine. Computer, 50 (9):112, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco2017090112.html.

Anonymous:2017:AYCb

[2397] Anonymous. Achieve your career goals with the fit that’s right for you. Computer, 50(9):c2, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco20170900c2.pdf.

Anonymous:2017:FCh

[2398] Anonymous. Front cover. Computer, 50(9):c1, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/09/mco20170900c1.pdf.

Anonymous:2017:MYCh

[2399] Anonymous. Move your career forward IEEE Computer Society membership. Computer, 50(9):c4, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco20170900c4.pdf.

Anonymous:2017:OMUc

[2400] Anonymous. One membership. Unlimited knowledge. Computer, 50(9): c3, September 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
09/mco20170900c3.pdf.

Anonymous:2017:TCj

[2401] Anonymous. Table of contents.

Computer, 50(10):1–2, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/10/mco2017100001.
pdf.

Anonymous:2017:Mm

[2402] Anonymous. Masthead. Computer, 50(10):3, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/10/mco2017100003.pdf.

Anonymous:2017:CHSg

[2403] Anonymous. Computer highlights Society magazines. Computer, 50(10): 4–7, October 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        10/mco2017100004.html.

Shneiderman:2017:RAG

[2404] Ben Shneiderman. Revisiting the astonishing growth of human–computer interaction research. Computer, 50 (10):8–11, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/10/mco2017100008.html.

Lomet:2017:FDM

[2405] David B. Lomet. The future of data management. Computer, 50(10):12– 13, October 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        10/mco2017100012.html.

Anonymous:2017:YAf

[2406] Anonymous. 50 & 25 years ago. Computer, 50(10):14–15, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        10/mco2017100014.html.

Oulasvirta:2017:SI

[2407] Antti Oulasvirta. On-skin interfaces. Computer, 50(10):16–18, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        10/mco2017100016.html.

Steimle:2017:SIU

[2408] Jurgen Steimle, Joanna Bergstrom-Lehtovirta, Martin Weigel, Aditya Shekhar Nittala, Sebastian Boring, Alex Olwal, and Kasper Hornbak. On-skin interaction using body landmarks. Computer, 50(10):19–27, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/10/mco2017100019-abs.html.

Lopes:2017:ISB

[2409] Pedro Lopes and Patrick Baudisch.

Interactive systems based on electrical muscle stimulation. Computer, 50 (10):28–35, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/10/mco2017100028-abs.html.

Satyanarayanan:2017:EC

[2410] Mahadev Satyanarayanan. Edge computing. Computer, 50(10):36– 38, October 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
10/mco2017100036.html.

Anonymous:2017:OMU

[2411] Anonymous. One membership. Unlimited knowledge. Computer, 50 (10):39, October 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
10/mco2017100039.pdf.

Klas:2017:ECR

[2412] Guenter Klas. Edge computing and the role of cellular networks. Computer, 50(10):40–49, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/10/mco2017100040-abs.html.

Liu:2017:LMN

[2413] Peng Liu, Lance Hartung, and Suman Banerjee. Lightweight multitenancy at the network’s extreme edge. Computer, 50(10):50–57, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/10/mco2017100050-abs.html.

Ananthanarayanan:2017:RTV

[2414] Ganesh Ananthanarayanan, Paramvir Bahl, Peter Bodik, Krishna Chintalapudi, Matthai Philipose, Lenin Ravindranath, and Sudipta Sinha. Realtime video analytics: The killer app for edge computing. Computer, 50 (10):58–67, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/10/mco2017100058-abs.html.

Berghel:2017:NNR

[2415] Hal Berghel. Net neutrality reloaded. Computer, 50(10):68–72, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/10/mco2017100068.
html.

Johnson:2017:WAG

[2416] Brian David Johnson. The weaponization of AI: A glimpse into future threats. Computer, 50(10):73, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
10/mco2017100073.html.

DeBenedictis:2017:SNR

[2417] Erik P. DeBenedictis. 3D software: A new research imperative. Computer, 50(10):74–77, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/10/mco2017100074-abs.html.

Voas:2017:HT

[2418] Jeffrey Voas and Nir Kshetri. Human tagging. Computer, 50(10):78– 85, October 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
10/mco2017100078-abs.html.

Charette:2017:LCA

[2419] Robert N. Charette. Low clearance ahead: Can predictable IT crashes be avoided? Computer, 50(10): 86–91, October 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
10/mco2017100086.html.

Tang:2017:EDL

[2420] Jie Tang, Dawei Sun, Shaoshan Liu, and Jean-Luc Gaudiot. Enabling deep learning on IoT devices. Computer, 50(10):92–96, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/10/mco2017100092-abs.html.

Anonymous:2017:ICSl

[2421] Anonymous. IEEE Computer Society. Computer, 50(10):97, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        10/mco2017100097.pdf.

Hislop:2017:HOS

[2422] Gregory W. Hislop and Heidi J. C. Ellis. Humanitarian open source software in computing education. Computer, 50 (10):98–101, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/10/mco2017100098-abs.html.

Schmidt:2017:TAM

[2423] Albrecht Schmidt. Technologies to amplify the mind. Computer, 50 (10):102–106, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/10/mco2017100102-abs.html.

Anonymous:2017:ITC

[2424] Anonymous. IEEE Transactions on Cloud Computing. Computer, 50(10): 107, October 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        10/mco2017100107.pdf.

Grier:2017:KIG

[2425] David Alan Grier. Keeping the Internet global. Computer, 50(10): 108, October 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        10/mco2017100108.html.

Anonymous:2017:FCi

[2426] Anonymous. Front cover. Computer, 50(10):c1, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/10/mco20171000c1.pdf.

Anonymous:2017:ICSm

[2427] Anonymous. IEEE Computer Society:

Be at the center of it all. Computer, 50(10):c2, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/10/mco20171000c2.pdf.

Anonymous:2017:LBTc

[2428] Anonymous. Looking for the BEST tech job for you? Computer, 50 (10):c3, October 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
10/mco20171000c3.pdf.

Anonymous:2017:PC

[2429] Anonymous. Prepare to connect. Computer, 50(10):c4, October 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
10/mco20171000c4.pdf.

Anonymous:2017:TCk

[2430] Anonymous. Table of contents. Computer, 50(11):1–2, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/11/mco2017110001.
pdf.

Anonymous:2017:Mn

[2431] Anonymous. [Masthead]. Computer, 50(11):3, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/11/mco2017110003.pdf.

Anonymous:2017:HSM

[2432] Anonymous. Highlights society magazines. Computer, 50(11):4–7, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        11/mco2017110004.html.

Leveson:2017:TYL

[2433] Nancy G. Leveson. The Therac-25: 30 years later. Computer, 50 (11):8–11, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/11/mco2017110008.html.

Anonymous:2017:YAg

[2434] Anonymous. 50 & 25 years ago. Computer, 50(11):12–13, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        11/mco2017110012.html.

Beigl:2017:PW

[2435] Michael Beigl, Florian Michahelles, Hide Tokuda, and Steve Hodges. Programming the world. Computer, 50 (11):14–17, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/11/mco2017110014.html.

Corno:2017:SWA

[2436] Fulvio Corno, Luigi De Russis, and Alberto Monge Roffarello. A Semantic Web approach to simplifying trigger-action programming in the IoT. Computer, 50(11):18–24, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
11/mco2017110018-abs.html.

Anonymous:2017:ICSn

[2437] Anonymous. IEEE Computer Society. Computer, 50(11):25, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
11/mco2017110025.pdf.

Hodges:2017:SDI

[2438] Jack Hodges, Kimberly Garcia, and Steven Ray. Semantic development and integration of standards for adoption and interoperability. Computer, 50 (11):26–36, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/11/mco2017110026-abs.html.

Anonymous:2017:ITSa

[2439] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 50(11):37, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/11/mco2017110037.pdf.

Kubitza:2017:MPP

[2440] Thomas Kubitza and Albrecht Schmidt. meSchup: A platform for programming interconnected smart things. Computer, 50(11):38–49, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        11/mco2017110038-abs.html.

Wolfe:2017:MPS

[2441] Henry B. Wolfe. The mobile phone as surveillance device: Progress, perils, and protective measures. Computer, 50 (11):50–58, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/11/mco2017110050-abs.html.

Anonymous:2017:SCb

[2442] Anonymous. Stay connected. Computer, 50(11):59, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2017/11/mco2017110059.
                                        pdf.

Bandari:2017:RSR

[2443] Roja Bandari, Zicong Zhou, Hai Qian, Timothy R. Tangherlini, and Vwani P. Roychowdhury. A resistant strain: Revealing the online grassroots rise of the antivaccination movement. Computer, 50(11):60–67, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/11/mco2017110060-abs.html.

Lam:2017:HST

[2444] Daniel Lam, Ethan Stillman, Anthony Cariello, Andrew McCorkle, and Ahmed Fouad. A high school team’s hard road to success in an international engineering competition. Computer, 50 (11):68–71, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/11/mco2017110068.html.

Berghel:2017:NSW

[2445] Hal Berghel. The new science wars. Computer, 50(11):72–76, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
11/mco2017110072.html.

Grier:2017:MNT

[2446] David Alan Grier. A more national technology. Computer, 50(11): 77, November 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
11/mco2017110077.html.

Grassi:2017:GAI

[2447] Paul Grassi, Mike Garcia, and Katie Boeckl. Government adopts an industry approach to open source collaboration. Computer, 50(11):78– 81, November 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
11/mco2017110078.html.

Young:2017:REE

[2448] Adam L. Young and Moti Yung. On ransomware and envisioning the enemy of tomorrow. Computer, 50 (11):82–85, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/11/mco2017110082-abs.html.

Endo:2017:MMC

[2449] Patricia Takako Endo, Guto Leoni Santos, Daniel Rosendo, Demis Moacir Gomes, Andre Moreira, Judith Kelner, Djamel Sadok, Glauco Estacio Goncalves, and Mozhgan Mahloo. Minimizing and managing cloud failures. Computer, 50(11):86–90, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
11/mco2017110086-abs.html.

Johnson:2017:FSE

[2450] Brian David Johnson. The future of shopping: An eerie silence? Computer, 50(11):91, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/11/mco2017110091.html.

Anonymous:2017:CCl

[2451] Anonymous. CS connection. Computer, 50(11):92–93, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/11/mco2017110092.
html.

Anonymous:2017:ITB

[2452] Anonymous. IEEE Transactions on Big Data. Computer, 50(11): 94, November 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
11/mco2017110094.pdf.

Anonymous:2017:ITM

[2453] Anonymous. IEEE Transactions on Multi-Scale Computing Systems. Computer, 50(11):95, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/11/mco2017110095.
pdf.

Brusilovsky:2017:TFM

[2454] Peter Brusilovsky. Toward flexible MOOCs: Student-sourcing of learning content at scale. Computer, 50 (11):96, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/11/mco2017110096.html.

Anonymous:2017:ICSo

[2455] Anonymous. IEEE Computer Society:

Be at the center of it all. Computer, 50(11):c2, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/11/mco20171100c2.pdf.

Anonymous:2017:LBTd

[2456] Anonymous. Looking for the BEST tech job for you? Computer, 50 (11):c4, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/11/mco20171100c4.pdf.

Anonymous:2017:Mo

[2457] Anonymous. myCS. Computer, 50 (11):c3, November 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/11/mco20171100c3.pdf.

Anonymous:2017:TCl

[2458] Anonymous. Table of contents. Computer, 50(12):1–2, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2017/12/mco2017120001.
                                        pdf.

Anonymous:2017:Mp

[2459] Anonymous. Masthead. Computer, 50(12):3, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/12/mco2017120003.pdf.

Gaudiot:2017:EC

[2460] Jean-Luc Gaudiot. Embracing changes. Computer, 50(12):4–6, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        12/mco2017120004.html.

Muller:2017:RIC

[2461] Hausi A. Muller. The rise of intelligent cyber-physical systems. Computer, 50 (12):7–9, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/12/mco2017120007.html.

Patterson:2017:RIS

[2462] David Patterson. Reduced instruction set computers then and now. Computer, 50(12):10–12, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
12/mco2017120010.html.

Anonymous:2017:CHSh

[2463] Anonymous. Computer highlights Society magazines. Computer, 50(12): 13–15, December 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
12/mco2017120013.html.

Anonymous:2017:YAh

[2464] Anonymous. 50 & 25 years ago. Computer, 50(12):16–17, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
12/mco2017120016.html.

Daily:2017:SDC

[2465] Mike Daily, Swarup Medasani, Reinhold Behringer, and Mohan Trivedi. Self-driving cars. Computer, 50(12): 18–23, December 2017. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
12/mco2017120018.html.

Dickmanns:2017:DSV

[2466] Ernst D. Dickmanns. Developing the sense of vision for autonomous road vehicles at UniBwM. Computer, 50 (12):24–31, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/12/mco2017120024-abs.html.

Hanna:2017:CUF

[2467] Mina J. Hanna and Shawn C. Kimmel. Current US Federal policy framework for self-driving vehicles: Opportunities and challenges. Computer, 50 (12):32–40, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/12/mco2017120032-abs.html.

Anonymous:2017:ICSp

[2468] Anonymous. IEEE Computer Society. Computer, 50(12):41, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        12/mco2017120041.pdf.

Liu:2017:UCP

[2469] Shaoshan Liu, Jie Tang, Chao Wang, Quan Wang, and Jean-Luc Gaudiot. A unified cloud platform for autonomous driving. Computer, 50 (12):42–49, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/12/mco2017120042-abs.html.

Ajmeri:2017:AEM

[2470] Nirav Ajmeri, Chung-Wei Hang, Simon D. Parsons, and Munindar P. Singh. Aragorn: Eliciting and maintaining secure service policies. Computer, 50(12):50–58, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
12/mco2017120050-abs.html.

Anonymous:2017:C

[2471] Anonymous. COMPSAC 2018. Computer, 50(12):59, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2017/12/mco2017120059.
pdf.

Knowles:2017:AGG

[2472] William Knowles, Jose M. Such, Antonios Gouglidis, Gaurav Misra, and Awais Rashid. All that glitters is not gold: On the effectiveness of cybersecurity qualifications. Computer, 50 (12):60–71, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/12/mco2017120060-abs.html.

Berghel:2017:ELR

[2473] Hal Berghel. Equifax and the latest round of identity theft roulette. Computer, 50(12):72–76, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
12/mco2017120072.html.

Arkenberg:2017:FIA

[2474] Chris Arkenberg. The future of intelligent assistants. Computer, 50 (12):77, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/12/mco2017120077.html.

Karakonstantis:2017:ERS

[2475] Georgios Karakonstantis, Dimitrios S. Nikolopoulos, Dimitris Gizopoulos, Pedro Trancoso, Yiannakis Sazeides, Christos D. Antonopoulos, Srikumar Venugopal, and Shidhartha Das. Errorresilient server ecosystems for edge and cloud datacenters. Computer, 50 (12):78–81, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/12/mco2017120078-abs.html.

Stankovic:2017:CCP

[2476] John A. Stankovic, James W. Sturges, and Jon Eisenberg. A 21st century cyber-physical systems education. Computer, 50(12):82–85, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2017/
                                        12/mco2017120082-abs.html.

DeBenedictis:2017:PSR

[2477] Erik P. DeBenedictis. Plotting a socially responsible course for computers using cybersecurity as an example. Computer, 50(12):86–90, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
12/mco2017120086-abs.html.

Kshetri:2017:HPG

[2478] Nir Kshetri and Jeffrey Voas. Hacking power grids: A current problem. Computer, 50(12):91–95, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
12/mco2017120091-abs.html.

Kott:2017:HDY

[2479] Alexander Kott and David S. Alberts.

How do you command an army of intelligent things? Computer, 50 (12):96–100, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2017/12/mco2017120096-abs.html.

Anonymous:2017:MVC

[2480] Anonymous. Metra voted 2018 Computer Society President-Elect. Computer, 50(12):101–103, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
12/mco2017120101.html.

Grier:2017:SC

[2481] David Alan Grier. The subway to China. Computer, 50(12):104, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2017/
12/mco2017120104.html.

Anonymous:2017:FCj

[2482] Anonymous. Front cover. Computer, 50(12):c1, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/12/mco20171200c1.pdf.

Anonymous:2017:ICSq

[2483] Anonymous. IEEE Computer Society: Be at the center of it all. Computer, 50(12):c2, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/12/mco20171200c2.pdf.

Anonymous:2017:ITSb

[2484] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 50(12):c3, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/12/mco20171200c3.pdf.

Anonymous:2017:LBTe

[2485] Anonymous. Looking for the BEST tech job for you? Computer, 50 (12):c4, December 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2017/12/mco20171200c4.pdf.

Anonymous:2018:TCa

[2486] Anonymous. Table of contents. Computer, 51(1):1–2, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2018/01/mco2018010001.
                                        pdf.

Anonymous:2018:Ma

[2487] Anonymous. Masthead. Computer, 51(1):3, January 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010003.pdf.

Kasahara:2018:SS

[2488] Hironori Kasahara. Satisfaction and sustainability. Computer, 51(1):4– 6, January 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010004.html.

Anonymous:2018:ITBa

[2489] Anonymous. IEEE Transactions on Big Data. Computer, 51(1):7, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010007.pdf.

Anonymous:2018:HSMa

[2490] Anonymous. Highlights Society magazines. Computer, 51(1):8–10, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010008.html.

Anonymous:2018:ICCa

[2491] Anonymous. IEEE Cloud Computing call for papers. Computer, 51 (1):11, January 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010011.pdf.

Anonymous:2018:YAa

[2492] Anonymous. 50 & 25 years ago. Computer, 51(1):12–13, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        01/mco2018010012.html.

Helal:2018:OEW

[2493] Sumi Helal. 2018 outlook: The emerging Web and information sciences. Computer, 51(1):14–16, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        01/mco2018010014.html.

Anonymous:2018:Mb

[2494] Anonymous. myCS. Computer, 51 (1):17, January 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        01/mco2018010017.pdf.

Hall:2018:BTS

[2495] Wendy Hall, Ramine Tinati, and Will Jennings. From Brexit to Trump: Social media’s role in democracy. Computer, 51(1):18–27, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        01/mco2018010018-abs.html.

Szpankowski:2018:FSI

[2496] Wojciech Szpankowski and Ananth Grama. Frontiers of science of information: Shannon meets Turing. Computer, 51(1):28–38, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010028-abs.html.

Anonymous:2018:ICSa

[2497] Anonymous. IEEE Computer Society. Computer, 51(1):39, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010039.pdf.

Bell:2018:RSC

[2498] C. Gordon Bell and Ike Nassi. Revisiting scalable coherent shared memory. Computer, 51(1):40–49, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010040-abs.html.

Mei:2018:TUO

[2499] Hong Mei and Yao Guo. Toward ubiquitous operating systems: A softwaredefined perspective. Computer, 51 (1):50–56, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/01/mco2018010050-abs.html.

Bruneo:2018:OAI

[2500] Dario Bruneo, Salvatore Distefano, Francesco Longo, Giovanni Merlino, and Antonio Puliafito. I/Ocloud: Adding an IoT dimension to cloud infrastructures. Computer, 51(1):57– 65, January 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010057-abs.html.

Chatterjee:2018:SIN

[2501] Tanusree Chatterjee, Sushmita Ruj, and Sipra Das Bit. Security issues in named data networks. Computer, 51(1):66–75, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/01/mco2018010066-abs.html.

Tingley:2018:TIC

[2502] Connor Tingley, George Serwin, Rohan Tuli, Jenya Kirsch Posner, Celvi Lisy, Lily Bresee, Keaton Viadro, and Chip Morimoto. Turning an international competition into a valuable learning experience. Computer, 51 (1):76–79, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/01/mco2018010076.html.

Voas:2018:ENG

[2503] Jeffrey Voas, Rick Kuhn, Celia Paulsen, and Kim Schaffer. Educating next-gen computer scientists. Computer, 51(1):80–88, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010080.html.

Berghel:2018:DDB

[2504] Hal Berghel. Disinformatics: The discipline behind grand deceptions. Computer, 51(1):89–93, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/01/mco2018010089.
html.

Vassilev:2018:PCY

[2505] Apostol Vassilev, Nicky Mouha, and Luis Brandao. Psst, can you keep a secret? Computer, 51(1):94– 97, January 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010094-abs.html.

Anonymous:2018:CCa

[2506] Anonymous. CS connection. Computer, 51(1):98–102, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/01/mco2018010098.
html.

Anonymous:2018:GF

[2507] Anonymous. Got flaws? Computer, 51(1):103, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/01/mco2018010103.pdf.

Shedroff:2018:SMR

[2508] Nathan Shedroff. Self-managing real estate. Computer, 51(1):104, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
01/mco2018010104.html.

Anonymous:2018:FCa

[2509] Anonymous. Front cover. Computer, 51(1):c1, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/01/mco20180100c1.pdf.

Anonymous:2018:ICSb

[2510] Anonymous. IEEE Computer Society: Be at the center of it all. Computer, 51(1):c2, January 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/01/mco20180100c2.pdf.

Anonymous:2018:ITS

[2511] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 51 (1):c3, January 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        01/mco20180100c3.pdf.

Anonymous:2018:LBTa

[2512] Anonymous. Looking for the BEST tech job for you? Computer, 51 (1):c4, January 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        01/mco20180100c4.pdf.

Anonymous:2018:TCb

[2513] Anonymous. Table of contents. Computer, 51(2):1–2, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2018/02/mco2018020001.
                                        pdf.

Anonymous:2018:Mc

[2514] Anonymous. Masthead. Computer, 51 (2):3, February 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
02/mco2018020003.pdf.

Anonymous:2018:HSMb

[2515] Anonymous. Highlights Society magazines. Computer, 51(2):4–7, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
02/mco2018020004.html.

Anonymous:2018:YAb

[2516] Anonymous. 50 & 25 years ago. Computer, 51(2):8–9, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
02/mco2018020008.html.

McMillin:2018:SE

[2517] Bruce McMillin. Software engineering. Computer, 51(2):10–12, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
02/mco2018020010.html.

Harel:2018:WCT

[2518] David Harel, Guy Katz, Rami Marelly, and Assaf Marron. Wise computing: Toward endowing system development with proactive wisdom. Computer, 51 (2):14–26, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/02/mco2018020014-abs.html.

Egyed:2018:MCA

[2519] Alexander Egyed, Klaus Zeman, Peter Hehenberger, and Andreas Demuth. Maintaining consistency across engineering artifacts. Computer, 51 (2):28–35, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/02/mco2018020028-abs.html.

Handler:2018:DSD

[2520] Ivan Handler. Data sharing defined — really! Computer, 51(2):36– 42, February 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        02/mco2018020036-abs.html.

Papadimitriou:2018:MVF

[2521] Stergios Papadimitriou and Lefteris Moussiades. Mac OS versus FreeBSD: A comparative evaluation. Computer, 51(2):44–53, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/02/mco2018020044-abs.html.

Beck:2018:BBR

[2522] Roman Beck. Beyond Bitcoin: The rise of blockchain world. Computer, 51(2):54–58, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/02/mco2018020054.html.

McGraw:2018:NKA

[2523] Gary McGraw. The new killer app for security: Software inventory. Computer, 51(2):60–62, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
02/mco2018020060-abs.html.

Berghel:2018:BSF

[2524] Hal Berghel. Bruce Schneier on future digital threats. Computer, 51 (2):64–67, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/02/mco2018020064.html.

DeBenedictis:2018:FQM

[2525] Erik P. DeBenedictis. A future with quantum machine learning. Computer, 51(2):68–71, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/02/mco2018020068-abs.html.

Raj:2018:TSU

[2526] Rajendra K. Raj and Allen Parrish. Toward standards in undergraduate cybersecurity education in 2018. Computer, 51(2):72–75, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
02/mco2018020072-abs.html.

Pal:2018:IBS

[2527] Amitangshu Pal and Krishna Kant.

IoT-based sensing and communications infrastructure for the fresh food supply chain. Computer, 51(2):76– 80, February 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        02/mco2018020076-abs.html.

Wright:2018:SEF

[2528] F. D. Wright. Standards: Engaging for the future. Computer, 51 (2):81–83, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/02/mco2018020081.html.

Anonymous:2018:CCb

[2529] Anonymous. CS connection. Computer, 51(2):84–87, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2018/02/mco2018020084.
                                        html.

Anonymous:2018:CWUa

[2530] Anonymous. Computer WRITE FOR US. Computer, 51(2):88, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        02/mco2018020088.pdf.

Anonymous:2018:FCb

[2531] Anonymous. Front cover. Computer, 51(2):c1, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/02/mco20180200c1.pdf.

Anonymous:2018:ICSc

[2532] Anonymous. IEEE Computer Society information. Computer, 51(2): c3, February 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
02/mco20180200c3.pdf.

Anonymous:2018:TCc

[2533] Anonymous. Table of contents. Computer, 51(3):1–2, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/03/mco2018030001.pdf.

Anonymous:2018:Md

[2534] Anonymous. Masthead. Computer, 51(3):3, March 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
03/mco2018030003.pdf.

Anonymous:2018:HSMc

[2535] Anonymous. Highlights Society magazines. Computer, 51(3):4–6, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
03/mco2018030004.html.

Anonymous:2018:ICSd

[2536] Anonymous. IEEE Computer Society. Computer, 51(3):7, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/03/mco2018030007.
pdf.

Sorin:2018:LPC

[2537] Daniel J. Sorin. Low-power content addressable memory. Computer, 51 (3):8–9, March 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
03/mco2018030008.html.

Anonymous:2018:YAc

[2538] Anonymous. 50 & 25 years ago. Computer, 51(3):10–11, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/03/mco2018030010.
html.

Banos:2018:CH

[2539] Oresti Banos and Christopher Nugent. E-coaching for health. Computer, 51(3):12–15, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/03/mco2018030012.html.

Ochoa:2018:ACS

[2540] Sergio F. Ochoa and Francisco J. Gutierrez. Architecting e-coaching systems: A first step for dealing with their intrinsic design complexity. Computer, 51(3):16–23, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/03/mco2018030016.html.

Robillard:2018:EMC

[2541] Julie M. Robillard and Jesse Hoey. Emotion and motivation in cognitive assistive technologies for dementia. Computer, 51(3):24–34, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
03/mco2018030024.html.

Anonymous:2018:IBAa

[2542] Anonymous. Impact a broader audience. Computer, 51(3):35, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
03/mco2018030035.pdf.

Pilloni:2018:RSL

[2543] Paolo Pilloni, Luca Piras, Salvatore Carta, Gianni Fenu, Fabrizio Mulas, and Ludovico Boratto. Recommender system lets coaches identify and help athletes who begin losing motivation. Computer, 51(3): 36–42, March 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
03/mco2018030036.html.

Parra:2018:MEB

[2544] Mario O. Parra, Jesus Favela, Luis A.

Castro, and Arturo Morales. Monitoring eating behaviors for a nutritionist e-assistant using crowdsourcing. Computer, 51(3):43–51, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
03/mco2018030043.html.

Lin:2018:PSN

[2545] Chit-Kwan Lin, Andreas Wild, Gautham N. Chinya, Yongqiang Cao, Mike Davies, Daniel M. Lavery, and Hong Wang. Programming spiking neural networks on Intel’s Loihi. Computer, 51(3):52–61, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/03/mco2018030052.html.

Willis:2018:MR

[2546] Scooter Willis. The maker revolution. Computer, 51(3):62–65, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        03/mco2018030062.html.

Berghel:2018:TP

[2547] Hal Berghel. Trolling pathologies. Computer, 51(3):66–69, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2018/03/mco2018030066.
                                        html.

Serpanos:2018:CPS

[2548] Dimitrios Serpanos. The cyber-physical systems revolution. Computer, 51(3): 70–73, March 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        03/mco2018030070.html.

Kshetri:2018:HBE

[2549] Nir Kshetri and Jeffrey Voas. Hacking’s brand-equity nexus. Computer, 51(3):74–77, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/03/mco2018030074.html.

Anonymous:2018:CCc

[2550] Anonymous. CS connection. Computer, 51(3):78–80, March 2018.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/03/mco2018030078.
html.

Anonymous:2018:CIa

[2551] Anonymous. Connect on interface. Computer, 51(3):c4, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/03/mco20180300c4.
pdf.

Anonymous:2018:FCc

[2552] Anonymous. Front cover. Computer, 51(3):c1, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/03/mco20180300c1.pdf.

Anonymous:2018:ICSe

[2553] Anonymous. IEEE Computer Society: Be at the center of it all. Computer, 51(3):c3, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/03/mco20180300c3.pdf.

Anonymous:2018:SGKa

[2554] Anonymous. Share the gift of knowledge: Give your favorite student a membership to the IEEE Computer Society! Computer, 51(3): c2, March 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
03/mco20180300c2.pdf.

Anonymous:2018:TCd

[2555] Anonymous. Table of contents. Computer, 51(4):1–2, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/04/mco2018040001.pdf.

Anonymous:2018:Me

[2556] Anonymous. Masthead. Computer, 51(4):3, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        04/mco2018040003.pdf.

Anonymous:2018:L

[2557] Anonymous. Letters. Computer, 51 (4):4–7, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        04/mco2018040004.html.

Anonymous:2018:CHSa

[2558] Anonymous. Computer highlights Society magazines. Computer, 51 (4):8–10, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        04/mco2018040008.html.

Anonymous:2018:ICSf

[2559] Anonymous. IEEE Computer Society information. Computer, 51 (4):11, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        04/mco2018040011.pdf.

DeMara:2018:NVM

[2560] Ronald F. DeMara and Paolo Montuschi. Non-volatile memory trends: Toward improving density and energy profiles across the system stack. Computer, 51(4):12–13, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
04/mco2018040012.html.

Anonymous:2018:YAd

[2561] Anonymous. 50 & 25 years ago.

Computer, 51(4):14–15, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/04/mco2018040014.
html.

Prince:2018:CSP

[2562] Daniel Prince. Cybersecurity: The security and protection challenges of our digital world. Computer, 51(4): 16–19, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
04/mco2018040016.html.

Bhattarai:2018:EET

[2563] Sulabh Bhattarai and Yong Wang.

End-to-end trust and security for Internet of Things applications. Computer, 51(4):20–27, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/04/mco2018040020-abs.html.

Ni:2018:GMB

[2564] Zhen Ni, Qianmu Li, and Gang Liu. Game-model-based network security risk control. Computer, 51 (4):28–38, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        04/mco2018040028-abs.html.

Anonymous:2018:IBAb

[2565] Anonymous. Impact a broader audience. Computer, 51(4):39, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        04/mco2018040039.pdf.

Lee:2018:DCR

[2566] Yongsuk Lee and Gyungho Lee. Detecting code reuse attacks with branch prediction. Computer, 51(4):40– 47, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        04/mco2018040040-abs.html.

Anonymous:2018:REHa

[2567] Anonymous. Recognizing excellence in high performance computing nominations are solicited for the Seymour Cray, Sidney Fernbach, and Ken Kennedy awards. Computer, 51 (4):48, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        04/mco2018040048.pdf.

Anonymous:2018:TCL

[2568] Anonymous. Take the CS Library wherever you go! Computer, 51 (4):49, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
04/mco2018040049.pdf.

Bouchard:2018:INI

[2569] Kristofer E. Bouchard, James B.

Aimone, Miyoung Chun, Thomas Dean, Michael Denker, Markus Diesmann, David D. Donofrio, Loren M. Frank, Narayanan Kasthuri, Christof Koch, Oliver Rubel, Horst D. Simon, F. T. Sommer, and Prabhat. International neuroscience initiatives through the lens of highperformance computing. Computer, 51(4):50–59, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/04/mco2018040050-abs.html.

Rubinov:2018:WWM

[2570] Konstantin Rubinov and Luciano Baresi. What are we missing when testing our Android apps? Computer, 51(4):60–68, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/04/mco2018040060-abs.html.

Anonymous:2018:ITBb

[2571] Anonymous. IEEE Transactions on Big Data. Computer, 51(4): 69, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
04/mco2018040069.pdf.

Berghel:2018:WTL

[2572] Hal Berghel. Weaponizing Twitter litter: Abuse-forming networks and social media. Computer, 51(4): 70–73, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
04/mco2018040070.html.

Conte:2018:RCA

[2573] Thomas M. Conte, Erik P. DeBenedictis, Avi Mendelson, and Dejan Milojicic. Rebooting computers to avoid Meltdown and Spectre. Computer, 51(4):74–77, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/04/mco2018040074-abs.html.

Hanna:2018:NNB

[2574] Mina J. Hanna. Net neutrality: A brief overview of the policy and the FCC’s ruling to upend it. Computer, 51(4):78–81, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/04/mco2018040078.html.

Chen:2018:PTI

[2575] Chung-Kuan Chen, Zhi-Kai Zhang, Shan-Hsin Lee, and Shiuhpyng Shieh. Penetration testing in the IoT Age. Computer, 51(4):82–85, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/04/mco2018040082-
abs.html.

Voas:2018:ICQ

[2576] Jeffrey Voas and Phillip A. Laplante. IoT’s certification quagmire. Computer, 51(4):86–89, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/04/mco2018040086-
abs.html.

Rosdahl:2018:CSS

[2577] Jon Rosdahl. Computer Society standards drive industry. Computer, 51 (4):90–94, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
04/mco2018040090.html.

Anonymous:2018:LBTb

[2578] Anonymous. Looking for the BEST tech job for you? Computer, 51 (4):95, April 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
04/mco2018040095.pdf.

Anonymous:2018:CCd

[2579] Anonymous. CS connection. Computer, 51(4):96, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/04/mco2018040096.html.

Anonymous:2018:CIb

[2580] Anonymous. Connect on interface. Computer, 51(4):c4, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/

                                        csdl/mags/co/2018/04/mco20180400c4.
                                        pdf.

Anonymous:2018:FCd

[2581] Anonymous. Front cover. Computer, 51(4):c1, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/04/mco20180400c1.pdf.

Anonymous:2018:ICSg

[2582] Anonymous. IEEE Computer Society: Be at the center of it all. Computer, 51(4):c3, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/04/mco20180400c3.pdf.

Anonymous:2018:SGKb

[2583] Anonymous. Share the gift of knowledge: Give your favorite student a membership to the IEEE Computer Society! Computer, 51(4): c2, April 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        04/mco20180400c2.pdf.

Anonymous:2018:TCe

[2584] Anonymous. Table of contents. Computer, 51(5):1–2, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/05/mco2018050001.pdf.

Anonymous:2018:Mf

[2585] Anonymous. Masthead. Computer, 51(5):3, May 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco2018050003.pdf.

Anonymous:2018:HSMd

[2586] Anonymous. Highlights Society magazines. Computer, 51(5):4–6, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco2018050004.html.

Anonymous:2018:CICa

[2587] Anonymous. 2018 CSI-IEEE CS Joint Education Award. Computer, 51(5):7, May 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco2018050007.pdf.

DeFloriani:2018:GBO

[2588] Leila De Floriani. Globe browsing in OpenSpace. Computer, 51 (5):8–9, May 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco2018050008.html.

Neuhold:2018:C

[2589] Erich Neuhold. Computer, 1968 and 1993. Computer, 51(5):10– 11, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco2018050010.html.

Lane:2018:DLT

[2590] Nicholas D. Lane and Pete Warden. The deep (learning) transformation of mobile and embedded computing. Computer, 51(5):12– 16, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco2018050012.html.

Anonymous:2018:RMAa

[2591] Anonymous. 2019 Richard E. Merwin Award for Distinguished Service. Computer, 51(5):17, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/05/mco2018050017.pdf.

Moshovos:2018:ETV

[2592] Andreas Moshovos, Jorge Albericio, Patrick Judd, Alberto Delmas Lascorz, Sayeh Sharify, Zissis Poulos, Tayler Hetherington, Tor Aamodt, and Natalie Enright Jerger. Exploiting typical values to accelerate deep learning. Computer, 51 (5):18–30, May 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco2018050018-abs.html.

Anonymous:2018:ICSh

[2593] Anonymous. IEEE Computer Society Harlan D. Mills Award. Computer, 51(5):31, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/05/mco2018050031.pdf.

Yao:2018:DLI

[2594] Shuochao Yao, Yiran Zhao, Aston Zhang, Shaohan Hu, Huajie Shao, Chao Zhang, Lu Su, and Tarek Abdelzaher. Deep learning for the Internet of Things. Computer, 51(5): 32–41, May 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco2018050032-abs.html.

Osia:2018:PSP

[2595] Seyed Ali Osia, Ali Shahin Shamsabadi, Ali Taheri, Hamid R. Rabiee, and Hamed Haddadi. Private and scalable personal data analytics using hybrid edge-to-cloud deep learning. Computer, 51(5):42–49, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/05/mco2018050042-abs.html.

Plotz:2018:DLH

[2596] Thomas Plotz and Yu Guan. Deep learning for human activity recognition in mobile computing. Computer, 51(5):50–59, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/05/mco2018050050-abs.html.

Chauhan:2018:BBA

[2597] Jagmohan Chauhan, Suranga Seneviratne, Yining Hu, Archan Misra, Aruna Seneviratne, and Youngki Lee. Breathing-based authentication on resource-constrained IoT devices using recurrent neural networks. Computer, 51(5):60–67, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/05/mco2018050060-abs.html.

Ahn:2018:FSB

[2598] Jungmo Ahn, Huynh Nguyen Loc, Rajesh Krishna Balan, Youngki Lee, and JeongGil Ko. Finding smallbowel lesions: Challenges in endoscopyimage-based learning systems. Computer, 51(5):68–76, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/05/mco2018050068-abs.html.

Anonymous:2018:CNEa

[2599] Anonymous. Call for nominees education awards nominations. Computer, 51(5):77, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/05/mco2018050077.pdf.

Charette:2018:PPP

[2600] Robert N. Charette. Puncturing pernicious project pufferies. Computer, 51(5):78–83, May 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        05/mco2018050078.html.

Berghel:2018:MDC

[2601] Hal Berghel. Malice domestic: The Cambridge Analytica dystopia. Computer, 51(5):84–89, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/05/mco2018050084.html.

Wolf:2018:CRW

[2602] Marilyn Wolf. Computing in the real world is the grandest of challenges. Computer, 51(5):90–91, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco2018050090-abs.html.

Kshetri:2018:CUB

[2603] Nir Kshetri and Jeffrey Voas. Cyberthreats under the bed. Computer, 51(5):92–95, May 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco2018050092-abs.html.

Anonymous:2018:ICSi

[2604] Anonymous. IEEE Computer Society. Computer, 51(5):96, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/05/mco2018050096.
pdf.

Anonymous:2018:FCe

[2605] Anonymous. Front cover. Computer, 51(5):c1, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/05/mco20180500c1.pdf.

Anonymous:2018:ICCb

[2606] Anonymous. 2019 IEEE-CS Charles Babbage Award. Computer, 51(5): c4, May 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
05/mco20180500c4.pdf.

Anonymous:2018:Mg

[2607] Anonymous. myCS. Computer, 51 (5):c2, May 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        05/mco20180500c2.pdf.

Anonymous:2018:REHb

[2608] Anonymous. Recognizing excellence in high performance computing nominations are solicited for the Seymour Cray Sidney Fernbach & Ken Kennedy Awards. Computer, 51 (5):c3, May 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        05/mco20180500c3.pdf.

Anonymous:2018:TCf

[2609] Anonymous. Table of contents. Computer, 51(6):1–2, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/06/mco2018060001.pdf.

Anonymous:2018:Mh

[2610] Anonymous. Masthead. Computer, 51(6):3, June 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        06/mco2018060003.pdf.

Anonymous:2018:HSMe

[2611] Anonymous. Highlights Society magazines. Computer, 51(6):4–6, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        06/mco2018060004.html.

Schuller:2018:WAC

[2612] Bjorn Schuller. What affective computing reveals about autistic children’s facial expressions of joy or fear. Computer, 51(6):7–8, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco2018060007.html.

Anonymous:2018:YAe

[2613] Anonymous. 50 & 25 years ago. Computer, 51(6):9–10, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco2018060009.html.

Anonymous:2018:LBTc

[2614] Anonymous. Looking for the BEST tech job for you? Computer, 51 (6):11, June 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
06/mco2018060011.pdf.

Hendler:2018:WSN

[2615] James Hendler, Wendy Hall, and Noshir Contractor. Web science: Now more than ever. Computer, 51(6): 12–17, June 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
06/mco2018060012.html.

Han:2018:UUP

[2616] Kyungsik Han, Hyunggu Jung, Jin Yea Jang, and Dongwon Lee. Understanding users’ privacy attitudes through subjective and objective assessments: An Instagram case study. Computer, 51(6):18–28, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco2018060018-abs.html.

Anonymous:2018:ICCc

[2617] Anonymous. 2019 IEEE-CS Charles Babbage Award. Computer, 51(6): 29, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
06/mco2018060029.pdf.

Gastil:2018:BCP

[2618] John Gastil and Sascha D. Meinrath. Bringing citizens and policymakers together online: Imagining the possibilities and taking stock of privacy and transparency hazards. Computer, 51(6):30–40, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco2018060030-abs.html.

Anonymous:2018:RMAb

[2619] Anonymous. 2019 Richard E. Merwin Award for Distinguished Service. Computer, 51(6):41, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco2018060041.pdf.

Nam:2018:HDO

[2620] Daye Nam and Mayank Kejriwal. How do organizations publish semantic markup? Three case studies using public Schema.org crawls. Computer, 51(6):42–51, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco2018060042-abs.html.

Sun:2018:USN

[2621] Jun Sun, Steffen Staab, and Jerome Kunegis. Understanding social networks using transfer learning. Computer, 51(6):52–60, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco2018060052-abs.html.

Anonymous:2018:ILCa

[2622] Anonymous. IEEE letters of the Computer Society. Computer, 51 (6):61, June 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
06/mco2018060061.pdf.

Maitland:2018:NYS

[2623] Carleen F. Maitland. Now you see it, now you don’t: Digital connectivity in marginalized communities. Computer, 51(6):62–71, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco2018060062-abs.html.

Ishida:2018:LSI

[2624] Toru Ishida, Yohei Murakami, Donghui Lin, Takao Nakaguchi, and Masayuki Otani. Language service infrastructure on the Web: The language grid. Computer, 51(6):72–81, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco2018060072-abs.html.

DeBenedictis:2018:AAC

[2625] Erik P. DeBenedictis. Accelerated architectures create programming opportunities. Computer, 51(6):82– 85, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        06/mco2018060082-abs.html.

Greene:2018:NPB

[2626] Kristen Greene, Michelle Steves, and Mary Theofanos. No phishing beyond this point. Computer, 51(6): 86–89, June 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        06/mco2018060086-abs.html.

Pering:2018:TIO

[2627] Trevor Pering, Kathy Farrington, and Thorsten Dahm. Taming the IoT: Operationalized testing to secure connected devices. Computer, 51(6): 90–94, June 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        06/mco2018060090-abs.html.

Anonymous:2018:ICSj

[2628] Anonymous. IEEE Computer Society. Computer, 51(6):95, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2018/06/mco2018060095.
                                        pdf.

Vedder:2018:BCD

[2629] Richard Vedder. Benefits of college— by degrees. Computer, 51(6):96– 99, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
06/mco2018060096.html.

Wright:2018:SRC

[2630] F. D. Wright and Thomas M.

Conte. Standards: Roadmapping computer technology trends enlightens industry. Computer, 51(6):100– 103, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
06/mco2018060100.html.

Anonymous:2018:CCe

[2631] Anonymous. CS connection. Computer, 51(6):104, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco2018060104.html.

Anonymous:2018:CNEb

[2632] Anonymous. Call for nominees education awards nominations. Computer, 51(6):c3, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco20180600c3.pdf.

Anonymous:2018:FCf

[2633] Anonymous. Front cover. Computer, 51(6):c1, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/06/mco20180600c1.pdf.

Anonymous:2018:ICSk

[2634] Anonymous. IEEE Computer Society Harlan D. Mills Award. Computer, 51(6):c4, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/06/mco20180600c4.pdf.

Anonymous:2018:SGKc

[2635] Anonymous. Share the gift of knowledge: Give your favorite student a membership to the IEEE Computer Society! Computer, 51(6): c2, June 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        06/mco20180600c2.pdf.

Anonymous:2018:TCg

[2636] Anonymous. Table of contents. Computer, 51(7):1–2, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/07/mco2018070001.pdf.

Anonymous:2018:Mi

[2637] Anonymous. Masthead. Computer, 51(7):3, July 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        07/mco2018070003.pdf.

Akmandor:2018:DAS

[2638] Ayten Ozge Akmandor and Niraj K. Jha. Detecting and alleviating stress with SoDA. Computer, 51(7):4– 5, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
07/mco2018070004.html.

Anonymous:2018:HSMf

[2639] Anonymous. Highlights Society magazines. Computer, 51(7):6– 8, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
07/mco2018070006.html.

Anonymous:2018:ICSl

[2640] Anonymous. IEEE Computer Society information. Computer, 51(7): 9, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
07/mco2018070009.pdf.

Anonymous:2018:YAf

[2641] Anonymous. 50 & 25 years ago. Computer, 51(7):10–11, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/07/mco2018070010.html.

Voas:2018:CIA

[2642] Jeffrey Voas, Rick Kuhn, Constantinos Kolias, Angelos Stavrou, and Georgios Kambourakis. Cybertrust in the IoT age. Computer, 51(7): 12–15, July 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
07/mco2018070012.html.

Roman-Castro:2018:ETI

[2643] Rodrigo Roman-Castro, Javier Lopez, and Stefanos Gritzalis. Evolution and trends in IoT security. Computer, 51(7):16–25, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/07/mco2018070016-abs.html.

Vlajic:2018:ILO

[2644] Natalija Vlajic and Daiwei Zhou. IoT as a land of opportunity for DDoS hackers. Computer, 51(7): 26–34, July 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
07/mco2018070026-abs.html.

Anonymous:2018:ICSm

[2645] Anonymous. IEEE Computer Society Harlan D. Mills Award. Computer, 51(7):35, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/07/mco2018070035.pdf.

Meng:2018:IDE

[2646] Weizhi Meng. Intrusion detection in the era of IoT: Building trust via traffic filtering and sampling. Computer, 51(7):36–43, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/07/mco2018070036-abs.html.

deFuentes:2018:ABC

[2647] Jose Maria de Fuentes, Lorena Gonzalez-Manzano, Agusti Solanas, and Fatbardh Veseli. Attributebased credentials for privacy-aware smart health services in IoT-based smart cities. Computer, 51(7):44– 53, July 2018. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
07/mco2018070044-abs.html.

Singh:2018:AIS

[2648] Jatinder Singh, Christopher Millard, Chris Reed, Jennifer Cobbe, and Jon Crowcroft. Accountability in the IoT: Systems, law, and ways forward. Computer, 51(7):54–65, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/07/mco2018070054-abs.html.

Kim:2018:MLM

[2649] Ryan Gary Kim, Janardhan Rao Doppa, Partha Pratim Pande, Diana Marculescu, and Radu Marculescu. Machine learning and manycore systems design: A serendipitous symbiosis. Computer, 51(7): 66–77, July 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
07/mco2018070066-abs.html.

Buzzelli:2018:UCF

[2650] Timothy Buzzelli. University of Central Florida takes the bronze at Annual International Programming Competition. Computer, 51(7):78– 81, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
07/mco2018070078.html.

Kuhn:2018:CRF

[2651] Rick Kuhn, Mohammad Raunak, and Raghu Kacker. Can reducing faults prevent vulnerabilities? Computer, 51(7):82–85, July 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        07/mco2018070082-abs.html.

Berghel:2018:NGI

[2652] Hal Berghel. Newton’s great insight. Computer, 51(7):86–90, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        07/mco2018070086.html.

Shiroishi:2018:SHS

[2653] Yoshihiro Shiroishi, Kunio Uchiyama, and Norihiro Suzuki. Society 5.0: For human security and well-being. Computer, 51(7):91–95, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/07/mco2018070091-abs.html.

Anonymous:2018:CCf

[2654] Anonymous. CS connection. Computer, 51(7):96, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/07/mco2018070096.html.

Anonymous:2018:FCg

[2655] Anonymous. Front cover. Computer, 51(7):c1, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/07/mco20180700c1.pdf.

Anonymous:2018:RMAc

[2656] Anonymous. 2019 Richard E. Merwin Award for Distinguished Service. Computer, 51(7):c4, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/07/mco20180700c4.pdf.

Anonymous:2018:SGKd

[2657] Anonymous. Share the gift of knowledge: Give your favorite student a membership to the IEEE Computer Society! Computer, 51(7): c2, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
07/mco20180700c2.pdf.

Anonymous:2018:TIC

[2658] Anonymous. TCHPC IEEE Computer Society Technical Consortium on High Performance Computing. Computer, 51(7):c3, July 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/07/mco20180700c3.pdf.

Anonymous:2018:TCh

[2659] Anonymous. Table of contents.

Computer, 51(8):1–2, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/08/mco2018080001.
pdf.

Anonymous:2018:Mj

[2660] Anonymous. Masthead. Computer, 51(8):3, August 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
08/mco2018080003.pdf.

Schaumont:2018:RHS

[2661] Patrick Schaumont and Paolo Montuschi. The rise of hardware security in computer architectures. Computer, 51(8):4–5, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/08/mco2018080004.html.

Anonymous:2018:CHSb

[2662] Anonymous. Computer highlights Society magazines. Computer, 51(8): 6–8, August 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
08/mco2018080006.html.

Anonymous:2018:CNEc

[2663] Anonymous. Call for nominees education awards nominations. Computer, 51(8):9, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/08/mco2018080009.pdf.

Anonymous:2018:CA

[2664] Anonymous. Computer, August 1968 and 1993. Computer, 51(8):10–12, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
08/mco2018080010.html.

Anonymous:2018:RMAd

[2665] Anonymous. 2019 Richard E. Merwin Award for Distinguished Service. Computer, 51(8):13, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/08/mco2018080013.pdf.

Banos:2018:MCT

[2666] Oresti Banos and Chris Nugent. Mcoaching: Towards the next generation of mobile-driven healthcare support services. Computer, 51(8): 14–17, August 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
08/mco2018080014.html.

Tschanz:2018:UEM

[2667] Mauro Tschanz, Tim Lucas Dorner, Jurgen Holm, and Kerstin Denecke. Using eMMA to manage medication. Computer, 51(8):18–25, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
08/mco2018080018-abs.html.

Wang:2018:SMB

[2668] Haolin Wang, Qingpeng Zhang, Mary Ip, and Joseph Tak Fai Lau. Social media-based conversational agents for health management and interventions. Computer, 51(8):26–33, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
08/mco2018080026-abs.html.

Hosio:2018:MDS

[2669] Simo Hosio, Jaro Karppinen, Niels van Berkel, Jonas Oppenlaender, and Jorge Goncalves. Mobile decision support and data provisioning for low back pain. Computer, 51(8): 34–43, August 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        08/mco2018080034-abs.html.

Berghel:2018:OTE

[2670] Hal Berghel and Daniel Berleant. The online trolling ecosystem. Computer, 51(8):44–51, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/08/mco2018080044.html.

DeBenedictis:2018:RIQ

[2671] Erik P. DeBenedictis. A role for IEEE in quantum computing. Computer, 51(8):52–55, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/08/mco2018080052-abs.html.

Isaak:2018:UDP

[2672] Jim Isaak and Mina J. Hanna. User data privacy: Facebook, Cambridge Analytica, and privacy protection. Computer, 51(8):56–59, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        08/mco2018080056.html.

Bloomfield:2018:SIS

[2673] Robin Bloomfield, Peter Bishop, Eoin Butler, and Robert Stroud. Securityinformed safety: Supporting stakeholders with codes of practice. Computer, 51(8):60–65, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/08/mco2018080060-abs.html.

Want:2018:AIL

[2674] Roy Want, Wei Wang, and Stan Chesnutt. Accurate indoor location for the IoT. Computer, 51(8): 66–70, August 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
08/mco2018080066-abs.html.

Anonymous:2018:ICSn

[2675] Anonymous. IEEE Computer Society. Computer, 51(8):71, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
08/mco2018080071.pdf.

Anonymous:2018:ICSo

[2676] Anonymous. IEEE Computer Society election. Computer, 51(8):72– 85, August 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
08/mco2018080072.html.

Anonymous:2018:IDV

[2677] Anonymous. IEEE Division V Delegate/Director candidates. Computer, 51(8):86–87, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/08/mco2018080086.html.

Anonymous:2018:IPE

[2678] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 51(8):88– 93, August 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        08/mco2018080088.html.

Anonymous:2018:CCg

[2679] Anonymous. CS connection. Computer, 51(8):94–96, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2018/08/mco2018080094.
                                        html.

Anonymous:2018:FCh

[2680] Anonymous. Front cover. Computer, 51(8):c1, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/08/mco20180800c1.pdf.

Anonymous:2018:ICCd

[2681] Anonymous. 2019 IEEE-CS Charles Babbage Award. Computer, 51(8): c4, August 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        08/mco20180800c4.pdf.

Anonymous:2018:ICSp

[2682] Anonymous. IEEE Computer Society: Be at the center of it all. Computer, 51(8):c2, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/08/mco20180800c2.pdf.

Anonymous:2018:ICSq

[2683] Anonymous. IEEE Computer Society Harlan D. Mills Award. Computer, 51(8):c3, August 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/08/mco20180800c3.pdf.

Anonymous:2018:TCi

[2684] Anonymous. Table of contents. Computer, 51(9):1–2, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/09/mco2018090001.
pdf.

Anonymous:2018:Mk

[2685] Anonymous. Masthead. Computer, 51(9):3, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/09/mco2018090003.pdf.

Anonymous:2018:CHS

[2686] Anonymous. Computer highlights Society magazines. Computer, 51(9): 4–6, September 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
09/mco2018090004.html.

Anonymous:2018:SC

[2687] Anonymous. Stay connected. Computer, 51(9):7, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2018/09/mco2018090007.
                                        pdf.

Vaidya:2018:PPS

[2688] Jaideep Vaidya. Privacy-preserving statistics. Computer, 51(9):8–9, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        09/mco2018090008.html.

Anonymous:2018:ILCb

[2689] Anonymous. IEEE letters of the Computer Society. Computer, 51(9): 10, September 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        09/mco2018090010.pdf.

Anonymous:2018:S

[2690] Anonymous. September, 1968 and 1993. Computer, 51(9):11–12, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        09/mco2018090011.html.

Anonymous:2018:CNEd

[2691] Anonymous. Call for nominees Education Awards nominations. Computer, 51(9):13, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/09/mco2018090013.pdf.

Helal:2018:EFA

[2692] Sumi Helal. The expanding frontier of artificial intelligence. Computer, 51 (9):14–17, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/09/mco2018090014.html.

Angelov:2018:TAM

[2693] Plamen P. Angelov and Xiaowei Gu. Toward anthropomorphic machine learning. Computer, 51(9):18– 27, September 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
09/mco2018090018-abs.html.

Hagras:2018:THU

[2694] Hani Hagras. Toward humanunderstandable, explainable AI. Computer, 51(9):28–36, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/09/mco2018090028-
abs.html.

Anonymous:2018:CICb

[2695] Anonymous. 2018 CSI-IEEE CS Joint Education Award. Computer, 51(9): 37, September 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
09/mco2018090037.pdf.

Schuller:2018:AAE

[2696] Dagmar Schuller and Bjorn W.

Schuller. The age of artificial emotional intelligence. Computer, 51(9): 38–46, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/09/mco2018090038-abs.html.

Anonymous:2018:ICSr

[2697] Anonymous. IEEE Computer Society Harlan D. Mills Award. Computer, 51(9):47, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/09/mco2018090047.pdf.

Dinh:2018:ABD

[2698] Thang N. Dinh and My T. Thai. AI and blockchain: A disruptive integration. Computer, 51(9):48– 53, September 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
09/mco2018090048-abs.html.

Bhat:2018:DCG

[2699] Wasim Ahmad Bhat. Is a datacapacity gap inevitable in big data storage? Computer, 51(9):54–62, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
09/mco2018090054-abs.html.

Anonymous:2018:LBTd

[2700] Anonymous. Looking for the BEST tech job for you? Computer, 51 (9):63, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/09/mco2018090063.pdf.

Hassan:2018:KPH

[2701] Mahbub Hassan, Wen Hu, Guhao Lan, Aruna Seneviratne, Sara Khalifa, and Sajal K. Das. Kineticpowered health wearables: Challenges and opportunities. Computer, 51(9): 64–74, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/09/mco2018090064-abs.html.

Anonymous:2018:ITB

[2702] Anonymous. IEEE Transactions on Big Data. Computer, 51(9): 75, September 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
09/mco2018090075.pdf.

Ernst:2018:ADC

[2703] Rolf Ernst. Automated driving: The cyber-physical perspective. Computer, 51(9):76–79, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/09/mco2018090076-abs.html.

Hurlburt:2018:TAC

[2704] George Hurlburt. Toward applied cyberethics. Computer, 51(9):80– 84, September 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
09/mco2018090080-abs.html.

Anonymous:2018:ICSs

[2705] Anonymous. IEEE Computer Society. Computer, 51(9):85, September 2018. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        09/mco2018090085.pdf.

Hudson:2018:WMI

[2706] Florence Hudson and Chris Clark. Wearables and medical interoperability: The evolving frontier. Computer, 51(9):86–90, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/09/mco2018090086.html.

Anonymous:2018:CWUb

[2707] Anonymous. Computer WRITE FOR US. Computer, 51(9):91, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        09/mco2018090091.pdf.

Anonymous:2018:CCh

[2708] Anonymous. CS connection. Computer, 51(9):92–96, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.org/
                                        csdl/mags/co/2018/09/mco2018090092.
                                        html.

Anonymous:2018:FCi

[2709] Anonymous. Front cover. Computer, 51(9):c1, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/09/mco20180900c1.pdf.

Anonymous:2018:ICCe

[2710] Anonymous. 2019 IEEE-CS Charles Babbage Award. Computer, 51(9): c4, September 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
09/mco20180900c4.pdf.

Anonymous:2018:ICSt

[2711] Anonymous. IEEE COMPUTER SOCIETY: Be at the center of it all. Computer, 51(9):c2, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
09/mco20180900c2.pdf.

Anonymous:2018:RMA

[2712] Anonymous. 2019 Richard E. Merwin Award for Distinguished Service. Computer, 51(9):c3, September 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
09/mco20180900c3.pdf.

Anonymous:2018:TCj

[2713] Anonymous. Table of contents. Computer, 51(11):1–2, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.org/
csdl/mags/co/2018/11/08625921.pdf.

Anonymous:2018:SL

[2714] Anonymous. Staff list. Computer, 51(11):3, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/11/08625937.pdf.

Cameron:2018:ECa

[2715] Lori Cameron. [Elsewhere in the CS]. Computer, 51(11):4–6, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
11/08625934.pdf.

Neuhold:2018:YA

[2716] Erich Neuhold. [50 & 25 years ago]. Computer, 51(11):7–9, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
11/08625940.pdf.

Sokolsky:2018:RCP

[2717] Oleg Sokolsky and Nicola Bezzo. Resiliency in cyber-physical systems [Guest Editors’ introduction]. Computer, 51(11):10–12, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
11/08625939.pdf.

Cofer:2018:FAC

[2718] Darren Cofer, Andrew Gacek, John Backes, Michael W. Whalen, Lee Pike, Adam Foltzer, Michal Podhradsky, Gerwin Klein, Ihor Kuz, June Andronick, Gernot Heiser, and Douglas Stuart. A formal approach to constructing secure air vehicle software. Computer, 51 (11):14–23, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/11/08625938-abs.html.

Abdelzaher:2018:TIB

[2719] Tarek Abdelzaher, Nora Ayanian, Tamer Basar, Suhas Diggavi, Jana Diesner, Deepak Ganesan, Ramesh Govindan, Susmit Jha, Tancrede Lepoint, Benjamin Marlin, Klara Nahrstedt, David Nicol, Raj Rajkumar, Stephen Russell, Sanjit Seshia, Fei Sha, Prashant Shenoy, Mani Srivastava, Gaurav Sukhatme, Ananthram Swami, Paulo Tabuada, Don Towsley, Nitin Vaidya, and Venu Veeravalli. Toward an Internet of Battlefield Things: A resilience perspective. Computer, 51 (11):24–36, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/11/08625936-abs.html.

Krishna:2018:CPF

[2720] Varun Badrinath Krishna, Ziping Wu, Vaidehi V. Ambardekar, Richard Macwan, and William H. Sanders. Cyberattacks on primary frequency response mechanisms in power grids. Computer, 51(11):37–45, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
11/08625915-abs.html.

Won:2018:ARC

[2721] Yuchang Won, Buyeon Yu, Jaegeun Park, In-Hee Park, Haegeon Jeong, Jeanseong Baik, Kyungtae Kang, Insup Lee, Sang Hyuk Son, Kyung-Joon Park, and Yongsoon Eun. An attack-resilient CPS architecture for hierarchical control: A case study on train control systems. Computer, 51 (11):46–55, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/11/08625932-abs.html.

Haque:2018:CBH

[2722] Mohammad Shihabul Haque, Daniel Jun Xian Ng, Arvind Easwaran, and Karthikeyan Thangamariappan. Contract-based hierarchical resilience management for cyber-physical systems. Computer, 51(11):56–65, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        11/08625911-abs.html.

Kumar:2018:SGF

[2723] Tanesh Kumar, Pawani Porambage, Ijaz Ahmad, Madhusanka Liyanage, Erkki Harjula, and Mika Ylianttila. Securing gadget-free digital services. Computer, 51(11):66–77, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        11/08625918-abs.html.

Halle:2018:SIC

[2724] Sylvain Halle, Raphael Khoury, and Mewena Awesso. Streamlining the inclusion of computer experiments in a research paper. Computer, 51(11):78– 89, November 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        11/08625929-abs.html.

Bergel:2018:EIW

[2725] Hal Bergel. Election integrity 2018: Are we better off than we were in 2016? [Aftershock]. Computer, 51 (11):90–94, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//www.computer.org/csdl/mags/co/
2018/11/08625933.pdf.

Koulamas:2018:CPS

[2726] Christos Koulamas and Athanasios Kalogeras. Cyber-physical systems and digital twins in the industrial Internet of Things [cyber-physical systems]. Computer, 51(11):95–98, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
11/08625931.pdf.

Kshetri:2018:CTV

[2727] Nir Kshetri. Cryptocurrencies: Transparency versus privacy [cybertrust]. Computer, 51(11):99–111, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
11/08625935.pdf.

Berghel:2018:GPD

[2728] Hal Berghel. The genesis of postmodern disinformatics [out of band]. Computer, 51(11):102–105, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

computer.org/csdl/mags/co/2018/
11/08625927.pdf.

Lima:2018:DOI

[2729] Claudio Lima. Developing open and interoperable DLT/blockchain standards [standards]. Computer, 51(11):

106–111, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/11/08625908.pdf.

Anonymous:2018:ICSu

[2730] Anonymous. IEEE Computer Society information. Computer, 51(11): 112, November 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https://www.

                                        computer.org/csdl/mags/co/2018/
                                        11/08625930.pdf.

Anonymous:2018:FCj

[2731] Anonymous. Front cover. Computer, 51(11):C1, November 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/mags/co/
                                        2018/11/08626741.pdf.

Anonymous:2018:FCk

[2732] Anonymous. Front cover. Computer, 51 (12):??, December 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2018:TCk

[2733] Anonymous. Table of contents. Computer, 51(12):1–2, December 2018. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2018:M

[2734] Anonymous. Masthead. Computer, 51 (12):??, December 2018. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2018:ECb

[2735] Lori Cameron. [Elsewhere in the CS].

Computer, 51(12):4–6, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2018:CDY

[2736] Erich Neuhold. Computer, December 1968 and 1993 [50 & 25 years ago]. Computer, 51(12):7–9, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:2018:GAB

[2737] Yasha Wang and Katsuyuki Yamazaki.

Governments in the age of big data and smart cities [Guest Editors’ introduction]. Computer, 51(12):10–12, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8637078/.

Anonymous:2018:ICSv

[2738] Anonymous. IEEE Computer Society information. Computer, 51(12):??, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8636450/.

Sumitomo:2018:PIO

[2739] Takahiro Sumitomo and Noboru Koshizuka. Progress and initiatives for open data policy in Japan. Computer, 51(12):14–23, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8636472/.

Gao:2018:HVC

[2740] Hui Gao, Chi Harold Liu, and Wendong Wang. Hybrid vehicular crowdsourcing with driverless cars: Challenges and a solution. Computer, 51 (12):24–31, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8636464/.

Wang:2018:SCD

[2741] Leye Wang, Bin Guo, and Qiang Yang. Smart city development with urban transfer learning. Computer, 51 (12):32–41, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8636587/.

Wang:2018:EHC

[2742] Jiangtao Wang, Junfeng Zhao, Yong Zhang, Xin Peng, Ying Li, and Yun Xie. Enabling human-centric smart cities: Crowdsourcing-based practice in China. Computer, 51(12):42–49, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8636468/.

Koshizuka:2018:CIE

[2743] Noboru Koshizuka, Stephan Haller, and Ken Sakamura. CPaaS.io: An EU– Japan collaboration on open smartcity platforms. Computer, 51(12): 50–58, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8636448/.

Kshetri:2018:BEH

[2744] Nir Kshetri. Blockchain and electronic healthcare records [Cybertrust]. Computer, 51(12):59–63, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8636458/.

Wang:2018:MIT

[2745] Ju Wang, Xiaojiang Chen, Dingyi Fang, Jie Xiong, Hongbo Jiang, and Rajesh Krishna Balan. Material identification and target imaging with RFIDs [IoT connection]. Computer, 51 (12):64–68, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8636463/.

DeBenedictis:2018:NQI

[2746] Erik P. DeBenedictis and Michael P.

Frank. The national quantum initiative will also benefit classical computers [rebooting computing]. Computer, 51(12):69–73, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8636459/.

Anonymous:2018:FVI

[2747] Anonymous. De Floriani voted 2019 IEEE Computer Society President-Elect [report to members]. Computer, 51(12):74–76, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8636484/.

Anonymous:2018:CC

[2748] Anonymous. [CS connection]. Computer, 51(12):78–80, December 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        https://ieeexplore.ieee.org/document/
                                        8636476/.

Anonymous:2019:TCa

[2749] Anonymous. Table of contents. Computer, 52(1):1–2, January 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Ma

[2750] Anonymous. Masthead. Computer, 52(1):3, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //ieeexplore.ieee.org/document/
                                        8666443/.

Metra:2019:ICSa

[2751] Cecilia Metra. The 2019 IEEE Computer Society: Targeting member satisfaction and technical excellence. Computer, 52(1):4–6, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL https://ieeexplore.
                                        ieee.org/document/8666651/.

Cameron:2019:CHSa

[2752] Lori Cameron. Computer highlights Society magazines. Computer, 52(1): 7–9, January 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2019:YAa

[2753] Erich Neuhold. 50 & 25 years ago. Computer, 52(1):10–12, January 2019.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:ROE

[2754] David Alan Grier. A revolutionary outlook [EIC’s message]. Computer, 52 (1):13–14, January 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISCa

[2755] Anonymous. IEEE Software call for articles. Computer, 52(1):14, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSa

[2756] Anonymous. IEEE Computer Society. Computer, 52(1):15, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8666656/.

Bertino:2019:RDT

[2757] Elisa Bertino, Shawn Merrill, Alina Nesen, and Christine Utz. Redefining data transparency: A multidimensional approach. Computer, 52(1):16–26, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8666655/.

Shneiderman:2019:IVE

[2758] Ben Shneiderman and Catherine Plaisant. Interactive visual event analytics: Opportunities and challenges. Computer, 52(1):27–35, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8666650/.

Wang:2019:ASS

[2759] Dong Wang, Boleslaw K. Szymanski, Tarek Abdelzaher, Heng Ji, and Lance Kaplan. The age of social sensing. Computer, 52(1):36–45, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL https://ieeexplore.
                                        ieee.org/document/8666667/.

Wang:2019:CAM

[2760] Jiangtao Wang, Yasha Wang, and Qin Lv. Crowd-assisted machine learning: Current issues and future directions. Computer, 52(1):46–53, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL https://ieeexplore.
                                        ieee.org/document/8666665/.

Wolf:2019:CSC

[2761] Marilyn Wolf. Computer security as civil defense. Computer, 52(1): 54–57, January 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //ieeexplore.ieee.org/document/
                                        8666652/.

Anonymous:2019:ETH

[2762] Anonymous. Erratum [for “Toward Human-Understandable, Explainable AI”]. Computer, 52(1):57, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL https://ieeexplore.
                                        ieee.org/document/8666659/.

Kshetri:2019:OAF

[2763] Nir Kshetri and Jeffrey Voas. Online advertising fraud. Computer, 52 (1):58–61, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8666661/.

Anonymous:2019:ITBa

[2764] Anonymous. IEEE Transactions on Big Data. Computer, 52(1):61, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dai:2019:UCR

[2765] Lu Dai and Stan Krolikoski. A unique collaboration: revitalizing existing standards while preparing for the future. Computer, 52(1):62–65, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8666653/.

Anonymous:2019:ITSa

[2766] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 52 (1):65, January 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2019:VTG

[2767] Hal Berghel. Vehicle telematics: The good, bad and ugly. Computer, 52 (1):66–70, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8666649/.

Anonymous:2019:BRRa

[2768] Anonymous. B. Ramakrishna Rau Award. Computer, 52(1):71, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kasahara:2019:CF

[2769] Hironori Kasahara. Collaboration for the future. Computer, 52(1): 72–76, January 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //ieeexplore.ieee.org/document/
                                        8666648/.

Anonymous:2019:FCa

[2770] Anonymous. Front cover. Computer, 52(1):C1, January 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:AIC

[2771] Anonymous. ACM/IEEE CS Eckert– Mauchly Award. Computer, 52(1): C2, January 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSb

[2772] Anonymous. IEEE Computer Society interface. Computer, 52(1):C3, January 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSc

[2773] Anonymous. IEEE Computer Society jobs board. Computer, 52(1): C4, January 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TCb

[2774] Anonymous. Table of contents. Computer, 52(2):1–2, February 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Mb

[2775] Anonymous. Masthead. Computer, 52 (2):3, February 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schuller:2019:MCC

[2776] Bjoern Schuller. Microexpressions:

A chance for computers to beat humans at detecting hidden emotions? Computer, 52(2):4–5, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISPa

[2777] Anonymous. IEEE Security & Privacy subscription. Computer, 52(2): 5, February 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2019:CHSb

[2778] Lori Cameron. Computer highlights Society magazines. Computer, 52(2): 6–8, February 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2019:YAb

[2779] Erich Neuhold. 50 & 25 years ago.

Computer, 52(2):9–11, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITBb

[2780] Anonymous. IEEE Transactions on Big Data. Computer, 52(2):11, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:HUT

[2781] David Alan Grier. Holding us together. Computer, 52(2):12–13, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CAS

[2782] Anonymous. Call for articles software magazine. Computer, 52(2): 13, February 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Antinyan:2019:PVC

[2783] Vard Antinyan, Anna B. Sandberg, and Miroslaw Staron. A pragmatic view on code complexity management. Computer, 52(2):14–22, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8672426/.

Anonymous:2019:IPCa

[2784] Anonymous. IT Professional call for articles. Computer, 52(2):22, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hole:2019:SSA

[2785] Kjell J. Hole and Christian Otterstad. Software systems with antifragility to downtime. Computer, 52 (2):23–31, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8672419/.

Anonymous:2019:IPCb

[2786] Anonymous. IEEE Pervasive Computing call for articles. Computer, 52 (2):31, February 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Treleaven:2019:ALR

[2787] Philip Treleaven, Jeremy Barnett, and Adriano Koshiyama. Algorithms: Law and regulation. Computer, 52 (2):32–40, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8672418/.

Siddique:2019:KCD

[2788] K. Siddique, Z. Akhtar, F. Aslam Khan, and Y. Kim. KDD Cup 99 data sets: A perspective on the role of data sets in network intrusion detection research. Computer, 52(2): 41–51, February 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICGa

[2789] Anonymous. IEEE Computer Graphics & Applications subscription. Computer, 52(2):51, February 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bresniker:2019:RSC

[2790] Kirk M. Bresniker, Paolo Faraboschi, Avi Mendelson, Dejan Milojicic, Timothy Roscoe, and Robert N. M. Watson. Rack-scale capabilities: Finegrained protection for large-scale memories. Computer, 52(2):52–62, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8672401/.

Anonymous:2019:IAHa

[2791] Anonymous. IEEE Annals of the History of Computing. Computer, 52 (2):62, February 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pal:2019:NCS

[2792] Amitangshu Pal and Krishna Kant. NFMI: Connectivity for short-range IoT applications. Computer, 52(2): 63–67, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //ieeexplore.ieee.org/document/
                                        8672420/.

Anonymous:2019:IISa

[2793] Anonymous. IEEE Intelligent Systems. Computer, 52(2):67, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IMa

[2794] Anonymous. IEEE Multimedia. Computer, 52(2):67, February 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhn:2019:RDL

[2795] Rick Kuhn, Dylan Yaga, and Jeffrey Voas. Rethinking distributed ledger technology. Computer, 52 (2):68–72, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //ieeexplore.ieee.org/document/
                                        8672407/.

Anonymous:2019:FCb

[2796] Anonymous. Front cover. Computer, 52(2):C1, February 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:BRRb

[2797] Anonymous. B. Ramakrishna Rau Award. Computer, 52(2):C2, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSd

[2798] Anonymous. IEEE Computer Society information. Computer, 52 (2):C3, February 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8672409/.

Anonymous:2019:ICSe

[2799] Anonymous. IEEE Computer Society job boards. Computer, 52(2): C4, February 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TCc

[2800] Anonymous. Table of contents. Computer, 52(3):1–2, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Mc

[2801] Anonymous. Masthead. Computer, 52(3):3, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8677348/.

Anonymous:2019:TCd

[2802] Anonymous. Table of contents. Computer, 52(3):1–2, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Md

[2803] Anonymous. Masthead. Computer, 52(3):3, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zomaya:2019:NHT

[2804] Albert Y. Zomaya. A novel hierarchical two-tier node-deployment strategy for sustainable wireless sensor networks. Computer, 52(3):4–5, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CHSa

[2805] Anonymous. Computer highlights Society magazines. Computer, 52(3): 6–8, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPCc

[2806] Anonymous. IT Professional call for articles. Computer, 52(3):8, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPCd

[2807] Anonymous. IT Professional call for articles. Computer, 52(3):8, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2019:YAc

[2808] Erich Neuhold. 50 & 25 years ago. Computer, 52(3):9–11, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:LLI

[2809] David Alan Grier. Learning to live in an insecure world. Computer, 52(3):12–13, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8677335/.

Anonymous:2019:SPS

[2810] Anonymous. Security & Privacy subscribe. Computer, 52(3):13, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sobel:2019:CFC

[2811] Ann Sobel, Allen Parrish, and Rajendra K. Raj. Curricular foundations for cybersecurity. Computer, 52 (3):14–17, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITBc

[2812] Anonymous. IEEE Transactions on Big Data. Computer, 52(3):17, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hudnall:2019:EWC

[2813] Matthew Hudnall. Educational and workforce cybersecurity frameworks: Comparing, contrasting, and mapping. Computer, 52(3):18–28, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8677327/.

Burley:2019:CCB

[2814] Diana L. Burley and Alfred H. Lewis Jr. Cybersecurity curricula 2017 and Boeing: Linking curricular guidance to professional practice. Computer, 52(3):29–37, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8678341/.

Gibson:2019:AUC

[2815] David Gibson, Vijay Anand, Josh Dehlinger, Charles Dierbach, Tracy Emmersen, and Andrew Phillips. Accredited undergraduate cybersecurity degrees: Four approaches. Computer, 52(3):38–47, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8677345/.

Anonymous:2019:ITSb

[2816] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 52 (3):47, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITSc

[2817] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 52 (3):47, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Emmersen:2019:UIA

[2818] Tracy Emmersen, Joseph M. Hatfield, Jeff Kosseff, and Steven R. Orr IV. The USNA’s interdisciplinary approach to cybersecurity education. Computer, 52(3):48–57, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8677342/.

Blair:2019:EFM

[2819] Jean R. S. Blair, Andrew O. Hall, and Edward Sobiesk. Educating future multidisciplinary cybersecurity teams. Computer, 52(3):58–66, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8677338/.

Ivy:2019:SCW

[2820] Jessica Ivy, Sarah B. Lee, Dana Franz, and Joseph Crumpton. Seeding cybersecurity workforce pathways with secondary education. Computer, 52 (3):67–75, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/document/
8677341/.

Anonymous:2019:IPCe

[2821] Anonymous. IEEE Pervasive Computing call for articles. Computer, 52 (3):75, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPCf

[2822] Anonymous. IEEE Pervasive Computing call for articles. Computer, 52 (3):75, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2019:CEP

[2823] Hal Berghel. Codes of ethics in a posttruth world. Computer, 52(3):76–80, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8677356/.

Charette:2019:EKH

[2824] Robert N. Charette. Evolving from know-how to know-what. Computer, 52(3):81–86, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //ieeexplore.ieee.org/document/
                                        8677333/.

Anonymous:2019:IAHb

[2825] Anonymous. IEEE Annals of the History of Computing. Computer, 52 (3):86, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IAHc

[2826] Anonymous. IEEE Annals of the History of Computing. Computer, 52 (3):86, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2019:TPS

[2827] Nir Kshetri and Jeffrey Voas. Trusting pirated software. Computer, 52 (3):87–90, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //ieeexplore.ieee.org/document/
                                        8677353/.

Anonymous:2019:FCc

[2828] Anonymous. Front cover. Computer, 52(3):C1, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:FCd

[2829] Anonymous. Front cover. Computer, 52(3):C1, March 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:LCSa

[2830] Anonymous. Letters of the Computer Society. Computer, 52(3):C2, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:LCSb

[2831] Anonymous. Letters of the Computer Society. Computer, 52(3):C2, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSf

[2832] Anonymous. IEEE Computer Society information. Computer, 52(3):C3, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://ieeexplore.
ieee.org/document/8677313/.

Anonymous:2019:ICSg

[2833] Anonymous. IEEE Computer Society jobs board. Computer, 52(3):C4, March 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TCe

[2834] Anonymous. Table of contents. Computer, 52(4):1–2, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Me

[2835] Anonymous. Masthead. Computer, 52(4):3, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CHSb

[2836] Anonymous. Computer highlights Society magazines. Computer, 52(4):4–6, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISPb

[2837] Anonymous. IEEE Software & Privacy. Computer, 52(4):6, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2019:YAd

[2838] E. Neuhold. 50 & 25 years ago. Computer, 52(4):7–9, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISCb

[2839] Anonymous. IEEE Software call for articles. Computer, 52(4):9, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:PAG

[2840] D. A. Grier. The path across the great deep. Computer, 52(4):10–11, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITBd

[2841] Anonymous. IEEE Transactions on Big Data. Computer, 52(4):11, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nag:2019:NAH

[2842] N. Nag and R. Jain. A navigational approach to health: Actionable guidance for improved quality of life. Computer, 52(4):12–20, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Such:2019:BCH

[2843] J. M. Such, P. Ciholas, A. Rashid, J. Vidler, and T. Seabrook. Basic cyber hygiene: Does it work? Computer, 52(4):21–31, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITSd

[2844] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 52(4):31, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFrancesco:2019:GDP

[2845] G. P. De Francesco. The General Data Protection Regulation’s practical impact on software architecture. Computer, 52(4):32–39, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fawaz:2019:SPI

[2846] K. Fawaz and K. G. Shin. Security and privacy in the Internet of Things. Computer, 52(4):40–49, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IAHd

[2847] Anonymous. IEEE Annals of the History of Computing. Computer, 52 (4):49, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Farahmand:2019:PDM

[2848] F. Farahmand and F. Farahmand. Privacy decision making: The brain approach. Computer, 52(4):50–58, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICGb

[2849] Anonymous. IEEE Computer Graphics & Applications subscribe. Computer, 52(4):58, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2019:IOS

[2850] D. Riehle. The innovations of open source. Computer, 52(4):59–63, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPCg

[2851] Anonymous. IT Professional call for articles. Computer, 52(4):63, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2019:TC

[2852] N. Kshetri and J. Voas. Thoughts on cyberbullying. Computer, 52(4): 64–68, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPCh

[2853] Anonymous. IEEE Persuasive Computing call for articles. Computer, 52(4):68, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Havens:2019:PSC

[2854] J. C. Havens and A. Hessami. From principles and standards to certification. Computer, 52(4):69–72, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Ea

[2855] Anonymous. Erratum. Computer, 52(4):72, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IMb

[2856] Anonymous. IEEE MultiMedia. Computer, 52(4):72, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:FCe

[2857] Anonymous. Front cover. Computer, 52(4):C1, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IWCa

[2858] Anonymous. IEEE World Congress on Services 2019. Computer, 52 (4):C2, April 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSh

[2859] Anonymous. IEEE Computer Society information. Computer, 52(4):C3, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:C

[2860] Anonymous. COMPSAC 2019. Computer, 52(4):C4, April 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TCf

[2861] Anonymous. Table of contents. Computer, 52(5):1–2, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Mf

[2862] Anonymous. Masthead. Computer, 52(5):3, May 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CHSc

[2863] Anonymous. Computer highlights Society magazines. Computer, 52(5):4–6, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IWCb

[2864] Anonymous. IEEE World Congress on Services 2019. Computer, 52(5):6, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2019:YAe

[2865] E. Neuhold. 50 & 25 years ago. Computer, 52(5):7–10, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CGA

[2866] Anonymous. Computer Graphics & Applications subscribe. Computer, 52(5):10, May 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:FV

[2867] D. A. Grier. A familiar vision. Computer, 52(5):11–12, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IAHe

[2868] Anonymous. IEEE Annals of the History of Computing. Computer, 52(5):12, May 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gudivada:2019:CCS

[2869] V. N. Gudivada, S. Pankanti, G. Seetharaman, and Y. Zhang. Cognitive computing systems: Their potential and the future. Computer, 52(5):13–18, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPCi

[2870] Anonymous. IEEE Pervasive Computing call for articles. Computer, 52(5):18, May 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:LCSc

[2871] Anonymous. Letters of the Computer Society. Computer, 52(5):19, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeBole:2019:TAZ

[2872] M. V. DeBole, B. Taba, A. Amir, F. Akopyan, A. Andreopoulos, W. P. Risk, J. Kusnitz, C. Ortega Otero, T. K. Nayak, R. Appuswamy, P. J. Carlson, A. S. Cassidy, P. Datta, S. K. Esser, G. J. Garreau, K. L. Holland, S. Lekuch, M. Mastro, J. McKinstry, C. di Nolfo, B. Paulovicks, J. Sawada, K. Schleupen, B. G. Shaw, J. L. Klamo, M. D. Flickner, J. V. Arthur, and D. S. Modha. TrueNorth: Accelerating from zero to 64 million neurons in 10 years. Computer, 52(5):20–29, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pyle:2019:LSS

[2873] S. D. Pyle, J. D. Sapp, and R. F. De-Mara. Leveraging stochasticity for in situ learning in binarized deep neural networks. Computer, 52(5):30–39, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:SSP

[2874] Anonymous. Subscribe to Security & Privacy magazine. Computer, 52 (5):39, May 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rasiwasia:2019:PBA

[2875] N. Rasiwasia. Perspectives on becoming an applied machine learning scientist. Computer, 52(5):40–47, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ma:2019:HEM

[2876] Z. Ma, J. Wu, S. Zhong, J. Jiang, and S. J. Heinen. Human eye movements reveal video frame importance. Computer, 52(5):48–57, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hwang:2019:BBR

[2877] S. O. Hwang and A. Mehmood. Blockchain-based resource syndicate. Computer, 52(5):58–66, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rinner:2019:CWT

[2878] B. Rinner. Can we trust smart cameras? Computer, 52(5):67–70, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITSe

[2879] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 52(5):70, May 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:SPMa

[2880] Anonymous. Security & Privacy magazine. Computer, 52(5):71, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morman:2019:HDW

[2881] L. M. Morman. How do we prepare the next generation for a career in our digital era? Computer, 52(5):72–74, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IISb

[2882] Anonymous. IEEE Intelligent Systems magazine. Computer, 52(5):74, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IMMa

[2883] Anonymous. IEEE MultiMedia magazine. Computer, 52(5):74, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITBe

[2884] Anonymous. IEEE Transactions on Big Data. Computer, 52(5):75, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CCa

[2885] Anonymous. CS connection. Computer, 52(5):76, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:FCf

[2886] Anonymous. Front cover. Computer, 52(5):C1, May 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSi

[2887] Anonymous. IEEE Computer Society Website launch. Computer, 52 (5):C2, May 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSj

[2888] Anonymous. IEEE Computer Society information. Computer, 52(5):C3, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IJBa

[2889] Anonymous. IEEE Jobs Board. Computer, 52(5):C4, May 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TCg

[2890] Anonymous. Table of contents. Computer, 52(6):1–2, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Mg

[2891] Anonymous. Masthead. Computer, 52(6):3, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaidya:2019:AHT

[2892] J. Vaidya. Advances in Hardware Trojan detection. Computer, 52(6):4–5, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITSf

[2893] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 52(6):5, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CHSd

[2894] Anonymous. Computer highlights Society magazines. Computer, 52(6):6–8, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISPc

[2895] Anonymous. IEEE Security & Privacy magazine subscribe. Computer, 52(6):8, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2019:YAf

[2896] E. Neuhold. 50 & 25 years ago. Computer, 52(6):9–11, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISMa

[2897] Anonymous. IEEE Software magazine call for articles. Computer, 52 (6):11, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:ECS

[2898] D. A. Grier. Exploring the current state of quantum computing. Computer, 52(6):12, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Humble:2019:QR

[2899] T. S. Humble and E. P. DeBenedictis. Quantum realism. Computer, 52 (6):13–17, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITBf

[2900] Anonymous. IEEE Transactions on Big Data. Computer, 52(6):17, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shaydulin:2019:HAS

[2901] R. Shaydulin, H. Ushijima-Mwesigwa, C. F. A. Negre, I. Safro, S. M. Mniszewski, and Y. Alexeev. A hybrid approach for solving optimization problems on small quantum computers. Computer, 52(6):18–26, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paler:2019:RSS

[2902] A. Paler, D. Herr, and S. J. Devitt. Really small shoe boxes: On realistic quantum resource estimation. Computer, 52(6):27–37, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPCj

[2903] Anonymous. IEEE Pervasive Computing call for articles. Computer, 52(6):37, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McGeoch:2019:PAB

[2904] C. C. McGeoch, R. Harris, S. P.

Reinhardt, and P. I. Bunyk. Practical annealing-based quantum computing. Computer, 52(6):38–46, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hu:2019:RBP

[2905] S. Hu, P. Liu, C. R. Chen, M. Pistoia, and J. Gambetta. Reduction-based problem mapping for quantum computing. Computer, 52(6):47–57, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2019:SOQ

[2906] P. Liu, S. Hu, M. Pistoia, C. R. Chen, and J. M. Gambetta. Stochastic optimization of quantum programs. Computer, 52(6):58–67, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IIPa

[2907] Anonymous. IEEE IT Professionals call for articles. Computer, 52 (6):67, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurlburt:2019:TU

[2908] G. F. Hurlburt. The tyranny of urgency. Computer, 52(6):68–72, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICGc

[2909] Anonymous. IEEE Computer Graphics & Applications magazine. Computer, 52(6):72, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2019:NIT

[2910] P. Laplante and S. Applebaum. NIST’s 18 Internet of Things trust concerns. Computer, 52(6):73–76, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:SN

[2911] Anonymous. Social networking. Computer, 52(6):77, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serpanos:2019:TNS

[2912] D. Serpanos. There is no safety without security and dependability. Computer, 52(6):78–81, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IAHf

[2913] Anonymous. IEEE Annals of the History of Computing. Computer, 52 (6):81, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ballhausen:2019:FOS

[2914] M. Ballhausen. Free and open source software licenses explained. Computer, 52(6):82–86, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IISc

[2915] Anonymous. IEEE Intelligent Systems magazine. Computer, 52(6):86, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IJBb

[2916] Anonymous. IEEE Jobs Board. Computer, 52(6):87, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wright:2019:MIS

[2917] D. Wright. Meeting industry’s standardization needs. Computer, 52(6): 88–92, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IMMb

[2918] Anonymous. IEEE MultiMedia magazine. Computer, 52(6):92, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ILC

[2919] Anonymous. IEEE letters of the Computer Society. Computer, 52(6):93, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CCb

[2920] Anonymous. CS connections. Computer, 52(6):94–96, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:FCg

[2921] Anonymous. Front cover. Computer, 52(6):C1, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IWCc

[2922] Anonymous. IEEE World Congress on Services 2019. Computer, 52 (6):C2, June 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSk

[2923] Anonymous. IEEE Computer Society information. Computer, 52(6):C3, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:SPMb

[2924] Anonymous. Security & Privacy magazine. Computer, 52(6):C4, June 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TCh

[2925] Anonymous. Table of contents. Computer, 52(7):1–2, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Mh

[2926] Anonymous. Masthead. Computer, 52(7):3, July 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CHSe

[2927] Anonymous. Computer highlights Society magazines. Computer, 52(7):4–6, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2019:YAg

[2928] E. Neuhold. 50 & 25 years ago. Computer, 52(7):7–9, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:TF

[2929] D. A. Grier. Transforming the field.

Computer, 52(7):10–11, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITSg

[2930] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 52(7):11, July 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Youssef:2019:TCC

[2931] M. Youssef and F. Kawsar. Transformative computing and communication. Computer, 52(7):12–14, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IAHg

[2932] Anonymous. IEEE Annals of the History of Computing. Computer, 52(7):14, July 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSl

[2933] Anonymous. IEEE Computer Society Technical Consortium on High Performance Computing. Computer, 52(7):15, July 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alloulah:2019:FMW

[2934] M. Alloulah and H. Huang. Future millimeter-wave indoor systems: A blueprint for joint communication and sensing. Computer, 52(7):16–24, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Savazzi:2019:USW

[2935] S. Savazzi, S. Sigg, F. Vicentini, S. Kianoush, and R. Findling. On the use of stray wireless signals for sensing: a look beyond 5G for the next generation of industry. Computer, 52 (7):25–36, July 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISCc

[2936] Anonymous. IEEE Software call for articles. Computer, 52(7):36, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Satyanarayanan:2019:ACT

[2937] M. Satyanarayanan and N. Davies. Augmenting cognition through edge computing. Computer, 52(7):37–46, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IMMc

[2938] Anonymous. IEEE Multimedia magazine. Computer, 52(7):46, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moreno:2019:HAS

[2939] L. Moreno and P. Martinez. The harmonization of accessibility standards for public policies. Computer, 52 (7):57–66, July 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITBg

[2940] Anonymous. IEEE Transactions on Big Data. Computer, 52(7):66, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowker:2019:CTC

[2941] G. C. Bowker. The computerization time crunch. Computer, 52 (7):67–71, July 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISPd

[2942] Anonymous. IEEE Security & Privacy subscription. Computer, 52(7):71, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khan:2019:SCS

[2943] M. T. Khan and H. Shrobe. Security of cyberphysical systems: Chaining induction and deduction. Computer, 52 (7):72–75, July 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roth:2019:STS

[2944] J. D. Roth, C. E. Lutton, and J. B.

Michael. Security through simplicity: a case study in logical segmentation inference. Computer, 52(7):76–79, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CCc

[2945] Anonymous. Cs connections. Computer, 52(7):80–83, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPCk

[2946] Anonymous. IT Professional call for papers. Computer, 52(7):83, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSm

[2947] Anonymous. IEEE Computer Society information. Computer, 52(7):84, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:FCh

[2948] Anonymous. Front cover. Computer, 52(7):C1, July 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSn

[2949] Anonymous. IEEE Computer Society Job Board. Computer, 52(7):C2, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSo

[2950] Anonymous. IEEE Computer Society social networking. Computer, 52 (7):C3, July 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICEa

[2951] Anonymous. IEEE Computing Edge magazine. Computer, 52(7):C4, July 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:FCi

[2952] Anonymous. Front cover. Computer, 52(8):C1, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IJBc

[2953] Anonymous. IEEE Job Boards. Computer, 52(8):C2, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TCi

[2954] Anonymous. Table of contents. Computer, 52(8):1–2, August 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Mi

[2955] Anonymous. Masthead. Computer, 52(8):3, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mueller:2019:AVR

[2956] K. Mueller. Advances in visualization recommender systems. Computer, 52 (8):4–5, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CHSf

[2957] Anonymous. Computer highlights Society magazines. Computer, 52(8): 6–8, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2019:YAh

[2958] E. Neuhold. 50 & 25 years ago.

Computer, 52(8):9–11, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2019:CTTa

[2959] E. Akleman. Computing through time. Computer, 52(8):10, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:GNG

[2960] D. A. Grier. Go/no go. Computer, 52 (8):12–13, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IAHh

[2961] Anonymous. IEEE Annals of the History of Computing. Computer, 52 (8):13, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Villari:2019:OOC

[2962] M. Villari, M. Fazio, S. Dustdar, O. Rana, D. N. Jha, and R. Ranjan. Osmosis: The osmotic computing platform for microelements in the cloud, edge, and Internet of Things. Computer, 52(8):14–26, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

SilvaBorges:2019:HDD

[2963] H. Silva Borges and M. T. Valente. How do developers promote open source projects? Computer, 52(8):27–33, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Figueroa:2019:DVL

[2964] A. Figueroa and J. Atkinson. Dualview learning for detecting web query intents. Computer, 52(8):34–42, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:2019:CRI

[2965] H. Kim, E. A. Lee, and S. Dustdar.

Creating a resilient IoT with edge computing. Computer, 52(8):43–53, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mcgraw:2019:SEM

[2966] G. Mcgraw, R. Bonett, H. Figueroa, and V. Shepardson. Security engineering for machine learning. Computer, 52 (8):54–57, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IMMd

[2967] Anonymous. IEEE Multimedia magazine. Computer, 52(8):57, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

HabibiGharakheili:2019:NLS

[2968] H. Habibi Gharakheili, A. Sivanathan, A. Hamza, and V. Sivaraman. Network-level security for the Internet of Things: Opportunities and challenges. Computer, 52(8):58–62, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISMb

[2969] Anonymous. IEEE Software magazine call for articles. Computer, 52 (8):62, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schoettle:2019:OSL

[2970] H. Schoettle. Open source license compliance — why and how? Computer, 52 (8):63–67, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICGd

[2971] Anonymous. IEEE Computer Graphics applications call for articles. Computer, 52(8):67, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Simpson:2019:ISA

[2972] R. Simpson. The IEEE Standards Association Corporate Program: Engaging industry in standards development. Computer, 52(8):68–71, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IIPb

[2973] Anonymous. IEEE IT Professionals call for articles. Computer, 52 (8):71, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hole:2019:BDA

[2974] K. J. Hole and S. Ahmad. Biologically driven artificial intelligence. Computer, 52(8):72–75, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPCl

[2975] Anonymous. IEEE Pervasive Computing call for articles. Computer, 52 (8):75, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICM

[2976] Anonymous. IEEE Computer magazine call for 2019 major awards nominations. Computer, 52(8):76, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSp

[2977] Anonymous. IEEE Computer Society election. Computer, 52(8):77–90, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISPe

[2978] Anonymous. IEEE Security & Privacy subscribe. Computer, 52(8):90, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IDV

[2979] Anonymous. IEEE Division VIII delegate/director candidates. Computer, 52 (8):91–92, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPE

[2980] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 52(8):93–96, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSq

[2981] Anonymous. IEEE Computer Society information. Computer, 52(8): C3, August 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICEb

[2982] Anonymous. IEEE Computing Edge subscribe. Computer, 52(8):C4, August 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:FCj

[2983] Anonymous. Front cover. Computer, 52 (9):C1, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CCSa

[2984] Anonymous. Computer call for special issue proposals. Computer, 52(9): C2, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TCj

[2985] Anonymous. Table of contents. Computer, 52(9):1–2, September 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Mj

[2986] Anonymous. Masthead. Computer, 52 (9):3, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fettes:2019:DVF

[2987] Q. Fettes, M. Clark, R. Bunescu, A. Karanth, and A. Louri. Dynamic voltage and frequency scaling in NoCs with supervised and reinforcement learning techniques. Computer, 52(9):4–5, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CHSg

[2988] Anonymous. Computer highlights Society magazines. Computer, 52(9):

6–8, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IMMe

[2989] Anonymous. IEEE Multimedia Magazine. Computer, 52(9):8, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2019:YAi

[2990] E. Neuhold. 50 & 25 years ago. Computer, 52(9):9–11, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:TCW

[2991] D. A. Grier. Translational computing:

What needs the translation? Computer, 52(9):12–14, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSr

[2992] Anonymous. 2019 IEEE Computer Society election. Computer, 52(9): 14, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:RMA

[2993] Anonymous. Richard E. Merwin Award. Computer, 52(9):15, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abramson:2019:TRC

[2994] D. Abramson and M. Parashar. Translational research in computer science. Computer, 52(9):16–23, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IIPc

[2995] Anonymous. IEEE IT Professional call for papers. Computer, 52(9): 23, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhn:2019:TSS

[2996] R. Kuhn, J. Voas, and P. Laplante. Taking score on the success of blockchain, so far. Computer, 52(9): 24–34, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPCm

[2997] Anonymous. IEEE Pervasive Computing call for articles. Computer, 52 (9):34, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zamansky:2019:LMD

[2998] A. Zamansky, D. van der Linden, I. Hadar, and S. Bleuer-Elsner. Log my dog: Perceived impact of dog activity tracking. Computer, 52(9):35– 43, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITBh

[2999] Anonymous. IEEE Transactions on Big Data. Computer, 52(9):43, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Salisbury:2019:WCA

[3000] M. Salisbury. When computers advise us: How to represent the types of knowledge users seek for expert advice. Computer, 52(9):44–51, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSs

[3001] Anonymous. 2019 IEEE Computer Society election. Computer, 52(9): 51, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hukal:2019:BCW

[3002] P. Hukal, N. Berente, M. Germonprez, and A. Schecter. Bots coordinating work in open source software projects. Computer, 52(9):52– 60, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huang:2019:ERB

[3003] C. Huang, T. Wang, and C. Lin. Early results on the business value of Internet of Things initiatives. Computer, 52(9): 61–67, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Varghese:2019:CF

[3004] B. Varghese, P. Leitner, S. Ray, K. Chard, A. Barker, Y. Elkhatib, H. Herry, C. Hong, J. Singer, F. P. Tso, E. Yoneki, and M. Zhani. Cloud futurology. Computer, 52(9):68–77, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fournaris:2019:GAN

[3005] A. P. Fournaris, A. S. Lalos, and D. Serpanos. Generative adversarial networks in AI-enabled safety-critical systems: Friend or foe? Computer, 52 (9):78–81, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bloomfield:2019:DID

[3006] R. Bloomfield, H. Khlaaf, P. Ryan Conmy, and G. Fletcher. Disruptive innovations and disruptive assurance: Assuring machine learning and autonomy. Computer, 52(9):82– 89, September 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISNa

[3007] Anonymous. IEEE social networking. Computer, 52(9):90, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CCd

[3008] Anonymous. CS connection. Computer, 52(9):91–92, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSt

[3009] Anonymous. IEEE Computer Society. Computer, 52(9):C3, September 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IJBd

[3010] Anonymous. IEEE Jobs Board. Computer, 52(9):C4, September 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:FCk

[3011] Anonymous. [Front cover]. Computer, 52(10):C1, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICAa

[3012] Anonymous. IEEE Computer Architecture Letters. Computer, 52(10): C2, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TCk

[3013] Anonymous. Table of contents. Computer, 52(10):1–2, October 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Mk

[3014] Anonymous. Masthead. Computer, 52 (10):3, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CHSh

[3015] Anonymous. Computer highlights Society magazines. Computer, 52(10): 4–6, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Montuschi:2019:NET

[3016] P. Montuschi. New and emerging topics in computing. Computer, 52 (10):7–8, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IISd

[3017] Anonymous. IEEE Intelligent Systems. Computer, 52(10):8, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:YAa

[3018] Anonymous. 50 25 years ago. Computer, 52(10):9–11, October 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2019:CTTb

[3019] E. Akleman. Computing through time. Computer, 52(10):11, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:GP

[3020] D. A. Grier. A good party. Computer, 52(10):12–13, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Crocker:2019:AIS

[3021] S. D. Crocker. The Arpanet and its impact on the state of networking. Computer, 52(10):14–23, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICGe

[3022] Anonymous. IEEE Computer Graphics Applications. Computer, 52(10): 23, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stimers:2019:UNQ

[3023] P. Stimers. The U.S. National Quantum Initiative. Computer, 52(10): 24–29, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lu:2019:SWT

[3024] Y. Lu, G. K. Thiruvathukal, A. S.

Kaseb, K. Gauen, D. Rijhwani, R. Dailey, D. Malik, Y. Huang, S. Aghajanzadeh, and M. M. Guo. See the world through network cameras. Computer, 52(10):30–40, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IIP

[3025] Anonymous. IEEE IT Professionals call for articles. Computer, 52 (10):40, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jukan:2019:FCC

[3026] A. Jukan, F. Carpio, X. Masip, A. J.

Ferrer, N. Kemper, and B. U. Stetina. Fog-to-cloud computing for farming: Low-cost technologies, data exchange, and animal welfare. Computer, 52(10): 41–51, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bertini:2019:CIH

[3027] F. Bertini, S. G. Rizzo, and D. Montesi. Can information hiding in social media posts represent a threat? Computer, 52(10):52–60, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IPC

[3028] Anonymous. IEEE Pervasive Computing call for papers. Computer, 52 (10):60, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:2019:STC

[3029] X. Wang, W. Sun, L. Hu, Y. Zhao, W. E. Wong, and Z. Chen. Softwaretesting contests: Observations and lessons learned. Computer, 52(10): 61–69, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISPf

[3030] Anonymous. IEEE Security & Privacy magazine subscribe. Computer, 52 (10):69, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:2019:AIA

[3031] T. W. Kim and S. Mejia. From artificial intelligence to artificial wisdom: What Socrates teaches us. Computer, 52(10): 70–74, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ID

[3032] Anonymous. IEEE DataPort. Computer, 52(10):75, October 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vassilev:2019:CDL

[3033] A. Vassilev. Can deep learning save us and itself from the avalanche of threats in cyberspace? Computer, 52(10): 76–82, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CCSb

[3034] Anonymous. Computer call for special issue proposals. Computer, 52 (10):83, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:2019:SSI

[3035] P. Zhang, S. Pan, M. Mirshekari, J. Fagert, and H. Y. Noh. Structures as sensors: Indirect sensing for inferring users and environments. Computer, 52(10):84–88, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITBi

[3036] Anonymous. IEEE Transactions on Big Data. Computer, 52(10):88, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2019:SDH

[3037] M. Campbell. Synthetic data: How AI is transitioning from data consumer to data producer … and why that’s important. Computer, 52(10): 89–91, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McAffer:2019:GSO

[3038] J. McAffer. Getting started with open source governance. Computer, 52(10): 92–96, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITSh

[3039] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 52 (10):96, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sobel:2019:NEV

[3040] A. E. K. Sobel. A new (?) educational view of computing. Computer, 52(10): 97–98, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISMc

[3041] Anonymous. IEEE Software magazine call for articles. Computer, 52 (10):98, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2019:HBI

[3042] H. Berghel. Huawei, BIS, and the IEEE: It’s déjà vu all over again. Computer, 52(10):99–104, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISNb

[3043] Anonymous. IEEE social networking. Computer, 52(10):105, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thompson:2019:EOD

[3044] G. Thompson. Ethernet: From office to data center to IoT. Computer, 52(10): 106–109, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IAHi

[3045] Anonymous. IEEE Annals of the History of Computing. Computer, 52 (10):109, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeBenedictis:2019:PQS

[3046] E. P. DeBenedictis. Powerball and quantum supremacy. Computer, 52 (10):110–112, October 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSu

[3047] Anonymous. IEEE Computer Society information. Computer, 52(10): C3, October 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IJBe

[3048] Anonymous. IEEE Job Board. Computer, 52(10):C4, October 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:FCl

[3049] Anonymous. Front cover. Computer, 52 (11):C1, November 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISNc

[3050] Anonymous. IEEE social networking.

Computer, 52(11):C2, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TCl

[3051] Anonymous. Table of contents. Computer, 52(11):1–2, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Ml

[3052] Anonymous. Masthead. Computer, 52 (11):3, November 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:YAb

[3053] Anonymous. 50 25 years ago. Computer, 52(11):4–6, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IMMf

[3054] Anonymous. IEEE Multimedia Magazine. Computer, 52(11):6, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parashar:2019:RI

[3055] M. Parashar. The reproducibility initiative. Computer, 52(11):7–8, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITBj

[3056] Anonymous. IEEE Transactions on Big Data. Computer, 52(11):8, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:EC

[3057] Anonymous. Elsewhere in the CS. Computer, 52(11):9–11, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IISe

[3058] Anonymous. IEEE Intelligent Systems Magazine. Computer, 52(11): 11, November 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:THL

[3059] D. A. Grier. Toward healthier lifestyles.

Computer, 52(11):12–13, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISPg

[3060] Anonymous. IEEE Security and Privacy Magazine subscribe. Computer, 52 (11):13, November 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Helal:2019:DHA

[3061] S. Helal and R. Jain. Digital healthactive and healthy living. Computer, 52 (11):14–17, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gadaleta:2019:EDR

[3062] M. Gadaleta, M. Rossi, E. J. Topol, S. R. Steinhubl, and G. Quer. On the effectiveness of deep representation learning: The atrial fibrillation case. Computer, 52(11):18–29, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Homdee:2019:AMP

[3063] N. Homdee, R. Alam, J. A. Hayes, T. Hamid, J. Park, S. Wolfe, H. Goins, N. Fyffe, T. Newbold, T. Smith-Jackson, A. Bankole, M. S. Anderson, and J. Lach. Agitation monitoring and prevention system for dementia caregiver empowerment. Computer, 52(11): 30–39, November 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tseng:2019:CBA

[3064] C. Tseng, C. Lin, H. Chang, C. Liu, B. M. F. Serafico, L. Wu, C. Lin, T. Hsu, C. Huang, and M. Lo. Cloudbased artificial intelligence system for large-scale arrhythmia screening. Computer, 52(11):40–51, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:2019:AIT

[3065] W. C. Chu, C. Shih, W. Chou, S. I. Ahamed, and P. Hsiung. Artificial intelligence of things in sports science: Weight training as an example. Computer, 52(11):52–61, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pinarer:2019:AAD

[3066] O. Pinarer and A. Ozgovde. Applicationaware dynamic energy management for portable devices. Computer, 52(11): 62–72, November 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shiroishi:2019:BAS

[3067] Y. Shiroishi, K. Uchiyama, and N. Suzuki. Better actions for Society 5.0: Using AI for evidence-based policy making that keeps humans in the loop. Computer, 52(11):73–78, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICAb

[3068] Anonymous. IEEE Computer Architecture Letters. Computer, 52(11): 79, November 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2019:ATE

[3069] J. B. Michael. Assessing the trustworthiness of electronic systems. Computer, 52(11):80–83, November 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITSi

[3070] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 52 (11):83, November 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CC

[3071] Anonymous. CS connection. Computer, 52(11):84, November 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISMd

[3072] Anonymous. IEEE Software Magazine call for articles. Computer, 52 (11):84, November 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSv

[3073] Anonymous. IEEE Computer Society information. Computer, 52(11): C3, November 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IJBf

[3074] Anonymous. IEEE Jobs Board. Computer, 52(11):C4, November 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:FC

[3075] Anonymous. [Front cover]. Computer, 52(12):C1, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITC

[3076] Anonymous. IEEE Transactions on Computers. Computer, 52(12):C2, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:TC

[3077] Anonymous. Table of contents. Computer, 52(12):1–2, December 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:M

[3078] Anonymous. Masthead. Computer, 52 (12):3, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Metra:2019:ICSb

[3079] C. Metra. The 2019 IEEE Computer Society: Hit target on member satisfaction and technical excellence. Computer, 52(12):4–11, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSw

[3080] Anonymous. IEEE Computer Society. Computer, 52(12):11, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:CHSi

[3081] Anonymous. Computer highlights Society magazines. Computer, 52(12):

12–14, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IIC

[3082] Anonymous. IEEE Internet Computing magazine. Computer, 52(12): 15, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lei:2019:WBE

[3083] H. Lei, C. Qiu, H. Yao, and S. Guo.

When blockchain-enabled Internet of Things meets cloud computing. Computer, 52(12):16–17, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISM

[3084] Anonymous. IEEE Software magazine call for articles. Computer, 52 (12):17, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2019:YA

[3085] E. Neuhold. 50 & 25 years ago. Computer, 52(12):18–21, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IMMg

[3086] Anonymous. IEEE Multimedia magazine. Computer, 52(12):19, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2019:CTT

[3087] E. Akleman. Computing through time.

Computer, 52(12):20, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITS

[3088] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 52 (12):21, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSx

[3089] Anonymous. IEEE Computer Society Awards. Computer, 52(12):22– 27, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2019:ICS

[3090] D. A. Grier. The IEEE Computer Society and the tools of the past. Computer, 52(12):28–30, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:IIS

[3091] Anonymous. IEEE Intelligent Systems magazine. Computer, 52(12): 30, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2019:TP

[3092] D. Milojicic. Technology predictions. Computer, 52(12):31–33, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Faraboschi:2019:TPA

[3093] P. Faraboschi, E. Frachtenberg, P. Laplante, K. Mansfield, and D. Milojicic. Technology predictions: Art, science, and fashion. Computer, 52(12): 34–38, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lange:2019:CRM

[3094] D. Lange. Cognitive robotics: Making robots sense, understand, and interact. Computer, 52(12):39–44, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bresniker:2019:GCA

[3095] K. Bresniker, A. Gavrilovska, J. Holt, D. Milojicic, and T. Tran. Grand challenge: Applying artificial intelligence and machine learning to cybersecurity. Computer, 52(12):45– 52, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Frachtenberg:2019:PDD

[3096] E. Frachtenberg. Practical drone delivery. Computer, 52(12):53–57, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Saracco:2019:DTB

[3097] R. Saracco. Digital twins: Bridging physical space and cyberspace. Computer, 52(12):58–64, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mendelson:2019:SPA

[3098] A. Mendelson. Security and privacy in the age of big data and machine learning. Computer, 52(12):65– 70, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coughlin:2019:NME

[3099] T. Coughlin. Nonvolatile memory express: The link that binds them. Computer, 52(12):71–79, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2019:PP

[3100] J. Voas. Predictions on predictions. Computer, 52(12):80–82, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoffman:2019:MHC

[3101] A. Hoffman. Moral hazards in cyber vulnerability markets. Computer, 52 (12):83–88, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Black:2019:OWP

[3102] P. E. Black and M. Singh. Opaque wrappers and patching: Negative results. Computer, 52(12):89–93, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pal:2019:UBP

[3103] A. Pal and K. Kant. Using blockchain for provenance and traceability in Internet of Things-integrated food logistics. Computer, 52(12):94–98, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2019:SEE

[3104] M. Campbell. Smart edge: The effects of shifting the center of data gravity out of the cloud. Computer, 52(12):

99–102, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spinellis:2019:HSO

[3105] D. Spinellis. How to select open source components. Computer, 52(12):103– 106, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sobel:2019:RWE

[3106] A. E. K. Sobel. Revisiting whether everyone should go to college. Computer, 52(12):107–108, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:Eb

[3107] Anonymous. Erratum. Computer, 52 (12):108, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hough:2019:ISO

[3108] David G. Hough. The IEEE Standard 754: One for the history books. Computer, 52(12):109–112, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://grouper.ieee.org/groups/
msc/ANSI_IEEE-Std-754-2019/background/
ieee-computer.pdf;   https://www.
computer.org/csdl/magazine/co/2019/
12/08909942/1f8KFWxbTCU.

Anonymous:2019:ISPh

[3109] Anonymous. IEEE Security Privacy magazine subscribe. Computer, 52(12): 113, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeBenedictis:2019:MLH

[3110] E. P. DeBenedictis. Moore’s Law: A hard act to follow. Computer, 52(12): 114–117, December 2019. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ITB

[3111] Anonymous. IEEE Transactions on Big Data. Computer, 52(12):117, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:SVI

[3112] Anonymous. Shull voted 2020 IEEE Computer Society President-Elect. Computer, 52(12):118–120, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ICSy

[3113] Anonymous. IEEE Computer Society. Computer, 52(12):C3, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2019:ISNd

[3114] Anonymous. IEEE Social Networking. Computer, 52(12):C4, December 2019. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).